### v1

Abstract
This paper develops a theory of bargaining in firm-to-firm trade with two-sided market power. The framework accommodates flexible market structures, yielding analytical expressions for pair-specific markups and pass-through elasticities. In U.S. import data, we estimate strong importer bargaining power and an upward-sloping export supply curve, consistent with oligopsony power. Pass-through of the 2018 tariffs in firm-to-firm relationships is incomplete, in contrast to product-level studies, primarily due to exporter cost reductions driven by falling demand from dominant buyers. Our study highlights how bargaining and network rigidities shape price outcomes, with implications for markup dispersion and shock propagation in global value chains.
1 Introduction
The recent resurgence of protectionist trade policies has reignited debate over how tariffs affect international prices. This question is particularly salient in today’s global economy, where production is fragmented across global value chains and most trade occurs within ongoing relationships between importers and exporters. These firm-to-firm linkages shape the cross-border transmission of shocks (Boehm et al., 2019); however, their implications for price setting and tariff pass-through remain less well understood.
A growing body of literature highlights the limitations of standard trade models in this context. Rather than engaging in anonymous spot-market transactions, firms often trade within long-term relationships shaped by contract incompleteness, relationship-specific investments, and lock-in frictions (Antràs, 2015). In this environment, prices arise from bilateral negotiations between exporters and importers, both of whom may hold significant market power (Antràs and Staiger, 2012; Morlacco, 2019; Dhyne et al., 2022).
This paper develops a theory of bargaining in firm-to-firm trade that incorporates two-sided market power and network rigidities. The framework accommodates flexible upstream and downstream market structures and delivers closed-form expressions for markups and pass-through elasticities as functions of market shares and a small set of structural parameters. We estimate these parameters using U.S. import data, validate the model’s cross-sectional predictions against observed pricing patterns, and use the estimated framework to quantify aggregate pass-through in firm-to-firm trade, highlighting how bargaining and network structure shape both the level and sources of pass-through in response to trade shocks.
Section 2 introduces the theory. We consider a fixed network of importer-exporter (or buyer-supplier) relationships, where each pair negotiates bilaterally over the price of an intermediate input. Given the price, the importer unilaterally chooses the traded quantity to minimize costs, making prices allocative. In the event of a disagreement, each party falls back on trade with its other existing partners. As a result, the network structure influences the value of each party’s outside option, and in turn, their effective bargaining power.
Our framework offers a unified theory of exporter and importer market power in firm-to-firm trade. Exporter rents arise from three sources: monopoly power through product differentiation, oligopoly power due to supplier concentration, and quasi-rents from an upward-sloping residual supply curve. When exporters hold full bargaining power, the model nests standard trade settings in which prices reflect an oligopoly markup over marginal cost (Atkeson and Burstein, 2008; Dhyne et al., 2022). This markup increases with the exporter’s share of the buyer’s total input purchases—a measure we refer to as the exporter’s supplier share.
Our first contribution is to characterize the role of importer market power in price-setting. This power arises when suppliers earn economic rents, creating scope for buyers to extract surplus by negotiating lower prices. It also reflects oligopsony power, as concentrated buyers can negotiate lower prices. When importers hold full bargaining power, prices can fall below marginal cost, generating a negative price–cost margin akin to an oligopsony markdown. This markdown declines with the importer’s share of the supplier’s total output–a measure we refer to as the importer’s buyer share. [^1] Our definition of the importer’s markdown differs from the standard price–marginal revenue wedge. See Section 2 for details. In general, the negotiated price reflects a convex combination of oligopoly and oligopsony forces, with weights determined by the firms’ outside options and other exogenous determinants of bargaining power.
Our second contribution is to characterize how importer market power shapes the pass-through of tariff shocks to bilateral prices. We focus on short-run pass-through elasticities, derived as local perturbations around the static price equilibrium, while holding the network and general equilibrium conditions fixed. This isolates the intensive margin of price adjustment from longer-run responses driven by entry, reallocation, or macroeconomic feedback.
Importer market power affects pass-through through two distinct mechanisms: a markup channel and a cost channel. The markup channel captures how bilateral markups respond to price changes. When exporters hold full bargaining power, markups exhibit strategic complementarities, a well-known source of incomplete pass-through (Amiti et al., 2014, 2019a). With two-sided market power, the markup elasticity becomes a convex combination of strategic complementarities and strategic substitutabilities, as importers reduce markdowns in response to higher prices, amplifying pass-through.
The cost channel captures how tariffs affect the exporter’s marginal cost through scale effects. While prior work links pass-through to cost adjustment under decreasing returns to scale (e.g., Burstein and Gopinath 2015), our framework introduces a new source of heterogeneity: the elasticity of residual supply increases with an importer’s buyer share. As a result, larger buyers face more incomplete pass-through. Although both channels interact to shape price responses, the cost channel dominates when bargaining power is two-sided and returns to scale are decreasing. In this setting, markup adjustments are modest and offset by opposing strategic forces, while cost-side responses amplify the heterogeneity of pass-through.
Section 3 describes the data and presents preliminary empirical evidence. Our main data source is the U.S. Census Longitudinal Firm Trade Transaction Database (LFTTD), which reports prices and quantities for each import transaction involving a U.S. importer, along with identifiers for the foreign exporter and the 10-digit product code. These features allow us to construct bilateral market shares on both sides of the transaction, which are the key sufficient statistics in the model governing the distribution of markups and pass-through elasticities. Our main sample focuses on repeated, arm’s-length relationships involving the exchange of intermediate inputs, excluding related-party links as flagged in the LFTTD.
Pricing and pass-through patterns in the data closely align with the model’s predictions under two-sided market power. Controlling for supplier–product–time fixed effects to proxy for marginal costs, we find that prices rise with the exporter’s supplier share and fall with the importer’s buyer share, consistent with the presence of both oligopoly and oligopsony forces. In contrast, tariff pass-through declines with buyer share and is non-decreasing in supplier share, suggesting that price adjustments are primarily driven by the cost channel, with a more limited role for the markup channel, as expected under two-sided market power.
These empirical comovements reinforce the model’s central mechanisms and are difficult to reconcile with conventional pricing frameworks that assume one-sided or purely competitive behavior. Nonetheless, reduced-form estimates may conflate the effects of competitor behavior, input price variation, or broader macroeconomic shocks, limiting their interpretability. A structural approach is therefore needed to isolate the underlying mechanisms. [^2] For a related discussion in a different context, see Berger et al. (2022).
To do so, Section 4 develops a structural estimation strategy around two key parameters: the importer’s bargaining power and the exporter’s returns to scale, which determines the slope of the input supply curve. We estimate that U.S. importers hold bargaining power of approximately 0.8, roughly four times greater than that of their foreign suppliers, and that the returns to scale parameter is about 0.45, consistent with upward-sloping supply. Identification exploits cross-sectional price variation across U.S. importers within a supplier–product–year. As the estimating equations hold independently of the general equilibrium environment, the approach remains valid without explicitly modeling broader equilibrium feedback.
We assess the model’s validity by evaluating its ability to replicate both the average and heterogeneous effects of the 2018 tariffs on prices and quantities. First, the model closely matches the average price response as well as its variation with buyer and supplier shares, consistent with the role of oligopoly, oligopsony, and cost-side adjustments. Second, applying the IV-based test proposed by Adão et al. (2023), we find that our model outperforms alternatives that either impose constant returns to scale or exclude bargaining. Third, while the model underestimates the magnitude of quantity changes, it successfully captures their direction, lending support to its core allocative mechanism. Together, these results validate the model’s key assumptions and suggest that general equilibrium feedback or omitted strategic interactions are unlikely to materially bias its pass-through predictions.
In Section 5, we use the estimated model to assess how the 2018 Trump tariffs affected aggregate U.S. import prices through firm-to-firm relationships. Rather than providing a full accounting of tariff incidence, which lies beyond the scope of our framework, we focus on one key margin: price adjustment within ongoing matches, and its decomposition into changes in markups and marginal costs.
We find that aggregate pass-through in firm-to-firm trade is substantially incomplete, with elasticities ranging from 65 to 71%. The model sheds light on the mechanisms behind this muted response. The key insight is that most of the adjustment occurs on the cost side: exporters facing weaker demand from powerful buyers move down their marginal cost curves, absorbing a sizable share of the tariff burden. Markup changes occur but contribute little to the aggregate response due to offsetting effects of strategic complementarities and substitutabilities. This distinction is important. While prior work often attributes incomplete pass-through to variable markups and strategic pricing, our results highlight the dominant role of cost adjustment and supply-side forces.
Our results stand in contrast to studies documenting near-complete pass-through of the 2018 tariffs (e.g., Amiti et al., 2019b, 2020; Cavallo et al., 2020; Fajgelbaum et al., 2020; Flaaen et al., 2020). We reconcile this divergence on two grounds. First, our analysis focuses on intermediate inputs and arm’s-length transactions, where pass-through may be lower due to input specificity, contractual frictions, or capacity constraints. Second, we isolate pricing within repeated matches between importers and exporters. By contrast, product-level unit values used in aggregate studies conflate within-match price changes with compositional shifts in trading partners, notably including one-off exchanges. Such one-off transactions are more likely to reflect marginal cost pricing and exhibit full pass-through. [^3] See Cajal-Grossi et al. (2023) for evidence that markups are lower and prices more competitive in spot transactions than in relational matches.
Related Literature
Our paper contributes to the literature on pricing and shock transmission in firm-to-firm trade, with a particular focus on the role of importer market power. A growing body of empirical work highlights the influence of dominant buyers in shaping supplier outcomes. Bernard et al. (2019) and Bernard et al. (2022) document the pivotal role of large buyers in determining supplier performance and market access. Using French customs data, Fontaine et al. (2020) show substantial variation in unit values across importers transacting with the same exporter, consistent with buyer-specific pricing. Similar patterns are found by Huang et al. (2021) in France, Chile, and China. Among the few studies to model buyer power directly, Morlacco (2019) estimates substantial oligopsony power among French importers, while Atkin et al. (2024) show that bargaining between Argentinian importers and foreign suppliers affects the price impact of trade policy.
Formal models of pricing in firm-to-firm trade remain limited. Dhyne et al. (2022) develops a network-based model of oligopoly in which suppliers’ markups increase with their relationship-specific market share. Our framework nests this model as a special case but extends it to incorporate bilateral bargaining and importer market power. Grossman et al. (2024) also study firm-to-firm pricing under Nash-in-Nash bargaining, emphasizing how supply-chain links adjust in response to trade shocks. In contrast, we take the trade network as fixed and focus on how bargaining and network rigidities interact to determine equilibrium prices and short-run pass-through.
We also contribute to the literature on the firm-level determinants of cost shock pass-through. A large body of work attributes incomplete pass-through to variable markups and strategic complementarities in price-setting (e.g., Amiti et al., 2014, 2019a). [^4] See also Berman et al. (2012), Auer and Schoenle (2016), and Garetto (2016) for firm-level evidence consistent with this class of models. We extend these insights to a bilateral oligopoly setting that reflects the structure of a large share of international trade. Our framework nests these existing mechanisms but allows for a richer characterization of both markup and cost channels, shaped by bargaining power and network position. Related work includes Gopinath and Itskhoki (2011) and Goldberg and Tille (2013), who model bargaining in firm-to-firm trade but abstract from importer market power and cost-side adjustment. Empirically, we find that these cost-based responses, rather than variable markups, are the primary driver of incomplete pass-through in these settings.
Finally, we contribute to the literature on shock propagation in production networks. A large body of work demonstrates that shocks can spread through input–output linkages and affect aggregate outcomes (Acemoglu et al., 2012; Di Giovanni et al., 2014; Grassi, 2018), often using natural experiments to trace transmission along supply chains (Barrot and Sauvagnat, 2016; Boehm et al., 2019; Carvalho et al., 2021). Recent work by Acemoglu and Tahbaz-Salehi (2025) highlights how bilateral bargaining can amplify supply chain disruptions through the price channel. We complement this perspective by modeling how idiosyncratic shocks to individual relationships affect prices under two-sided market power. While we abstract from general equilibrium feedback, our focus is on the within-network transmission of trade shocks through pricing, a key but underexplored margin in the propagation of shocks across production networks.
2 Theoretical Framework
This section develops a bargaining theory of firm-to-firm trade with two-sided market power. The model links markups and pass-through elasticities to importer and exporter market shares, along with a small set of structural parameters that capture key features of market structure upstream and downstream. The model yields analytic structural equations and testable predictions, which we later bring to the data using U.S. customs records.
We impose the following assumptions to keep the model tractable. First, we treat the trade network as fixed. This assumption captures the lock-in effects in firm-to-firm trade and allows us to focus on the price-setting problem while abstracting from firms’ decisions to form or sever links. [^5] This separation between extensive and intensive margin decisions is consistent with prior work showing that dynamic sourcing choices, while important for long-run outcomes, are not necessary to analyze the short-run price effects of trade shocks (Blaum et al., 2018). Second, we abstract from nominal rigidities, such as fixed-price contracts or currency denomination, as these are unlikely to materially affect tariff pass-through. [^6] Evidence from Amiti et al. (2020) supports this view, showing similar short- and long-run pass-through rates of 2018 U.S. import tariffs. See Goldberg and Tille (2013) for a theory of importer-exporter bargaining over the transaction price and exchange rate exposure. Third, we consider a static framework of single-product negotiations, despite our data reflecting repeated, multi-product interactions between firms. [^7] Repeated relationships affect contract structure and trade volumes (Macchiavello and Morjaria, 2015), exchange rate pass-through (Heise, 2024), and the propagation of shocks (Martin et al., 2023). We return to these features and discuss how we incorporate information on relationship duration in Section 3 and 4.
2.1 Environment
We focus on the relationship between exporter $i$ and importer $j$ of an intermediate input. We denote by $\mathcal{Z}_{i}$ the set of importers connected to exporter $i,$ and by $\mathcal{Z}_{j}$ the set of exporters connected to importer $j$. These sets vary across firms and are treated as given.
Exporters and Supply
Exporter $i$ produces $q_{i}$ units of the unique input variety and sells them to all importers in $\mathcal{Z}_{i}$, where total output satisfies $q_{i}\equiv\sum_{j\in\mathcal{Z}_{i}}q_{ij}$, and $q_{ij}$ denotes the quantity of the intermediate input purchased by importer $j$.
We assume that exporter $i$ operates a short-run production technology with returns to scale parameter $\theta\in(0,1]$. Marginal cost (or equivalently, the short-run supply curve) is given by:
$c_{i}\equiv MC_{i}(q_{i})=k_{i}\,q_{i}^{\frac{1-\theta}{\theta}},$ (2.1)
where $k_{i}$ captures exogenous factors such as the exporter productivity or foreign wages.
This cost structure implies that the average cost of $q_{i}$ units is $\theta$-times the marginal cost. When $\theta<1$, indicating decreasing returns to scale, the short-run average cost lies below the marginal cost, and the gap between them widens as $q_{i}$ increases. In contrast, $\theta=1$ indicates a constant returns technology, with marginal and average costs constant and equal to $k_{i}$.
Importers and Demand
Importer $j$ combines domestic and foreign inputs, denoted by $q_{j}^{d}$ and $q_{j}^{f}$ respectively, to produce $q_{j}$ units of a final good, which is subsequently sold in a downstream market. The production technology of importer $j$ is given by:
$q_{j}=\varphi_{j}\left(q_{j}^{f}\right)^{\gamma}\left(q_{j}^{d\vphantom{f}}\right)^{\varrho-\gamma},$ (2.2)
where $\varphi_{j}$ is the importer’s productivity, $\gamma$ and $\varrho-\gamma$ are the output elasticities of foreign and domestic inputs, respectively, with $\varrho$ governing the degree of returns to scale in production. [^8] The domestic input $q_{j}^{d}$ can be interpreted as a constant returns to scale aggregator of primary factors, such as labor and domestic intermediates. Due to the lack of data on domestic input purchases, we model foreign and domestic intermediates as distinct rather than symmetric or substitutable.
The foreign input $q_{j}^{f}$ is a constant elasticity of substitution (CES) composite of differentiated input varieties sourced from exporters in the set $\mathcal{Z}_{j}$. Each variety $q_{ij}$ is weighted by a demand shifter $\varsigma_{ij}$, and the elasticity of substitution across varieties is given by $\rho>1$:
$q_{j}^{f}=\left(\sum_{i\in\mathcal{Z}_{j}}\varsigma_{ij}\,q_{ij}^{\frac{\rho-1}{\rho}}\right)^{\frac{\rho}{\rho-1}}.$ (2.3)
Downstream, each importer operates in a monopolistically competitive market, facing CES demand with constant elasticity $\nu>1$ and an importer-specific demand shifter $D_{j}$.
2.2 Bargaining Protocol and Gains From Trade
Each $i-j$ vertical interaction proceeds in two stages. In the first stage, the importer chooses input quantity $q_{ij}$ to minimize total input costs, taking the price $p_{ij}$ as given:
$\min_{q_{ij}}\;TC_{j}(q_{ij};p_{ij})=p_{ij}q_{ij}+\sum_{\ell\in\mathcal{Z}_{j}\setminus i}p_{\ell j}q_{\ell j}+p^{d}_{j}q^{d}_{j}.$
The solution to this problem yields importer $j$’s input demand:
$q_{ij}=q_{j}^{f}\,\varsigma_{ij}^{\rho}\left(\frac{p_{ij}}{p_{j}^{f}}\right)^{-\rho}\quad\text{with}\quad q_{j}^{f}=\gamma\,c_{j}\,q_{j}\left(p_{j}^{f}\right)^{-1},$ (2.4)
where $p_{j}^{f}=\left(\sum_{i}\varsigma_{ij}^{\rho}p_{ij}^{1-\rho}\right)^{\frac{1}{1-\rho}}$ is the foreign input price index.
In the second stage, the importer and exporter determine the negotiated price $p_{ij}$ through Nash bargaining, taking the importer’s demand in equation (2.4) as given:
$\max_{p_{ij}}\left(\underbrace{\pi^{i}(p_{ij};q_{ij})-\tilde{\pi}^{i}_{(-j)}}_{GFT_{ij}^{i}}\right)^{1-\phi}\left(\underbrace{\pi^{j}(p_{ij};q_{ij})-\tilde{\pi}_{(-i)}^{j}}_{GFT_{ij}^{j}}\right)^{\phi}.$ (2.5)
In equation (2.5), $\phi\in(0,1)$ denotes the importer’s bargaining leverage, and $GFT_{ij}^{k}$ for $k\in\{i,j\}$ captures the gains from trade, defined as the difference between a firm’s profits with and without the bilateral relationship. In case of disagreement, each party falls back on trade with its other existing partners. All bilateral matches generate strictly positive surplus, and trade occurs in equilibrium across all relationships.
To solve equation (2.5), we adopt the Nash-in-Nash bargaining solution concept, in which each bilateral negotiation takes as given the outcomes of all other matches in the network (Horn and Wolinsky, 1988; Collard-Wexler et al., 2019). [^9] A common alternative is the sequential bargaining model of Stole and Zwiebel (1996), in which disagreement with one partner triggers renegotiation with others. This framework is suited to one-to-many settings such as labor markets, but is less appropriate for global supply chains, where firms engage in many-to-many relationships and isolated breakdowns rarely affect unrelated contracts. Moreover, it introduces additional complexity without yielding clear empirical gains in our context. For this reason, we leave the dependence on prices and quantities in other links in the network implicit throughout the analysis.
This setup implies two key properties. First, prices are allocative: the negotiated price $p_{ij}$ pins down the traded quantity $q_{ij}$. Second, the equilibrium allocation $(p_{ij},q_{ij})$ satisfies the importer’s demand in equation (2.4). We discuss the allocative implications of these assumptions, along with alternative bargaining protocols, in Section 2.5.
2.3 Equilibrium Prices and Allocations
The solution to (2.5) yields a bilateral price of the form $p_{ij}=\mu_{ij}\cdot c_{i},$ where $c_{i}$ is exporter $i$’s marginal cost, and $\mu_{ij}$ is a pair-specific markup.
We define two bilateral market shares that serve as key sufficient statistics in the model:
$s_{ij}$ $\equiv\frac{p_{ij}q_{ij}}{\sum_{k\in\mathcal{Z}_{j}}p_{kj}q_{kj}}\in[0,1],\quad\quad x_{ij}\equiv\frac{q_{ij}}{\sum_{k\in\mathcal{Z}_{i}}q_{ik}}\in[0,1],$
where $s_{ij}$ denotes exporter $i$’s supplier share, defined as its sales to importer $j$ as a fraction of $j$’s total foreign input expenditures, and $x_{ij}$ denotes importer $j$’s buyer share, equal to its purchases as a share of exporter $i$’s total output.
The following proposition characterizes the equilibrium markup as a convex combination of two limiting pricing regimes:
Proposition 1.
The bilateral markup $\mu_{ij}$ can be expressed as
$\mu_{ij}=(1-\omega_{ij})\cdot\mu_{ij}^{\textnormal{oligopoly}}+\omega_{ij}\cdot\mu_{ij}^{\textnormal{oligopsony}},$ (2.6)
where the weight
$\omega_{ij}=\frac{\frac{\phi}{1-\phi}\lambda_{ij}}{1+\frac{\phi}{1-\phi}\lambda_{ij}}\in(0,1)$ (2.7)
represents the importer’s effective bargaining power, and $\lambda_{ij}\equiv-\frac{d\ln\pi_{j}}{d\ln(p_{ij}\,q_{ij})}\cdot\frac{\pi_{j}}{GFT_{ij}^{j}}$ captures the strength of the importer’s outside option.
Proof: See Appendix A.1.
Equation (2.6) shows that the equilibrium markup lies between two polar cases. When the exporter holds all the bargaining power ($\phi\to 0$), the markup converges to the oligopoly case $\mu_{ij}^{\text{oligopoly}}$. When the importer holds full bargaining power ($\phi\to 1$), it converges to the oligopsony case $\mu_{ij}^{\text{oligopsony}}$. For intermediate values, the outcome reflects the influence of both parties, with weight $\omega_{ij}$ determined by bargaining power and the importer’s outside option (see Section 2.3.3).
We now examine these limiting cases in more detail and characterize how the bilateral markup depends on market shares and model primitives.
2.3.1 Oligopoly Markup
Under full exporter bargaining power ($\phi=0$), the markup takes the standard form:
$\mu_{ij}^{\text{oligopoly}}$ $=\frac{\varepsilon_{ij}}{\varepsilon_{ij}-1}\geq 1,$ (2.8)
where $\varepsilon_{ij}$ denotes the residual demand elasticity faced by exporter $i$, given by:
$\varepsilon_{ij}=(1-s_{ij})\cdot\rho+s_{ij}\cdot\eta.$ (2.9)
This elasticity is a weighted average of two components: the elasticity of substitution across foreign suppliers, $\rho$, and the elasticity $\eta$ of the importer’s foreign input bundle $q_{j}^{f}$ with respect to its price index $p_{j}^{f}$:
$\eta\equiv-\frac{d\ln q_{j}^{f}}{d\ln p_{j}^{f}}=\frac{(\varrho-\gamma)+\nu\cdot(1-(\varrho-\gamma))}{\varrho+\nu\cdot(1-\varrho)}.$
Provided that $\rho>\eta$, a standard parameter condition, the exporter’s markup (2.8) increases with its supplier share $s_{ij}$, reflecting greater oligopoly power.
This case corresponds to the firm-to-firm trade setting in Dhyne et al. (2022). It is also closely related to the oligopolistic competition model of Atkeson and Burstein (2008), but differs in three key respects. First, supplier shares are defined at the match level, rather than at the firm or industry level, reflecting our assumption of a fixed network. Second, whereas the outer nest elasticity $\eta$ is typically treated as a fixed preference parameter, here it summarizes how downstream market structure–determined by both technology ($\gamma$, $\varrho$) and demand ($\nu$)–shapes the bargaining environment. Third, we allow for decreasing returns to scale in production, in contrast to the constant returns assumption common in related models.
Figure 1 illustrates the equilibrium outcome for a representative $i-j$ match. The downward-sloping residual demand and marginal revenue curves (black and gray) intersect with the upward-sloping residual marginal and average cost curves (blue and red).
Panel (A) depicts the case with $\phi=0$. The equilibrium quantity is determined by the intersection of marginal cost and marginal revenue, and the price lies on the residual demand curve. Exporter rents have two components: the oligopoly rent (red), which corresponds to the markup in equation (2.8), and the quasi-rent (purple), which arises from the gap between marginal and average cost under decreasing returns to scale.
Figure 1: Illustration of Equilibrium Allocations With Different Bargaining Power
(a) $\phi=0$
(b) $\phi\in(0,1)$
(c) $\phi=1$
Notes: Each panel illustrates the negotiated price and quantity under a different value of $\phi$. The exporter’s residual demand function $p(q_{ij})$ is shown in black; the importer’s residual supply function $MC_{i}(q_{ij})$ is shown in blue and the average cost function $AC_{i}(q_{ij})$ is in red. All panels assume $\theta<1$.
2.3.2 Oligopsony Markdown
Under full importer bargaining power ($\phi=1$), the bilateral markup simplifies to:
$\mu_{ij}^{\text{oligopsony}}=\theta\cdot\frac{1-(1-x_{ij})^{\frac{1}{\theta}}}{x_{ij}}\leq 1,$ (2.10)
which we refer to as an oligopsony markup, or markdown, since it lies weakly below one. [^10] Although $\mu_{ij}^{\text{oligopsony}}$ denotes a price–cost ratio, we refer to it as a markdown because it falls below one. This differs from standard oligopsony models, where markdowns typically reflect a wedge between input prices and marginal revenue product. See Section 2.5 for further discussion.
Expression (2.10) characterizes the lowest price at which the exporter is willing to supply, with the importer extracting the exporter’s entire surplus through bargaining. [^11] Formally, the markdown is proportional to the percentage increase in quasi-rents attributable to the $i-j$ match. See equation (A.9) in Appendix A.1. The markdown declines with the buyer’s share $x_{ij}$: larger buyers induce greater quasi-rents and can negotiate lower prices. As $x_{ij}\to 0$, the importer behaves atomistically, generates no quasi-rents, and pays a price equal to marginal cost.
This dependence of the price–cost ratio on $x_{ij}$ reflects a form of oligopsony power that arises only under decreasing returns to scale. When $\theta=1$, marginal and average costs coincide, quasi-rents vanish, and the importer can extract only oligopoly rents. In this case, the price cannot fall below marginal cost, and the markdown remains fixed at one.
Panel (C) of Figure 1 illustrates the equilibrium under $\phi=1$. Compared to Panel (A), the price lies at the intersection of the downstream demand curve and the exporter’s residual average cost curve, leaving no surplus to the exporter. The result is a lower price and higher quantity relative to the oligopoly case.
Clarification of Terminology
The terms “buyer power” and “oligopsony power” are often used interchangeably in the literature, but rarely precisely defined (Noll, 2005). We distinguish between these concepts in our framework.
We use buyer power or importer market power to refer broadly to the importer’s ability to influence prices through bargaining (i.e., any $\phi>0$). In contrast, we reserve oligopsony power for the specific case when a higher buyer share $x_{ij}$ enables the importer to negotiate lower prices, holding $\phi$ fixed. Importantly, oligopsony power requires both $x_{ij}>0$ and $\theta<1$, as it depends on the scale-dependent gap between marginal and average cost.
2.3.3 Two-Sided Market Power and Bargaining Weights
Panel (B) of Figure 1 shows the intermediate case with $\phi\in(0,1)$. The equilibrium price lies on the thick blue curve and reflects a weighted average of the outcomes in Panels (A) and (C), consistent with Proposition 1. The resulting price–marginal cost ratio can exceed or fall below one, depending on the relative bargaining power of the exporter and importer.
Determinants of $\omega_{ij}$.
The bargaining weight $\omega_{ij}$ in equation (2.6), defined in equation (2.7), governs the balance between oligopoly and oligopsony forces in price setting. It depends both on the exogenous bargaining power $\phi$ and an endogenous term $\lambda_{ij}$, defined as:
[…]
To interpret $\lambda_{ij}$, we decompose it into two terms. The cost exposure component, $\lambda_{ij}^{C}=\frac{(\eta-1)s_{ij}}{\varepsilon_{ij}-1}$, captures how sensitive importer $j$’s profits are to price changes from supplier $i$ and increases with the supplier share $s_{ij}$. The network dependence term, $\lambda_{ij}^{N}=\left(1-(1-s_{ij})^{\frac{\eta-1}{\rho-1}}\right)^{-1}$, reflects how much of importer $j$’s profits rely on the match with $i$, and declines with $s_{ij}$ as reduced diversification weakens the importer’s outside option.
Taken together, $\lambda_{ij}$ captures both the importer’s incentive to negotiate aggressively (through cost exposure) and the strength of its fallback option (through network dependence). It follows a hump-shaped pattern in $s_{ij}$, increasing at low values and declining at higher ones, and converges to 1 as $s_{ij}\to 0$ or $s_{ij}\to 1$.
This interpretation clarifies the meaning of $\omega_{ij}$ as a measure of the importer’s effective bargaining power. While $\phi$ governs baseline influence in Nash bargaining, $\lambda_{ij}$ endogenously adjusts it based on the structure of the trading relationship.
Testable Implications.
Equation (2.6) writes the bilateral markup $\mu_{ij}$ as a convex combination of the oligopoly markup and the oligopsony markdown, with bargaining power as weight. Since these components scale with bilateral market shares, the model delivers clear predictions on how markups co-move with $s_{ij}$ and $x_{ij}$ under two-sided market power.
Proposition 2.
The bilateral markup $\mu_{ij}$ exhibits the following properties:
  1. 1.
    Markup and Exporter’s Share. If $\mu_{ij}$ increases with $s_{ij}$, then $\phi<1$.
  2. 2.
    Markup and Importer’s Share. If $\mu_{ij}$ decreases with $x_{ij}$, then $\phi>0$ and $\theta<1$.
Proof: See Appendix A.2.
Proposition 2 provides sufficient conditions for detecting two-sided market power. A positive relationship between $\mu_{ij}$ and supplier share $s_{ij}$ implies $\phi<1$, since exporters retain pricing power and $\mu_{ij}^{\text{oligopoly}}$ increases with $s_{ij}$. If $\phi=1$, prices are fully disciplined by buyers, and no such pattern should emerge. While the dependence of $\omega_{ij}$ on $s_{ij}$ may attenuate this effect, a positive slope remains inconsistent with full buyer power.
Conversely, a negative relationship between $\mu_{ij}$ and buyer share $x_{ij}$ signals oligopsony power. This requires both positive bargaining leverage for importers ($\phi>0$) and decreasing returns to scale ($\theta<1$). When both conditions are met, markups reflect the joint influence of importer and exporter market power.
2.4 Equilibrium: Tariff Pass-Through Elasticities
Our pricing framework provides a basis for analyzing the short-run impact of an unanticipated tariff imposed on imports from country $c$. Let $T_{c}$ denote the gross tariff rate. The (log) price that exporter $i$ from country $c$ charges importer $j$ is given by:
$\ln p_{ij}=\ln\mu_{ij}+\ln c_{i}+\ln T_{c}.$
While a tariff on country $c$ may, in principle, affect all trade relationships involving exporters from $c$, we focus on the short-run, direct effect at the relationship level. We interpret the tariff shock as a small, unanticipated perturbation around the initial equilibrium. Accordingly, we treat $T_{c}$ as a pair-specific shock and hold constant prices in all other relevant matches such that $dp_{rj}=0$ for all $r\in\mathcal{Z}_{j}\setminus{i}$ and $dp_{i\ell}=0$ for all $\ell\in\mathcal{Z}_{i}\setminus{j}$.
The following proposition characterizes this direct component of tariff pass-through.
Proposition 3.
The tariff pass-through elasticity into the bilateral import price $p_{ij}$, holding fixed all other prices in the network and general equilibrium variables, is:
$\Phi_{ij}\equiv\frac{d\ln p_{ij}}{d\ln T_{c}}=\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}},$ (2.11)
where:
$\Gamma_{ij}\equiv-\frac{d\ln\mu_{ij}}{d\ln p_{ij}}\quad\textit{and}\quad\Lambda_{ij}\equiv-\frac{d\ln c_{i}}{d\ln p_{ij}}$
denote the partial elasticities of the equilibrium markup $\mu_{ij}$ and the exporter’s marginal cost $c_{i}$ to changes in the bilateral price $p_{ij}$, respectively.
Proof: See Appendix A.3.
Proposition 3 highlights two distinct mechanisms shaping tariff pass-through in firm-to-firm trade: a markup channel, capturing strategic pricing responses, and a cost channel, capturing how the exporter’s marginal cost adjusts in response to price changes. In models with CES demand, monopolistic competition, and constant marginal costs, both $\Gamma_{ij}$ and $\Lambda_{ij}$ are zero, yielding full pass-through ($\Phi_{ij}=1$). More generally, $\Phi_{ij}$ may lie above, below, or at one depending on the magnitudes of $\Gamma_{ij}$ and $\Lambda_{ij}$.
2.4.1 Markup Elasticity
The markup elasticity $\Gamma_{ij}$ measures how the equilibrium markup $\mu_{ij}$ responds to changes in the bilateral price $p_{ij}$. It is given by:
[…]
where $\omega_{ij}^{\Gamma}\equiv\omega_{ij}\cdot\frac{\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\in[0,1]$, $\Gamma_{ij}^{\text{oligopoly}}$ denotes the oligopoly markup elasticity, $\Gamma_{ij}^{\text{oligopsony}}$ the oligopsony markup elasticity, and $\Gamma_{ij}^{\omega}$ the elasticity of the bargaining weight with respect to the bilateral price.
The term in square brackets captures the direct elasticity of the markup, expressed as a convex combination of the oligopoly and oligopsony components. The second term reflects how the bargaining weight $\omega_{ij}$ itself responds to price changes. We discuss each component in turn.
Oligopoly Markup Elasticity
The oligopoly markup elasticity dominates when $\phi\to 0$, in which case $\Gamma_{ij}\to\Gamma_{ij}^{\text{oligopoly}}$. It is given by
$\Gamma_{ij}^{\text{oligopoly}}\equiv-\frac{d\ln\mu_{ij}^{\text{oligopoly}}}{d\ln p_{ij}}=\frac{1}{\varepsilon_{ij}-1}\cdot\frac{\rho-\varepsilon_{ij}}{\varepsilon_{ij}}\cdot(\rho-1)(1-s_{ij})\geq 0.$
This elasticity reflects the standard logic of strategic complementarities in price-setting: when a tariff raises the bilateral price $p_{ij}$, exporter $i$ reduces its markup to limit trade diversion, resulting in incomplete pass-through (see Amiti et al., 2014; Auer and Schoenle, 2016; Garetto, 2016; Amiti et al., 2019a). The response is strongest at intermediate supplier shares, giving rise to a U-shaped relationship between $\Gamma_{ij}^{\text{oligopoly}}$ and $s_{ij}$. [^12] Amiti et al. (2014) shows that, to a first-order approximation, pass-through decreases with the exporter’s share through strategic complementarities. However, we cannot rely on the same approximation, given our focus on bilateral markets, where both very low and very high market shares are observed in the data.
Oligopsony Markdown Elasticity
The oligopsony markdown elasticity dominates when $\phi\to 1$, in which case $\Gamma_{ij}\to\Gamma_{ij}^{\text{oligopsony}}$. It is given by
$\Gamma_{ij}^{\text{oligopsony}}\equiv-\frac{d\ln\mu_{ij}^{\text{oligopsony}}}{d\ln p_{ij}}=\left(\frac{x_{ij}\cdot(1-x_{ij})^{\frac{1}{\theta}-1}}{\theta\cdot(1-(1-x_{ij})^{\frac{1}{\theta}})}-1\right)(1-x_{ij})\,\varepsilon_{ij}\leq 0.$
This elasticity captures the logic of strategic substitutabilities among importers. [^13] Strategic substitutabilities arise because a decline in demand by other buyers lowers marginal costs and prices, encouraging importer $j$ to expand purchases. As the price $p_{ij}$ rises, importer $j$ reduces demand, shrinking the exporter’s quasi-rents. This weakens the importer’s bargaining position, reducing the markdown and amplifying the price response, potentially resulting in more-than-complete pass-through.
The elasticity $\Gamma_{ij}^{\text{oligopsony}}$ depends on both the importer’s buyer share $x_{ij}$ and the exporter’s supplier share $s_{ij}$. It is U-shaped in $x_{ij}$: the elasticity vanishes when $x_{ij}\to 0$ (atomistic buyer) or $x_{ij}\to 1$ (monopsonist), and reaches its peak at intermediate values. It also declines with $s_{ij}$, since a higher supplier share reduces the demand elasticity $\varepsilon_{ij}$, limiting changes in $x_{ij}$ and weakening the markdown response.
The Role of the Endogenous Bargaining Weight
The elasticity $\Gamma_{ij}^{\omega}\equiv\frac{d\ln\omega_{ij}}{d\ln p_{ij}}$ captures how the bargaining weight $\omega_{ij}$ responds to price changes. Its sign and magnitude depend on the exporter’s supplier share $s_{ij}$ and are derived in Appendix A.3. For empirically relevant values of $s_{ij}$ and parameter ranges, this elasticity is typically small. In particular, as $s_{ij}$ approaches zero or one, where $\omega_{ij}$ converges to the importer’s exogenous bargaining strength $\phi$, the elasticity $\Gamma_{ij}^{\omega}$ goes to zero.
2.4.2 Cost Elasticity
The cost elasticity $\Lambda_{ij}$ measures how exporter $i$’s marginal cost $c_{i}$ responds to changes in the bilateral price $p_{ij}$, via adjustments in traded quantity. While prior work emphasizes the role of the cost channel in shaping pass-through (e.g., Burstein and Gopinath, 2015; Amiti et al., 2019a), our contribution is to show that this elasticity varies systematically across matches through both demand and supply channels.
Formally:
$\Lambda_{ij}\equiv-\frac{d\ln c_{i}}{d\ln p_{ij}}$ $=\frac{d\ln c_{i}}{d\ln q_{ij}}\cdot\left(-\frac{d\ln q_{ij}}{d\ln p_{ij}}\right)$
$=\frac{1-\theta}{\theta}\cdot x_{ij}\cdot\varepsilon_{ij}\geq 0.$
The elasticity declines with the exporter’s supplier share $s_{ij}$, which reduces the residual demand elasticity $\varepsilon_{ij}$, and increases with the importer’s buyer share $x_{ij}$, which amplifies the residual supply elasticity $\frac{d\ln c_{i}}{d\ln q_{ij}}$. Importantly, this mechanism operates independently of bargaining power ($\phi$), and instead reflects the concentrated nature of the trade network.
2.4.3 Pass-Through Elasticity and Bilateral Market Shares
Figure 2: Pass-Through Elasticity and Bilateral Market Shares
Notes: The figure presents heatmaps of the pass-through elasticity $\Phi_{ij}$ across combinations of $s_{ij}$ (x-axis) and $x_{ij}$ (y-axis), under alternative assumptions about $\phi$ and $\theta$. We set $\theta=0.5$ in the decreasing returns case and $\phi=0.5$ in the intermediate bargaining case. Other parameters are fixed at $\gamma=0.5$, $\varrho=1$, $\nu=4$, and $\rho=10$.
The interaction between markup and cost channels generates rich heterogeneity in pass-through elasticities across matches, shaped by bargaining power ($\phi$), returns to scale ($\theta$), and bilateral market shares ($s_{ij}$ and $x_{ij}$). Figure 2 illustrates these patterns using heatmaps of the pass-through elasticity $\Phi_{ij}$ as a function of supplier share ($s_{ij}$) and buyer share ($x_{ij}$), under three bargaining regimes—$\phi\to 0$, $\phi\in(0,1)$, and $\phi\to 1$—with decreasing returns to scale ($\theta<1$) in the top row and constant returns to scale ($\theta=1$) in the bottom row.
Several key insights emerge. First, pass-through $\Phi_{ij}$ increases with importer bargaining power $\phi$. As $\phi$ rises, the markup elasticity shifts toward the oligopsony markdown elasticity, increasing pass-through through strategic substitutabilities. This is most visible in the shift from the left to the right columns within each row.
Second, when $\theta<1$, pass-through is incomplete across most of the $(s_{ij},x_{ij})$ space. This is because the cost elasticity $\Lambda_{ij}$, which is always non-negative, dominates the overall pass-through elasticity $(\Gamma_{ij}+\Lambda_{ij})$ when $\Gamma_{ij}$ is negative due to strong buyer power. Moreover, pass-through values under $\phi\in(0,1)$ and $\phi=1$ are nearly indistinguishable, indicating that pass-through is relatively insensitive to the exact level of $\phi$ in this regime. This reflects two forces: (i) the cost channel is strong when returns to scale are decreasing and importers’ bargaining power is high, and (ii) the markup elasticity $\Gamma_{ij}$ is either muted by offsetting effects of strategic complementarities and substitutabilities or generally low in values. [^14] For more details, see Appendix A.4.
Third, pass-through $\Phi_{ij}$ depends on $x_{ij}$ only when $\theta<1$, in which case it declines with $x_{ij}$. This yields a sharp empirical prediction: a negative relationship between pass-through and buyer share $x_{ij}$ implies decreasing returns to scale, regardless of the level of bargaining power. This result is formalized below.
Proposition 4.
If pass-through $\Phi_{ij}$ decreases with the importer’s buyer share $x_{ij}$, then $\theta<1$.
Proof: See Appendix A.4.
By contrast, the relationship between $\Phi_{ij}$ and the supplier share $s_{ij}$ is less clear-cut. As shown in Figure 2, pass-through declines with $s_{ij}$ primarily when $\phi$ is low and $\theta$ is near one, a setting in which the markup channel dominates and cost-based adjustments are limited. In this region, strategic complementarities give rise to a U-shaped relationship between $\Phi_{ij}$ and $s_{ij}$. As either $\phi$ or $x_{ij}$ increases, the cost channel becomes more prominent, and pass-through tends to rise with $s_{ij}$. While we do not formally characterize the conditions under which this reversal occurs, we show below that $\phi$ is weakly increasing in $s_{ij}$. This monotonicity supports interpreting the empirical evidence as inconsistent with the low-$\phi$, high-$\theta$ case.
2.5 Discussion
This section reviews key modeling assumptions and discusses potential extensions.
2.5.1 Bargaining Protocol and Quantities
The baseline model assumes demand-determined quantities, whereby the importer chooses input quantities to minimize total cost, taking the price as given. This assumption provides analytical tractability, yields closed-form solutions for markups and pass-through, and nests standard models of international trade. An important implication is that prices are allocative, a feature that aligns with our empirical findings and with firm-level evidence in related settings (Gopinath and Itskhoki, 2011).
We consider two alternative bargaining protocols as benchmarks for understanding the role of quantity determination. The first is efficient bargaining, discussed more formally in Appendix B.1. This case corresponds to the vertically integrated benchmark, where importer and exporter jointly negotiate over both price and quantity to maximize total surplus. While theoretically appealing, this setup implies that prices are non-allocative transfers, inconsistent with the empirical evidence. Moreover, the vertically integrated case may be unrealistic in the context of arm’s-length firm-to-firm trade, where limited commitment and contracting frictions are prevalent (Antràs, 2020).
The second alternative, detailed in Appendix B.2, is supply-driven bargaining: the exporter sets quantity for a given price, and price is then negotiated based on the resulting supply curve. As $\phi\to 1$, this nests the classic monopsony benchmark commonly used in labor markets (e.g., Berger et al., 2022). The key distinction between this setup and our baseline lies in their welfare implications: in the supply-driven case, buyer power lowers both prices and quantities, while supplier power helps restore efficiency. In contrast, under demand-driven bargaining, buyer power mitigates upstream distortions and improves efficiency. [^15] See Avignon et al. (2024) and Demirer and Rubens (2025) for recent discussions of how quantity-setting assumptions affect welfare outcomes.
While both alternatives offer useful benchmarks, they are less suited to the goals of this paper. The supply-driven model, in particular, does not provide closed-form expressions for key variables, which makes it challenging to generate clear predictions or link the model to data compared to our baseline setup. Still, it remains a valuable direction for future work, especially for studying welfare effects when buyer power leads to inefficiencies.
2.5.2 Outside Options
In our model, each firm’s outside option reflects payoffs from trade with all other existing partners, excluding the focal match. This assumption allows us to express markups and pass-through in terms of market shares and a small set of parameters, facilitating structural estimation and counterfactual analysis.
While analytically convenient, the assumption may appear restrictive if disagreement leads a firm to form or sever other relationships. Appendix B.3 explores a more flexible setup in which disagreement affects the importer’s cost and the exporter’s revenue non-parametrically. Although more realistic, this extension introduces an identification problem: the parameters $\phi$ and $\theta$ can no longer be separately identified from the outside option, which limits their interpretability and empirical tractability.
2.5.3 General Equilibrium Forces
We conclude the theoretical section by noting that all results are derived under a partial equilibrium approach, holding fixed general equilibrium variables such as wages, demand shifters, and competitor behavior. This approach simplifies the analysis but also raises questions about the role of general equilibrium adjustments in the results.
General equilibrium considerations are critical for interpreting pass-through elasticities. The structural pass-through elasticity derived in Section 2.4 is a direct or partial elasticity, holding constant aggregate variables such as wages, demand conditions, and, importantly, competitors’ prices and sourcing decisions. In reality, tariff shocks may also affect variables such as foreign wages or domestic export prices. If these effects are not fully observed or controlled for, reduced-form estimates may conflate the direct impact with general equilibrium responses, making it harder to align empirical pass-through coefficients with model-based elasticities (Burstein and Gopinath, 2015). [^16] In Appendix B.4, we extend equation (2.11) to account for indirect effects, such as how a shock to exporter $i$ influences other prices and quantities, which may in turn affect $p_{ij}$. While we do not model full general equilibrium dynamics, this extension illustrates how spillovers across relationships may lead to reduced-form estimates diverging from structural ones. Therefore, our goal is not to recover structural elasticities from reduced-form regressions, which would be inappropriate in this context (Berger et al., 2022).
Instead, we pursue a structural approach, which helps mitigate similar concerns. Our approach has three advantages. First, the theoretical relationship between bilateral shares, markups, and pass-through holds parametrically regardless of the specific general equilibrium environment. While the parameter values depend on how aggregate variables such as tariffs, demand shifters, and wages co-move in equilibrium, we do not need to model their relationships explicitly. Instead, identification is based on the cross-sectional variation of prices across importers within exporter–product–year cells, without making assumptions on the general equilibrium environment.
Second, we can also isolate and test the cross-sectional predictions of the model independently of general equilibrium forces. In particular, even if aggregate variables shift over time, the model predicts specific patterns between firm-to-firm shares, markups, and pass-through at a given point in time. These predictions can be assessed empirically using within-period variation across matches, which we exploit in the next sections.
Third, we can evaluate the performance of the estimated model by comparing its ability to replicate observed price changes in response to tariff shocks, thereby gauging the significance of general equilibrium forces. As shown in the empirical analysis, the estimated model fits the observed price changes well. This suggests that the short-run effects of tariffs on prices can be understood primarily through the lens of partial equilibrium mechanisms, and supports the usefulness of our framework for studying firm-to-firm pricing and pass-through.
3 Data and Stylized Facts
This section describes the data and preliminary empirical analysis. Section 3.1 outlines the main data sources. Section 3.2 discusses how we adapt the baseline model to the data to construct key variables. Section 3.3 details the sample selection and provides summary statistics. Finally, Section 3.4 presents evidence testing the model’s predictions.
3.1 Data Sources
Our main dataset is the U.S. Census Bureau’s Linked/Longitudinal Firm Trade Transaction Database (LFTTD), which covers the universe of U.S. import transactions from 2001 to 2018. Each observation corresponds to a shipment from a foreign exporter to a U.S. importer and includes the transaction date, product classification at the 10-digit Harmonized System (HS10) level, FOB import value in U.S. dollars, physical quantity, transportation mode, and country of origin. Exporters are identified using a manufacturer ID (MID) constructed by the Census Bureau from the exporter’s name, street address, city, and country. [^17] The MID combines the country code, (elements from) the firm name, city, and address (Kamal and Monarch, 2018). Because the algorithm is not standardized, it may generate inconsistent identifiers due to misspellings or minor location changes, leading to one firm having multiple MIDs or several firms sharing one. Following Kamal and Monarch (2018), we construct a robustness version that truncates location fields to improve consistency. Our baseline uses the full MID, and results are robust to this alternative.
To focus on arm’s-length trade, we exclude related-party transactions from the baseline sample. The LFTTD includes a related-party indicator based on a mandatory field in U.S. Customs forms, flagging relationships with ownership stakes of at least five percent. While widely used, this measure may misclassify firms due to its reliance on self-reporting and a low reporting threshold (Ruhl, 2015). To improve accuracy, we construct an alternative indicator using ORBIS, which provides firm-level cross-border ownership links. We merge ORBIS to the LFTTD as described in Appendix C.1.
We supplement the transaction-level data with information on statutory U.S. import tariffs introduced during the 2018 trade war. We use the dataset from Fajgelbaum et al. (2020), which records the timing, product coverage, and country-specific scope of these measures at the HS8–month level. The tariffs averaged 25 percentage points and were imposed on top of existing rates, targeting selected goods. They were implemented in phases over the course of the year, beginning with imports from China and later expanding to goods from other trade partners, including Canada, Mexico, and the European Union. Tariff changes are annualized based on the number of months each measure was in effect.
3.2 Measuring Key Variables of the Model
To construct the key variables of interest, we extend the model to include multiple foreign inputs, indexed by $h$. Each input corresponds to an HS10 product category. We model the foreign input bundle as a Cobb-Douglas composite of individual product quantities:
$q_{j}^{f}$ $=\prod_{h\in\mathcal{H}_{j}}\left(q_{jh}^{f}\right)^{\alpha_{jh}},\ \ \ \text{where}\ \ \ q_{jh}^{f}=\Big{(}\sum_{i\in\mathcal{Z}_{j}^{h}}\varsigma_{ijh}\cdot\left(q_{ijh}\right)^{\frac{\rho-1}{\rho}}\Big{)}^{\frac{\rho}{\rho-1}},$
and $\alpha_{jh}\in(0,1)$ denotes the (observed) Cobb-Douglas share of input $h$ in firm $j$’s total imports of foreign intermediates. This formulation implies that the elasticity of the importer’s marginal cost with respect to the price of foreign input $h$ is $\frac{d\ln c_{j}}{d\ln p_{jh}^{f}}=\alpha_{jh}\gamma\in(0,1]$.
We construct the exporter’s supplier share as $s_{ijh}\equiv\frac{p_{ijh}q_{ijh}}{\sum_{k\in\mathcal{Z}_{j}^{h}}p_{kjh}q_{kjh}}$, where $\mathcal{Z}_{j}^{h}$ denotes the set of firm $j$’s foreign suppliers of input $h$. The numerator captures the total value of imports of product $h$ from exporter $i$ (a MID in our data) to firm $j$ in a given year. The denominator aggregates imports of product $h$ from all foreign suppliers to $j$.
In contrast, the importer’s buyer share is constructed as $x_{ijh}\equiv\frac{q_{ijh}}{\sum_{k\in\mathcal{Z}_{i}^{h}}q_{ihk}}$, where $\mathcal{Z}_{i}^{h}$ is the set of all U.S. importers buying product $h$ from exporter $i$. Since our dataset only includes U.S. importers, we assume that exporter $i$ operates product- and destination-specific production lines. Under this assumption, the denominator of $x_{ijh}$, which captures the total quantity of product $h$ from exporter $i$, includes only those sold to U.S. buyers. This restriction reflects a data limitation, as we do not observe importer destinations beyond the U.S. and thus cannot account for the full set of an exporter’s buyers. [^18] To address the possibility that the importer’s buyer share $x_{ij}$ may be overstated due to unobserved sales to other destinations, we replicate the analysis using only exporters from Canada and Mexico. These countries direct the majority of their exports to the U.S. (71% for Canada and 73% for Mexico in 2019), making the assumption of destination-specific production less restrictive. Reassuringly, our estimates remain stable in this subsample. Full results are available upon request.
3.3 Sample Construction and Summary Statistics
We apply a series of restrictions to the LFTTD to align the empirical sample with the model’s focus on decentralized bargaining over intermediate inputs. Full details are provided in Appendix C.2. The selection criteria are designed to ensure that we observe relationship-level price changes, exclude related-party transactions, and maintain sufficient variation in buyer–supplier matches for identification purposes. Appendix Table C.1 summarizes how the sample evolves with each restriction across key dimensions.
We begin by restricting the sample to importer–exporter–product triples observed in two consecutive years, allowing us to compute relationship-level price changes, which are central to our analysis. While this requirement eliminates over half of the raw matches, it retains a substantial share of trade, covering roughly 88% of import value between 2001–2016 and 80% between 2017–2018, highlighting the importance of repeated relationships.
Next, we restrict the sample to capital and intermediate inputs by excluding products classified as consumption goods under the Broad Economic Categories (BEC) system. This step reduces the number of suppliers and relationships, as shown in Table C.1.191919As a robustness check, we also consider a broader sample that includes consumption goods. Appendix Table C.2 reports the composition of this sample and Table C.3 reports the corresponding summary statistics. These statistics are broadly similar to the baseline, indicating that including consumption goods does not significantly alter the data composition. To further mitigate measurement error, we apply three filters: (i) exclude transactions involving energy goods, (ii) drop observations with unit values outside the 1st–99th percentile within product, and (iii) remove transactions with absolute log price changes above four.
We then exclude related-party transactions, which are less likely to reflect decentralized bargaining and may involve internal pricing strategies such as transfer pricing. [^20] Bernard et al. (2006) document that prices in related-party trade differ systematically from arm’s-length transactions, with lower average prices and distinct pass-through behavior. In our baseline definition, a buyer–supplier pair is considered related if ORBIS identifies a shared corporate parent. [^21] We retain all observations not flagged as related in either ORBIS or the LFTTD. This approach preserves sample size while ensuring that limited ORBIS coverage does not unduly constrain the selection. See Appendix C.1 for details. For robustness, we consider two alternative definitions: one based solely on the LFTTD’s related-party flag, and another that combines this flag with ORBIS data identifying the U.S. importer as a multinational (domestic or foreign-owned).
Lastly, we impose restrictions based on our identification strategy. Since the latter relies on observing the variation in prices across U.S. buyers for the same supplier-product-year combination, we restrict our sample to supplier–product pairs in which the supplier transacts with at least two U.S. buyers in consecutive years. Table C.1 reports that after these restrictions, the sample accounts for approximately \$160 billion in import value and 250 thousand buyer-supplier-product-year combinations between 2017-2018.
Summary Statistics
Table 1 reports summary statistics for our final sample. Panel A shows that the concentration of importers and exporters is substantial. On average, an exporter supplies 32% of an importer’s total imports of a given HS10 product, with a median share of 15%. The average buyer share is lower, at 25%, with a median of 10%. The two shares are highly dispersed and largely uncorrelated, with a correlation of 0.04.
Table 1: Summary Statistics for Main Estimation Sample (2001–2018)
Variable Mean Std. Dev. P25 Median P75
Panel A: Characteristics of Trade Relationships
$s_{ijh}$: Supplier share 0.32 0.35 0.03 0.15 0.57
$x_{ijh}$: Buyer share 0.25 0.29 0.02 0.10 0.40
 Relationship length (product $h$) 4.00 2.80 2.50 3.50 5.50
 Relationship length (all products) 4.80 3.30 2.50 4.50 6.50
 # Transactions (product $h$) 120 1100 6.50 16 50
 # Transactions (all products) 360 3000 11 36 140
 # Products per pair 3.80 7.30 1.50 2.50 4.50
 Multi-HS10 dummy 0.59 0.49 0.00 1.00 1.00
 # Suppliers per buyer (HS10) 1.80 3.20 1.50 2.50 5.50
 Buyer tenure (all products) 9.90 5.00 6.50 10.00 14.00
 Buyer tenure (product $h$) 6.90 4.40 3.50 6.50 10.00
 # Buyers per supplier (HS10) 3.20 3.90 2.50 3.50 7.50
 Supplier tenure (all products) 8.00 4.60 4.50 8.50 12.00
 Supplier tenure (product $h$) 6.40 4.00 3.50 6.50 9.50
 Corr. between $s_{ijh}$ and $x_{ijh}$ 0.041
Panel B: Prices
$\log p$ (pre-duty) 3.50 2.80 1.40 3.10 5.40
$\log p$ (pre-duty, excl. charges) 3.40 2.80 1.30 3.00 5.40
$\log p^{\text{duty}}$ (post-duty) 3.50 2.80 1.40 3.10 5.40
Notes: This table reports summary statistics for the estimation sample used in the empirical analysis. The data span 2001–2016 and include importer–supplier–product matches observed in two consecutive years. The sample excludes consumption goods (based on BEC), energy products, statistical outliers, and related-party trade, and is restricted to suppliers trading with at least two U.S. buyers in two consecutive years. This sample corresponds to the “+ Supplier Multi-Buyer” row in Panel B of Table C.1. Columns report the mean, standard deviation, and selected quantiles (25th, 50th, 75th percentile) for each variable. Prices in Panel B are log unit values (FOB value over quantity), with variants including charges or duties. $s_{ijh}$ denotes exporter $i$’s share in buyer $j$’s imports of product $h$; $x_{ijh}$ denotes buyer $j$’s share in exporter $i$’s U.S. exports of the same product. Relationship length and tenure are in years; concentration is measured at the HS10–year level. Counts of buyers, suppliers, and origin countries are per product per firm. Statistics are based on confidential LFTTD data and rounded to four significant digits per U.S. Census Bureau Disclosure Guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Long-term relationships are a hallmark of intermediate input trade (Antràs and Chor, 2013; Monarch, 2022). In our data, importer–exporter pairs trade the same product for an average of four years and remain connected across all products for nearly five years. Pairs transact frequently (median: 16 times per product), often spanning multiple products (mean: 3.8), and typically involve durable links, with average tenures ranging from 7 to 10 years.
Panel B of Table 1 reports descriptive statistics for three bilateral (log) price measures based on FOB unit values: baseline pre-duty prices, pre-duty prices excluding insurance and other ancillary charges, and post-duty prices. All three exhibit wide dispersion across importer–exporter–product matches, with interquartile ranges exceeding four log points.
To assess the sources of this variation, we perform a variance decomposition exercise, as described in Appendix D.1. Table D.1 shows that product-year fixed effects explain about 50% of the total variance, while match-specific residuals account for 4%. Crucially, when we focus only on variation within supplier–product–year cells, 77% of the remaining dispersion is explained by match-specific factors. This pattern holds across price definitions, underscoring the importance of relationship-specific forces in pricing. [^22] See Fontaine et al. (2020) for related evidence in French data.
3.4 Test of Model Predictions
We now examine how the comovements between markups, pass-through, and bilateral market shares align with the model’s predictions in Propositions 2 and 4. While not a formal test, this analysis provides supporting evidence for the mechanisms emphasized in the theory. Section 4 then develops a structural approach to quantify these forces more directly.
3.4.1 Test of Proposition 2: Markups and Bilateral Market Shares
We begin by testing Proposition 2, which predicts that with two-sided market power ($\phi\in(0,1)$ and $\theta<1$), bilateral markups increase with the exporter’s supplier share ($s_{ijht}$) and decrease with the importer’s buyer share ($x_{ijht}$). Since markups are not observed, we use log prices, equal to log markups plus log marginal costs, and include supplier–product–time fixed effects to absorb cost variation and isolate the markup component.
We estimate the following specification:
$\ln p_{ijht}$ $=\alpha_{s}\,s_{ijht}+\alpha_{x}\,x_{ijht}+\mathbf{X}_{ijht}\boldsymbol{\gamma}+\mathbf{FE}+\upsilon_{ijht},$ (3.1)
where the coefficients of interest are $\alpha_{s}$ and $\alpha_{x}$, which we expect to be positive and negative, respectively.
To address endogeneity concerns, we construct leave-one-out instruments that isolate variation in market structure plausibly exogenous to the pricing decision of a given buyer–supplier pair. Specifically, we use the average supplier share (excluding $i$) among other buyers of exporter $i$ (excluding $j$) to instrument for $s_{ijh}$. Similarly, to instrument for $x_{ijh}$, we use the average buyer share (excluding $j$) among other suppliers to importer $j$ (excluding $i$).
Table 2 shows the results. Columns (1)–(2) include exporter ($\text{FE}_{i}$), importer ($\text{FE}_{j}$), and product–year ($\text{FE}_{ht}$) fixed effects. Column (3)–(4) replaces exporter fixed effects with exporter–product–year ($\text{FE}_{iht}$) to account for unobserved marginal costs. Columns (5) and (6) further replace importer fixed effects with importer–product–year ($\text{FE}_{jht}$) to capture buyer-specific demand shocks. All regressions control for the relationship duration, measured as years since the first shipment of product $h$ between $i$ and $j$.
Table 2: Prices and Bilateral Concentration
Dependent Variable: $\ln p_{ijht}$
(1) (2) (3) (4) (5) (6)
OLS IV OLS IV OLS IV
$s_{ijht}$ 0.183 0.174 0.251 0.169 0.269 0.681
(0.0033) (0.0296) (0.0036) (0.029) (0.0065) (0.0147)
$x_{ijht}$0.537 -0.077 -0.594 -0.186 -0.533 -0.777
(0.0030) (0.0175) (0.0028) (0.0249) (0.0061) (0.0112)
$\text{FE}_{i}+\text{FE}_{j}+\text{FE}_{ht}$ Yes Yes No No No No
$\text{FE}_{iht}+\text{FE}_{j}$ No No Yes Yes No No
$\text{FE}_{iht}+\text{FE}_{jht}$ No No No No Yes Yes
Observations 1,200,000 1,200,000 1,200,000 1,200,000 1,200,000 1,200,000
R-squared 0.957 0.010 0.976 0.032 0.991 0.032
First-stage F stat. 5,270 3,485 19,760
SW F stat ($s_{ijht}$) 10,710 7,464 39,830
SW F stat ($x_{ijht}$) 21,120 7,197 43,550
Notes: This table reports OLS and IV estimates of equation (3.1), where the dependent variable is the log FOB unit value of product $h$ imported by buyer $j$ from supplier $i$ in year $t$. Columns alternate between OLS and IV specifications. All regressions control for log relationship length (in years) within HS10 products. Columns (1)–(2) include buyer ($\text{FE}_{j}$), supplier ($\text{FE}_{i}$), and product-year ($\text{FE}_{ht}$) fixed effects. Columns (3)–(4) use supplier–product–year ($\text{FE}_{iht}$) and buyer ($\text{FE}_{j}$) fixed effects. Columns (5)–(6) include fully interacted buyer–product–year and supplier–product–year fixed effects ($\text{FE}_{jht}$, $\text{FE}_{iht}$), flexibly controlling for sourcing and pricing patterns. IV estimates (even-numbered columns) use leave-one-out instruments: $s_{ijht}$ is instrumented with the average share of other suppliers across buyers of $i$ (excluding $j$), and $x_{ijht}$ with the average share of other buyers across suppliers to $j$ (excluding $i$). Because the model includes multiple endogenous regressors, we report both first-stage and conditional F-statistics from Sanderson and Windmeijer (2016), which assess instrument strength for each endogenous regressor conditional on the others, addressing limitations of standard first-stage tests in multi-equation IV settings. Standard errors are robust. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
We estimate $\alpha_{s}>0$ and $\alpha_{x}<0$, with both coefficients statistically and economically significant. A positive $\alpha_{s}$ indicates oligopoly power, requiring exporter bargaining power ($\phi<1$), while a negative $\alpha_{x}$ reflects oligopsony power, requiring importer bargaining power ($\phi>0$) and upward-sloping supply ($\theta<1$). These findings are consistent with the model’s core assumption of two-sided market power, i.e., $\phi\in(0,1)$ and $\theta<1$.
3.4.2 Test of Proposition 4: Pass-Through and Bilateral Market Shares
We next test Proposition 4, which links tariff pass-through to the importer’s buyer share. Under decreasing returns ($\theta<1$), the model predicts that pass-through declines with the buyer share via the cost channel, providing a direct test for $\theta<1$.
For this analysis, we focus on 2017–2018, when U.S. imports experienced sharp and unanticipated tariff increases under the Trump administration. We estimate the following regression specification:
$\Delta\ln p_{ijht}=$ $\alpha_{0}+\alpha_{1}\,\Delta\ln(1+\tau_{cht})+\alpha_{s}\,\Delta\ln(1+\tau_{cht})\cdot s_{ijh,t-1}+\alpha_{x}\,\Delta\ln(1+\tau_{cht})\cdot x_{ijh,t-1}$
$+\alpha_{2}\,s_{ijh,t-1}+\alpha_{3}\,x_{ijh,t-1}+\textbf{X}_{ijht}\boldsymbol{\gamma}\,+\mathbf{FE}+\epsilon_{ijht}.$ (3.2)
where we use the change in the duty-exclusive price as the dependent variable, defined as $\Delta\ln p_{ijht}=\Delta\ln p_{ijht}^{\text{duty}}-\Delta\ln(1+\tau^{\text{app}}_{cht})$, where $\tau^{\text{app}}_{cht}$ is the applied ad-valorem tariff. This transformation isolates price changes net of applied duties, mitigating measurement error.
The interaction terms capture heterogeneity in pass-through with respect to bilateral market shares, $s_{ijh,t-1}$ and $x_{ijh,t-1}$, measured at the beginning of the period. The vector $\textbf{X}_{ijht}$ includes controls for changes in exporter $i$’s sales to other U.S. buyers and the average price change faced by importer $j$ from alternative suppliers, helping to isolate bilateral responses emphasized in the model.
We estimate two specifications. The first includes product–time ($\text{FE}_{ht}$) and exporting country–sector ($\text{FE}_{cs}$) fixed effects, following standard practice in the pass-through literature. This is our baseline. The second is more demanding, adding importer–time ($\text{FE}_{jt}$) and exporting country–time ($\text{FE}_{ct}$) fixed effects.
Table 3: Pass-Through and Relationship Heterogeneity
Dependent variable: $\Delta\ln p_{ijht}$
(1) (2) (3) (4) (5) (6)
$\Delta\ln(1+\tau_{cht})$0.151 -0.188 -0.171 -0.045 -0.066 -0.123
(0.093) (0.105) (0.095) (0.099) (0.093) (0.104)
$\Delta\ln(1+\tau_{cht})\cdot\ln\text{longevity}_{ijht}$ 0.026 0.042
(0.019) (0.018)
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.050 0.054 0.053
(0.075) (0.072) (0.071)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.403 -0.403 -0.411
(0.113) (0.113) (0.114)
$\text{FE}_{ht}+\text{FE}_{cs}$ Yes Yes Yes Yes Yes Yes
Observations 249,000 249,000 249,000 249,000 249,000 249,000
R-squared 0.04 0.04 0.04 0.04 0.04 0.04
Notes: This table reports estimates of the pass-through of statutory tariffs, $\Delta\ln(1+\tau_{cht})$, to duty-exclusive prices at the exporter–importer–product–year level, $\Delta\ln p_{ijht}$. Columns (2) and (6) interact tariffs with the log of relationship longevity, measured as the number of years that buyer $j$ and supplier $i$ have transacted in product $h$. Columns (3) and (5) interact tariffs with the lagged supplier share, $s_{ijht-1}$, defined as supplier $i$’s share in buyer $j$’s imports of product $h$. Columns (4) and (5) interact tariffs with the lagged buyer share, $x_{ijht-1}$, defined as buyer $j$’s share in supplier $i$’s exports of product $h$. All regressions include product–year and exporter country–sector fixed effects ($\text{FE}_{ht}+\text{FE}_{cs}$). Controls include: (i) $\ln\text{longevity}_{ijht}$; (ii) $\Delta\ln q_{i(-j)ht}$, the change in exporter $i$’s total sales of $h$ to U.S. buyers other than $j$; and (iii) $\Delta\ln p_{(-i)jht}$, the weighted average price change charged by other suppliers of $h$ to buyer $j$, using lagged shares as weights. Standard errors are clustered at the HS8 product and exporter-country level. The sample corresponds to the ”+ Supplier Multi-Buyer” definition in Table C.1. Observation counts are rounded to four significant digits per U.S. Census Bureau disclosure guidelines. See Table D.11 for results using an alternative definition of arm’s-length trade based on LFTTD. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Table 3 presents the results using the baseline fixed effects. Column (1) shows that, on average, pass-through into duty-exclusive prices is incomplete: a 10% tariff increase reduces exporter prices by 1.5% , corresponding to 85% pass-through rate. Column (2) adds an interaction with relationship age to account for the role of match longevity, which has been shown to influence price adjustments. The results confirm that pass-through rises with relationship length, consistent with the results in Heise (2024).
Columns (3)–(6) show that the coefficient on supplier share ($\alpha_{s}$) is positive but insignificant, suggesting limited pass-through heterogeneity on the exporter side. In contrast, the coefficient on buyer share ($\alpha_{x}$) is consistently negative and statistically significant, indicating that importers with greater buyer share face lower pass-through. Table D.2 in Appendix D.2 confirms that these patterns hold under more demanding fixed effects. Table D.3 shows they are robust to alternative price definitions and general equilibrium controls.
To assess nonlinearities, we interact tariff changes with quartiles of lagged supplier and buyer shares. Figure D.1 shows no systematic pattern across supplier share quartiles but a clear, monotonic decline across buyer share quartiles, regardless of fixed effects.
Together, these results suggest that pass-through is largely unresponsive to supplier concentration but declines strongly with buyer concentration. This pattern aligns with the model’s predictions under two-sided market power: the strong buyer share gradient provides direct evidence of decreasing returns, as formalized in Proposition 4, while the weak supplier share gradient reflects the dominance of cost-channel, which suggests a high-$\phi$, low-$\theta$ environment.
4 Structural Estimation
The patterns documented above are consistent with the model’s predictions under two-sided market power. However, as discussed, they do not warrant direct inference on the structural parameters. We now turn to a structural estimation approach to quantify the role of two-sided market power in shaping international prices.
Our estimation targets two key parameters: the importer’s bargaining power, $\phi$, and the returns to scale parameter, $\theta$, which governs the elasticity of foreign export supply. The remaining parameters are taken from the literature or directly measured from the data.
We set the elasticity of substitution across foreign varieties to $\rho=10$, consistent with Anderson and van Wincoop (2004) and Edmond et al. (2023), who adopt similar values to match observed U.S. markups. The downstream demand elasticity faced by importers is set to $\nu=4$, based on the estimates in Broda and Weinstein (2006). [^23] Appendix E.1 provides further discussion. The elasticity of importer $j$’s marginal cost with respect to the foreign input price index is set to $\gamma=0.5$, following Eldridge and Powers (2018), who document the share of imported inputs in total material costs for U.S. manufacturers. Input cost shares $\alpha_{jht}$ are directly measured from the data.
Finally, we set the importer’s returns to scale parameter to $\varrho=1$ in the baseline specification. As discussed in Section 2.3.1, the parameters $\nu$, $\gamma$, and $\varrho$ jointly determine the elasticity $\eta$ of importer $j$’s foreign input bundle $q^{f}_{j}$ with respect to its price index $p^{f}_{j}$. Since $\varrho$ enters the model only through $\eta$, fixing it to one does not restrict generality. We verify that our results are robust to alternative values of $\eta$.
4.1 Identification and Estimation of the Parameters $\theta$ and $\phi$
Let $\Omega_{ijt}$ denote the information set available to a given $i-j$ pair during negotiations. This includes observed market shares ($s_{ijht}$, $x_{ijht}$, $\alpha_{jht}$) and calibrated parameters ($\nu$, $\gamma$, $\rho$, $\varrho$). As shown in equation (2.6), the bilateral markup depends only on the model primitives $(\phi,\theta)$ conditional on $\Omega_{ijt}$, i.e., $\mu_{ij}=\mu\left(\phi,\theta\mid\Omega_{ijt}\right).$ The log price of product $h$ transacted between exporter $i$ and importer $j$ in year $t$ can thus be written as:
$\ln p_{ijht}=\ln\mu\left(\phi,\theta\mid\Omega_{ijt}\right)+\ln c_{iht},$
where $\ln c_{iht}$ denotes the exporter’s marginal cost. From equation (2.1), this is given by:
$\ln c_{ijht}=\frac{1-\theta}{\theta}\ln q_{iht}+\ln k_{iht}.$
In the theoretical model, the term $k_{iht}$ captured exporter-level cost shifters. In the empirical implementation, we generalize this term to allow for match-specific cost components, such as relationship-specific know-how or specialization, by letting it vary flexibly at the match-year level, $k_{ijht}$, thereby absorbing all $(i,j,h,t)$-specific variation. In contrast, the term $\frac{1-\theta}{\theta}\ln q_{iht}$ is constant across importer matches for a given supplier–product–year. Identification of $(\phi,\theta)$ thus relies on cross-sectional variation in bilateral market shares, as discussed next.
Consider an exporter $i$ matched with two importers, $j$ and $\ell$. Conditional on the joint information set $\boldsymbol{\Omega}_{ij\ell t}\equiv(\Omega_{ijt},\Omega_{i\ell t})$, we assume that the unobserved component of marginal cost is mean-independent of the buyer identity: $\Delta k_{ij\ell ht}\equiv\mathbb{E}_{k}\left[k_{ijht}-k_{i\ell ht}\mid\boldsymbol{\Omega}_{ij\ell t}\right]=0.$ [^24] Omitted variables may induce $\Delta k_{ij\ell ht}\neq 0$, raising endogeneity concerns. One possible issue is endogenous network formation: unobserved factors may simultaneously affect both match formation and pricing, biasing estimates of $\hat{\phi}$ and $\hat{\theta}$. We address this by differencing across buyers of the same exporter, which removes exporter-level shocks common to all matches. To address remaining concerns, we further implement an instrumental variable strategy. Taking log price differences across buyers $j$ and $\ell$ served by the same exporter $i$ in year $t$ yields the following moment condition:
$g(\phi,\theta\mid\boldsymbol{\Omega}_{ij\ell t})\equiv\mathbb{E}_{k}\Big{[}\ln p_{ijht}-\ln p_{i\ell ht}-\big{(}\ln\mu(\phi,\theta\mid\Omega_{ijt})-\ln\mu(\phi,\theta\mid\Omega_{i\ell t})\big{)}\Big{|}\boldsymbol{\Omega}_{ij\ell t}\Big{]}=0.$ (4.1)
Identification requires that equation (4.1) does not hold for two pairs ($\phi,\theta)$ such that $(\phi^{A},\theta^{A})\neq(\phi^{B},\theta^{B})$. Since the oligopoly markup is independent of the returns to scale parameter $\theta$, identification of $\theta$ requires that the oligopsony channel plays a role in price determination, i.e., $\phi>0$, a condition supported by the reduced-form evidence in Tables 2 and 3.
We therefore focus on the empirically relevant case of bilateral bargaining power $\phi\in(0,1)$. The markup function $\mu(\phi,\theta\mid\Omega)$ is strictly monotonic in both parameters and thus invertible in each. It follows that the moment condition in equation (4.1) is also invertible in $\phi$ and $\theta$. Identification then relies on observing multiple importer–exporter pairs in the same year, or multiple matches for a given exporter over time, under the assumption that bargaining weights remain constant across matches. [^25] Formally, identification relies on the nonlinearity of the markup equation (2.6) in $s_{ijht}$ and $x_{ijht}$. Consider moment conditions from two periods $t$ and $t-1$: the associated derivatives with respect to $(\phi,\theta)$ are not collinear, satisfying the full-rank condition. Similar variation across multiple matches for the same exporter in a given year (e.g., $i$$j$$k$ vs. $i$$j$$\ell$) also secures identification.
Importantly, our identification strategy does not hinge on strong assumptions about the exogeneity of tariffs or other aggregate variables. While the estimated parameters may be shaped by general equilibrium forces, identification relies solely on cross-sectional variation across buyers within supplier–product–year cells. As a result, we do not need to specify or model the broader general equilibrium environment.
Estimation
We estimate equation (4.1) via generalized method of moments (GMM),
$\min_{\{\mathbf{{\phi}},\theta\}}\,\,\mathbf{g}(\mathbf{{\phi}},\theta)\ \mathbf{Z}^{\prime}\ \mathbf{W}\ \mathbf{Z}\ \mathbf{g}(\mathbf{{\phi}},\theta)^{\prime},$ (4.2)
where $\mathbf{g}(\mathbf{\phi},\theta)$ stacks all moment conditions in equation (4.1) across all $i$$j$$\ell$ pairs and years and $\mathbf{W}$ is the optimal weighting matrix. [^26] Appendix E.2 presents Monte Carlo simulations based on a data-generating process that mirrors the setup in Section 2. The results confirm that our estimators is consistent.
To address endogeneity concerns, we first include fixed effects by demeaning $\mathbf{g}({\phi},\theta)$ at the HS10 product, year, and buyer level. This removes average variation across those dimensions, so that only time-varying, pair-specific shocks could bias $\Delta k_{ij\ell ht}$. In addition, we employ instrumental variables ($\mathbf{Z}$) that are plausibly exogenous with respect to the network formation process and other omitted variables.
In particular, the vector $\mathbf{Z}$ includes the total number of importers and exporters in each HS10 product-year, which we interpret as proxies for the pool of potential US buyers and foreign suppliers in a given variety. We also include in $\mathbf{Z}$ the mean and median of the distributions of the two bilateral shares within each year, excluding the focal pairs $i-j$ and $i-\ell$ to preserve over-identification. These instruments vary with the competitive structure within each HS10 product-year and are correlated with the endogenous variables through market structure, but, by construction, are not correlated with the idiosyncratic shocks affecting individual matches.
Extension: pair-specific bargaining weights
While our baseline assumes a constant bargaining weight $\phi$ across all importer–exporter pairs, we also consider an extension to allow $\phi$ to vary at the pair level.
Given the large number of trade pairs in the data, estimating a separate $\phi_{ij}$ for each is computationally burdensome. Moreover, our identification strategy does not allow bargaining weights to vary both across pairs and over time. We therefore model bargaining power as a function of observable characteristics:
$\phi_{ijt}=\frac{\exp\left({\bf X}_{ijht}\,\mathbf{\boldsymbol{\kappa}}\right)}{1+\exp\left({\bf X}_{ijht}\,\mathbf{\boldsymbol{\kappa}}\right)}\in[0,1],$ (4.3)
where $\boldsymbol{\kappa}$ is a parameter vector to be estimated and $\mathbf{X}_{ijht}$ includes covariates that plausibly influence bargaining outcomes but are not direct determinants of gains from trade in our model. Specifically, we include: (i) the longevity of the $i-j$ relationship, (ii) the number of transactions between $i-j$ in a year, (iii) the relative outside option of the two, measured by the ratio of the quantity of the exporter $i$’s sales to buyers other than $j$ in year $t-1$ over the quantity of the importer $j$’ purchases from suppliers other than $i$ in year $t-1$, and (iv) an indicator variable of whether the buyer and supplier transact multiple HS10 products.
4.2 Estimation Results
We estimate equation (4.2) using data from 2001 to 2016. We exclude 2017 and 2018, as these years will be used to validate the model out-of-sample in Section 4.3, leveraging the tariff shocks that occurred during this period. To avoid convergence issues when $\phi$ is near one, we estimate the transformed parameter $\bar{\phi}\equiv\ln\frac{\phi}{1-\phi}$, which enters the markup equation linearly.
Table 4: Estimated Model Primitives
Panel A: Calibrated Parameters $\hat{\nu}$ $\hat{\gamma}$ $\hat{\rho}$ 4 0.5 10 Panel B: Estimated Parameters (GMM) (1) (2) (3) (4) Rel. bargaining power: $\ln\widehat{\frac{\phi}{1-\phi}}$ 1.565 0.863 (0.055) (0.043) Returns to scale ($\hat{\theta}$) 0.454 0.497 0.383 0.502 (0.004) (0.006) (0.006) (0.007) Constant 4.118 1.454 (0.428) (0.180) Longevity -0.360 0.332 (0.062) (0.064) Number of HS10 transactions -0.264 -0.003 (0.029) (0.014) Multiple HS10 dummy -0.180 0.131 (0.047) (0.034) Lagged outside option -0.235 -0.230 (0.031) (0.030) None Yes Yes No No $\text{FE}_{h}+\text{FE}_{t}+\text{FE}_{j}$ No No Yes Yes Observations 3,120,000 Panel C: Implied Bargaining Powers ($\hat{\phi}$) Mean 0.827 0.922 0.703 0.860 (0.008) (0.074) (0.009) (0.099) Median 0.945 0.886 (0.074) (0.099) Notes: This table presents model estimates based on our main estimation sample, which focuses on U.S. imports of intermediate inputs and capital goods for the period 2001-2016. Panel A reports calibrated parameters: the elasticity of demand ($\nu$), the elasticity of costs with respect to foreign input prices ($\gamma$), and the elasticity of substitution across foreign varieties ($\rho$). We set $\varrho=1$, so that $\eta=2.5$. Panel B presents GMM estimates. Columns (1) and (3) impose a constant $\phi$ across bilateral pairs, while Columns (2) and (4) estimate the full vector $\boldsymbol{\kappa}$ to allow for heterogeneity in bargaining power. Specifications differ in the inclusion of fixed effects. Controls include: (i) the log of relationship longevity between exporter $i$ and importer $j$; (ii) the log of the number of transactions between $i$ and $j$ in a given year; (iii) the log of the relative outside option, defined as the ratio of exporter $i$’s sales to other U.S. buyers (excluding $j$) over importer $j$’s purchases from other suppliers (excluding $i$), both in year $t{-}1$; and (iv) a dummy variable equal to one if the $i-j$ pair transacts in more than one HS10 product. Panel C reports the mean and median of the implied bargaining power. Standard errors are robust; those in Panel C are computed using the delta method. The set of instruments includes the number of exporters and importers at the HS10 level, as well as lagged bilateral shares (excluding the focal pair). The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Table 4 presents the estimation results. Panel B reports the GMM estimates. Columns (1) and (3) assume a constant $\phi$, while Columns (2) and (4) allow $\phi_{ij}$ to vary by trade pair as specified in equation (4.3). The specifications in Columns (1) and (2) are estimated without fixed effects; those in Columns (3) and (4) include year, product, and importer fixed effects. Panel C shows the implied values of $\phi$ or $\phi_{ijt}$.
The parameters are precisely estimated. Across specifications, U.S. importers appear to wield substantial bargaining power, with estimated values of $\phi$ ranging from 0.70 to 0.92. Our preferred estimate, reported in Column (1), is $\hat{\phi}=0.83$, implying that U.S. importers have, on average, roughly four times the bargaining power of their foreign suppliers. [^27] These findings are consistent with evidence from related settings. Morlacco (2019) documents significant markdowns by French importers in input trade, while Atkin et al. (2024) show that Argentine importers often exercise considerable bargaining power.
The returns to scale parameter $\hat{\theta}$ is consistently estimated below one, ranging from 0.40 to 0.50 across specifications, with a preferred estimate of 0.45.282828While standard estimates of returns to scale often cluster near one, $\theta$ here captures a short-run returns to scale elasticity, i.e., the slope of the exporter’s marginal cost curve during bargaining. This implies a residual export supply elasticity between 0.25 and 0.375 for the average importer, indicating relatively steep foreign supply curves. [^29] The implied residual supply elasticity is computed from the marginal cost slope $c^{\prime}_{q_{ijh}}=\frac{1-\theta}{\theta}\cdot x_{ijh}$, using the average buyer share of 0.25 from Table 1. These values are consistent with evidence from U.S. manufacturing under short-run constraints: Boehm and Pandalai-Nayar (2022) report median inverse elasticities around 0.3 at typical capacity levels, and Broda et al. (2008) document similarly low elasticities across many traded goods.
Moving to the estimates of the vector $\hat{{\boldsymbol{\kappa}}}$, we find that the coefficients on relationship longevity, frequency of transactions, and the multiple-product indicator are highly significant, although their sign varies depending on the set of fixed effects included. By contrast, the coefficient on the relative outside option is stable across specifications: importers hold less bargaining power when their supplier has a stronger outside option. Specifically, an increase in the supplier’s past sales to other buyers relative to the importer’s purchases from other suppliers is consistently associated with lower bargaining power for the importer.
Robustness
We assess the robustness of our structural estimates to alternative sample definitions and model calibrations. Appendix D.3 first considers a broader sample that includes all products in the BEC classification, notably extending the baseline by adding consumption goods. We also examine an alternative sample that uses related-party trade indicators from the LFTTD instead of ORBIS. On the calibration side, we consider a lower elasticity of substitution across foreign varieties ($\rho=5$ instead of 10) and introduce decreasing returns to scale in downstream production by setting $\varrho=0.5$ rather than 1. This choice aligns with the estimated returns to scale on the exporter side and allows us to test the robustness of the estimates to alternative values of the importer’s downstream demand elasticity, $\eta$. Across all variations, the structural estimates remain highly stable.
Implied Markups
Using equation (2.6), we compute markups for all buyer–supplier–product matches given the estimated parameters and the observed distribution of market shares.
The resulting markup distribution is clustered near the competitive benchmark. Our preferred estimates yield a mean markup of 0.94, with the median even closer to competitive levels. These low markups reflect the strong countervailing power of importers. The model implies that observed markups are a convex combination of oligopoly and oligopsony markups, which average 1.34 and 0.87, respectively. The large estimated bargaining power of buyers ($\hat{\phi}\approx 0.8$) shifts weight toward the oligopsony case, allowing importers to extract a substantial share of the surplus from exporters. [^30] The effective bargaining weight $\omega_{ijh}$ averages 0.77 (standard deviation 0.05), slightly below $\hat{\phi}$. This indicates that network effects, on average, dampen the importers’ effective bargaining power relative to $\phi$, although the gap is small.
4.3 Model Validation
We assess the model’s empirical validity by testing its ability to predict both the level and heterogeneity of price and quantity changes following the 2017–2018 tariff increases. These moments were not targeted in the estimation.
Model-predicted price changes are computed as:
$\widehat{\Delta\ln p_{ijht}}=\Phi_{ijht}(s_{ijht},x_{ijht}\mid\boldsymbol{\hat{\Theta}})\cdot\Delta\ln(1+\tau_{cht}),$ (4.4)
where $\tau_{cht}$ denotes the ad-valorem tariff on product $h$ from country $c$, and $\Phi_{ijht}$ is the model-implied pass-through elasticity defined in equation (2.11), which depends on bilateral shares and the estimated parameter vector $\boldsymbol{\hat{\Theta}}$.
Although the model is primarily designed to explain prices, it embeds a demand-driven allocation rule that links prices to quantities (and sales) via the importer’s demand curve. This mapping imposes a specific quantity-setting structure, enabling a direct comparison between predicted price responses and observed adjustments in quantities and trade values. Predicted quantity changes are given by $\widehat{\Delta\ln q_{ijht}}=-\hat{\varepsilon}_{ijht}\cdot\widehat{\Delta\ln p_{ijht}},$ where $\hat{\varepsilon}_{ijht}$ denotes the match-specific residual demand elasticity implied by the model. For sales, the corresponding mapping is $\widehat{\Delta\ln r_{ijht}}=-(1-\hat{\varepsilon}_{ijht})\cdot\widehat{\Delta\ln p_{ijht}}$.
Table 5: Price Responses and Relationship Heterogeneity: Data vs. Model
Panel A: Data
(1) (2) (3) (4)
$\Delta\ln(1+\tau_{cht})$0.151 -0.066 -0.223 -0.163
(0.093) (0.093) (0.109) (0.107)
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.054 0.029
(0.072) (0.157)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.403 -0.271
(0.113) (0.135)
R-squared 0.04 0.04 0.31 0.31
Panel B: Model
(1) (2) (3) (4)
$\Delta\ln(1+\tau_{cht})$0.248 -0.144 -0.249 -0.136
(0.008) (0.009) (0.010) (0.011)
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.104 0.091
(0.011) (0.010)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.475 -0.486
(0.026) (0.029)
R-squared 0.32 0.46 0.50 0.59
$\text{FE}_{ht}+\text{FE}_{cs}$ Yes Yes No No
$\text{FE}_{ht}+\text{FE}_{ct}+\text{FE}_{jt}$ No No Yes Yes
Observations 249,000
Notes: This table reports the pass-through of tariffs to duty-exclusive prices at the exporter–importer–product level. Panel A presents reduced-form estimates from the data. Panel B shows corresponding pass-through estimates generated by the model. Columns (2)–(4) interact tariff changes with lagged supplier share ($s_{ijht-1}$) and lagged buyer share ($x_{ijht-1}$). Columns (1) and (2) use baseline fixed effects ($\text{FE}_{ht}+\text{FE}_{cs}$), while Columns (3) and (4) employ a more stringent specification with product–year, country–year, and buyer–year fixed effects ($\text{FE}_{ht}+\text{FE}_{ct}+\text{FE}_{jt}$). Standard errors are clustered at the HS8 product and exporter-country level. Observation counts are rounded to four significant digits per U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
4.3.1 Price Predictions
We begin with price outcomes. Table 5 compares tariff pass-through elasticities in the data (Panel A) and in the model (Panel B), in terms of average effects and their heterogeneity with respect to buyer and supplier shares. Columns (1) and (2) adopt a baseline specification with product–time and country–sector fixed effects, while Columns (3) and (4) adopt the more demanding specification with buyer–time and country–time fixed effects.
The model predicts an average pass-through elasticity on duty-exclusive prices of $-0.25$, which translates to a 75% pass-through rate (Columns (1) and (3) of Panel B). These values fall within the 95% confidence intervals of the corresponding reduced-form estimates in the data (Columns (1) and (3) of Panel A).
Columns (2) and (4) of Panel A, consistent with Column (6) of Table 3, show that pass-through declines significantly with buyer share ($x_{ijht}$) but exhibits no robust relationship with supplier share ($s_{ijht}$). The model closely replicates these patterns: as shown in Panel B, pass-through declines steeply in $x_{ijht}$ and increases mildly in $s_{ijht}$, with magnitudes comparable to those in the data. This alignment supports our interpretation of Table 3 as consistent with a low-$\theta$, high-$\phi$ environment, which is also consistent with our structural estimates.
4.3.2 Goodness-of-fit Test
Having shown that the model replicates average and heterogeneous price responses to tariff shocks, we formally test its predictive performance by evaluating how well model-implied price changes explain observed variation. This exercise complements Table 5 by providing a direct measure of goodness of fit relative to standard alternatives.
Specifically, we estimate:
$\Delta\ln p_{ijht}=\beta\,\,\widehat{\Delta\ln p_{ijht}}\,+\,\mathbf{FE}\,+\,u_{ijht},$ (4.5)
where $\Delta\ln p_{ijht}$ is the observed change in the duty-inclusive price for product $h$ between exporter $i$ and importer $j$, and $\widehat{\Delta\ln p_{ijht}}$ is the corresponding model-predicted change.
To benchmark the results, we compare the full model to three nested alternatives that sequentially shut down bilateral bargaining ($\phi=0$), decreasing returns to scale ($\theta=1$), or both. The fully restricted case ($\phi=0$, $\theta=1$) corresponds to a standard Nash-Bertrand model with constant marginal costs (e.g., Dhyne et al., 2022); the intermediate case with $\phi>0$ and $\theta=1$ mirrors the bargaining framework of Gopinath and Itskhoki (2011), which overlooks the heterogeneous buyer dimension. In all cases, we use the same parameter values from Column (1) of Table 4, without re-estimating. [^31] As discussed in Section 4.1, $\theta$ is not identified when $\phi=0$. Moreover, fixing $\theta=1$ and estimating equation (4.5) using the value of $\phi$ jointly obtained with $\theta$ from equation (4.2) yields an upper bound on the value of $\hat{\beta}$ attainable with a re-estimated $\phi$. Appendix E.3 provides formal proof and supporting simulation evidence.
A key challenge is that observed price changes may reflect shocks unrelated to tariffs, increasing the risk of rejecting a valid model for reasons unrelated to its tariff predictions. In addition, while the model uses statutory tariff changes, observed duty-inclusive prices reflect actual duties paid, potentially introducing measurement error. To address both issues, we follow the logic of the IV-based goodness of fit test in Adão et al. (2023) and estimate equation (4.5) using two-stage least squares, instrumenting $\widehat{\Delta\ln p_{ijht}}$ with statutory tariff changes. This isolates variation in predicted prices that is directly attributable to tariff shocks, which is the one targeted by the model. Under the null that the model accurately captures pass-through, the IV coefficient $\hat{\beta}$ should equal one.
Figure 3: IV-Based Goodness-of-Fit Test
Notes: Each point reports the coefficient from an IV regression of observed log price changes on model-predicted changes $\widehat{\Delta\ln p_{ijht}}$, using statutory tariffs as instruments. Lines show 95% confidence intervals. Blue and red denote regressions with $\mathrm{FE}_{ht}+\mathrm{FE}_{cs}$ and $\mathrm{FE}_{ht}+\mathrm{FE}_{ct}+\mathrm{FE}_{jt}$, respectively. Standard errors are clustered by product and exporter-country. Observation counts (249,000) are rounded per Census disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Figure 3 presents the results across the four model variants, each estimated under two alternative fixed effects specifications. Blue points and lines correspond to regressions with product–time and country–sector fixed effects, while red ones use the more stringent specification with product–time, country–time, and buyer–time fixed effects. In all cases, the baseline model provides the best fit: the estimated coefficient is very close to one and not statistically different from it under both specifications.
Models with constant returns to scale ($\theta=1$), whether or not they include bargaining, perform noticeably worse, although their coefficients are not rejected at conventional significance levels. These variants lack a cost channel, and while the case with $\phi=0$ features strategic complementarities that generate some degree of incomplete pass-through, this mechanism alone does not produce sufficient variation to match the data. Similarly, the fourth model with decreasing returns but no bargaining ($\phi=0$, $\theta<1$), which was previously rejected based on price-level evidence (Table 2), also underperforms relative to the baseline model. In this case, markup responses driven by strategic complementarities amplify cost-based adjustments, but in a way that also fails to replicate the observed pass-through patterns. [^32] Appendix D.4 (Table D.6) presents additional robustness checks using alternative calibrations, including $\rho=5$ instead of 10 and $\varrho=0.5$ instead of 1. The model’s predictive performance remains stable across these variations, lending further support to our main findings.
Overall, these results suggest that cost adjustments, along with weak strategic complementarities, are necessary to account for the empirical evidence. This is consistent with a setting in which oligopsony forces are dominant.
4.3.3 Quantity Predictions
Next, we assess the model’s ability to predict changes in bilateral quantities. Table D.7 in Appendix D.4 reports relationship-level quantity responses to tariff changes. Columns (3)–(4) of Panel A show that tariff increases reduce traded volumes, with larger declines under more stringent fixed effects. However, the interaction terms with supplier and buyer shares are statistically imprecise, suggesting inconclusive evidence of heterogeneity in the data. [^33] In robustness exercises (not shown), the signs on the interaction coefficients with supplier and buyer shares occasionally flip, though they largely remain statistically insignificant. We therefore conclude that the data do not offer robust evidence of heterogeneous quantity responses by relationship structure.
Panel B shows that the model generates sizable average quantity declines and predicts heterogeneity across relationships. In particular, quantity responses become less negative with higher buyer shares, consistent with the model’s allocative logic: stronger buyers face smaller price increases and thus smaller quantity reductions. In contrast, the positive interaction with supplier share is not explained by price adjustments because the model predicts little variation in pass-through along this dimension. Instead, it reflects the curvature of the importer’s demand curve embedded in the model.
Table D.8 in Appendix D.4 evaluates model fit for quantities (Panel A) and sales (Panel B) across the four alternative parameterizations discussed above. Since all models share the same demand system, performance differences reflect variation in price predictions rather than differences in quantity mechanisms. While all specifications are formally rejected, the baseline model and the specification with decreasing returns to scale perform best, mirroring their superior performance in prices.
Despite differences in magnitude and precision, the model captures the broad directional patterns in quantity responses, lending support to its core allocative mechanism. However, its weaker quantitative fit suggests that additional forces beyond price-based allocation shape the observed quantity adjustments. The model’s strong performance on prices highlights this asymmetry and points to potential gains from extending the framework to incorporate supply-side responses.
5 Aggregate Implications of Bargaining in Firm-to-Firm Trade
This section examines the impact of the Trump tariffs on aggregate import prices through the lens of our bargaining model. We begin by situating our analysis within the broader literature on tariff pass-through. We then use the model to simulate the aggregate effects of the tariffs and assess the mechanisms driving price adjustments.
5.1 Comparison with Existing Pass-Through Estimates
Several recent studies find near-complete pass-through of the 2018 Trump tariffs to U.S. import prices (e.g., Fajgelbaum et al., 2020; Amiti et al., 2019b, 2020). These analyses primarily rely on product-level data. In contrast, our match-level estimates indicate incomplete pass-through. For instance, in a standard specification similar to those used in the literature, Column (1) of Table 5 (Panel A) reports a pass-through elasticity on duty-exclusive prices of $-0.15$, implying an 85% pass-through rate.
This divergence partly reflects differences in sample composition. As discussed in Section 3.3, our estimates are based on repeated firm-to-firm relationships where bilateral prices can be observed over time. They therefore capture within-relationship price changes net of compositional shifts across firms or products, rather than on changes in product-level unit values. In addition, we restrict attention to arm’s-length transactions involving intermediate goods and suppliers with two or more buyers. Within this sample, we estimate pass-through rates ranging from 78 to 85%, depending on the fixed effects used, as shown in Table 5. The model’s predicted pass-through closely matches these empirical patterns.
To further assess the role of sample composition, Table D.9 in Appendix D.5 reports pass-through estimates across alternative sample definitions. Expanding the baseline to include single-buyer relationships raises the estimate to approximately 83%. Adding matches involving related parties, energy goods, or extreme price changes increases it to roughly 90%. The most inclusive specification, which further incorporates final consumption goods, yields pass-through elasticities between 93% and 95%, depending on the fixed effects. Although these broader samples imply higher pass-through, they still fall short of full pass-through, in contrast to product-level studies. [^34] Using the monthly LFTTD data, we show in Table D.10 in Appendix D.5 that pass-through remains incomplete in product-level regressions when restricting to the firms and products used in our baseline sample, confirming that this is a feature of repeated firm-to-firm relationships.
These product-level estimates are based on aggregated unit values that combine within-match price changes with shifts in the composition of transactions, particularly the inclusion of one-off (spot) exchanges. Such transactions may differ systematically from repeated firm-to-firm relationships. In particular, incomplete pass-through may be a feature of relational trade, where prices reflect bilateral bargaining and upward-sloping residual supply. Spot exchanges, by contrast, are more likely to reflect marginal cost pricing and thus exhibit full pass-through. While speculative, this interpretation is consistent with Cajal-Grossi et al. (2023), who find lower markups and more competitive pricing in spot relationships.
5.2 Tariffs and Aggregate Import Prices
We use the estimated model to assess the impact of the 2018 tariffs on aggregate import prices and decompose the underlying contributions of markup and marginal cost changes. A full evaluation of the tariffs’ macroeconomic effects would require a general equilibrium framework that incorporates changes in expenditure, wages, export prices, and retaliation, as in Fajgelbaum et al. (2020). While such elements are essential for welfare analysis, they lie beyond the scope of our model.
Instead, we focus on one central component of tariff incidence: the change in aggregate import prices within ongoing firm-to-firm relationships. We compute model-implied bilateral price changes using equation (4.4) and estimate aggregate pass-through by regressing these predicted changes on the tariff shock, weighting each observation by its initial-period import value. The approach connects the model’s micro-level predictions to aggregate outcomes and provides a framework for decomposing the contribution of distinct adjustment channels.
To that end, we separately isolate the roles of markup and marginal cost adjustments. For each mechanism, we compute the model-implied price changes while holding the other channel constant, and re-estimate the weighted regression. The predicted price change due solely to markup adjustment is defined as:
$\widehat{\Delta^{\Gamma}\ln p_{ijht}}=\Phi^{\Gamma}_{ijht}(s_{ijht},\,x_{ijht}\,\mid\,\hat{\boldsymbol{\Theta}})\cdot\Delta\ln T_{cht},\quad\text{where}\quad\Phi^{\Gamma}_{ijht}\equiv\frac{1}{1+\Gamma_{ij}}.$
Similarly, the predicted price change driven solely by cost adjustment is given by:
$\widehat{\Delta^{\Lambda}\ln p_{ijht}}=\Phi^{\Lambda}_{ijht}(s_{ijht},\,x_{ijht}\,\mid\,\hat{\boldsymbol{\Theta}})\cdot\Delta\ln T_{cht},\quad\text{where}\quad\Phi^{\Lambda}_{ijht}\equiv\frac{1}{1+\Lambda_{ij}}.$
Table 6: Aggregate Tariff Pass-Through and Decomposition
Baseline FE Stringent FE
(1) (2)
Panel A: Aggregate Passthrough ()
Agg. pass-through elasticity 70.7 64.6
Cost channel only: $1/(1+\Lambda_{ij})$ 69.3 62.7
Markup channel only: $1/(1+\Gamma_{ij})$ 97.1 88.0
Panel B: Variance Decomposition of $\Lambda_{ij}+\Gamma_{ij}$
Cost Elasticity: $\Lambda_{ij}$ 1.01 0.99
Markup Elasticity: $\Gamma_{ij}$0.01 0.01
Notes: This table reports model-implied aggregate pass-through estimates following the 2018 U.S. tariff increases. Column (1) includes product–time and country–sector fixed effects. Column (2) includes product–time, country–time, and buyer–time fixed effects. The overall pass-through elasticity is computed as $1+$ the estimated coefficient on $\Delta\ln(1+\tau)$, and decomposed into contributions from the cost channel ($\Lambda_{ij}$) and the markup channel ($\Gamma_{ij}$). The counterfactual “Cost channel only” row shows the predicted pass-through when markup elasticities are set to zero, while the “Markup channel only” row sets cost elasticities to zero. Panel B reports the relative contribution of each channel to the cross-sectional variance of $\Lambda_{ij}+\Gamma_{ij}$, the total elasticity governing pass-through. These shares sum to one and are derived from a variance decomposition. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Panel A of Table 6 reports the aggregate tariff pass-through predicted by the model, which incorporates both markup and marginal cost adjustments: 71% under the baseline specification (Column (1)) and 65% under the alternative using product–time, country–time, and buyer–time fixed effects (Column (2)). These estimates are lower than the average pass-through of 75% obtained from unweighted regressions (Table 5), reflecting the more incomplete pass-through associated with larger, high-share buyers who account for a greater share of aggregate trade.
Nearly all of the predicted price response is driven by the cost channel. When markup elasticities are shut down (i.e., setting $\Gamma_{ij}=0$), the model yields pass-through rates between 69 and 63%, closely matching the full-model estimates. By contrast, shutting down the cost elasticities results in much higher pass-through, ranging from 88 to 97%.
Panel B offers an alternative approach to quantify the contribution of each channel by decomposing the variance of the term $\Lambda_{ij}+\Gamma_{ij}$, which governs the pass-through elasticity $1/(1+\Lambda_{ij}+\Gamma_{ij})$. This decomposition attributes the cross-sectional variation in pass-through to cost and markup elasticities, with their shares summing to one by construction. The results reinforce the earlier findings: nearly all of the variation is explained by the cost elasticity $\Lambda_{ij}$, underscoring its dominant role in shaping price responses.
Together, the results suggest a clear conclusion: tariff pass-through is incomplete, reflecting the combined effects of cost-side adjustments and bargaining. The underlying mechanism is intuitive. When tariffs reduce demand from large U.S. buyers, exporters move down their marginal cost curves, dampening the price impact of the shock. Although markups adjust endogenously, the opposing forces of strategic complementarities and substitutabilities limit their quantitative contribution to aggregate price change. This is a robust implication of our framework. As illustrated in Figure 2, when returns to scale are decreasing ($\theta<1$) and buyer power is high (large $\phi$), most of the variation in pass-through across the $(s_{ij},\,x_{ij})$ space is driven by cost adjustments, with a comparatively smaller role for markup responses.
5.3 The Importance of Bargaining in Firm-to-Firm Trade
Although bargaining plays a central role in our model, it appears to have limited influence on tariff pass-through, which primarily reflects cost-side adjustments. This is not because bargaining is unimportant, but because it endogenously mutes the markup channel. When buyer power is strong ($\phi$ high), the markup elasticity $(\Gamma_{ij})$ tends toward zero. As a result, pass-through is driven almost entirely by cost adjustments and appears relatively unresponsive to markup variation. In this sense, the weak role of the markup channel in shaping pass-through is itself a consequence of strong importer market power.
In addition, bargaining remains essential to understand price levels, markups, and the welfare consequences of market power. As shown in Table 2, two-sided market power is critical to explaining the cross-sectional variation in prices and markups. Moreover, as discussed in Section 2.5, $\phi$ governs how markup dispersion maps into misallocation. When exporters hold bargaining power, markup heterogeneity leads to underproduction. But when bargaining power shifts toward importers, the same dispersion can improve efficiency by reallocating output toward lower-cost suppliers. Ignoring bargaining would therefore yield misleading conclusions about the allocative effects of firm-to-firm trade.
6 Conclusions
Firm-to-firm relationships are a central feature of international trade. These relationships often involve market power on both sides, with prices determined through bilateral negotiations rather than market-clearing conditions. We develop a framework that departs from standard models of price-taking buyers by allowing both importers and exporters to influence price formation through bargaining. This approach yields analytical and empirical tools to study how market power and network frictions shape prices, markups, and the transmission of shocks in firm-to-firm trade.
Using transaction-level data, we show that U.S. importers wield substantial bargaining power and face upward-sloping residual supply, consistent with oligopsonistic behavior. Within ongoing firm-to-firm relationships, tariff pass-through during the 2018 trade war was incomplete, with exporters absorbing much of the tariff through cost-side adjustments. This muted price response is primarily driven by cost-side adjustments: while markups do adjust, their contribution to pass-through is limited because strong buyer power endogenously flattens the markup elasticity. In this sense, the limited role of markups in shaping tariff responses is itself a consequence of strong importer power.
Although not the primary focus of our analysis, the results highlight the importance of bargaining for understanding the allocative implications of market power in firm-to-firm trade. In our model, when supplier power dominates, markup dispersion reflects inefficiencies and underproduction. By contrast, when buyer power is strong, the same dispersion can lead to efficient reallocation and higher aggregate output. These findings suggest that markup heterogeneity does not uniformly signal misallocation, and that its welfare consequences depend critically on the distribution of bargaining power. Extending the framework to study these welfare implications more formally remains a promising avenue for future work.
More broadly, our study offers a foundation for analyzing how market power shapes the incidence and transmission of shocks through the price channel. While we focus on detailed price-setting within firm-to-firm relationships, Acemoglu and Tahbaz-Salehi (2025) emphasize the general equilibrium implications of market power in production networks in shaping aggregate fluctuations. A promising direction for future research is to integrate these perspectives by combining micro-level bargaining dynamics with macro-level spillovers to study the broader implications of buyer and supplier power, including in domestic supply chains.
Appendix A Mathematical Derivations
This appendix derives the key equations from the main text: the bilateral markup $\mu_{ij}$ (equation (2.6)) and the bilateral pass-through elasticity $\Phi_{ij}$ (equation (2.11)).
A.1 Proof of Proposition 1
To derive $\mu_{ij}$, we first obtain the demand elasticity $\varepsilon_{ij}$ from the importer’s cost minimization problem, then solve the bilateral bargaining problem.
Demand Elasticity $\varepsilon_{ij}$
Our baseline model assumes that the importer first chooses the input quantity to minimize costs given a price, then negotiates the price bilaterally with the exporter. The importer’s nested CES production structure (equations (2.2)–(2.3)) yields the demand for input $q_{ij}$ in (2.4):
$q_{ij}=q_{j}^{f}\varsigma_{ij}^{\rho}\left(\frac{p_{ij}}{p_{j}^{f}}\right)^{-\rho},$
where $p_{j}^{f}=\left(\sum_{i}\varsigma_{ij}^{\rho}p_{ij}^{1-\rho}\right)^{\frac{1}{1-\rho}}$ is the price index (i.e., shadow cost) of imported inputs. Solving the outer CES aggregator for total input demand yields:
$q_{j}^{f}$ $=\gamma c_{j}q_{j}\left(p_{j}^{f}\right)^{-1},$ (A.1)
$q_{j}^{d}$ $=(\varrho-\gamma)c_{j}q_{j}\left(p_{j}^{d}\right)^{-1},$ (A.2)
where $c_{j}$ is the unit cost of output $q_{j}$, given by:
$c_{j}=\left[\varphi_{j}^{-1}\left(\frac{p_{j}^{f}}{\gamma}\right)^{\gamma}\left(\frac{p_{j}^{d}}{\varrho-\gamma}\right)^{\varrho-\gamma}\right]^{\frac{1}{\varrho}}q_{j}^{\frac{1-\varrho}{\varrho}}\equiv k_{j}q_{j}^{\frac{1-\varrho}{\varrho}}.$ (A.3)
This form illustrates symmetry with exporter-side technology. Finally, note that the cost share of foreign inputs is constant and equal to
$\frac{p_{j}^{f}q_{j}^{f}}{p_{j}^{f}q_{j}^{f}+p_{j}^{d}q_{j}^{d}}=\frac{\gamma}{\varrho}.$
A relevant object for our derivations will be $\frac{d\ln c_{j}}{d\ln p_{j}^{f}},$ namely, the elasticity of the marginal cost $c_{j}$ with respect to $p_{j}^{f}.$ To find this elasticity, we first use the demand function downstream to write $q_{j}$ as
$q_{j}=\left(\frac{\nu}{\nu-1}\right)^{-\nu}c_{j}^{-\nu}D_{j},$ (A.4)
where $D_{j}$ is the firm-level demand shifter. Substituting equation (A.4) into equation (A.3) and rearranging, we can write:
$c_{j}$ $=\left[z_{j}^{-1}\left(\frac{p_{j}^{f}}{\gamma}\right)^{\gamma}\left(\frac{p_{j}^{d}}{\varrho-\gamma}\right)^{(\varrho-\gamma)}\right]^{\frac{1}{\varrho+\nu-\nu\varrho}}\left(\frac{\nu}{\nu-1}\right)^{-\nu\frac{1-\varrho}{\varrho+\nu-\nu\varrho}}\left(D_{j}\right)^{\frac{1-\varrho}{\varrho+\nu-\nu\varrho}},$
which implies
$\frac{d\ln c_{j}}{d\ln p_{j}^{f}}=\frac{\gamma}{\varrho+\nu-\nu\varrho}.$
Armed with these equations, we proceed to find the elasticity of interest. Given the log demand:
$\ln q_{ij}=\ln q_{j}^{f}+\rho\ln\varsigma_{ij}-\rho\left(\ln p_{ij}-\ln p_{j}^{f}\right),$
and equation (A.1), we find equation (2.9):
$\varepsilon_{ij}=-\frac{d\ln q_{ij}}{d\ln p_{ij}}$ $=-\left(\frac{d\ln q_{j}^{f}}{d\ln p_{j}^{f}}+\rho\right)\frac{d\ln p_{j}^{f}}{d\ln p_{ij}}+\rho$
$=\left((\nu-1)\frac{d\ln c_{j}}{d\ln p_{j}^{f}}+1-\rho\right)s_{ij}+\rho$
$=\eta s_{ij}^{f}+(1-s_{ij}^{f})\rho,$ (A.5)
where we defined
$\eta$ $\equiv-\frac{d\ln q_{j}^{f}}{d\ln p_{j}^{f}}=\frac{(\nu-1)\gamma}{\varrho+\nu(1-\varrho)}+1.$
$=\frac{(\varrho-\gamma)+\nu(1-(\varrho-\gamma))}{\varrho+\nu(1-\varrho)}.$
Equilibrium Price
The problem of the $i-j$ pair is to choose a bilateral price $p_{ij}$ that solves the following problem:
$\max_{p}\Big{(}\underbrace{\pi_{i}\left(p\right)-\tilde{\pi}_{i\left(-j\right)}}_{GFT_{ij}^{i}\left(p\right)}\Big{)}^{1-\phi}\Big{(}\underbrace{\pi_{j}\left(p\right)-\tilde{\pi}_{j\left(-i\right)}}_{GFT_{ij}^{j}\left(p\right)}\Big{)}^{\phi},$ (A.6)
where $\phi\in(0,1)$ is $j$’s bargaining power, and the terms inside parentheses are the gains from trade for exporter $i$ ($GFT_{ij}^{i}\left(p\right)$) and importer $j$ ($GFT_{ij}^{j}\left(p\right)$), written as a function of $p_{ij}$.
The FOC associated with the problem (A.6) can be written as:
00 0 $=\frac{d\ln\pi_{i}}{d\ln p_{ij}}+\frac{\phi}{1-\phi}\cdot\frac{GFT_{ij}^{i}}{\pi_{i}}\frac{\pi_{j}}{GFT_{ij}^{j}}\cdot\frac{d\ln\pi_{j}}{d\ln p_{ij}},$ (A.7)
where we used the fact that $\frac{dGFT_{ij}^{k}}{dp}=\frac{d\pi_{k}}{dp}$ for $k=\{i,j\}.$ In what follows, we derive expressions for $\frac{d\ln\pi_{i}}{d\ln p_{ij}},\frac{d\ln\pi_{j}}{d\ln p_{ij}},GFT_{ij}^{i}$ and $GFT_{ij}^{j}.$
Exporter $i$’s Profits and Gains from Trade–Firm $i$’s profit under a successful negotiation can be expressed as
$\pi_{i}=p_{ij}q_{ij}+\sum_{k\neq j}p_{ik}q_{ik}-\theta c_{i}q_{i},$
where $c_{i}=k_{i}q_{i}^{\frac{1-\theta}{\theta}}.$ The elasticity of the profit $\pi_{i}$ with respect to $p_{ij}$ can be found as:
$\frac{d\ln\pi_{i}}{d\ln p_{ij}}$ $=\frac{q_{ij}}{\pi_{i}}(\varepsilon_{ij}-1)\left(\mu_{ij}^{\text{oligopoly}}c_{i}-p_{ij}\right),$ (A.8)
where $\varepsilon_{ij}\equiv-\frac{d\ln q_{ij}}{d\ln p_{ij}}$ is defined in equation (A.5) and
$\mu^{\text{oligopoly}}:=\frac{\varepsilon_{ij}}{\varepsilon_{ij}-1}.$
The “outside” profit of firm $i$ are
$\tilde{\pi}_{i\left(-j\right)}=\sum_{k\neq j}p_{ik}q_{ik}-\theta\widetilde{c_{i}q_{i}}.$
For a constant $k_{i}$ and given $q_{i}=\sum_{j\in\mathcal{Z}_{i}}q_{ij}$, the total cost in case of failed agreement can be found as:
$\theta\widetilde{c_{i}q_{i}}$ $=\theta c_{i}q_{i}\left(\frac{\widetilde{q_{i}}}{q_{i}}\right)^{\frac{1}{\theta}}.$
$=\theta c_{i}q_{i}\left(1-x_{ij}\right)^{\frac{1}{\theta}}.$
The exporter’s gains from trade $GFT_{ij}^{i}$ are thus given by:
$GFT_{ij}^{i}\left(p_{ij}\right)$ $\equiv\pi_{i}\left(p\right)-\tilde{\pi}_{i\left(-j\right)}$ (A.9)
$=q_{ij}\left(p_{ij}-c_{i}\mu_{ij}^{\text{oligopsony}}\right),$
where $\mu_{ij}^{\text{oligopsony}}:=\theta\left[\frac{\Delta_{ij}^{x}}{x_{ij}}\right]$ and $\Delta_{ij}^{x}:=\left[1-\left(1-x_{ij}\right)^{\frac{1}{\theta}}\right].$
Importer $j$’s Profits and Gains from Trade–Firm $j$’s profit under a successful negotiation can be expressed as
$\pi_{j}=p_{j}q_{j}-\varrho c_{j}q_{j}=\frac{\gamma}{\eta-1}c_{j}q_{j},$
where $c_{j}$ is defined in equation (A.3), and $q_{j}$ in equation (A.4). The elasticity of the profit $\pi_{j}$ with respect to $p_{ij}$ can be found as:
$-\frac{d\ln\pi_{j}}{d\ln p_{ij}}=(\eta-1)s_{ij}.$ (A.10)
The outside profit of firm $j$ under a failed negotiation is
$\tilde{\pi}_{j\left(-i\right)}=\frac{\gamma}{\eta-1}\widetilde{c_{j}q_{j}},$
where $\varrho\widetilde{c_{j}q_{j}}$ denotes the total “outside” cost of importer $j$. The assumptions on technology and demand downstream imply that we can write:
$\widetilde{c_{j}q_{j}}$ $=\left(1-s_{ij}\right)^{\frac{\eta-1}{\rho-1}}c_{j}q_{j}.$
Putting things together, the importer’s gains from trade can be written as:
$GFT_{j}(p_{ij})$ $=\pi_{j}\Delta_{ij}^{s}$ (A.11)
where we defined $\Delta_{ij}^{s}:=\left(1-\left(1-s_{ij}\right)^{\frac{\eta-1}{\rho-1}}\right).$
We’re now ready to solve for the bilateral price $p_{ij}.$ Substituting equations (A.8), (A.9), (A.10), and (A.11) into the FOC in equation (A.7), we obtain:
$\mu_{ij}:=\frac{p_{ij}}{c_{i}}=(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}+\omega_{ij}\mu_{ij}^{\text{oligopsony}}$ (A.12)
where the weighting factor is
$\omega_{ij}:=\frac{\frac{\phi}{1-\phi}\lambda_{ij}}{1+\frac{\phi}{1-\phi}\lambda_{ij}}\in(0,1)$ (A.13)
and where
$\lambda_{ij}$ $:=\frac{\pi_{j}}{GFT_{ij}^{j}}\cdot\left(-\frac{d\ln\pi_{j}}{d\ln p_{ij}}\cdot(\varepsilon_{ij}-1)^{-1}\right)$ (A.14)
$=\underbrace{\frac{\pi_{j}}{GFT_{ij}^{j}}}_{\lambda_{ij}^{N}}\cdot\underbrace{-\frac{d\ln\pi_{j}}{d\ln p_{ij}q_{ij}}}_{\lambda_{ij}^{I}}$
$=\left(\Delta_{ij}^{s}\right)^{-1}\cdot\frac{\eta-1}{\varepsilon_{ij}-1}s_{ij}>0.$
Equations (A.12)-(A.14) forms the basis for Proposition 1 in Section 2. ∎
A.2 Proof of Proposition 2
Proposition 2 characterizes how the bilateral markup $\mu_{ij}$ co-moves with the exporter’s and importer’s bilateral market shares, $s_{ij}$ and $x_{ij}$.
Part (i): $\mu_{ij}$ and Exporter’s supplier share $s_{ij}$
To study how $\mu_{ij}$ responds to the exporter’s supplier share $s_{ij}$, note that the oligopoly markup $\mu^{\text{oligopoly}}_{ij}$ is strictly increasing in $s_{ij}$ under standard assumptions ($\eta<\rho<\infty$). Differentiating the overall markup $\mu_{ij}$ in equation (A.12) with respect to $s_{ij}$ yields:
$\frac{\partial\mu_{ij}}{\partial s_{ij}}=(1-\omega_{ij})\cdot\underbrace{\frac{\partial\mu^{\text{oligopoly}}_{ij}}{\partial s_{ij}}}_{>0}+\frac{\partial\omega_{ij}}{\partial s_{ij}}\cdot\underbrace{\left(\mu^{\text{oligopsony}}_{ij}-\mu^{\text{oligopoly}}_{ij}\right)}_{<0}.$
The first term is strictly positive when $\omega_{ij}<1$ (i.e., $\phi<1$), reflecting the direct effect of $s_{ij}$ on the oligopoly markup. The second term captures how changes in $s_{ij}$ affect the bargaining weight $\omega_{ij}$ and thus the relative influence of the oligopsony markdown. Although $\omega_{ij}$ is hump-shaped in $s_{ij}$, increasing at low values and decreasing at high values, its sensitivity to $s_{ij}$ is limited in most of the parameter space. As a result, $\frac{\partial\omega_{ij}}{\partial s_{ij}}$ is typically small, and the first (positive) term generally dominates.
In contrast, when $\phi=1$ (full importer bargaining power), the bilateral markup reduces to the oligopsony markdown: $\mu_{ij}=\mu^{\text{oligopsony}}_{ij}$, and $\omega_{ij}=1$. In this case,
$\frac{\partial\mu_{ij}}{\partial s_{ij}}=0.$
Thus, we have:
$\frac{\partial\mu_{ij}}{\partial s_{ij}}=\begin{cases}\gtrless 0&\text{if }\phi<1\text{ (typically positive)},\\ =0&\text{if }\phi=1.\end{cases}$
Thus, $\mu_{ij}$ can increase with $s_{ij}$ in our theory only when $\phi<1$, which proves the first part of Proposition 1.
Part (ii): $\mu_{ij}$ and Importer’s buyer share $x_{ij}$
The effect of $x_{ij}$ on $\mu_{ij}$ is limited to the oligopsony markdown term and is therefore easier to characterize. Differentiating the bilateral markup with respect to $x_{ij}$ yields:
$\frac{\partial\mu_{ij}}{\partial x_{ij}}=\omega_{ij}\cdot\underbrace{\frac{\partial\mu^{\text{oligopsony}}_{ij}}{\partial x_{ij}}}_{<0\ \text{if}\ \theta<1}.$
The markdown $\mu^{\text{oligopsony}}_{ij}$ decreases with $x_{ij}$ when marginal cost is increasing (i.e., $\theta<1$), as larger buyers elicit stronger cost reductions and thus negotiate lower prices. This implies:
$\frac{\partial\mu_{ij}}{\partial x_{ij}}<0\iff\omega_{ij}>0\text{ and }\theta<1.$
Therefore, observing a negative slope of $\mu_{ij}$ with respect to the importer’s share $x_{ij}$ provides evidence of both importer bargaining power ($\phi>0$ so that $\omega_{ij}>0$) and decreasing returns to scale ($\theta<1$). This completes the proof of Proposition 2. ∎
A.3 Proof of Proposition 3
The log (tariff-inclusive) price is given by:
$\ln p_{ij}=\ln\mu_{ij}+\ln c_{i}+\ln T_{c},$
where $c_{i}$ and $\mu_{ij}$ are as in equations (2.1) and (2.6), respectively.
Taking a full log-differential and rearranging terms yields:
$d\ln p_{ij}$ $=$ $-\Gamma_{ij}\cdot d\ln p_{ij}-\Lambda_{ij}d\ln p_{ij}+d\ln T_{c}$
$\frac{d\ln p_{ij}}{d\ln T_{c}}$ $=$ $\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}},$
where $\Gamma_{ij}\equiv-\frac{d\ln\mu_{ij}}{d\ln p_{ij}}$ and $\Lambda_{ij}\equiv-\frac{d\ln c_{i}}{d\ln p_{ij}}$ are the partial markup and cost elasticities, respectively.
The Cost Elasticity
Taking the logarithm of equation (2.1), we obtain:
$\ln c_{i}$ $=\ln k_{i}+\frac{1-\theta}{\theta}\ln q_{i}.$
It immediately follows that:
$\Lambda_{ij}\equiv-\frac{d\ln c_{i}}{d\ln p_{ij}}$ $=\frac{1-\theta}{\theta}\frac{d\ln q_{i}}{d\ln q_{ij}}\left(-\frac{d\ln q_{ij}}{d\ln p_{ij}}\right)$
$=\frac{1-\theta}{\theta}\cdot x_{ij}\cdot\varepsilon_{ij}\geq 0.$
Moreover, the comparative statics with respect to the bilateral shares are easy to compute as:
$\frac{d\Lambda_{ij}}{dx_{ij}}=\frac{1-\theta}{\theta}\cdot\varepsilon_{ij}\geq 0,$
with strict inequality whenever $\theta<1,$ whereas:
$\frac{d\Lambda_{ij}}{ds_{ij}}=\frac{1-\theta}{\theta}\cdot x_{ij}\cdot(\eta-\rho)<0.$
Thus, the cost elasticity weakly increases with the importer’s buyer share $x_{ij},$ and it decreases with the exporter’s supplier share $s_{ij}.$
Markup Elasticity
Taking logs of equation (2.6) and differentiating, we obtain:
$d\ln\mu_{ij}$ $=\frac{(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\,d\ln\mu_{ij}^{\text{oligopoly}}+\frac{\omega_{ij}\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\,d\ln\mu_{ij}^{\text{oligopsony}}$
$\ \ +\frac{\omega_{ij}\left(\mu_{ij}^{\text{oligopsony}}-\mu_{ij}^{\text{oligopoly}}\right)}{\mu_{ij}}\,d\ln\omega_{ij}.$
Rearranging terms, the price elasticity of the bilateral markup can be expressed as:
$\Gamma_{ij}\equiv-\frac{d\ln\mu_{ij}}{d\ln p_{ij}}=\frac{(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\,\Gamma_{ij}^{\text{oligopoly}}+\frac{\omega_{ij}\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\,\Gamma_{ij}^{\text{oligopsony}}+\left(1-\frac{\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\right)\Gamma_{ij}^{\omega},$
where: $\Gamma_{ij}^{\text{oligopoly}}\equiv-\frac{d\ln\mu_{ij}^{\text{oligopoly}}}{d\ln p_{ij}},\ \ \Gamma_{ij}^{\text{oligopsony}}\equiv-\frac{d\ln\mu_{ij}^{\text{oligopsony}}}{d\ln p_{ij}},$ and $\Gamma_{ij}^{\omega}\equiv-\frac{d\ln\omega_{ij}}{d\ln p_{ij}}.$
Oligopoly Markup Elasticity–The oligopoly markup elasticity is given by:
$\Gamma_{ij}^{\text{oligopoly}}\equiv$ $-\frac{d\ln\mu_{ij}^{\text{oligopoly}}}{d\ln p_{ij}}=-\frac{d\ln\mu_{ij}^{\text{oligopoly}}}{d\ln s_{ij}}\cdot\frac{d\ln s_{ij}}{d\ln p_{ij}}.$
From the definition of $s_{ij},$ the last term is:
$\frac{d\ln s_{ij}}{d\ln p_{ij}}=-\left(\rho-1\right)\left(1-s_{ij}\right).$
Given $\mu_{ij}^{oligopoly}=\frac{\varepsilon_{ij}}{\varepsilon_{ij}-1}$ , we find:
$-\frac{d\ln\mu_{ij}^{\text{oligopoly}}}{d\ln s_{ij}}$ $=-\frac{1}{(\varepsilon_{ij}-1)}\cdot\frac{\rho-\varepsilon_{ij}}{\varepsilon_{ij}},$
which implies:
$\Gamma_{ij}^{\text{oligopoly}}=-\frac{d\ln\mu_{ij}^{\text{oligopoly}}}{d\ln s_{ij}}\cdot\frac{d\ln s_{ij}}{d\ln p_{ij}}=\frac{1}{\varepsilon_{ij}-1}\frac{\rho-\varepsilon_{ij}}{\varepsilon_{ij}}\left(\rho-1\right)\left(1-s_{ij}\right)\geq 0.$
Oligopsony Markup Elasticity– The oligopsony markup elasticity is given by:
$\Gamma_{ij}^{\text{oligopsony}}\equiv$ $-\frac{d\ln\mu_{ij}^{\text{oligopsony}}}{d\ln p_{ij}}=-\left(\frac{d\ln\mu_{ij}^{\text{oligopsony}}}{d\ln x_{ij}}\right)\left(\frac{d\ln x_{ij}}{d\ln p_{ij}}\right).$
From the definition of $x_{ij},$ the last term is:
$\frac{d\ln x_{ij}}{d\ln p_{ij}}=-(1-x_{ij})\varepsilon_{ij}.$
Given $\mu_{ij}^{oligopsony}:=\theta\left(\frac{1-\left(1-x_{ij}\right)^{\frac{1}{\theta}}}{x_{ij}}\right),$ we find:
$\frac{d\ln\mu_{ij}^{\text{oligopsony}}}{d\ln x_{ij}}=\left(\frac{x_{ij}\left(1-x_{ij}\right)^{\frac{1}{\theta}-1}}{\theta\left(1-\left(1-x_{ij}\right)^{\frac{1}{\theta}}\right)}-1\right)$
which implies:
$\Gamma_{ij}^{\text{oligopsony}}=-\left(\frac{d\ln\mu_{ij}^{\text{oligopsony}}}{d\ln x_{ij}}\right)\left(\frac{d\ln x_{ij}}{d\ln p_{ij}}\right)=\left(\frac{x_{ij}\left(1-x_{ij}\right)^{\frac{1}{\theta}-1}}{\theta\left(1-\left(1-x_{ij}\right)^{\frac{1}{\theta}}\right)}-1\right)(1-x_{ij})\varepsilon_{ij}.$
with
$\frac{\partial\Gamma_{ij}^{\text{oligopsony}}}{\partial x_{ij}}={}$ $\Bigg{[}\frac{\partial}{\partial x_{ij}}\left(\frac{x_{ij}(1-x_{ij})^{\frac{1}{\theta}-1}}{\theta\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]}\right)(1-x_{ij})$
$\quad-\left(\frac{x_{ij}(1-x_{ij})^{\frac{1}{\theta}-1}}{\theta\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]}-1\right)\Bigg{]}\varepsilon_{ij}$
[…]
$\quad-\left(\frac{x_{ij}(1-x_{ij})^{\frac{1}{\theta}-1}}{\theta\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]}-1\right)$
[…]
$\quad\left.-\left(\frac{x_{ij}(1-x_{ij})^{\frac{1}{\theta}-1}}{\theta\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]}-1\right)\right]$
Omega Elasticity
The elasticity of the weight $\omega_{ij}$ with respect to price is:
$\Gamma_{ij}^{\omega}\equiv-\frac{d\ln\omega_{ij}}{d\ln p_{ij}}=$ $\frac{d\ln\omega_{ij}}{d\ln s_{ij}}\left(-\frac{d\ln s_{ij}}{d\ln p_{ij}}\right).$
Given
$\frac{d\ln\omega_{ij}}{d\ln s_{ij}}=\left(1-\omega_{ij}\right)\frac{d\ln\lambda_{ij}}{d\ln s_{ij}},$
the above elasticity becomes:
$\Gamma_{ij}^{\omega}=\frac{d\ln\lambda_{ij}}{d\ln s_{ij}}\left(1-\omega_{ij}\right)\left(\rho-1\right)\left(1-s_{ij}\right),$
Since $\lambda_{ij}=\frac{(\eta-1)s_{ij}}{\varepsilon_{ij}-1}\cdot\left(1-(1-s_{ij})^{\frac{\eta-1}{\rho-1}}\right)^{-1}$ , we find:
$\frac{d\ln\lambda_{ij}}{d\ln s_{ij}}$ $=1-\frac{\varepsilon_{ij}-\rho}{\varepsilon_{ij}-1}-\frac{\eta-1}{\rho-1}\cdot\frac{(1-s_{ij})^{\frac{\eta-1}{\rho-1}}}{1-(1-s_{ij})^{\frac{\eta-1}{\rho-1}}}\cdot\frac{s_{ij}}{(1-s_{ij})}.$
Thus:
$\Gamma_{ij}^{\omega}$ $=\left(1-\frac{\varepsilon_{ij}-\rho}{\varepsilon_{ij}-1}-\frac{\eta-1}{\rho-1}\cdot\frac{(1-s_{ij})^{\frac{\eta-1}{\rho-1}}}{1-(1-s_{ij})^{\frac{\eta-1}{\rho-1}}}\cdot\frac{s_{ij}}{(1-s_{ij})}\right)\left(1-\omega_{ij}\right)\left(\rho-1\right)\left(1-s_{ij}\right).$
A.4 Proof of Proposition 4
We now prove Proposition 4, which characterizes the equilibrium comovement between pass-through $\Phi_{ij}$ and the bilateral market shares $x_{ij}$ and $s_{ij}$.
Comparative Statics with Respect to $x_{ij}$
We begin by analyzing how pass-through $\Phi_{ij}$ responds to the importer’s buyer share $x_{ij}$. As discussed in the main text, $\Phi_{ij}$ depends on $x_{ij}$ through two channels: (i) the oligopsony markdown elasticity $\Gamma_{ij}^{\text{oligopsony}}$, which is non-monotonic (U-shaped) in $x_{ij}$, and (ii) the cost elasticity $\Lambda_{ij}$, which increases linearly in $x_{ij}$ when $\theta<1$.
When $\theta=1$, both elasticities are zero: $\Gamma_{ij}^{\text{oligopsony}}=\Lambda_{ij}=0$. Therefore, pass-through is constant in $x_{ij}$:
$\left.\frac{\partial\Phi_{ij}}{\partial x_{ij}}\right|_{\theta=1}=0.$
When $\theta<1$, however, both elasticities are active. The pass-through elasticity can be written as:
$\Phi_{ij}=\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}}=\frac{1}{1+\cdots+\omega_{ij}\cdot\frac{\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\cdot\Gamma_{ij}^{\text{oligopsony}}+\Lambda_{ij}},$
where $\cdots$” denotes terms that do not depend on $x_{ij}$. Differentiating with respect to $x_{ij}$:
$\frac{\partial\Phi_{ij}}{\partial x_{ij}}=-\frac{1}{(1+\Gamma_{ij}+\Lambda_{ij})^{2}}\cdot\frac{\partial}{\partial x_{ij}}\left(\omega_{ij}\cdot\frac{\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\cdot\Gamma_{ij}^{\text{oligopsony}}+\Lambda_{ij}\right).$
Let:
$T_{1}\equiv\omega_{ij}\cdot\frac{\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\cdot\Gamma_{ij}^{\text{oligopsony}},\quad T_{2}\equiv\Lambda_{ij}.$
We can expand these as:
$T_{1}$ $=\omega_{ij}\cdot\left(\frac{(1-x_{ij})^{\frac{1}{\theta}-1}-\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\right)(1-x_{ij})\varepsilon_{ij},$
$T_{2}$ $=\frac{1-\theta}{\theta}\cdot x_{ij}\cdot\varepsilon_{ij}.$
While $T_{1}$ is non-monotonic in $x_{ij}$, $T_{2}$ increases linearly in $x_{ij}$ and dominates $T_{1}$ for all $x_{ij}$. This can be verified numerically; for instance, when $\theta=\frac{1}{2}$:
$T_{1}+T_{2}=x_{ij}\cdot\varepsilon_{ij}\left(1-\frac{\omega_{ij}}{\mu_{ij}}\cdot\left(\frac{5}{2}-\frac{3}{2}x_{ij}\right)\right),$
which increases in $x_{ij}$ as the term in parentheses remains positive.
Hence, the term in the numerator of $\frac{\partial\Phi_{ij}}{\partial x_{ij}}$ increases in $x_{ij}$, implying:
$\operatorname{sign}\left(\frac{\partial\Phi_{ij}}{\partial x_{ij}}\right)=-\operatorname{sign}\left(\frac{\partial\Lambda_{ij}}{\partial x_{ij}}\right),$
so that:
$\frac{\partial\Phi_{ij}}{\partial x_{ij}}\leq 0.$
This proves that, under $\theta<1$, pass-through decreases in $x_{ij}$, a robust and testable implication of the model.
Comparative Statics with Respect to $s_{ij}$
We now turn to how pass-through $\Phi_{ij}$ varies with the exporter’s supplier share $s_{ij}$. Unlike the case of $x_{ij}$, this relationship is more complex, as all components, namely, $\Gamma_{ij}^{\text{oligopoly}}$, $\Gamma_{ij}^{\text{oligopsony}}$, $\Gamma_{ij}^{\omega}$, and $\Lambda_{ij}$, may vary with $s_{ij}$. Specifically,
$\Phi_{ij}=\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}}=\frac{1}{1+\frac{(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopoly}}+\frac{\omega_{ij}\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopsony}}+\left(1-\frac{\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\right)\Gamma_{ij}^{\omega}+\Lambda_{ij}}.$
Differentiating with respect to $s_{ij}$ gives:
$\frac{\partial\Phi_{ij}}{\partial s_{ij}}$ $=-\frac{\partial}{\partial s_{ij}}\bigg{(}\frac{(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopoly}}+\frac{\omega_{ij}\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopsony}}$
$\quad+\left(1-\frac{\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\right)\Gamma_{ij}^{\omega}+\Lambda_{ij}\bigg{)}\cdot\left(1+\Gamma_{ij}+\Lambda_{ij}\right)^{-2}.$
When $\theta=1$, both $\Gamma_{ij}^{\text{oligopsony}}=0$ and $\Lambda_{ij}=0$, and the expression simplifies to:
$\frac{\partial\Phi_{ij}}{\partial s_{ij}}=-\frac{\partial}{\partial s_{ij}}\left(\frac{(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopoly}}+\left(1-\frac{\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\right)\Gamma_{ij}^{\omega}\right)\cdot(1+\Gamma_{ij})^{-2}.$
To build intuition, consider the limit $s_{ij}\to 1$, where $\omega_{ij}\to\phi$ and $\Gamma_{ij}^{\omega}\to 0$. Then:
$\frac{\partial\Phi_{ij}}{\partial s_{ij}}\propto\begin{cases}-\frac{\partial}{\partial s_{ij}}\left(\frac{(1-\phi)\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopoly}}+\frac{\phi\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopsony}}+\Lambda_{ij}\right)&\text{if }\theta<1,\\ -\frac{\partial}{\partial s_{ij}}\left(\frac{(1-\phi)\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopoly}}\right)&\text{if }\theta=1.\end{cases}$
The term involving $\Gamma_{ij}^{\text{oligopoly}}$ is hump-shaped in $s_{ij}$, while the remaining components, particularly $\Lambda_{ij}$, decline monotonically with $s_{ij}$ due to $\frac{\partial\varepsilon_{ij}}{\partial s_{ij}}=-(\rho-\eta)<0$. Quantitatively, these latter terms dominate in most parameter ranges.
Thus, we have:
$\operatorname{sign}\left(\frac{\partial\Phi_{ij}}{\partial s_{ij}}\right)\propto\begin{cases}(+)&\text{if }\theta<1\text{ or }s_{ij}\text{ sufficientlylarge},\\ (-)&\text{if }\theta=1\text{ and }s_{ij}\text{ sufficiently small}.\end{cases}$
Due to the non-monotonicity and interaction of multiple channels, the relationship between $\Phi_{ij}$ and $s_{ij}$ does not yield a clean, general prediction. While pass-through may increase with $s_{ij}$ under decreasing returns to scale and sufficient buyer power, the sign of this relationship depends on the relative strength of strategic complementarities, oligopsony markdowns, and cost elasticity. As a result, we omit this relationship from Proposition 3 and focus instead on the more robust prediction involving $x_{ij}$.
Appendix B Theory: Discussion and Extensions
B.1 Efficient Bargaining
In the efficient bargaining setup, the importer and exporter negotiate over a two-part tariff $(p,q)$ by maximizing a Nash product of generalized firm-specific gains from trade (GFT):
$\max_{p,q}\left[GFT_{ij}^{i}(p,q)\right]^{1-\phi}\left[GFT_{ij}^{j}(p,q)\right]^{\phi}\quad\text{s.t.}\quad GFT_{ij}^{i}\geq 0,\quad GFT_{ij}^{j}\geq 0$
where:
$GFT_{ij}^{i}(p_{ij},q_{ij})$ $\equiv\pi_{i}(p)-\tilde{\pi}_{i(-j)}=p_{ij}q_{ij}-\left(\theta c_{i}q_{i}-\theta\widetilde{c_{i}q_{i}}\right),$
$GFT_{ij}^{j}(p_{ij},q_{ij})$ $\equiv\pi_{j}(p)-\tilde{\pi}_{j(-i)}=\left(p_{j}q_{j}-\tilde{p}_{j}\tilde{q}_{j}\right)-p_{ij}q_{ij}.$
This formulation corresponds to a setting in which the firm pair first selects the input quantity $q_{ij}$ to maximize joint surplus, and then negotiates over the price $p_{ij}$ to determine how surplus is split.
Figure B.1: Efficient Bargaining
The efficient quantity solves:
$q_{ij}^{*}$ $:\quad\arg\max_{q}\left[GFT_{ij}^{i}(q)+GFT_{ij}^{j}(q)\right]$
$:\quad\frac{d(p_{j}q_{j})}{dq_{ij}}=\frac{d(\theta c_{i}q_{i})}{dq_{ij}}$
That is, the efficient quantity $q^{*}_{ij}$ equates the exporter’s marginal cost to the importer’s marginal revenue, similar to the outcome under vertical integration. Solving yields:
$q_{ij}^{*}:\quad p_{ij}=c_{i}(q_{ij}^{*}),$
so that the negotiated price equals the exporter’s marginal cost.
Given $q_{ij}^{*}$, the price $p_{ij}$ solves the Nash bargaining problem:
$\max_{p_{ij}}\left(p_{ij}q^{*}_{ij}-\Delta C_{i}(q^{*}_{ij})\right)^{1-\phi}\left(\Delta R_{j}(q^{*}_{ij})-p_{ij}q^{*}_{ij}\right)^{\phi},$
where $\phi\in(0,1)$ denotes the importer’s bargaining power. The terms $\Delta R_{j}(q_{ij})$ and $\Delta C_{i}(q_{ij})$ represent the additional revenue for the importer and the additional cost for the exporter attributable to the match:
$\Delta R_{j}(q_{ij})$ $=p_{j}q_{j}-\widetilde{p_{j}q_{j}},$
$\Delta C_{i}(q_{ij})$ $=\theta c_{i}q_{i}-\theta\widetilde{c_{i}q_{i}}.$
Standard derivations lead to the equilibrium price:
$p_{ij}$ $=(1-\phi)\cdot\frac{\Delta R_{j}(q_{ij})}{q_{ij}}+\phi\cdot\frac{\Delta C_{i}(q_{ij})}{q_{ij}}$
$=(1-\phi)\cdot MC_{i}(q^{*}_{ij})+\phi\cdot AC(q^{*}_{ij})$
a weighted average of the per-unit downstream revenue gain and per-unit upstream cost increase from the match, with bargaining weights given by $\phi$.
Figure B.1 illustrates the set of feasible equilibria. The efficient quantity is $q^{*}$, while the negotiated price ranges between $MC(q^{*})$ and $AC(q^{*})$ depending on $\phi$.
B.2 Supply-Driven Quantity Bargaining
In the case of supply-driven quantity bargaining, the exporter first chooses the quantity $q_{ij}$ for a given price $p_{ij}$ to maximize profits, and bargaining occurs over the price, holding the induced supply curve fixed. For tractability, we solve the dual problem where the exporter selects a price $p_{ij}$ for a given quantity $q_{ij}$, and bargaining takes place over the quantity. Formally, this is expressed as:
$\begin{cases}\max_{p_{ij}}\pi_{i}(p_{ij};q_{ij})\\ \max_{q_{ij}}\left[GFT_{ij}^{i}(q_{ij},p_{ij})\right]^{1-\phi}\left[GFT_{ij}^{j}(q_{ij},p_{ij})\right]^{\phi},\end{cases}$
where the dependence on price and quantity in other parts of the network is left implicit. The solution to the exporter’s problem yields the supply function, which we write as:
$p_{ij}(q_{ij})=k_{i}q_{i}^{\frac{1-\theta}{\theta}},\quad\text{where}\quad q_{i}=q_{i(-j)}+q_{ij},$
as in equation (2.1). Solving the bargaining problem using derivations similar to those in Section A.1 of this Appendix yields the following expression for the bilateral price:
$p_{ij}=(1-\omega_{ij}^{S})\cdot c_{i}+\omega_{ij}^{S}\cdot\psi_{ij}^{-1}MRP_{j},$
where
$\psi_{ij}=1+c^{\prime}_{i,q_{ij}}=1+\frac{1-\theta}{\theta}x_{ij}.$
Here, $\omega_{ij}^{S}$ is given by:
$\omega_{ij}^{S}\equiv\frac{\frac{\phi}{1-\phi}\cdot\lambda^{S}_{ij}}{1+\frac{\phi}{1-\phi}\cdot\lambda^{S}_{ij}}\quad\text{and}\quad\lambda^{S}_{ij}\equiv\frac{GFT_{ij}^{i}}{GFT_{ij}^{j}}\psi_{ij}=\frac{p_{ij}q_{ij}-\theta c_{i}q_{i}\Delta_{ij}^{x}}{p_{j}q_{j}\Delta_{ij}^{s}-p_{ij}q_{ij}}\psi_{ij}.$
Figure B.2: Supply-Driven Quantity Bargaining
Hence, the bilateral price is a convex combination of the exporter’s marginal cost ($c_{i}$) and a term reflecting the markdown ($\psi_{ij}$) below the importer’s marginal revenue product ($MRP_{j}\equiv\frac{dp_{j}q_{j}}{dq_{ij}}$). As in the demand-driven quantity bargaining case, the weight $\omega_{ij}^{S}$ has an intuitive interpretation as the importer’s effective bargaining power.
For $\phi=0$, $\omega_{ij}^{S}=0$, and the allocation is efficient with $p_{ij}=c_{i}$. For $\phi=1$, $\omega_{ij}^{S}=1$, and we obtain the standard oligopsony benchmark where the input price is a markdown below the importer’s marginal revenue product, with the markdown equal to one plus the importer’s residual supply elasticity. This reflects the importer’s view, where the cost of increasing quantity is driven by raising the unit price for all inframarginal units when supply curves are upward-sloping.
Figure B.2 plots the set of feasible allocations. The red line indicates the pure oligopsony markdown, and the pink area represents the set of markdowns for $\phi\in(0,1)$.
While this case is intuitive, it is more complex to characterize due to the term $\lambda^{S}_{ij}$, which is itself a function of prices and quantities and thus no longer a simple function of market shares as in the baseline case. As a result, bringing this model to the data is less straightforward than our baseline model.
B.3 Generalized Outside Option
In the baseline model, we assume that the importer’s (exporter’s) gains from trade are given by the firm’s total payoff from trading with all partners, minus the payoff from trading with all partners except exporter $i$ (importer $j$). In simpler terms, each importer (exporter) treats itself as the marginal buyer (supplier). This implies that, in case of disagreement, neither party considers the possibility of forming new relationships, which substantially influences the definition of outside profits for both importers and exporters.
We now consider a more general specification that imposes less structure on outside options. Let
$\Delta_{ij}^{c_{i}}=\frac{\tilde{c}_{i}}{c_{i}}\qquad\text{and}\qquad\Delta_{ij}^{c_{j}}=\frac{\tilde{c}_{j}}{c_{j}}$
denote the percentage change in exporter $i$’s and importer $j$’s marginal cost in the event of a failed negotiation. Under this generalization, the gains from trade for firms $i$ and $j$ are given by:
$GFT_{ij}^{i}(p_{ij})$ $=q_{ij}\left(p_{ij}-c_{i}\mu_{ij}^{\text{oligopsony},G}\right),$
$GFT_{ij}^{j}(p_{ij})$ $=\pi_{j}\Delta_{ij}^{s,G},$
where we define:
$\mu_{ij}^{\text{oligopsony},G}$ $:=\theta\left[\frac{\Delta_{ij}^{x}}{x_{ij}}\right],\qquad\Delta_{ij}^{x}:=\left[\frac{1-\Delta_{ij}^{c_{i}}(1-x_{ij})}{x_{ij}}\right],$
$\Delta_{ij}^{s,G}$ $:=1-\left(\Delta_{ij}^{c_{j}}\right)^{1-\nu}.$
The first-order condition under this generalized outside option implies:
$\mu_{ij}:=\frac{p_{ij}}{c_{i}}=(1-\omega^{G}_{ij})\mu_{ij}^{\text{oligopoly}}+\omega^{G}_{ij}\mu_{ij}^{\text{oligopsony},G},$ (B.1)
where the weighting factor is given by:
$\omega^{G}_{ij}$ $:=\frac{\frac{\phi}{1-\phi}\lambda_{ij}^{G}}{1+\frac{\phi}{1-\phi}\lambda_{ij}^{G}}\in(0,1),$
$\lambda_{ij}^{G}$ $:=\frac{\gamma s_{ij}}{\Delta_{ij}^{s,G}}\cdot\frac{\nu-1}{\varepsilon_{ij}-1}>0.$
Equation (B.1) shares the structure of the markup equation in equation (2.6), but with two key differences. First, the term $\lambda_{ij}$ now depends on $\Delta_{ij}^{s,G}$, which cannot be expressed solely as a function of the supplier share. Second, the oligopsony markdown depends on $\Delta_{ij}^{x}$, which likewise cannot be written as a function of the buyer share alone.
To summarize, generalizing the structure of outside options preserves the overall structure of the markup equation and yields similar comparative statics. However, it introduces an identification problem: the effective bargaining power and oligopsony markdown now depend on unobserved terms, such as $\Delta_{ij}^{s,G}$ and $\Delta_{ij}^{x}$, that cannot be expressed solely as functions of market shares. In the absence of external data to estimate these terms, empirical implementation becomes infeasible without further assumptions. Imposing the more restrictive baseline assumption enables us to express both components as functions of observable bilateral market shares and a small set of structural parameters, allowing us to implement the model using available data and avoid excessive computational complexity.
B.4 Full Pass-Through Elasticity
In deriving Proposition 3, we assumed that the shock is applied at the firm-to-firm level and that prices and quantities in other relationships remain fixed. These assumptions allow us to isolate the direct, short-run effects of a shock. However, actual trade policy shocks, such as the Trump tariffs, often apply at the exporter- or product-level and may induce broader adjustments that our static, partial equilibrium model does not capture.
This section extends the analysis to incorporate certain indirect effects by capturing how a shock to exporter $i$ influences prices and quantities in other relationships, and how those changes feed back into the bilateral price $p_{ij}$. While we continue to abstract from full general equilibrium forces, this exercise clarifies how network spillovers may cause reduced-form pass-through estimates to deviate from structural ones, as emphasized by Berger et al. (2022).
We generalize the cost shock to an exporter-level shock, denoted $\vartheta_{i}$, and re-derive the relevant elasticities to allow for cross-relationship spillovers. First, the impact on the importer’s buyer share becomes:
$\frac{d\ln x_{ij}}{d\ln\vartheta_{i}}$ $=-\varepsilon_{ij}(1-x_{ij})\frac{d\ln p_{ij}}{d\ln\vartheta_{i}}+\sum_{z\in\mathcal{Z}_{i},\,z\neq j}x_{iz}\varepsilon_{iz}\frac{d\ln p_{iz}}{d\ln\vartheta_{i}}.$
The effect on the exporter’s marginal cost is:
$\frac{d\ln c_{i}}{d\ln\vartheta_{i}}$ $=\frac{1-\theta}{\theta}\left(-\varepsilon_{ij}x_{ij}\frac{d\ln p_{ij}}{d\ln\vartheta_{i}}-\sum_{z\in\mathcal{Z}_{i},\,z\neq j}x_{iz}\varepsilon_{iz}\frac{d\ln p_{iz}}{d\ln\vartheta_{i}}\right).$
We also account for the fact that $p_{ij}$ affects rival suppliers’ shares and prices. The elasticity of supplier share with respect to the shock is:
$\frac{d\ln s_{ij}}{d\ln\vartheta_{i}}=(1-\rho)\frac{d\ln p_{ij}}{d\ln\vartheta_{i}}\left((1-s_{ij})+s_{ij}(1-\rho)\sum_{k\in\mathcal{Z}_{j},\,k\neq i}s_{kj}\Gamma_{kj}^{s}\right).$
Incorporating these indirect effects, the full pass-through elasticity $\Psi_{ij}=\frac{d\ln p_{ij}}{d\ln\vartheta_{i}}$ is implicitly defined by:
$\Psi_{ij}=\tilde{\Phi}_{ij}+\tilde{\Phi}_{ij}\left(\Gamma_{ij}^{x}-\frac{1-\theta}{\theta}\right)\sum_{z\in\mathcal{Z}_{i},\,z\neq j}x_{iz}\varepsilon_{iz}\Psi_{iz},$
where:
$\tilde{\Phi}_{ij}=\left[1+\Gamma_{ij}^{s}(\rho-1)\left((1-s_{ij})-s_{ij}(\rho-1)\sum_{k\in\mathcal{Z}_{j},\,k\neq i}s_{kj}\Gamma_{kj}^{s}\right)+\Gamma_{ij}^{x}\varepsilon_{ij}(1-x_{ij})+\frac{1-\theta}{\theta}\varepsilon_{ij}x_{ij}\right]^{-1}.$
This elasticity $\Psi_{ij}$ embeds two key indirect effects beyond the direct elasticity $\Phi_{ij}$ derived earlier.
First, an increase in $p_{ij}$ may cause rivals’ supplier shares (e.g., $s_{kj}$) to rise, increasing their prices $p_{kj}$ via strategic interactions. These adjustments dampen the original substitution away from exporter $i$, raising $\tilde{\Phi}_{ij}$ relative to $\Phi_{ij}$.
Second, a cost shock to firm $i$ may propagate to other buyers $z\in\mathcal{Z}_{i}$, affecting $p_{iz}$, which then feeds back into $x_{ij}$ via firm $i$’s overall scale and market presence. These changes affect $p_{ij}$ through both the markup and cost channels, amplifying pass-through further.
Together, these network spillovers push the full pass-through elasticity $\Psi_{ij}$ away from the direct elasticity $\Phi_{ij}$. Whether the net effect is amplification or attenuation depends on the strength of substitution patterns and strategic responses,an empirical question. In the main text, we focus on the direct pass-through elasticity $\Phi_{ij}$, which is tightly grounded in our model’s Nash-in-Nash structure and match-level pricing assumptions.
Appendix C Data Appendix
C.1 Related-Party Trade Measured via Ownership Linkages
A key advantage of the ORBIS dataset is the breadth and detail of its ownership information. It provides comprehensive listings of both direct and indirect shareholders and subsidiaries, along with indicators of each firm’s independence, global ultimate ownership, and group affiliations. This enables us to identify corporate structures at the firm level, including ownership links between firms located in different countries. We define a parent–subsidiary relationship as one in which the parent firm holds at least a 50% ownership stake in the affiliate.
Linking U.S. Importers to Multinational Ownership
To identify U.S.-based multinational firms, we match firms in the Census Business Register to their ORBIS counterparts using names, addresses, and GPS coordinates. This linkage combines probabilistic record matching with manual validation, producing a high match rate. As a result, we can flag U.S. establishments that are either majority-owned affiliates of foreign multinationals or parent firms with majority-owned affiliates abroad. This information allows us to identify multinationals with operations in the U.S. without relying solely on the Related Party Trade (RPT) indicator reported in the LFTTD.
Identifying Cross-Border Ownership Links
To assess whether the foreign exporter also belongs to the same corporate group, we match the Manufacturer ID (MID) reported in LFTTD to firm records in ORBIS. The MID is constructed by U.S. Customs based on the exporter’s name, address, and country of origin using a set of formatting rules. The MID begins with a two-character country code (or a province code for Canada), followed by a name-based segment derived from the first three letters of the first and second words in the company’s name. If the company name consists of only one word, the first six letters are used. The next segment contains the first four digits from the address number, and the final three characters are the first three alphabetic characters of the city name. Standard formatting conventions apply, including the exclusion of punctuation, one-letter initials, and common stop words such as “the,” “and,” or “of.” Country-specific prefixes (e.g., “OAO” or “ZAO” in Russia or “PT” in Indonesia) are also omitted when constructing the MID.
Using these same rules, we replicate the MID structure for foreign firms in the ORBIS database. We then match each MID in the customs data to candidate firms in ORBIS based on the reconstructed name segment. We assess the quality of each potential match using two dimensions: location and product alignment. A location score is computed based on the match between city names in the MID and in ORBIS. A product match score is computed by comparing the NAICS6 industry code listed in ORBIS to the HS6 product code recorded in the customs data, using the concordance developed by Pierce and Schott (2009). We retain only those matches where both location and product scores exceed 90%. In addition, we drop from the matched dataset any ORBIS firm with fewer than five transactions to filter out spurious exporters and potential noise.
Another concern is that the MID may sometimes refer to intermediaries rather than manufacturers. Although U.S. customs rules require that the MID correspond to the producer or manufacturer, not to wholesalers or freight forwarders, compliance with this rule is imperfect. To mitigate this concern, we use ORBIS industry codes to exclude retailers, wholesalers, and logistics providers from the matched dataset.
Finally, another challenge with the MID is that it is not a unique firm identifier: a given MID can correspond to multiple legal entities. In our matched data, we address this issue directly by checking whether a MID maps to more than one firm in ORBIS. If multiple firms share the same MID but belong to the same corporate group based on majority ownership links reported in ORBIS, we retain the match. Otherwise, we exclude the ambiguous MID from the analysis.
Taken together, these steps yield a linked dataset that offers a more transparent and conservative definition of related-party trade, based on majority ownership (at least 50%). In contrast to the standard related-party trade (RPT) flag in customs data, which applies a lower threshold of 6% ownership for imports, this approach reduces false positives and more precisely captures transactions where ownership ties are likely to influence pricing. By combining MID-based matching with firm-level ownership structures, the final dataset is well-suited for analyzing pricing behavior in cross-border transactions.
C.2 Data Cleaning, Sample Construction and Summary Statistics
We construct the analysis sample in several steps to align the data with the model’s structure and requirements.
We begin by removing observations that are incomplete or inconsistent with the modeling framework. Specifically, we drop transactions with missing or zero values for import value or quantity, invalid exporter identifiers (e.g., strings with fewer than three characters or beginning with a number), or U.S. importers that cannot be linked to the Longitudinal Business Database (LBD). We also exclude transactions associated with special provisions or temporary classifications (HS codes 98–99).
Next, we restrict our attention to trade in intermediate and capital goods by removing HS10 products, which are classified as consumption goods under the BEC system.
Although the customs data are reported at the transaction level, we aggregate them to the annual level for each buyer–supplier–product triplet. We choose annual aggregation because few relationships appear in adjacent months or in the same month across years. We then retain only those triplets that are active in at least two consecutive years and where the supplier transacts the same HS10 product with more than one U.S. buyer. This ensures a panel structure with repeated observations, supporting the identification of the model’s parameters.
We also exclude related-party trade, which is less likely to reflect decentralized bargaining and more likely to involve internal pricing practices such as transfer pricing. [^35] Bernard et al. (2006) shows that related-party prices differ systematically from arm’s-length transactions, including lower average prices and different pass-through behavior. In our baseline, a buyer–supplier pair is flagged as related if ORBIS identifies a shared corporate parent. To preserve coverage, we retain all observations not flagged as related by either ORBIS or the LFTTD. For robustness, we also consider two alternative definitions: one based solely on the LFTTD flag, and another combining it with ORBIS data to identify the U.S. importer as a multinational. See Appendix C.1 for further details.
To address outliers, we follow Heise (2024) and apply two filters. First, we drop observations with extreme price levels, defined as log unit values below the 1st or above the 99th percentile of the HS10 product–country distribution. Second, we trim extreme price changes by excluding year-on-year log price differences smaller than $-4$ or greater than $+4$. These outliers are removed from estimation but retained when computing tenure and relationship length. We also exclude all HS10 products under HS chapter 27 (energy-related goods), which lie outside the scope of the model.
After applying these restrictions, which exclude outliers, energy products, and related-party trade, and focusing on repeated, arm’s-length relationships, the sample retains over 20% of U.S. imports by value and nearly 15% of buyer–supplier–product triplets.
Finally, we restrict attention to suppliers that sell the same HS10 product to more than one U.S. buyer. This condition is essential for the identification of the model’s firm-level parameters and ensures that the empirical setting aligns with the model’s structure oulined in Setion 4.
Table C.1 summarizes the cumulative impact of these steps. Panel A focuses on the 2001–2016 sample used for structural estimation. Panel B reports the corresponding summary for the 2017–2018 sample used in the pass-through analysis. The first row of each panel (“All Imports”) includes all U.S. import records for the relevant years. Subsequent rows show the effect of each restriction in turn, including the requirement that buyer–supplier pairs trade the same product in two consecutive calendar years, the exclusion of consumption and energy goods, and the restriction to arm’s-length relationships with sufficient variation for identification.
The final rows show the estimation samples used in the analysis. For 2001–2016 (Panel A), the data include approximately \$880 billion in import value, 480,000 buyer–supplier pairs, and 630,000 buyer–supplier–HS10 triplets. For 2017–2018 (Panel B), the final sample includes \$160 billion in imports, 190,000 pairs, and 250,000 triplets. [^36] Panel A omits intermediate sample steps and shows only the full and final samples for 2001–2016, as these are not reported in the analysis nor disclosed under Census data policies. The same applies to Table C.2. These samples correspond exactly to the data used in estimation and post-estimation analysis.
Table C.2 reports a similar analysis using a broader sample that includes capital, intermediate, and consumption goods (“BEC – All (Consec.)”), rather than limiting to consumption goods only. Analogous to Table 1, Table C.3 reports the summary statistics of this broader sample.
Table C.1: Sample Composition by Period – Excluding BEC-Classified Consumption Goods
Sample Import value (bn USD) Importers (th) Exporters (th) Pairs (th) Triplets (th)
Panel A: 2001–2016
All Imports 22,000 1,000 6,900 17,000 40,000
BEC – Non-Cons. (Consec.) - - - - -
+ No Energy/Outliers/RPT - - - - -
+ Supplier Multi-Buyer 880 70 100 480 630
Panel B: 2017–2018
All Imports 2,000 330 1,300 2,500 5,300
BEC – All (Consec.) 1,600 160 530 890 1,800
BEC – Non-Cons. (Consec.) 1,000 120 320 540 950
+ No Energy/Outliers/RPT 420 110 270 470 730
+ Supplier Multi-Buyer 160 71 43 190 250
Notes: This table reports sample characteristics for a series of progressively restricted datasets used in the empirical analysis. The first row (“All Imports”) includes all U.S. import records in the sample period. All subsequent rows restrict the sample to buyer–supplier pairs that trade the same HS-10 product in two consecutive calendar years. The “BEC – excl. Cons. (Consec.)” sample includes only capital and intermediate goods, excluding consumption goods as defined by the Broad Economic Categories (BEC) classification. The next sample (“+ No Energy/Outliers”) adds four filters: (i) transactions involving energy-sector goods are excluded; (ii) observations with price levels below the 1st percentile or above the 99th percentile of the within-product price distribution are removed; (iii) extreme log price changes (above 4 or below –4) are excluded; and (iv) related-party transactions, which are defined as trade between entities with ownership ties or corporate control, are dropped following U.S. Census Bureau classification. “+ Supplier Multi-Buyer” restricts to suppliers that trade with at least two different buyers in consecutive years for the same product. “Import value” denotes the total annual value of imports in billions of U.S. dollars. “Importers” and “Exporters” correspond to distinct U.S. buyers and foreign suppliers, respectively. “Pairs” refer to unique buyer–supplier–product matches. “Triplets” refer to unique buyer–supplier–product–year combinations. All figures are reported separately for the 2001–2016 and 2017–2018 periods and are rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. These samples are the exact ones used in the empirical analysis. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Table C.2: Sample Composition by Period – All BEC Categories
Sample Import value (bn USD) Importers (th) Exporters (th) Pairs (th) Triplets (th)
Panel A: 2001–2016
All Imports 22,000 1,000 6,900 17,000 40,000
BEC – All (Consec.) - - - - -
+ No Energy/Outliers/RPT - - - -
+ Supplier Multi-Buyer 1,600 110 210 990 1,500
Panel B: 2017–2018
All Imports 2,000 330 1,300 2,500 5,300
BEC – All (Consec.) 1,600 160 530 890 1,800
+ No Energy/Outliers/RPT 710 150 470 800 1,500
+ Supplier Multi-Buyer 260 99 79 330 470
Notes: This table reports sample characteristics for a series of progressively restricted datasets used in the empirical analysis. The first row (“All Imports”) includes all U.S. import records in the sample period. All subsequent rows restrict the sample to buyer–supplier pairs that trade the same HS-10 product in two consecutive calendar years. The “BEC – All Categories (Consec.)” sample retains all transactions in capital, intermediate, and consumption goods as defined by the Broad Economic Categories (BEC) system, subject to the consecutive-year condition. The next sample (“+ No Energy/Outliers”) adds four filters: (i) transactions involving energy-sector goods are excluded; (ii) observations with price levels below the 1st percentile or above the 99th percentile of the within-product price distribution are removed; (iii) extreme log price changes (above 4 or below –4) are excluded; and (iv) related-party transactions are dropped following U.S. Census Bureau classification. “+ Supplier Multi-Buyer” restricts to suppliers that trade with at least two different buyers in consecutive years for the same product. “Import value” denotes the total annual value of imports in billions of U.S. dollars. “Importers” and “Exporters” correspond to distinct U.S. buyers and foreign suppliers, respectively. “Pairs” refer to unique buyer–supplier–product matches. “Triplets” refer to unique buyer–supplier–product–year combinations. All figures are reported separately for the 2001–2016 and 2017–2018 periods and are rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Table C.3: Summary Statistics – All BEC Categories (2001–2018)
Variable Mean Std. Dev. P25 Median P75
Panel A: Characteristics of Trade Relationships
$s_{ijh}$: Supplier share 0.27 0.32 0.02 0.10 0.43
$x_{ijh}$: Buyer share 0.27 0.30 0.03 0.13 0.45
 Relationship length (product $h$) 3.90 2.60 2.50 3.50 5.50
 Relationship length (all products) 4.60 3.10 2.50 4.50 6.50
 # Transactions (product $h$) 100 890 6.50 16 50
 # Transactions (all products) 410 3200 13 45 180
 # Products per pair 5.60 12.00 1.50 2.50 5.50
 Multi-HS10 dummy 0.68 0.47 0.00 1.00 1.00
 # Suppliers per buyer (HS10) 2.00 3.60 1.50 2.50 6.50
 Buyer tenure (all products) 9.50 4.90 5.50 9.50 14.00
 Buyer tenure (product $h$) 6.80 4.30 3.50 6.50 10.00
 # Buyers per supplier (HS10) 3.00 3.40 2.50 3.50 5.50
 Supplier tenure (all products) 7.70 4.40 4.50 7.50 11.00
 Supplier tenure (product $h$) 6.00 3.80 3.50 5.50 9.50
 Corr. between $s_{ijh}$ and $x_{ijh}$ 0.053
Panel B: Prices
$\log p$ (pre-duty) 3.40 2.50 1.50 3.10 5.10
$\log p$ (pre-duty, excl. charges) 3.30 2.50 1.40 3.00 5.00
$\log p^{\text{duty}}$ (post-duty) 3.40 2.50 1.50 3.10 5.20
Notes: This table reports summary statistics for a sample that covers all BEC product categories except energy goods, and excludes statistical outliers and related-party trade. It further restricts to suppliers that trade with at least two different U.S. buyers in consecutive years. This corresponds to the cumulative sample underlying the “+ Supplier Multi-Buyer” row in Panel B of Table C.2. Columns report the mean, standard deviation, and selected quantiles (25th percentile, median, and 75th percentile) for each variable. Prices in Panel B are log unit values (FOB value over quantity), with variants including charges or duties. $s_{ijh}$ denotes exporter $i$’s share in buyer $j$’s imports of product $h$; $x_{ijh}$ denotes buyer $j$’s share in exporter $i$’s U.S. exports of the same product. Relationship length and tenure are in years; concentration is measured at the HS10–year level. Counts of buyers, suppliers, and origin countries are per product per firm. Statistics are based on confidential LFTTD data and rounded to four significant digits per U.S. Census Bureau Disclosure Guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Appendix D Additional Empirical Results
D.1 Decomposition of Price Dispersion
To explore the sources of price heterogeneity, we report in Table D.1 the results from OLS regressions decomposing price variation using the specification:
$\ln p_{ijht}=FE_{i}+FE_{j}+FE_{ht}+\beta\boldsymbol{X}_{ijht}+\varepsilon_{ijht},$
estimated over the period 2001-2016. We consider three alternative prices: prices that exclude both duties and charges ($\ln p_{ijht}$), prices that include charges but exclude duties ($\ln p^{c}_{ijht}$), and prices that include duties but exclude charges ($\ln p^{\text{duty}}_{ijht}$).
Table D.1 finds that controlling for product and year fixed effects explains approximately 50% of the overall price dispersion, while 4% is attributed to match-specific residuals. Notably, this figure changes substantially when isolating variation within supplier–product–year combinations (Panel B), with the buyer–supplier match accounting for 77% of the price variance. This emphasizes that a significant share of price heterogeneity stems from bilateral characteristics that are not solely attributable to either buyer or supplier individually.
Table D.1: Fixed-Effect Decomposition of Price Dispersion
Source of Variation $\ln p_{ijht}$ $\ln p^{c}_{ijht}$ $\ln p^{\text{duty}}_{ijht}$
Panel A: Overall price dispersion
$FE_{ht}$ 0.483 0.485 0.486
$FE_{i}$ 0.427 0.424 0.423
$FE_{j}$ 0.0452 0.0464 0.0463
Match residual 0.0444 0.0441 0.0442
Panel B: Within exporter–product dispersion
$FE_{j}$ 0.231 0.233 0.233
Match residual 0.768 0.765 0.765
Notes: The columns correspond to alternative price definitions: $\ln p_{ijht}$ excludes both duties and charges; $\ln p^{c}_{ijht}$ includes charges but excludes duties; $\ln p^{\text{duty}}_{ijht}$ includes duties but excludes charges. The estimation sample includes importer–exporter–product matches observed in two consecutive calendar years, and applies the following restrictions: (i) excludes transactions involving consumption goods (based on the BEC classification), energy-sector products, statistical outliers, and related-party trade; and (ii) retains only suppliers that trade with at least two distinct U.S. buyers in consecutive years. This corresponds to the cumulative sample underlying the “+ Supplier Multi-Buyer” row in Panel B of Table C.1. The control vector $\boldsymbol{X}_{ijht}$ includes the log of transaction value, the log of relationship longevity (years since the exporter first supplied the buyer with the given HS10 product), and the log of the relative number of partners (the supplier’s number of HS10-level buyers divided by the buyer’s number of HS10-level suppliers). The sample includes 1,2000,000 importer–exporter–product–year observations, which have been rounded to four significant digits per U.S. Census Bureau disclosure guidelines. $R^{2}=0.956$. All coefficients in a regression model are significantly different from zero at the 1% significance level. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
D.2 Pass-Through Heterogeneity
Table D.2 examines the heterogeneity in tariff pass-through through specification (3.2) by including buyer-by-year fixed effects ($FE_{jt}$). This more demanding specification accounts for time-varying shocks at the buyer level, while also controlling for product–year and exporter country–year fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Findings replicate the results of incomplete tariff pass-through reported in Table 3, suggesting that exporters adjust marginal costs in response to demand shifts from dominant importers, thereby absorbing a substantial fraction of tariff shocks. These results demonstrate the critical role of the cost channel as, by Proposition 3, the pass-through decreases with the buyer share $x_{ij}$.
Table D.2: Pass-Through and Relationship Heterogeneity, Stringent Fixed Effects
Dependent variable: $\Delta\ln p_{ijht}$
(1) (2) (3) (4) (5) (6)
$\Delta\ln(1+\tau_{cht})$0.223 -0.342 -0.230 -0.154 -0.163 -0.292
(0.109) (0.155) (0.096) (0.130) (0.107) (0.145)
$\Delta\ln(1+\tau_{cht})\cdot\ln\text{longevity}_{ijht}$ 0.086 0.097
(0.044) (0.044)
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.023 0.029 0.018
(0.141) (0.157) (0.157)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.271 -0.271 -0.280
(0.127) (0.135) (0.130)
$FE_{ht}+FE_{ct}+FE_{jt}$ Yes Yes Yes Yes Yes Yes
Observations 249,000 249,000 249,000 249,000 249,000 249,000
R-squared 0.31 0.31 0.31 0.31 0.31 0.31
Notes: This table reports estimates of the pass-through of statutory tariffs, $\Delta\ln(1+\tau_{cht})$, to duty-exclusive prices at the exporter–importer–product–year level, $\Delta\ln p_{ijht}$. Columns (2) and (6) interact tariffs with the log of relationship longevity, measured as the number of years that buyer $j$ and supplier $i$ have transacted in product $h$. Columns (3) and (5) interact tariffs with the lagged supplier share, $s_{ijht-1}$, defined as supplier $i$’s share in buyer $j$’s imports of product $h$. Columns (4) and (5) interact tariffs with the lagged buyer share, $x_{ijht-1}$, defined as buyer $j$’s share in supplier $i$’s exports of product $h$. All regressions include product–year, exporter country–year, and importer–year fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Controls include: (i) $\ln\text{longevity}_{ijht}$; (ii) $\Delta\ln q_{i(-j)ht}$, the change in exporter $i$’s total sales of $h$ to U.S. buyers other than $j$; and (iii) $\Delta\ln p_{(-i)jht}$, the weighted average price change charged by other suppliers of $h$ to buyer $j$, using lagged shares as weights. Standard errors are clustered at the HS8 product and exporter-country level. The sample corresponds to the ”+ Supplier Multi-Buyer” definition in Table C.1. Observation counts are rounded to four significant digits per U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Nonlinear Effects
To explore nonlinearities in tariff pass-through along the distribution of bilateral concentration, we interact the tariff change with quartile dummies of the lagged supplier share ($s_{ijh,t-1}$) and buyer share ($x_{ijh,t-1}$). Specifically, we estimate equation D.1, where $\mathbf{1}\{s_{ijh,t-1}\in Q_{q}\}$ and $\mathbf{1}\{x_{ijh,t-1}\in Q_{q}\}$ are indicator variables for quartiles $q=2,3,4$, with the first quartile serving as the omitted category. To separate level and interaction effects, the regression also includes the shares $s_{ijh,t-1}$ and $x_{ijh,t-1}$ themselves as covariates. This specification allows us to test whether pass-through varies nonlinearly across the concentration distribution, while flexibly controlling for underlying differences in market structure. Panel (A) of Figure D.1 includes product–time and exporting country–sector fixed effects, while Panel (B) features product–time, importer–time, and exporting country–time fixed effects.
$\Delta\ln p_{ijht}=$ $\alpha_{0}+\alpha_{1}\Delta\ln(1+\tau_{cht})+\sum_{q=2}^{Q}\alpha_{s,q}\cdot\Delta\ln(1+\tau_{cht})\cdot\mathbf{1}\{s_{ijh,t-1}\in Q_{q}\}$
$+\sum_{q=2}^{Q}\alpha_{x,q}\cdot\Delta\ln(1+\tau_{cht})\cdot\mathbf{1}\{x_{ijh,t-1}\in Q_{q}\}+\gamma^{\prime}\textbf{X}_{ijht}+\textbf{FE}+\epsilon_{ijht}.$ (D.1)
The results reveal no evidence of nonlinearities in pass-through with respect to supplier shares (in purple). Interaction coefficients across the upper quartiles of $s_{ijht}$ are uniformly positive, but small in magnitude and statistically insignificant, indicating that supplier concentration does not materially affect the degree of pass-through. In contrast, buyer shares exhibit a strong, monotonic relationship (in green): pass-through declines significantly at higher quartiles of $x_{ijht}$, consistent with the model’s prediction that dominant buyers constrain suppliers’ ability to shift cost shocks.
Figure D.1: Pass-Through by Bilateral Market Share Quartiles
(A) Fixed Effects: $FE_{ht}+FE_{cs}$
(B) Fixed Effects: $FE_{ht}+FE_{ct}+FE_{jt}$
Notes: The figure plots estimated coefficients from regressions of bilateral price changes on tariff changes interacted with quartiles of supplier share ($s_{ijht}$) and buyer share ($x_{ijht}$). The first quartile serves as the omitted category. The estimated coefficients correspond to equation (D.1), where $\alpha_{s,q}$ and $\alpha_{x,q}$ capture the interaction of the tariff term with the $q$th quartile of supplier and buyer shares, respectively. The top panel includes product-year and exporting country-sector fixed effects. The bottom panel includes product-year, importer-year, and exporting country-year fixed effects. Bars indicate 95% confidence intervals. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Additional Robustness: GE Controls and Price Definitions
Table D.3 presents robustness checks using alternative price definitions and specifications. Columns (1)-(2) exclude the general equilibrium controls from our baseline model. Columns (3)-(4) use duty-exclusive prices that include charges, while Columns (5)-(6) use tariff-inclusive prices. In all cases, we interact tariff changes with lagged supplier and buyer shares, and hold fixed effects constant across specifications for comparability. Across all variations, pass-through estimates and their interaction effects with bilateral market shares remain stable in sign and magnitude, supporting the robustness of the main findings.
Table D.3: Additional Robustness: GE Controls and Price Definitions
Dependent variable: $\ln p_{ijht}$ (excl. GE controls) $\ln p^{c}_{ijht}$ (before duty, incl. charges) $\ln p^{\text{duty}}_{ijht}$ (tariff inclusive)
(1) (2) (3) (4) (5) (6)
$\Delta\ln(1+\tau_{cht})$0.060 -0.153 -0.032 -0.114 0.506 0.401
(0.093) (0.106) (0.091) (0.104) (0.099) (0.126)
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.048 0.036 0.033 0.004 0.017 0.007
(0.069) (0.162) (0.080) (0.165) (0.068) (0.143)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.399 -0.278 -0.407 -0.280 -0.421 -0.287
(0.113) (0.136) (0.112) (0.134) (0.121) (0.145)
$FE_{ht}+FE_{cs}$ Yes No Yes No Yes No
$FE_{ht}+FE_{ct}+FE_{jt}$ No Yes No Yes No Yes
Observations 249,000 249,000 249,000 249,000 249,000 249,000
R-squared 0.04 0.31 0.04 0.31 0.05 0.31
Notes: This table reports robustness checks on tariff pass-through specifications using alternative price definitions and control sets. Columns (1)–(2) exclude general equilibrium controls; Columns (3)–(4) use pre-duty prices including charges; and Columns (5)–(6) use tariff-inclusive prices. In each case, we report specifications using either baseline fixed effects ($FE_{ht}+FE_{cs}$) or a more stringent set of fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Standard errors are clustered at the HS8 product and exporter-country level. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
D.3 Additional Results on Model Estimation
We assess the robustness of our structural estimates in Section 4.2. We first replicate the main GMM estimation using an expanded sample that includes all Broad Economic Categories (BECs), rather than excluding consumption goods as in the baseline. This extended product scope allows us to test whether our key parameter estimates, namely, returns to scale ($\theta$) and relative bargaining power ($\phi$), are sensitive to the exclusion of consumer-oriented products. As shown in Table D.4, the results remain fairly stable, suggesting that the baseline findings are not driven by product composition.
We then examine how sensitive the estimates are to alternative values of calibrated model parameters. First, we vary the elasticity of substitution across foreign varieties ($\rho$) by setting $\rho=5$ instead of 10, which is consistent with the lower end of the estimates in the literature. Second, we relax the assumption of constant returns to scale for the importers and set $\varrho=0.5$ instead of 1. We report the results in Table D.5. Column (1) shows the estimated values when setting $\rho=5$, and Column (2) shows the estimated values when setting $\varrho=0.5$. Furthermore, in Table D.5, we also estimate the parameters using an alternative sample constructed by utilizing the RPT indicator from LFTTD. [^37] See Section 3.3 for the discussion on selection through RPT indicators. Columns (3) and (4) report the estimated values for this set of sample. Throughout these alternative setups,the resulting estimates remain robust, suggesting that the estimated values are not sensitive to particular values of other parameters or set of sample.
Table D.4: Estimated Model Primitives - All BEC Categories
Panel A: Calibrated Parameters $\hat{\nu}$ $\hat{\gamma}$ $\hat{\rho}$ 4 0.5 10 Panel B: Estimated Parameters (GMM) (1) (2) (3) (4) Rel. bargaining power: $\ln\widehat{\frac{\phi}{1-\phi}}$ 1.162 0.558 (0.026) (0.022) Returns to scale ($\hat{\theta}$) 0.505 0.573 0.432 0.586 (0.003) (0.005) (0.004) (0.005) Constant 2.780 0.742 (0.190) (0.083) Longevity 0.061 0.742 (0.023) (0.069) Number of HS10 transactions -0.240 -0.019 (0.019) (0.011) Multiple HS10 dummy 0.080 0.190 (0.026) (0.025) Lagged outside option -0.188 -0.237 (0.017) (0.021) None Yes Yes No No $FE_{h}+FE_{t}+FE_{j}$ No No Yes Yes Observations 6,143,000 Panel C: Implied Bargaining Powers ($\hat{\phi}$) Mean 0.762 0.896 0.636 0.848 (0.005) (0.068) (0.005) (0.112) Median 0.913 0.878 (0.068) (0.112) Notes: This table presents model estimates based on a sample that includes all Broad Economic Categories (BEC), including consumption goods, for the period 2001-2016. Panel A reports calibrated parameters: the elasticity of demand ($\nu$), the elasticity of costs with respect to foreign input prices ($\gamma$), and the elasticity of substitution across foreign varieties ($\rho$). We set $\varrho=1$, so that $\eta=2.5$. Panel B presents GMM estimates. Columns (1) and (3) impose a constant $\phi$ across bilateral pairs, while Columns (2) and (4) estimate the full vector $\boldsymbol{\kappa}$ to allow for heterogeneity in bargaining power. Specifications differ in the inclusion of fixed effects. Controls include: (i) the log of relationship longevity between exporter $i$ and importer $j$; (ii) the log of the number of transactions between $i$ and $j$ in a given year; (iii) the log of the relative outside option, defined as the ratio of exporter $i$’s sales to other U.S. buyers (excluding $j$) over importer $j$’s purchases from other suppliers (excluding $i$), both in year $t{-}1$; and (iv) a dummy variable equal to one if the $i–j$ pair transacts in more than one HS10 product. Panel C reports the mean and median of the implied bargaining power. Standard errors are robust; those in Panel C are computed using the delta method. The set of instruments includes the number of exporters and importers at the HS10 level, as well as lagged bilateral shares (excluding the focal pair). The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Table D.5: Robustness of Model Estimates
$\rho=5$ $\varrho=0.5$ Related Party Trade LFTTD (1) (2) (3) (4) Rel. bargaining power: $\ln\widehat{\frac{\phi}{1-\phi}}$ 1.455 1.838 1.618 0.892 (0.038) (0.052) (0.062) (0.047) Returns to scale ($\theta$) 0.427 0.473 0.453 0.381 (0.004) (0.003) (0.005) (0.007) Mean $\hat{\phi}$ 0.811 0.863 0.835 0.709 (0.006) (0.006) (0.009) (0.010) Median $\hat{\phi}$ 0.811 0.863 0.835 0.709 (0.000) (0.000) (0.000) (0.000) None Yes Yes Yes No $FE_{h}+FE_{t}+FE_{j}$ No No No Yes Observations 3,120,000
Notes: This table reports robustness checks for the main model estimates. The columns explore sensitivity to changes in key calibrated parameters and sample definitions. Column (1) varies the elasticity of substitution across foreign varieties ($\rho$); Column (2) changes the downstream returns to scale parameter ($\varrho$). Columns (3)–(4) use a sample based on related-party indicators from LFTTD (RPT), without and with fixed effects for buyer, product, and time. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
D.4 Additional Results on Model Fit
Sensitivity to alternative parameter values.
Table D.6 presents IV-based goodness-of-fit tests as in equation (4.5) for an alternative set of calibrated parameters of the model and two alternative sets of fixed effects. The results show that models incorporating bargaining and decreasing returns to scale (Columns (1) and (5)) have coefficients closer to one, indicating empirical alignment, consistent with the results in Figure 3.
Quantity responses and relationship heterogeneity: data vs. model.
Table D.7 compares how quantities respond to tariffs in the data (Panel A) and in the model (Panel B). Columns (1) and (2) use baseline fixed effects; Columns (3) and (4) add more demanding fixed effects. Columns (2) and (4) include interactions with supplier and buyer shares.
The table shows that the model generates sizable average quantity declines and predicts heterogeneity across relationships. In contrast, the interaction terms with supplier and buyer shares for the quantity responses in the data are statistically imprecise, suggesting inconclusive evidence of heterogeneity in the data.
Testing the model-predicted quantity changes.
Panel A in Table D.8 presents IV-based tests comparing observed quantity changes to those predicted by the model under alternative parameterizations. The model-predicted change in quantity is given by:
$\widehat{\Delta\ln q_{ijht}}=-\varepsilon_{ijht}\widehat{\Delta\ln p_{ijht}},$
where $\varepsilon_{ijht}$ is the match-specific demand elasticity and $\widehat{\Delta\ln q_{ijht}}$ is as in equation (4.5). Although formal tests reject all models, the baseline model with bargaining and decreasing returns (Columns (1) and (2)) demonstrates the strongest fit, indicating that this specification best captures the underlying mechanisms of tariff-induced quantity adjustments.
Testing the model-predicted sales changes.
Panel B in Table D.8 provides analogous IV-based tests for observed sales changes computed as:
$\widehat{\Delta\ln r_{ijht}}=(1-\varepsilon_{ijht})\,\widehat{\Delta\ln p_{ijht}}.$
The baseline specification (Columns (1) and (2)) shows superior alignment between predicted and observed sales, reinforcing the conclusion that the bargaining model with decreasing returns most effectively matches empirical patterns in the data.
Table D.6: IV-Based Goodness-of-Fit Test, Alternative Parameters
Panel A: Baseline Fixed Effects
Dependent variable: $\Delta\ln p_{ijht}$
$\rho=5$ $\varrho=0.5$
Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$ Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$
(1) (2) (3) (4) (5) (6) (7) (8)
$\widehat{\Delta\ln p_{ijht}}$ 1.211 0.781 0.735 1.368 1.247 0.871 0.740 1.544
(0.343) (0.220) (0.207) (0.387) (0.352) (0.246) (0.209) (0.437)
$FE_{ht}+FE_{cs}$ Yes Yes Yes Yes Yes Yes Yes Yes
Panel B: Stringent Fixed Effects
Dependent variable: $\Delta\ln p_{ijht}$
Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$ Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$
(1) (2) (3) (4) (5) (6) (7) (8)
$\widehat{\Delta\ln p_{ijht}}$ 1.000 0.648 0.612 1.132 1.047 0.720 0.615 1.302
(0.351) (0.227) (0.214) (0.398) (0.368) (0.252) (0.216) (0.458)
$FE_{ht}+FE_{ct}+FE_{jt}$ Yes Yes Yes Yes Yes Yes Yes Yes
Observations 249,000
Notes: Each column reports the coefficient from an IV regression of the observed change in log price on the corresponding model-predicted change $\widehat{\Delta\ln p_{ijht}}$, using statutory tariff changes as instruments. Columns (1)-(4) use predicted prices using a $\rho=5$ instead of $\rho=10$ as in the baseline. Columns (5)-(8) use predicted prices using a $\varrho=0.5$ instead of $\varrho=1$ as in the baseline. Notice that in each case, we re-estimate the calibrated parameters $\phi$ and $\theta$, reported in Appendix D.3 In Panel A, all columns include product-time and country-sector fixed effects ($FE_{ht}+FE_{cs}$). In Panel B, all columns include product-time, country-time, and buyer-time fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Standard errors are clustered at the product and exporter-country level. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Table D.7: Quantity Responses and Relationship Heterogeneity: Data vs. Model
Panel A: Data
(1) (2) (3) (4)
$\Delta\ln(1+\tau_{cht})$0.568 0.006 -1.021 -0.761
(0.249) (0.331) (0.185) (0.237)
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$0.511 -0.278
(0.323) (0.276)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.201 -0.130
(0.229) (0.242)
R-squared 0.06 0.10 0.32 0.36
Panel B: Model
(1) (2) (3) (4)
$\Delta\ln(1+\tau_{cht})$1.962 -3.967 -2.169 -3.984
(0.080) (0.201) (0.108) (0.230)
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 2.441 2.420
(0.409) (0.551)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$ 2.993 3.270
(0.229) (0.316)
R-squared 0.28 0.38 0.44 0.51
$FE_{ht}+FE_{cs}$ Yes Yes No No
$FE_{ht}+FE_{ct}+FE_{jt}$ No No Yes Yes
Observations 249,000
Notes: This table reports the pass-through of tariffs to quantities at the exporter–importer–product level. Panel A presents reduced-form estimates from the data. Panel B shows corresponding pass-through estimates generated by the model. Columns (2)–(4) interact tariff changes with lagged supplier share ($s_{ijht-1}$) and lagged buyer share ($x_{ijht-1}$). Columns (1) and (2) use baseline fixed effects ($FE_{ht}+FE_{cs}$), while Columns (3) and (4) employ a more stringent specification with product–year, country–year, and buyer–year fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Standard errors are clustered at the HS8 product and exporter-country level. Observation counts are rounded to four significant digits per U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Table D.8: IV-Based Goodness-of-Fit Test for Quantities and Sales
Panel A: Quantities
Dependent Variable: $\Delta\ln q_{ijht}$
Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$
(1) (2) (3) (4) (5) (6) (7) (8)
$\widehat{\Delta\ln q_{ijht}}$ 0.349 0.525 0.205 0.311 0.185 0.282 0.420 0.627
(0.144) (0.142) (0.084) (0.084) (0.076) (0.076) (0.174) (0.170)
Panel B: Sales
Dependent Variable: $\Delta\ln(p_{ijht}\cdot q_{ijht})$
Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$
(1) (2) (3) (4) (5) (6) (7) (8)
$\widehat{\Delta\ln(p\cdot q)_{ijht}}$ 0.176 0.426 0.102 0.250 0.092 0.228 0.212 0.509
(0.157) (0.154) (0.091) (0.090) (0.082) (0.082) (0.189) (0.184)
$FE_{ht}+FE_{cs}$ Yes No Yes No Yes No Yes No
$FE_{ht}+FE_{ct}+FE_{jt}$ No Yes No Yes No Yes No Yes
Observations 249,000
Notes: Each column reports the coefficient from an IV regression of the observed change in log quantity (Panel A) or log sales (Panel B) on the corresponding model-predicted change, using statutory tariff changes as instruments. Columns (1), (3), (5), and (7) include product–time and country–sector fixed effects ($FE_{ht}+FE_{cs}$), while Columns (2), (4), (6), and (8) include product–time, country–time, and buyer–time fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$). Standard errors are clustered at the HS8 product and exporter-country level. Observation counts are rounded to four significant digits per U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
D.5 Additional Pass-Through Results
D.5.1 Pass-Through Across Samples
Panel A and Panel B in Table D.9 evaluate how tariff pass-through estimates vary under different sample restrictions using baseline fixed effects (product–time, country–sector) and more stringent fixed effects (including buyer-time), respectively. For ease of exposition, the first two columns of Table D.9 replicate the numbers of Columns (1) and (3) of Table 5. These two columns show that the baseline sample yields a pass-through estimate of around 78-85%, closely matched by the model-implied prediction with no statistically significant difference. As the sample is progressively broadened–from including single-buyer matches (Column (3)), to adding related-party transactions, energy goods, and outliers (Column (4)), and finally to the most inclusive specification (Column (5))–estimated pass-through increases steadily, reaching up to 93-95%. This pattern underscores the sensitivity of reduced-form estimates to sample composition and the role of relationship filtering in uncovering pricing patterns consistent with bilateral bargaining.
Table D.9: Tariff Pass-Through Across Different Samples
Dependent variable: Model Baseline +Suppliers + Energy/ + Final
$\Delta\ln p_{ijht}$ w/ RPT/ Goods
$<$ 2 Buyers Outliers
(1) (2) (3) (4) (5)
Panel A: Baseline Fixed Effects
$\Delta\ln(1+\tau_{cht})$0.248 -0.151 -0.168 -0.099 -0.066
(0.008) (0.093) (0.035) (0.050) (0.043)
$FE_{ht}+FE_{cs}$ Yes Yes Yes Yes Yes
R-squared 0.32 0.04 0.02 0.02 0.02
Panel B: Stringent Fixed Effects
$\Delta\ln(1+\tau_{cht})$0.249 -0.223 -0.171 -0.123 -0.047
(0.010) (0.109) (0.039) (0.059) (0.042)
$FE_{ht}+FE_{ct}+FE_{jt}$ Yes Yes Yes Yes Yes
R-squared 0.50 0.31 0.18 0.14 0.13
Observations 249,000 249,000 732,000 945,000 1,768,000
Notes: This table reports tariff pass-through estimates to duty-exclusive prices at the exporter–importer–product level. Panel A uses baseline fixed effects: product–time and country–sector ($FE_{ht}+FE_{cs}$). Panel B uses a more stringent specification: product–time, country–time, and buyer–time ($FE_{ht}+FE_{ct}+FE_{jt}$). Column (1) uses the model-predicted price change as the dependent variable. Column (2) uses the observed price change in the baseline sample and is identical to Column (1) of Table 3. Column (3) adds relationships in which the supplier trades with only one U.S. importer. Column (4) further expands the sample to include relationships that are either related parties, involve energy commodities, or exhibit extreme price levels or changes. Column (5) incorporates consumption goods, thus encompassing all consecutive exporter-importer-product combinations. Standard errors are clustered at the HS8 product and country level. Observation counts are rounded per U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
D.5.2 Pass-Through Using Data Aggregated at the Product-Level
To complement our main analysis, we replicate a standard pass-through specification using data aggregated at the product-country-month level. We construct the data directly from the buyer–supplier–product triplets that form the basis of our firm-level regressions. While much of the literature analyzes monthly price and tariff changes at the product–country level, such approaches reflect both intensive and extensive margin adjustments, including changes in trading partners or the entry and exit of relationships. In contrast, our aggregation focuses exclusively on consecutive transactions between the same buyer and supplier for a given product. This setup isolates price responses within ongoing relationships, capturing what is arguably the most direct expression of tariff pass-through at the micro level.
Data
We measure price changes at the HS10–country–month level, using the same subset of buyer–supplier–product links as in the baseline analysis. These are links with at least one transaction in both 2017 and 2018. For each product and country, we construct monthly prices by aggregating trade values and quantities. We then relate monthly price changes to changes in statutory tariffs, controlling for product-month, country-month, and country-sector fixed effects. Standard errors are clustered at the HS8–country level.
Models
Table D.10, reports results separately for all products and for the subset that excludes consumption goods. Within each group, columns reflect increasingly selective samples. Columns (1) and (4) includes buyer–supplier–product pairs observed in two consecutive calendar years. The next specification, Columns (2) and (5), excludes consumption goods and applies additional filters: energy-sector goods are dropped; transactions with extreme price levels and price changes are excluded; and related-party trade is removed. The final sample, Columns (3) and (6), is restricted to suppliers trading with at least two U.S. buyers for the same product in consecutive years.
Results
Across specifications, we find consistent evidence of incomplete tariff pass-through to U.S. import prices. As in our firm-level regressions, the degree of pass-through incompleteness increases as we move to more selective samples, particularly those that condition on firms with multiple trading partners over consecutive years. The estimated effects are generally larger in magnitude than at the match level, suggesting that relationship-level frictions may be amplified when observed in aggregated trade flows. While this exercise remains suggestive, it helps connect the mechanisms explored in the main analysis to pricing patterns in product-level data.
Table D.10: Tariff Pass-Through Using Data Aggregated at the Product-Level
Dependent Variable: $\Delta\ln p_{cht}$
All Products Excl. Consumption Goods
(1) (2) (3) (4) (5) (6)
$\Delta\ln(1+\tau_{cht})$0.115 -0.136 -0.321 -0.072 -0.119 -0.700
(0.078) (0.058) (0.186) (0.136) (0.101) (0.164)
Consecutive obs only Yes Yes Yes Yes Yes Yes
(-) Energy/RPT/Outliers No Yes Yes No Yes Yes
(-) Suppliers with $<$2 buyers No No Yes No No Yes
Observations 800,000 540,000 180,000 510,000 320,000 100,000
R-squared 0.13 0.14 0.22 0.13 0.14 0.22
Notes: This table reports regressions of month-over-month changes in log unit values (FOB, excluding charges) on corresponding changes in statutory tariffs ($\Delta\tau_{hct}$), measured at the HS10–country–month level. The data are aggregated from the firm-level sample used in our main analysis, retaining only buyer–supplier–product triplets with consecutive transactions. Columns (1)-(3) refer to all products; Columns (4)-(6) exclude consumption goods, based on the Broad Economic Categories (BEC) classification. Each column reflects a progressively more restricted sample: Columns (1) and (4) include all consecutive transactions; Columns (2) and (5) drop energy goods, extreme price levels and changes, and related-party trade; and Columns (3) and (6) restrict to suppliers with multiple buyers in consecutive years. All regressions include product–month, country–month, and country–sector fixed effects. Standard errors are clustered at the HS8–country level and reported in brackets. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
D.5.3 Pass-Through Results Using LFTTD RPT Indicators
As a robustness check, we re-estimate the baseline specifications using the related-party transaction (RPT) indicator provided in the LFTTD. Table D.11 shows that the pass-through estimates remain stable, and the interaction effects with supplier and buyer shares are qualitatively similar to those reported in Table 3.
Table D.11: Pass-Through and Relationship Heterogeneity–Alternative Related Party Trade
Dependent variable: $\Delta\ln p_{ijht}$
(1) (2) (3) (4)
$\Delta\ln(1+\tau_{cht})$0.163 -0.087 -0.304 -0.253
(0.096) (0.100) (0.111) (0.114)
$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.043 0.011
(0.070) (0.174)
$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.374 -0.222
(0.110) (0.148)
$FE_{ht}+FE_{cs}$ Yes Yes No No
$FE_{ht}+FE_{ct}+FE_{jt}$ No No Yes Yes
Observations 249,000 249,000 249,000 249,000
R-squared 0.04 0.05 0.32 0.32
Notes: This table reports estimates of the pass-through of statutory tariffs, $\Delta\ln(1+\tau_{cht})$, to duty-exclusive prices at the exporter–importer–product–year level, $\Delta\ln p_{ijht}$. Columns (2) and (4) include interactions between tariffs and lagged bilateral characteristics: supplier share ($s_{ijht-1}$), defined as supplier $i$’s share in buyer $j$’s imports of product $h$, and buyer share ($x_{ijht-1}$), defined as buyer $j$’s share in supplier $i$’s exports of product $h$. All regressions include the following controls: (i) $\ln\text{longevity}_{ijht}$, the number of years $i$ and $j$ have transacted in $h$; (ii) $\Delta\ln q_{i(-j)ht}$, exporter $i$’s sales of $h$ to U.S. buyers other than $j$; and (iii) $\Delta\ln p_{(-i)jht}$, the average price change charged by other suppliers of $h$ to buyer $j$, using lagged shares as weights. Columns (1)–(2) include product–year and exporter country–sector fixed effects ($FE_{ht}+FE_{cs}$), while Columns (3)–(4) include a more demanding set of fixed effects: product–year, importer–year, and exporter country–year ($FE_{ht}+FE_{ct}+FE_{jt}$). Standard errors are clustered at the HS8 product and exporter-country level. Arm’s length transactions are defined using LFTTD related party trade indicator. The number of observations is rounded to four significant digits in accordance with U.S. Census Bureau disclosure guidelines. Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
Appendix E Estimation Appendix
E.1 Downstream Demand Elasticity $\left(\nu\right)$
Consider a model where importer $j$ sells its output $q_{j}$ to downstream customers in different countries. A representative consumer in each country maximises utility by choosing a composite of domestic and imported goods. The sub-utility derived from the composite imported good will be given by a CES aggregation across imported varieties with a good-importer specific elasticity of substitution given by $\sigma_{g}$. Broda and Weinstein (2006) provide estimates of the elasticity $\sigma_{g}$ at the HS10 good $g$-level in U.S. import data. The plot below shows the distribution of these elasticities. We base the calibration of the elasticity $\nu$ in our model on these estimates. We consider a value of 4 for $\nu$, close to the mean value of 3.85, which we see as a conservative choice.
Figure E.1: Downstream Demand Elasticity
Notes: The figure displays the estimates of the import demand elasticity $\sigma_{g}$ from Broda and Weinstein (2006). The mean and median value of $\sigma_{g}^{US}$ is 3.85 and 2.8, respectively. Estimates are truncated above at 20, and below at 1.
E.2 Monte Carlo Simulation
Data for one replicate. Each exporter $i\in\{1,\dots,200\}$ belongs to a block with exactly two importers, labeled $j(i)$ and $\ell(i)$. Store the log‐price difference
$\Delta p_{i}\;=\;\ln p_{ij}\;-\;\ln p_{i\ell}\quad\text{and the pair }(j,\ell).$
We set the marginal cost to 1 for all pairs for simplicity and use the parameters of $\varrho=1,\nu=4,\gamma=0.5,\rho=10$. $\phi^{*}=0.827$ and $\theta^{*}=0.454$. $s_{ij}$ and $x_{ij}$ are drawn from a $U[0,1]$ so that all shares within a block sum to 1. All Monte-Carlo exercises use 501 random replicas.
Joint estimation of $\phi$ and $\theta$.
  1. 1.
    Candidate markups. For any $(\phi,\theta)\in(0,1)\times(0,1)$ compute the bilateral markup $\mu_{ij}(\phi,\theta)$ from the structural formula (2.6).
  2. 2.
    Model-implied gap for exporter $i$: ${\Delta\mu}_{i}(\phi,\theta)\;=\;\ln\mu_{ij}(\phi,\theta)\;-\;\ln\mu_{i\ell}(\phi,\theta).$
  3. 3.
    Non-linear least squares criterion.
    $Q(\phi,\theta)\;=\;\sum_{i=1}^{200}\Bigl{[}\,\Delta p_{i}-{\Delta\mu}_{i}(\phi,\theta)\Bigr{]}^{2}.$
  4. 4.
    Estimation. Minimize $Q(\phi,\theta)$ subject to the simple box constraints
    $0.01\leq\phi\leq 0.99,\qquad 0.01\leq\theta\leq 1$
    We record the resulting estimates $\hat{\phi},\hat{\theta}$ for each of the 500 Monte-Carlo replicates.
Results.
Figure E.2 reports the frequencies for the jointly estimated $(\phi,\theta)$ under the parameters noted above. The left panel refers to $\hat{\phi}$ and the right panel to $\hat{\theta}$. Similarly, Figure E.3 shows the analogous results for an alternative set of parameters of $\varrho=1,\nu=2.5,\gamma=0.5,\rho=5$. Across the figures, the true and average estimates are numerically close, and the distributions of the estimates are centered around the true parameters, showing that our estimator is consistent.
Figure E.2: Estimated $(\phi,\theta)$ when $\varrho=1,\nu=4,\gamma=0.5,\rho=10$
Figure E.3: Estimated $(\phi,\theta)$ when $\varrho=1,\nu=2.5,\gamma=0.5,\rho=5$
E.3 Estimation Under $\theta=1$
In Section E.3.1, we first demonstrate that assuming $\theta=1$ in the estimation leads to overestimating $\phi$ when the true parameters are $(\phi^{*},\theta^{*})\in(0,1)^{2}$. Further, in Section E.3.2, we show that using $\phi^{*}$ (instead of the overestimated $\phi$) in the model validation exercise in Section 4.3 where $\theta$ is imposed to be $1$ would yield an upper bound on the attainable correlation between model-predicted and observed price changes.
E.3.1 Estimation Bias When Fixing $\theta=1$ in the Estimation
In what follows, we maintain the following simplification assumptions. We assume $Cor(s_{ij},x_{ij})=0$, mirroring the low correlation between the two bilateral shares (Table 1). We also impose marginal costs to be constant across firms, implying $p_{ij}=\mu_{ij}$ (can be relaxed by assuming a distribution for $\Delta k_{ij\ell}$). Finally, we assume that $\omega(\phi)=\phi$ to keep the notation clean. This assumption is without loss since $\omega$ is increasing in $\phi$.
The objective function when jointly estimating $\phi$ and $\theta$ is:
$\arg\min_{\phi,\theta}R(\phi,\theta)$ $=\mathbb{E}\big{[}p_{ij}-p_{i\ell}-\mu_{ij}(\phi,\theta)+\mu_{i\ell}(\phi,\theta)\big{]}^{2}$ (E.1)
$=\mathbb{E}\big{[}\Delta p_{i}-\Delta\mu_{i}(\phi,\theta)\big{]}^{2}$
where $j$ and $\ell$ are two importers to exporter $i$ and we dropped the $\ln$ sign to simplify the notation. Under the full-rank condition, minimization of (E.1) leads to the estimation of the true parameters $(\theta^{*},\phi^{*})$.
Now define $\tilde{R}(\phi)=R(\phi,\theta)|_{\theta=1}$. When fixing $\theta=1$, the objective function (E.1) becomes:
$\arg\min_{\phi}\tilde{R}(\phi)=\mathbb{E}\big{[}\Delta p_{i}-\Delta\mu_{i}(\phi;\theta=1)\big{]}^{2}.$ (E.2)
To study the bias, we can replace $p_{ij}=(1-\phi^{*})\mu^{oligopoly}_{ij}+\phi^{*}\mu^{oligopsony}_{ij}(\theta^{*})$ and rearrange equation (E.2) as follows:
$\tilde{R}(\phi)=\mathbb{E}\left[(\phi-\phi^{*})\,(\mu^{oligopoly}_{ij}-\mu^{oligopoly}_{i\ell})+\phi^{*}\,(\mu^{oligopsony}_{ij}(\theta^{*})-\mu^{oligopsony}_{i\ell}(\theta^{*}))\right]^{2},$
where, in a slight abuse of notation, we denote the oligopsony markdown computed at the true $\theta^{*}$ as $\mu^{oligopsony}_{ij}(\theta^{*})$ to distinguish it from the oligopsony markdown when $\theta=1$, which is $\mu^{oligopsony}_{ij}(\theta=1)=1$.
We can use the price equation to transform this into a function of only prices (data) and $\mu^{oligopsony}(\theta^{*})$:
$\tilde{R}(\phi)=$ $\mathbb{E}\Bigg{[}(\phi-\phi^{*})\,\left(\frac{p_{ij}-\phi^{*}\mu^{oligopsony}_{ij}(\theta^{*})}{1-\phi^{*}}-\frac{p_{i\ell}-\phi^{*}\mu^{oligopsony}_{i\ell}(\theta^{*})}{1-\phi^{*}}\right)$
$\hskip 28.45274pt+\phi^{*}\,\left(\mu^{oligopsony}_{ij}(\theta^{*})-\mu^{oligopsony}_{i\ell}(\theta^{*})\right)\Bigg{]}^{2}$
$=$ $\mathbb{E}\Bigg{[}\frac{\phi-\phi^{*}}{1-\phi^{*}}\,(p_{ij}-p_{i\ell})-\frac{\phi-\phi^{*}}{1-\phi^{*}}\phi^{*}\left(\mu^{oligopsony}_{ij}(\theta^{*})-\mu^{oligopsony}_{i\ell}(\theta^{*})\right)$
$\hskip 28.45274pt+\phi^{*}\,\left(\mu^{oligopsony}_{ij}(\theta^{*})-\mu^{oligopsony}_{i\ell}(\theta^{*})\right)\Bigg{]}^{2}$
$=$ 1($\frac{1}{(1-\phi^{*})^{2}}\mathbb{E}\left[(\phi-\phi^{*})\,(p_{ij}-p_{i\ell})+\left(\mu^{oligopsony}_{ij}(\theta^{*})-\mu^{oligopsony}_{i\ell}(\theta^{*})\right)\,\Big{(}(1-\phi)\,\phi^{*}\Big{)}\right]^{2}$
$=$ 1($\frac{1}{(1-\phi^{*})^{2}}\mathbb{E}\left[(\phi-\phi^{*})\,\Delta p_{i}+\Delta\mu^{oligopsony}_{i}\Big{(}(1-\phi)\,\phi^{*}\Big{)}\right]^{2}.$
Notice that when $\phi=\phi^{*}$, the first term is zero, but the second term is not. The last equation above can be further rewritten as
$\tilde{R}(\phi)=$ $\left(\phi-\phi^{*}\right)^{2}\cdot\left(\sigma_{p}^{2}+\overline{\Delta p}^{2}\right)+\big{(}\left(1-\phi\right)\cdot\phi^{*}\big{)}^{2}\cdot\left(\sigma_{\Delta\mu^{oligopsony}}^{2}+\overline{\Delta\mu}_{oligopsony}^{2}\right)$ (E.3)
$+2\,\left(\phi-\phi^{*}\right)\cdot\left(1-\phi\right)\cdot\phi^{*}\cdot\bigg{(}\operatorname{Corr}\left(\Delta p_{i},\Delta\mu^{oligopsony}_{i}\right)\cdot\sigma_{p}\cdot\sigma_{\Delta\mu^{oligopsony}}\,+\overline{\Delta p}\cdot\overline{\Delta\mu}_{oligopsony}\bigg{)},$
where we denote the standard deviation and average of variable $z$ by $\sigma_{z}$ and $\overline{z}$ respectively.
Taking derivative with respect to $\phi$ and setting it equal to zero yields:
$\phi=\frac{A\cdot\phi^{*}+B\cdot(\phi^{*})^{2}-\phi^{*}\cdot C\cdot(1+\phi^{*})}{A+B\cdot(\phi^{*})^{2}-2\cdot\phi^{*}C}=\phi^{*}\cdot\,\underbrace{\frac{A+B\cdot\phi^{*}-C\cdot(1+\phi^{*})}{A+B\cdot(\phi^{*})^{2}-2\cdot\phi^{*}C}}_{>1}>\phi^{*}$
where $A=\sigma_{p}^{2}+\overline{\Delta p}^{2}>0,\;B=\sigma_{\mu^{oligopsony}}^{2}+\overline{\Delta\mu}_{oligopsony}^{2}>0,\;C=\operatorname{Corr}\left(\Delta p_{i},\Delta\mu^{oligopsony}_{i}\right)\cdot\sigma_{p}\cdot\sigma_{\mu^{oligopsony}}+\overline{\Delta p}\cdot\overline{\Delta\mu}_{oligopsony}>0$.
Therefore, the objective (E.2) would estimate $\phi>\phi^{*}$ when setting $\theta=1$. This argument can be extended to heterogeneous marginal costs across pairs. In that case, we would replace $\Delta\mu_{i}$ for $\Delta p_{i}$.
Simulation
We repeat the simulations from Appendix E.2, now estimating only $\phi$ (with $\theta$ fixed at 1). We draw $s_{ij}\sim U[0,1]$, injecting enough within-exporter variation to identify $\phi$ from differences in markups. If $s_{ij}$ varies very little for exporter $i$, then $\ln\mu(s_{ij},\phi)-\ln\mu(s_{i\ell},\phi)\approx 0,$ so $\phi$ cannot be identified as the markup difference cancels out in the moment (4.1).
Panel (A) of Figure E.4 maintains 2 importers and exporters in each market, showing a large bias. In Panel (B) of Figure E.4, we increase the number of buyers and suppliers to create more cross-sectional variation. The figure confirms our argument above: The estimated value of $\phi$ is larger than the true $\phi^{*}$ when the estimation imposes $\theta=1$.
Figure E.4: Estimation of the parameter $\phi$ while fixing $\theta=1$
(a) 2 importers and exporters
(b) 10 importers and exporters
Notes: In each similuation there are either 2 (Panel a) or 10 (Panel b) buyers and suppliers in each market. There are 100 markets. 501 simulations. We set $\varrho=1,\nu=4,\gamma=0.5,\rho=10$ and compute $\phi$ by minimum distance.
E.3.2 Correlation between Pass-Through and Price Changes
In the previous section we proved that the estimated $\phi$ is necessarily larger than the true value of $\phi^{*}$, when one imposes $\theta=1$ in the estimation. In the exercise of Section 4.3, we do not re-estimate $\phi$ and use $\phi^{*}$ when testing the model under $\theta=1$.
In this section, we show that the coefficients reported for $\theta-1$ of Figure 3 are upper bounds on the attainable correlations between predicted and observed price changes.
In what follows, we drop $\ln$ and subscripts to simplify the notation. Call the re-estimated $\phi$ when setting $\theta=1$ as $\phi^{R}$. Then, combining equations (4.5) and (4.4) in Section 4.3, the pass-through regression for this model is
$\Delta p\;=\;\beta^{R}\,\cdot\,\Phi^{R}\,\cdot\,\Delta T\,+\,u^{R},$ (E.4)
where we denoted $\Phi^{R}=\Phi(s,x;\phi^{R},\theta=1)$. [^38] We disregard the fixed effects and consider each variable as demeaned for simplicity. We also assume that the residual is independent from $\Phi^{R}\,\cdot\,\Delta T$.
From Appendix E.3.1, setting $\theta=1$ and re-estimating $\phi$, we would find $\phi^{R}>\phi^{*}$, so that
$\tilde{\Phi}\,=\,\Phi(s,x;\phi^{*},1)\,<\,\Phi(s,x;\phi^{R},1)\,=\,\Phi^{R},$
since the pass-through increases in $\phi$.
Using $\tilde{\Phi}$ as an independent variable instead of $\Phi^{R}$ in the OLS regression (E.4),
$\Delta p=\tilde{\beta}\,\cdot\,\tilde{\Phi}\,\cdot\,\Delta T+\tilde{u}$
yields the following estimated coefficient:
$\tilde{\beta}=\frac{\operatorname{Cov}(\tilde{\Phi}\,\Delta T,\;\Delta p)}{\operatorname{Var}(\tilde{\Phi}\,\Delta T)}=\beta^{R}\,\frac{\operatorname{Cov}(\tilde{\Phi}\,\Delta T,\;\Phi^{R}\,\Delta T+u^{R})}{\operatorname{Var}(\tilde{\Phi}\,\Delta T)}.$
From independence of $\Delta T$ from $\Phi^{R}$ and $\tilde{\Phi}$ we know that
$\operatorname{Cov}(\tilde{\Phi}\,\Delta T,\;\Phi^{R}\,\Delta T)=\operatorname{Var}(\tilde{\Phi}\,\Delta T)+\operatorname{Cov}(\Phi^{R},\tilde{\Phi})\,\cdot\,\operatorname{Var}(\Delta T),$
which implies
$\frac{\operatorname{Cov}(\tilde{\Phi}\,\Delta T,\;\Phi^{R}\,\Delta T)}{\operatorname{Var}(\tilde{\Phi}\,\Delta T)}=1+\frac{\operatorname{Cov}(\Phi^{R},\tilde{\Phi})\,\cdot\,\operatorname{Var}(\Delta T)}{\operatorname{Var}(\tilde{\Phi}\,\Delta T)}>1$
since $\operatorname{Cov}(\Phi^{R},\tilde{\Phi})>0$ and $\operatorname{Var}(\cdot)>0$. Hence,
$\tilde{\beta}=\beta^{R}\,\cdot\,\Bigl{(}1+\frac{\operatorname{Cov}(\Phi^{R},\tilde{\Phi})\,\operatorname{Var}(\Delta T)}{\operatorname{Var}(\tilde{\Phi}\,\Delta T)}\Bigr{)}\,>\,\beta^{R}.$
Thus, since $\tilde{\beta}$, estimated with $\phi^{*}$ and $\theta=1$, exceeds the pass-through coefficient $\beta^{R}$ that would be obtained by re-estimating with $\phi=\phi^{R}$ and $\theta=1$ (i.e. $\tilde{\beta}>\beta^{R}$), $\tilde{\beta}$ constitutes an upper bound on the pass-through coefficient in equation (E.4). Consequently, the estimates for $\theta-1$ of Figure 3—obtained under $\phi=\phi^{*}$ and $\theta=1$ via the pass-through formula—can be interpreted as upper bounds on the coefficients one would recover by re-estimating with $\phi=\phi^{R}$ at $\theta=1$.

### v2

@@
 Abstract
-This paper develops a theory of bargaining in firm-to-firm trade with two-sided market power.
-The framework accommodates flexible market structures, yielding analytical expressions for pair-specific markups and pass-through elasticities.
-In U.S.
-import data, we estimate strong importer bargaining power and an upward-sloping export supply curve, consistent with oligopsony power.
-Pass-through of the 2018 tariffs in firm-to-firm relationships is incomplete, in contrast to product-level studies, primarily due to exporter cost reductions driven by falling demand from dominant buyers.
-Our study highlights how bargaining and network rigidities shape price outcomes, with implications for markup dispersion and shock propagation in global value chains.
+We develop and estimate a structural model of bargaining in firm-to-firm trade to study the determinants of tariff pass-through.
+The model features oligopoly and oligopsony power and yields analytical expressions for bilateral markups and pass-through based on two sufficient statistics: the supplier’s share in the buyer’s purchases and the buyer’s share in the supplier’s output.
+Using U.S.
+import data, we find substantial importer bargaining power and steep export supply curves.
+These primitives imply that cost changes, rather than markup adjustments, dominate pass-through, accounting for the bulk of incomplete pass-through of the 2018 U.S.
+tariffs and its heterogeneity across buyer-supplier links.
 1 Introduction
-The recent resurgence of protectionist trade policies has reignited debate over how tariffs affect international prices.
-This question is particularly salient in today’s global economy, where production is fragmented across global value chains and most trade occurs within ongoing relationships between importers and exporters.
-These firm-to-firm linkages shape the cross-border transmission of shocks (Boehm et al., 2019); however, their implications for price setting and tariff pass-through remain less well understood.
-A growing body of literature highlights the limitations of standard trade models in this context.
-Rather than engaging in anonymous spot-market transactions, firms often trade within long-term relationships shaped by contract incompleteness, relationship-specific investments, and lock-in frictions (Antràs, 2015).
-In this environment, prices arise from bilateral negotiations between exporters and importers, both of whom may hold significant market power (Antràs and Staiger, 2012; Morlacco, 2019; Dhyne et al., 2022).
-This paper develops a theory of bargaining in firm-to-firm trade that incorporates two-sided market power and network rigidities.
-The framework accommodates flexible upstream and downstream market structures and delivers closed-form expressions for markups and pass-through elasticities as functions of market shares and a small set of structural parameters.
-We estimate these parameters using U.S.
-import data, validate the model’s cross-sectional predictions against observed pricing patterns, and use the estimated framework to quantify aggregate pass-through in firm-to-firm trade, highlighting how bargaining and network structure shape both the level and sources of pass-through in response to trade shocks.
-Section 2 introduces the theory.
-We consider a fixed network of importer-exporter (or buyer-supplier) relationships, where each pair negotiates bilaterally over the price of an intermediate input.
-Given the price, the importer unilaterally chooses the traded quantity to minimize costs, making prices allocative.
-In the event of a disagreement, each party falls back on trade with its other existing partners.
-As a result, the network structure influences the value of each party’s outside option, and in turn, their effective bargaining power.
+International trade is organized around a network of buyers and suppliers that transact repeatedly across borders and jointly generate rents along global supply chains.
+How these rents are divided, and who bears cost shocks, has first-order implications for the incidence of trade policy.
+Yet canonical trade models abstract from bilateral relationships, thereby missing key mechanisms of shock transmission.
+We develop and estimate a structural model of firm-to-firm trade where two-sided market power determines the level and heterogeneity of relationship-level markups and pass-through.
+Applied to the 2018 U.S.
+tariffs, our model explains within-relationship pass-through as the outcome of bilateral bargaining between importers and exporters.
+Pass-through of a cost shock reflects two distinct forces.
+The first is a markup channel, whereby the negotiated price-cost margin adjusts to the shock, reallocating surplus within the match.
+The second is a cost channel: at the higher effective price, large buyers contract demand, moving suppliers down their supply curves and lowering marginal cost.
+The relative strength of these forces is ultimately an empirical question that depends on underlying primitives.
+A central primitive is the inverse export supply elasticity, which not only governs oligopsony power but also shapes the balance between the two channels of transmission through the slope of exporter’s marginal cost curve.
+With bilateral bargaining, these elasticities are match-specific, generating systematic heterogeneity in pass-through across trading relationships that models with atomistic buyers or suppliers are unable to capture.
+We estimate the model using pre 2017 U.S.
+import data and apply it to analyze the 2018 tariff episode.
+The estimates indicate strong importer bargaining power and steep export supply curves, consistent with substantial oligopsony power.
+At these parameter values, the cost channel drives tariff pass-through, with marginal cost adjustments dominating markup responses.
+Pass-through varies systematically across relationships and declines with buyer share, so larger importers face lower pass-through through a higher residual inverse supply elasticity.
+In the aggregate, pass-through is incomplete at 67 to 73 percent, implying that exporters absorb roughly one third of the tariff burden, through reductions in scale and quasi-rents.
+This match-level perspective contrasts with product-level studies that find near complete pass-through of the 2018 tariffs, [^1] We reconcile the match level and product level evidence later in the introduction.
+while reduced-form evidence supports the model predictions relative to alternatives without bargaining or upward-sloping export supply.
+The model extends the network-based oligopoly framework of kikkawa2019imperfect to incorporate two-sided market power and decreasing returns to scale, which generate upward-sloping export supply and capture lock-in effects in global supply chains (antras2015global).
+Prices are set through bilateral Nash bargaining, with outside options given by each party’s alternative incumbent relationships.
+The equilibrium yields analytical expressions for bilateral markups and pass-through as functions of two sufficient statistics: the supplier share, measuring the exporter’s concentration among the buyer’s input expenditure, and the buyer share, measuring the importer’s concentration among the exporter’s total input supply.
 Our framework offers a unified theory of exporter and importer market power in firm-to-firm trade.
-Exporter rents arise from three sources: monopoly power through product differentiation, oligopoly power due to supplier concentration, and quasi-rents from an upward-sloping residual supply curve.
-When exporters hold full bargaining power, the model nests standard trade settings in which prices reflect an oligopoly markup over marginal cost (Atkeson and Burstein, 2008; Dhyne et al., 2022).
-This markup increases with the exporter’s share of the buyer’s total input purchases—a measure we refer to as the exporter’s supplier share.
-Our first contribution is to characterize the role of importer market power in price-setting.
-This power arises when suppliers earn economic rents, creating scope for buyers to extract surplus by negotiating lower prices.
-It also reflects oligopsony power, as concentrated buyers can negotiate lower prices.
-When importers hold full bargaining power, prices can fall below marginal cost, generating a negative price–cost margin akin to an oligopsony markdown.
-This markdown declines with the importer’s share of the supplier’s total output–a measure we refer to as the importer’s buyer share.
-[^1] Our definition of the importer’s markdown differs from the standard price–marginal revenue wedge.
+On the exporter side, rents arise from product differentiation and supplier concentration, expressed as an oligopoly markup over marginal cost that rises with the exporter’s supplier share, as in standard models (e.g.
+Atkeson2008).
+On the importer side, buyer power takes two distinct forms.
+Positive bargaining power gives a buyer countervailing power: the ability to extract a share of supplier rents, without affecting their scale.
+A positive inverse export supply elasticity further generates quasi-rents, resulting in an oligopsony markdown: a price-cost ratio below one that declines with the importer’s buyer share, as larger buyers capture a greater share of those quasi-rents through bargaining.
+[^2] In standard oligopsony models, oligopsony power operates through quantity distortions: buyers restrict demand to reduce the price they face.
+In our model, quantities always lie on buyers’ demand curves and there are no such distortions in equilibrium; the inverse supply elasticity nonetheless governs oligopsony power through the bargaining channel.
 See Section 2 for details.
-In general, the negotiated price reflects a convex combination of oligopoly and oligopsony forces, with weights determined by the firms’ outside options and other exogenous determinants of bargaining power.
-Our second contribution is to characterize how importer market power shapes the pass-through of tariff shocks to bilateral prices.
-We focus on short-run pass-through elasticities, derived as local perturbations around the static price equilibrium, while holding the network and general equilibrium conditions fixed.
-This isolates the intensive margin of price adjustment from longer-run responses driven by entry, reallocation, or macroeconomic feedback.
-Importer market power affects pass-through through two distinct mechanisms: a markup channel and a cost channel.
-The markup channel captures how bilateral markups respond to price changes.
-When exporters hold full bargaining power, markups exhibit strategic complementarities, a well-known source of incomplete pass-through (Amiti et al., 2014, 2019a).
-With two-sided market power, the markup elasticity becomes a convex combination of strategic complementarities and strategic substitutabilities, as importers reduce markdowns in response to higher prices, amplifying pass-through.
-The cost channel captures how tariffs affect the exporter’s marginal cost through scale effects.
-While prior work links pass-through to cost adjustment under decreasing returns to scale (e.g., Burstein and Gopinath 2015), our framework introduces a new source of heterogeneity: the elasticity of residual supply increases with an importer’s buyer share.
-As a result, larger buyers face more incomplete pass-through.
-Although both channels interact to shape price responses, the cost channel dominates when bargaining power is two-sided and returns to scale are decreasing.
-In this setting, markup adjustments are modest and offset by opposing strategic forces, while cost-side responses amplify the heterogeneity of pass-through.
-Section 3 describes the data and presents preliminary empirical evidence.
+Under two-sided market power, the equilibrium price-cost margin is a convex combination of the oligopoly markup and the oligopsony markdown, with bargaining power as weight.
+We study pass-through within existing buyer-supplier relationships: local price responses holding fixed all other match prices and general-equilibrium variables, and abstracting from longer-run adjustments driven by relationship entry and exit.
+A novel insight is that the two primitives that govern oligopsony power, namely the inverse export supply elasticity and bargaining power, also govern the pass-through determinants: the former determines whether pass-through reflects cost or markup adjustments; the latter shapes the nature of markup adjustment.
+When the inverse supply elasticity is high, demand contractions produce large movements in exporters’ marginal cost, and the cost channel drives most of the pass-through, regardless of bargaining power; when it is low, the cost channel is more muted and markup adjustments dominate.
+Within the markup channel, bargaining power governs the nature of strategic interaction: low buyer bargaining power implies strategic complementarities, which dampen pass-through (e.g., amiti2014importers; amiti2019international); high buyer bargaining power implies strategic substitutabilities arising from oligopsony power, which amplify it.
+We take the model to U.S.
+transaction-level customs data covering the 2018 tariff episode.
 Our main data source is the U.S.
-Census Longitudinal Firm Trade Transaction Database (LFTTD), which reports prices and quantities for each import transaction involving a U.S.
-importer, along with identifiers for the foreign exporter and the 10-digit product code.
-These features allow us to construct bilateral market shares on both sides of the transaction, which are the key sufficient statistics in the model governing the distribution of markups and pass-through elasticities.
-Our main sample focuses on repeated, arm’s-length relationships involving the exchange of intermediate inputs, excluding related-party links as flagged in the LFTTD.
-Pricing and pass-through patterns in the data closely align with the model’s predictions under two-sided market power.
-Controlling for supplier–product–time fixed effects to proxy for marginal costs, we find that prices rise with the exporter’s supplier share and fall with the importer’s buyer share, consistent with the presence of both oligopoly and oligopsony forces.
-In contrast, tariff pass-through declines with buyer share and is non-decreasing in supplier share, suggesting that price adjustments are primarily driven by the cost channel, with a more limited role for the markup channel, as expected under two-sided market power.
-These empirical comovements reinforce the model’s central mechanisms and are difficult to reconcile with conventional pricing frameworks that assume one-sided or purely competitive behavior.
-Nonetheless, reduced-form estimates may conflate the effects of competitor behavior, input price variation, or broader macroeconomic shocks, limiting their interpretability.
-A structural approach is therefore needed to isolate the underlying mechanisms.
-[^2] For a related discussion in a different context, see Berger et al.
-(2022).
-To do so, Section 4 develops a structural estimation strategy around two key parameters: the importer’s bargaining power and the exporter’s returns to scale, which determines the slope of the input supply curve.
-We estimate that U.S.
-importers hold bargaining power of approximately 0.8, roughly four times greater than that of their foreign suppliers, and that the returns to scale parameter is about 0.45, consistent with upward-sloping supply.
-Identification exploits cross-sectional price variation across U.S.
-importers within a supplier–product–year.
-As the estimating equations hold independently of the general equilibrium environment, the approach remains valid without explicitly modeling broader equilibrium feedback.
-We assess the model’s validity by evaluating its ability to replicate both the average and heterogeneous effects of the 2018 tariffs on prices and quantities.
-First, the model closely matches the average price response as well as its variation with buyer and supplier shares, consistent with the role of oligopoly, oligopsony, and cost-side adjustments.
-Second, applying the IV-based test proposed by Adão et al.
-(2023), we find that our model outperforms alternatives that either impose constant returns to scale or exclude bargaining.
-Third, while the model underestimates the magnitude of quantity changes, it successfully captures their direction, lending support to its core allocative mechanism.
-Together, these results validate the model’s key assumptions and suggest that general equilibrium feedback or omitted strategic interactions are unlikely to materially bias its pass-through predictions.
-In Section 5, we use the estimated model to assess how the 2018 Trump tariffs affected aggregate U.S.
-import prices through firm-to-firm relationships.
-Rather than providing a full accounting of tariff incidence, which lies beyond the scope of our framework, we focus on one key margin: price adjustment within ongoing matches, and its decomposition into changes in markups and marginal costs.
-We find that aggregate pass-through in firm-to-firm trade is substantially incomplete, with elasticities ranging from 65 to 71%.
-The model sheds light on the mechanisms behind this muted response.
-The key insight is that most of the adjustment occurs on the cost side: exporters facing weaker demand from powerful buyers move down their marginal cost curves, absorbing a sizable share of the tariff burden.
-Markup changes occur but contribute little to the aggregate response due to offsetting effects of strategic complementarities and substitutabilities.
-This distinction is important.
-While prior work often attributes incomplete pass-through to variable markups and strategic pricing, our results highlight the dominant role of cost adjustment and supply-side forces.
-Our results stand in contrast to studies documenting near-complete pass-through of the 2018 tariffs (e.g., Amiti et al., 2019b, 2020; Cavallo et al., 2020; Fajgelbaum et al., 2020; Flaaen et al., 2020).
-We reconcile this divergence on two grounds.
-First, our analysis focuses on intermediate inputs and arm’s-length transactions, where pass-through may be lower due to input specificity, contractual frictions, or capacity constraints.
-Second, we isolate pricing within repeated matches between importers and exporters.
-By contrast, product-level unit values used in aggregate studies conflate within-match price changes with compositional shifts in trading partners, notably including one-off exchanges.
-Such one-off transactions are more likely to reflect marginal cost pricing and exhibit full pass-through.
-[^3] See Cajal-Grossi et al.
-(2023) for evidence that markups are lower and prices more competitive in spot transactions than in relational matches.
+Census Longitudinal Firm Trade Transaction Database (LFTTD), which records values and quantities of import transactions at the importer-exporter-product level.
+Our main sample focuses on repeated, arm’s-length relationships involving the exchange of intermediate and capital goods.
+Identification exploits cross-sectional price variation across buyers sourcing from the same exporter, product, and year: comparing prices across buyers within an exporter-product-year cell differences out exporter-level cost variation, so that remaining price differences reflect bilateral price-cost margins, which the model maps to observed market shares and unobserved parameters.
+We estimate importer bargaining weights of approximately 0.8, four times those of foreign suppliers, and a (short-run) returns-to-scale parameter around 0.5, implying an average residual inverse export supply elasticity of 0.3 for the average link.
+[^3] This estimate is below standard returns-to-scale estimates from the production function literature, reflecting a fundamentally different object: short-run supply constraints within existing relationships rather than long-run technological returns to scale.
+We discuss this in Section 4.
+Under constant returns to scale, this elasticity would be zero and both oligopsony power and the cost channel would vanish.
+Price and pass-through patterns in the data closely align with the model’s predictions.
+Controlling for marginal costs, prices rise with the exporter’s supplier share and fall with the importer’s buyer share, consistent with two-sided market power.
+Out of sample, the model closely replicates the 2018 tariff episode, matching both the average level of pass-through (78% in the model, against 87% in the data) and its cross-sectional gradient: a one-unit increase in buyer share lowers pass-through by 0.42 in the model, against 0.41 in the data.
+Quantity responses point in the right direction but are estimated less precisely.
+Specifications imposing constant returns to scale generate a near-zero interaction between pass-through and buyer share and are readily rejected.
+A specification retaining decreasing returns but setting bargaining power to zero performs comparably on this gradient: consistent with the theory, under strong decreasing returns the cost channel dominates and the gradient is insensitive to bargaining power.
+The two specifications differ, however, on average pass-through.
+Strong importer bargaining power implies strategic substitutabilities among buyers that partially offset the cost channel, while zero bargaining power implies strategic complementarities that reinforce it, pushing pass-through further below one.
+The baseline model outperforms its no bargaining counterpart in matching average pass-through (78% vs.
+72%), as it lies closer to the empirical point estimate, although both fall within the confidence interval.
+We use the estimated model to study aggregate pass-through in firm-to-firm trade and decompose tariff incidence within these relationships.
+Weighting by initial import shares, aggregate pass-through is substantially incomplete, with elasticities between 66 and 73%, below the unweighted average of 87% because larger (high-share) buyers experience the lowest pass-through.
+The adjustment operates almost entirely through the cost channel: shutting down markup elasticities leaves aggregate pass-through largely unchanged, whereas eliminating cost elasticities raises it significantly.
+Roughly one third of the tariff burden is thus borne by foreign exporters, absorbed not through compressed margins but through lost quasi-rents as lower demand moves exporters down their marginal-cost schedules.
+A related literature documents near complete pass-through of the 2018 tariffs at the product level, implying that U.S.
+buyers bore most of the tariff burden (amiti2019impact; amiti2020s; flaaen2019production; fajgelbaum2020return; cavallo2020tariff).
+Our finding of incomplete pass-through within continuing relationships helps reconcile these results.
+Product-level estimates combine within-relationship price adjustment with changes in the set of active trading partners.
+Our results indicate that a full account of aggregate incidence requires understanding how the set of trading relationships responds to tariffs.
 Related Literature
-Our paper contributes to the literature on pricing and shock transmission in firm-to-firm trade, with a particular focus on the role of importer market power.
-A growing body of empirical work highlights the influence of dominant buyers in shaping supplier outcomes.
-Bernard et al.
-(2019) and Bernard et al.
-(2022) document the pivotal role of large buyers in determining supplier performance and market access.
-Using French customs data, Fontaine et al.
-(2020) show substantial variation in unit values across importers transacting with the same exporter, consistent with buyer-specific pricing.
-Similar patterns are found by Huang et al.
-(2021) in France, Chile, and China.
-Among the few studies to model buyer power directly, Morlacco (2019) estimates substantial oligopsony power among French importers, while Atkin et al.
-(2024) show that bargaining between Argentinian importers and foreign suppliers affects the price impact of trade policy.
-Formal models of pricing in firm-to-firm trade remain limited.
-Dhyne et al.
-(2022) develops a network-based model of oligopoly in which suppliers’ markups increase with their relationship-specific market share.
-Our framework nests this model as a special case but extends it to incorporate bilateral bargaining and importer market power.
-Grossman et al.
-(2024) also study firm-to-firm pricing under Nash-in-Nash bargaining, emphasizing how supply-chain links adjust in response to trade shocks.
-In contrast, we take the trade network as fixed and focus on how bargaining and network rigidities interact to determine equilibrium prices and short-run pass-through.
-We also contribute to the literature on the firm-level determinants of cost shock pass-through.
-A large body of work attributes incomplete pass-through to variable markups and strategic complementarities in price-setting (e.g., Amiti et al., 2014, 2019a).
-[^4] See also Berman et al.
-(2012), Auer and Schoenle (2016), and Garetto (2016) for firm-level evidence consistent with this class of models.
-We extend these insights to a bilateral oligopoly setting that reflects the structure of a large share of international trade.
-Our framework nests these existing mechanisms but allows for a richer characterization of both markup and cost channels, shaped by bargaining power and network position.
-Related work includes Gopinath and Itskhoki (2011) and Goldberg and Tille (2013), who model bargaining in firm-to-firm trade but abstract from importer market power and cost-side adjustment.
-Empirically, we find that these cost-based responses, rather than variable markups, are the primary driver of incomplete pass-through in these settings.
-Finally, we contribute to the literature on shock propagation in production networks.
-A large body of work demonstrates that shocks can spread through input–output linkages and affect aggregate outcomes (Acemoglu et al., 2012; Di Giovanni et al., 2014; Grassi, 2018), often using natural experiments to trace transmission along supply chains (Barrot and Sauvagnat, 2016; Boehm et al., 2019; Carvalho et al., 2021).
-Recent work by Acemoglu and Tahbaz-Salehi (2025) highlights how bilateral bargaining can amplify supply chain disruptions through the price channel.
-We complement this perspective by modeling how idiosyncratic shocks to individual relationships affect prices under two-sided market power.
-While we abstract from general equilibrium feedback, our focus is on the within-network transmission of trade shocks through pricing, a key but underexplored margin in the propagation of shocks across production networks.
+Our paper contributes to three main strands of the literature.
+First, the paper relates to the growing literature on importer market power.
+Early contributions emphasize the role of large buyers in shaping supplier outcomes (bernard2019production; bernard2022origins), while more recent firm-to-firm evidence documents systematic price discrimination across buyers, consistent with buyer power (Fontaine2019; huang2024firm; ignatenko2024market; macedoni2024pricing).
+Despite this evidence, formal theories of buyer power in international trade remain limited.
+[^4] Alternative explanations include endogenous matching, whereby more productive buyers source from a larger set of suppliers and intensify upstream competition (huang2024firm), and supplier-side price discrimination, under which buyers with stronger outside options obtain lower markups (ignatenko2024market).
+Exceptions include Morlacco2019, who estimates substantial oligopsony power among French importers, and atkin2024trade, who show that importer-supplier bargaining shapes the effects of trade policy.
+We develop a tractable framework that nests kikkawa2019imperfect and extends it to incorporate importer oligopsony power via bilateral bargaining and upward-sloping export supply.
+These ingredients are quantitatively central: together they account for cross-sectional markup comovements and the level and heterogeneity of pass-through.
+Second, the paper relates to the literature on determinants of incomplete pass-through using micro-level data.
+A large literature attributes incomplete pass-through to strategic markup adjustment under oligopoly (e.g., amiti2014importers; amiti2019international), while an alternative line emphasizes cost-side adjustment via nontraded distribution and retail costs in the importing country (goldberg2013structural).
+Our cost channel instead arises upstream, from exporters’ decreasing returns and demand contractions by large buyers.
+[^5] Prior work on pass-through under bargaining (gopinath2011search; goldberg2013bargaining) imposes constant marginal costs and therefore abstracts from oligopsony and the cost channel.
+Which channel prevails is state-dependent: in our setting, where short-run supply constraints bind, the cost channel dominates, while the standard markup approach is nested as a special case under constant marginal costs and applies naturally to longer-run horizons.
+Finally, we contribute to the macro literature on shock propagation in production networks.
+A large body of work shows how input-output linkages affect the propagation of idiosyncratic shocks (Acemoglu2012; Grassi2018).
+[^6] Empirical evidence confirms these mechanisms.
+See, for example, barrot2016input, boehm2019input, and carvalho2021supply.
+A common feature of this literature is that prices are either taken as given or governed by fixed markups, so shocks propagate primarily through quantities.
+acemoglu2025macroeconomics allow for bilateral bargaining to study the consequences of supply chain disruptions.
+In their framework, however, price distortions play no role.
+We complement this approach by characterizing how unit prices respond to cost shocks under bilateral bargaining with decreasing returns, and showing that price adjustment itself is a quantitatively important transmission channel.
+The remainder of the paper is organized as follows.
+Section 2 develops the theoretical framework and derives the results that guide the empirical analysis.
+Section 3 describes the data and presents reduced form evidence on prices and pass-through.
+Section 4 outlines the structural estimation strategy and reports the results.
+Section 5 examines aggregate pass-through in firm-to-firm trade.
+Section 6 concludes.
 2 Theoretical Framework
-This section develops a bargaining theory of firm-to-firm trade with two-sided market power.
-The model links markups and pass-through elasticities to importer and exporter market shares, along with a small set of structural parameters that capture key features of market structure upstream and downstream.
-The model yields analytic structural equations and testable predictions, which we later bring to the data using U.S.
-customs records.
-We impose the following assumptions to keep the model tractable.
-First, we treat the trade network as fixed.
-This assumption captures the lock-in effects in firm-to-firm trade and allows us to focus on the price-setting problem while abstracting from firms’ decisions to form or sever links.
-[^5] This separation between extensive and intensive margin decisions is consistent with prior work showing that dynamic sourcing choices, while important for long-run outcomes, are not necessary to analyze the short-run price effects of trade shocks (Blaum et al., 2018).
-Second, we abstract from nominal rigidities, such as fixed-price contracts or currency denomination, as these are unlikely to materially affect tariff pass-through.
-[^6] Evidence from Amiti et al.
-(2020) supports this view, showing similar short- and long-run pass-through rates of 2018 U.S.
-import tariffs.
-See Goldberg and Tille (2013) for a theory of importer-exporter bargaining over the transaction price and exchange rate exposure.
-Third, we consider a static framework of single-product negotiations, despite our data reflecting repeated, multi-product interactions between firms.
-[^7] Repeated relationships affect contract structure and trade volumes (Macchiavello and Morjaria, 2015), exchange rate pass-through (Heise, 2024), and the propagation of shocks (Martin et al., 2023).
-We return to these features and discuss how we incorporate information on relationship duration in Section 3 and 4.
+This section develops a tractable pricing theory of firm-to-firm trade.
+We focus on how bilateral bargaining shapes prices and quantities within existing trade relationships, taking the network structure as given to isolate short-run tariff transmission.
+The framework is static and considers single-product negotiations; repeated interactions and multi-product relationships are incorporated in reduced form through the estimated bargaining parameters (see Section 4).
+We also abstract from nominal rigidities, such as fixed-price contracts or currency denomination, which appear to play a limited role in tariff pass-through.
+[^7] For example, amiti2020s shows similar short- and long-run pass-through for the 2018 U.S.
+tariffs.
+All proofs and derivations are in Appendix A.
 2.1 Environment
@@
 We denote by $\mathcal{Z}_{i}$ the set of importers connected to exporter $i$, and by $\mathcal{Z}_{j}$ the set of exporters connected to importer $j$.
-These sets vary across firms and are treated as given.
-Exporters and Supply
-Exporter $i$ produces $q_{i}$ units of the unique input variety and sells them to all importers in $\mathcal{Z}_{i}$, where total output satisfies $q_{i}\equiv\sum_{j\in\mathcal{Z}_{i}}q_{ij}$, and $q_{ij}$ denotes the quantity of the intermediate input purchased by importer $j$.
-We assume that exporter $i$ operates a short-run production technology with returns to scale parameter $\theta\in(0,1]$. Marginal cost (or equivalently, the short-run supply curve) is given by:
-$c_{i}\equiv MC_{i}(q_{i})=k_{i}\,q_{i}^{\frac{1-\theta}{\theta}},$ (2.1)
-where $k_{i}$ captures exogenous factors such as the exporter productivity or foreign wages.
-This cost structure implies that the average cost of $q_{i}$ units is $\theta$-times the marginal cost.
-When $\theta<1$, indicating decreasing returns to scale, the short-run average cost lies below the marginal cost, and the gap between them widens as $q_{i}$ increases.
-In contrast, $\theta=1$ indicates a constant returns technology, with marginal and average costs constant and equal to $k_{i}$.
-Importers and Demand
-Importer $j$ combines domestic and foreign inputs, denoted by $q_{j}^{d}$ and $q_{j}^{f}$ respectively, to produce $q_{j}$ units of a final good, which is subsequently sold in a downstream market.
-The production technology of importer $j$ is given by:
-$q_{j}=\varphi_{j}\left(q_{j}^{f}\right)^{\gamma}\left(q_{j}^{d\vphantom{f}}\right)^{\varrho-\gamma},$ (2.2)
-where $\varphi_{j}$ is the importer’s productivity, $\gamma$ and $\varrho-\gamma$ are the output elasticities of foreign and domestic inputs, respectively, with $\varrho$ governing the degree of returns to scale in production.
-[^8] The domestic input $q_{j}^{d}$ can be interpreted as a constant returns to scale aggregator of primary factors, such as labor and domestic intermediates.
-Due to the lack of data on domestic input purchases, we model foreign and domestic intermediates as distinct rather than symmetric or substitutable.
-The foreign input $q_{j}^{f}$ is a constant elasticity of substitution (CES) composite of differentiated input varieties sourced from exporters in the set $\mathcal{Z}_{j}$.
-Each variety $q_{ij}$ is weighted by a demand shifter $\varsigma_{ij}$, and the elasticity of substitution across varieties is given by $\rho>1$:
-$q_{j}^{f}=\left(\sum_{i\in\mathcal{Z}_{j}}\varsigma_{ij}\,q_{ij}^{\frac{\rho-1}{\rho}}\right)^{\frac{\rho}{\rho-1}}.$ (2.3)
-Downstream, each importer operates in a monopolistically competitive market, facing CES demand with constant elasticity $\nu>1$ and an importer-specific demand shifter $D_{j}$.
+These sets vary across firms and are taken as given.
+Exporters and Supply.
+Exporter $i$ supplies a differentiated input to importers in $\mathcal{Z}_{i}$. Total exported quantity is $q_{i}=\sum_{j\in\mathcal{Z}_{i}}q_{ij}$, where $q_{ij}$ denotes importer $j$’s demand.
+The exporter’s marginal cost is
+$MC_{i}(q_{i})=k_{i}\,q_{i}^{\frac{1-\theta}{\theta}},$ (2.1)
+where $\theta\in(0,1]$ governs returns to scale and $k_{i}$ captures exogenous cost shifters, such as productivity or foreign wages.
+When $\theta<1$, the technology exhibits decreasing returns, so marginal cost rises with output.
+When $\theta=1$, marginal cost is constant and equal to $k_{i}$.
+The implied total cost function is $TC_{i}(q_{i})=\theta\,k_{i}\,q_{i}^{\frac{1}{\theta}}$, so average cost equals $\theta$ times marginal cost.
+Importers and Demand.
+Importer $j$ produces a final good $q_{j}$ by combining a foreign input ($q_{j}^{f}$) with a domestic input ($q_{j}^{d}$) according to the production function
+$q_{j}=\varphi_{j}\,(q_{j}^{f})^{\gamma}\,(q_{j}^{d})^{\varrho-\gamma},$ (2.2)
+where $\varphi_{j}$ denotes productivity, $\gamma$ the output elasticity of foreign inputs, and $\varrho$ the returns-to-scale parameter.
+[^8] Since domestic input purchases are not observed, we impose a unit elasticity of substitution between foreign and domestic inputs, which fixes the cost share of $q_{j}^{d}$ and $q_{j}^{f}$.
+Under this restriction, $q_{j}^{d}$ can be interpreted as a constant-returns aggregator of primary factors such as labor or domestic intermediates.
+The foreign input consists of a bundle of input varieties supplied by the exporters in $\mathcal{Z}_{j}$ and aggregated according to a CES technology:
+$q_{j}^{f}=\left(\sum_{i\in\mathcal{Z}_{j}}\varsigma_{ij}\,q_{ij}^{\,\frac{\rho-1}{\rho}}\right)^{\!\frac{\rho}{\rho-1}},$ (2.3)
+where $\varsigma_{ij}$ is a demand shifter, and $\rho>1$ is the substitution elasticity across input varieties.
+Downstream, importer $j$ sells its final good in a monopolistically competitive market and faces CES demand with elasticity $\nu>1$. Revenue is therefore given by
+$R_{j}(q_{j})=D_{j}\,q_{j}^{\frac{\nu-1}{\nu}},$ (2.4)
+where $D_{j}$ denote an exogenous demand shifter.
 2.2 Bargaining Protocol and Gains From Trade
-Each $i-j$ vertical interaction proceeds in two stages. In the first stage, the importer chooses input quantity $q_{ij}$ to minimize total input costs, taking the price $p_{ij}$ as given:
-$\min_{q_{ij}}\;TC_{j}(q_{ij};p_{ij})=p_{ij}q_{ij}+\sum_{\ell\in\mathcal{Z}_{j}\setminus i}p_{\ell j}q_{\ell j}+p^{d}_{j}q^{d}_{j}.$
-The solution to this problem yields importer $j$’s input demand:
-$q_{ij}=q_{j}^{f}\,\varsigma_{ij}^{\rho}\left(\frac{p_{ij}}{p_{j}^{f}}\right)^{-\rho}\quad\text{with}\quad q_{j}^{f}=\gamma\,c_{j}\,q_{j}\left(p_{j}^{f}\right)^{-1},$ (2.4)
-where $p_{j}^{f}=\left(\sum_{i}\varsigma_{ij}^{\rho}p_{ij}^{1-\rho}\right)^{\frac{1}{1-\rho}}$ is the foreign input price index.
-In the second stage, the importer and exporter determine the negotiated price $p_{ij}$ through Nash bargaining, taking the importer’s demand in equation (2.4) as given:
-$\max_{p_{ij}}\left(\underbrace{\pi^{i}(p_{ij};q_{ij})-\tilde{\pi}^{i}_{(-j)}}_{GFT_{ij}^{i}}\right)^{1-\phi}\left(\underbrace{\pi^{j}(p_{ij};q_{ij})-\tilde{\pi}_{(-i)}^{j}}_{GFT_{ij}^{j}}\right)^{\phi}.$ (2.5)
-In equation (2.5), $\phi\in(0,1)$ denotes the importer’s bargaining leverage, and $GFT_{ij}^{k}$ for $k\in\{i,j\}$ captures the gains from trade, defined as the difference between a firm’s profits with and without the bilateral relationship.
-In case of disagreement, each party falls back on trade with its other existing partners.
-All bilateral matches generate strictly positive surplus, and trade occurs in equilibrium across all relationships.
-To solve equation (2.5), we adopt the Nash-in-Nash bargaining solution concept, in which each bilateral negotiation takes as given the outcomes of all other matches in the network (Horn and Wolinsky, 1988; Collard-Wexler et al., 2019).
-[^9] A common alternative is the sequential bargaining model of Stole and Zwiebel (1996), in which disagreement with one partner triggers renegotiation with others.
-This framework is suited to one-to-many settings such as labor markets, but is less appropriate for global supply chains, where firms engage in many-to-many relationships and isolated breakdowns rarely affect unrelated contracts.
-Moreover, it introduces additional complexity without yielding clear empirical gains in our context.
-For this reason, we leave the dependence on prices and quantities in other links in the network implicit throughout the analysis.
-This setup implies two key properties.
-First, prices are allocative: the negotiated price $p_{ij}$ pins down the traded quantity $q_{ij}$.
-Second, the equilibrium allocation $(p_{ij},q_{ij})$ satisfies the importer’s demand in equation (2.4).
-We discuss the allocative implications of these assumptions, along with alternative bargaining protocols, in Section 2.5.
-2.3 Equilibrium Prices and Allocations
-The solution to (2.5) yields a bilateral price of the form $p_{ij}=\mu_{ij}\cdot c_{i},$ where $c_{i}$ is exporter $i$’s marginal cost, and $\mu_{ij}$ is a pair-specific markup.
-We define two bilateral market shares that serve as key sufficient statistics in the model:
-$s_{ij}$ $\equiv\frac{p_{ij}q_{ij}}{\sum_{k\in\mathcal{Z}_{j}}p_{kj}q_{kj}}\in[0,1],\quad\quad x_{ij}\equiv\frac{q_{ij}}{\sum_{k\in\mathcal{Z}_{i}}q_{ik}}\in[0,1],$
-where $s_{ij}$ denotes exporter $i$’s supplier share, defined as its sales to importer $j$ as a fraction of $j$’s total foreign input expenditures, and $x_{ij}$ denotes importer $j$’s buyer share, equal to its purchases as a share of exporter $i$’s total output.
-The following proposition characterizes the equilibrium markup as a convex combination of two limiting pricing regimes:
+Each bilateral interaction between importer $j$ and exporter $i$ unfolds in two stages.
+In the first stage, for a given output target $q_{j}$, importer $j$ chooses foreign inputs $\{q_{\ell j}\}_{\ell\in\mathcal{Z}_{j}}$ and domestic input $q_{j}^{d}$ to minimize total cost, taking input prices as given:
+$\min_{\{q_{\ell j}\},\,q_{j}^{d}}\quad\sum_{\ell\in\mathcal{Z}_{j}}p_{\ell j}q_{\ell j}+p_{j}^{d}q_{j}^{d}\quad\text{s.t.}\quad\varphi_{j}(q_{j}^{f})^{\gamma}(q_{j}^{d})^{\rho-\gamma}\geq q_{j}.$ (2.5)
+Solving this problem yields the standard CES demand for input varieties:
+$q_{ij}=\varsigma_{ij}^{\rho}\left(\frac{p_{ij}}{p_{j}^{f}}\right)^{-\rho}q_{j}^{f},$ (2.6)
+where $q_{j}^{f}$ is the foreign input aggregator and $p_{j}^{f}$ the associated price index, satisfying $p_{j}^{f}q_{j}^{f}=\gamma\,c_{j}\,q_{j}$, with $c_{j}=MC_{j}(q_{j})$ denoting importer $j$’s marginal cost.
+In the second stage, the two parties bargain over the price $p_{ij}$, taking the demand schedule (2.6) as given.
+[^9] This protocol is analogous to the employer “right-to-manage” model in labor economics (e.g., azkarate2025union), where wages are bargained while firms retain the choice of employment.
+Operating profits are given by $\pi^{i}_{ij}=\sum_{k\in\mathcal{Z}_{i}}p_{ik}q_{ik}-TC_{i}(q_{i})$ for the exporter and $\pi^{j}_{ij}=R_{j}(q_{j})-\sum_{\ell\in\mathcal{Z}_{j}}p_{\ell j}q_{\ell j}-p_{j}^{d}q_{j}^{d}$ for the importer, where $TC_{i}(q_{i})$ and $R_{j}(q_{j})$ are defined in Section 2.1.
+If bargaining fails and link $(i,j)$ is severed, each party falls back on profits from its remaining matches, denoted $\tilde{\pi}_{(-j)}^{i}$ and $\tilde{\pi}_{(-i)}^{j}$.
+We define the gains from trade as $GFT_{ij}^{k}\equiv\pi^{k}_{ij}-\tilde{\pi}_{(-\cdot)}^{k}$ for $k\in\{i,j\}$, which measure the value each firm derives from the match.
+For the exporter, these gains correspond to revenue from $j$ net of the incremental cost of serving that buyer.
+For the importer, losing variety $i$ raises the input price index $p_{j}^{f}$, increasing marginal costs, reducing optimal output, and lowering downstream revenues.
+In our setting, these gains are strictly positive.
+[^10] Appendix B.3 discusses an extension with more general nonparametric outside options.
+The negotiated price solves a Nash bargaining problem that splits these gains, with $\phi\in(0,1)$ denoting the importer’s exogenous bargaining weight:
+$\max_{p_{ij},\,q_{ij}}\;\left(GFT_{ij}^{i}\right)^{1-\phi}\;\left(GFT_{ij}^{j}\right)^{\phi}\quad\text{s.t.}\quad q_{ij}=q_{j}^{f}\varsigma_{ij}^{\rho}\left(\frac{p_{ij}}{p_{j}^{f}}\right)^{-\rho}.$ (2.7)
+We adopt the Nash-in-Nash solution concept, taking prices and quantities on all other links as given (Horn1988; collard2019nash).
+2.3 Price Characterization
+To characterize equilibrium prices, it is useful to summarize each bilateral relationship with two market shares that capture each party’s relative importance to the other.
+For a given link $(i,j)$, the exporter’s supplier share $s_{ij}$ is its share of the importer’s foreign input expenditure; the importer’s buyer share $x_{ij}$ measures its share of the exporter’s total output.
+Formally,
+$s_{ij}\equiv\frac{p_{ij}q_{ij}}{\sum_{k\in\mathcal{Z}_{j}}p_{kj}q_{kj}},\qquad\text{and}\qquad x_{ij}\equiv\frac{q_{ij}}{\sum_{k\in\mathcal{Z}_{i}}q_{ik}}.$ (2.8)
+We also define $\xi(GFT_{ij}^{j},p_{ij})$ as the elasticity of the importer’s gains from trade with respect to the bilateral price:
+$\xi(GFT_{ij}^{j},p_{ij})\equiv-\frac{d\ln GFT_{ij}^{j}}{d\ln p_{ij}}=\frac{(\eta-1)s_{ij}}{1-(1-s_{ij})^{\frac{\eta-1}{\rho-1}}}.$ (2.9)
+With these objects in hand, we can now present our first key theoretical result.
 Proposition 1.
-The bilateral markup $\mu_{ij}$ can be expressed as
-$\mu_{ij}=(1-\omega_{ij})\cdot\mu_{ij}^{\textnormal{oligopoly}}+\omega_{ij}\cdot\mu_{ij}^{\textnormal{oligopsony}},$ (2.6)
-where the weight
-$\omega_{ij}=\frac{\frac{\phi}{1-\phi}\lambda_{ij}}{1+\frac{\phi}{1-\phi}\lambda_{ij}}\in(0,1)$ (2.7)
-represents the importer’s effective bargaining power, and $\lambda_{ij}\equiv-\frac{d\ln\pi_{j}}{d\ln(p_{ij}\,q_{ij})}\cdot\frac{\pi_{j}}{GFT_{ij}^{j}}$ captures the strength of the importer’s outside option.
-Proof: See Appendix A.1.
-Equation (2.6) shows that the equilibrium markup lies between two polar cases.
-When the exporter holds all the bargaining power ($\phi\to 0$), the markup converges to the oligopoly case $\mu_{ij}^{\text{oligopoly}}$.
-When the importer holds full bargaining power ($\phi\to 1$), it converges to the oligopsony case $\mu_{ij}^{\text{oligopsony}}$.
-For intermediate values, the outcome reflects the influence of both parties, with weight $\omega_{ij}$ determined by bargaining power and the importer’s outside option (see Section 2.3.3).
-We now examine these limiting cases in more detail and characterize how the bilateral markup depends on market shares and model primitives.
-2.3.1 Oligopoly Markup
-Under full exporter bargaining power ($\phi=0$), the markup takes the standard form:
-$\mu_{ij}^{\text{oligopoly}}$ $=\frac{\varepsilon_{ij}}{\varepsilon_{ij}-1}\geq 1,$ (2.8)
-where $\varepsilon_{ij}$ denotes the residual demand elasticity faced by exporter $i$, given by:
-$\varepsilon_{ij}=(1-s_{ij})\cdot\rho+s_{ij}\cdot\eta.$ (2.9)
-This elasticity is a weighted average of two components: the elasticity of substitution across foreign suppliers, $\rho$, and the elasticity $\eta$ of the importer’s foreign input bundle $q_{j}^{f}$ with respect to its price index $p_{j}^{f}$:
-$\eta\equiv-\frac{d\ln q_{j}^{f}}{d\ln p_{j}^{f}}=\frac{(\varrho-\gamma)+\nu\cdot(1-(\varrho-\gamma))}{\varrho+\nu\cdot(1-\varrho)}.$
-Provided that $\rho>\eta$, a standard parameter condition, the exporter’s markup (2.8) increases with its supplier share $s_{ij}$, reflecting greater oligopoly power.
-This case corresponds to the firm-to-firm trade setting in Dhyne et al.
-(2022).
-It is also closely related to the oligopolistic competition model of Atkeson and Burstein (2008), but differs in three key respects.
-First, supplier shares are defined at the match level, rather than at the firm or industry level, reflecting our assumption of a fixed network.
-Second, whereas the outer nest elasticity $\eta$ is typically treated as a fixed preference parameter, here it summarizes how downstream market structure–determined by both technology ($\gamma$, $\varrho$) and demand ($\nu$)–shapes the bargaining environment.
-Third, we allow for decreasing returns to scale in production, in contrast to the constant returns assumption common in related models.
-Figure 1 illustrates the equilibrium outcome for a representative $i-j$ match.
-The downward-sloping residual demand and marginal revenue curves (black and gray) intersect with the upward-sloping residual marginal and average cost curves (blue and red).
-Panel (A) depicts the case with $\phi=0$.
-The equilibrium quantity is determined by the intersection of marginal cost and marginal revenue, and the price lies on the residual demand curve.
-Exporter rents have two components: the oligopoly rent (red), which corresponds to the markup in equation (2.8), and the quasi-rent (purple), which arises from the gap between marginal and average cost under decreasing returns to scale.
-Figure 1: Illustration of Equilibrium Allocations With Different Bargaining Power
-(a) $\phi=0$
-(b) $\phi\in(0,1)$
-(c) $\phi=1$
+The solution to (2.7) yields the price–cost ratio $\mu_{ij}\equiv p_{ij}/MC_{i}$, which satisfies
+$\mu_{ij}=(1-\omega_{ij})\,\mu^{\textnormal{oligopoly}}(s_{ij})+\omega_{ij}\,\mu^{\textnormal{oligopsony}}(x_{ij}),$ (2.10)
+with
+$\omega_{ij}=\frac{\phi\,\xi(GFT_{ij}^{j},p_{ij})}{\phi\,\xi(GFT_{ij}^{j},p_{ij})+(1-\phi)(\varepsilon_{ij}-1)}\in(0,1).$ (2.11)
+Proof. See Appendix A.1.
+The first component in equation (2.10), $\mu^{\textnormal{oligopoly}}_{ij}$, is the standard oligopoly markup arising under Bertrand competition with differentiated products:
+$\mu_{ij}^{\text{oligopoly}}=\frac{\varepsilon_{ij}}{\varepsilon_{ij}-1}\geq 1\qquad\text{where}\qquad\varepsilon_{ij}=(1-s_{ij})\rho+s_{ij}\eta.$ (2.12)
+This expression coincides with the equilibrium price-cost ratio when the exporter has full bargaining power ($\phi=0$), in which case $\omega_{ij}=0$.
+Here, the residual demand elasticity $\varepsilon_{ij}$ is a weighted average of the within-input elasticity $\rho$ and the elasticity
+$\eta\equiv-\frac{d\ln q_{j}^{f}}{d\ln p_{j}^{f}}=\frac{(\varrho-\gamma)+\nu(1-(\varrho-\gamma))}{\varrho+\nu(1-\varrho)},$ (2.13)
+which depends on technology ($\gamma,\varrho$) and demand ($\nu$).
+[^11] This is different than in standard nested-CES models with Bertrand competition, where $\eta$ is typically a primitive preference parameter (e.g., Atkeson2008).
+When $\rho>\eta$, the markup increases in the supplier share $s_{ij}$, reflecting greater oligopoly power.
+The second component in equation (2.10), $\mu^{\textnormal{oligopsony}}_{ij}$, is given by
+$\mu_{ij}^{\text{oligopsony}}=\theta\cdot\frac{1-(1-x_{ij})^{1/\theta}}{x_{ij}}\leq 1.$ (2.14)
+This expression corresponds to the equilibrium price–cost ratio when the importer has full bargaining power ($\phi=1$), in which case $\omega_{ij}=1$.
+Since $\mu_{ij}^{\text{oligopsony}}\leq 1$, we refer to it as an oligopsony markdown.
+[^12] Although we refer to $\mu_{ij}^{\text{oligopsony}}$ as a markdown, it is formally a price–cost ratio.
+This differs from the conventional definition of a markdown as the wedge between the input price and the input’s marginal revenue product.
+The oligopsony markdown has a natural quasi-rents interpretation.
+Under full importer bargaining power, the importer extracts all exporters’ surplus and bargain a price equals the exporter’s incremental cost of serving them.
+Expressed relative to current marginal cost, the corresponding price-cost ratio is
+$\mu_{ij}^{\text{oligopsony}}=\frac{TC_{i}(q_{i})-TC_{i}(q_{i}-q_{ij})}{MC_{i}(q_{i})\,q_{ij}}.$
+The numerator is the exporter’s actual cost saving if the relationship with buyer $j$ is severed.
+The denominator is what that cost would be if all units sold to $j$ were valued at current marginal cost.
+Under decreasing returns, the incremental cost of serving $j$ lies below current marginal cost, allowing the negotiated price to fall below marginal cost.
+This gap, which owes to quasi rents, captures oligopsony power in the model.
+This markdown decreases with the buyer share $x_{ij}$.
+Larger buyers account for a greater share of the exporter’s output, so walking away from the relationship generates larger cost savings for the exporter and strengthens the importer’s bargaining position.
+As $x_{ij}\to 0$, the buyer is atomistic, the incremental cost of serving the match converges to marginal cost, and $\mu_{ij}^{\text{oligopsony}}\to 1$, so that $p_{ij}=MC_{i}$.
+As $x_{ij}\to 1$, the importer is a monopsonist, the negotiated price converges to average cost, and $\mu_{ij}^{\text{oligopsony}}\to AC_{i}/MC_{i}=\theta$.
+[^13] Our oligopsony mechanism is observationally equivalent to quantity discounts across a supplier’s buyers: conditional on a supplier and product, marginal cost is common across buyers, yet buyers that account for a larger share of the supplier’s output $(x_{ij})$ obtain higher markdowns and pay lower negotiated prices.
+This mechanism operates only under decreasing returns to scale.
+When $\theta=1$, marginal and average costs coincide, leaving no surplus to extract.
+As a result, the negotiated price cannot fall below marginal cost, and the markdown in (2.14) equals one for all $x_{ij}$.
+Bargaining Weights.
+Equation (2.10) expresses the bilateral markup as a weighted average of the oligopoly markup (2.12) and the oligopsony markdown (2.14).
+The weight $\omega_{ij}$ governs the relative importance of these two forces.
+It is increasing in the exogenous bargaining parameter $\phi$, but it is also endogenous because it depends on both the importer’s gains-from-trade elasticity $\xi(GFT_{ij}^{j},p_{ij})$ and the residual demand elasticity $\varepsilon_{ij}$.
+Thus, even for a given value of $\phi$, effective bargaining power varies systematically across matches.
+In particular, for $\phi\in(0,1)$, the weight $\omega_{ij}$ is hump-shaped in the supplier share $s_{ij}$.
+When $s_{ij}$ is small, the bilateral price has little effect on the importer’s overall costs, so the importer has little to gain from bargaining aggressively.
+When $s_{ij}$ is very large, the importer is highly exposed to the relationship but also depends heavily on it, which weakens its outside option.
+Effective bargaining power is therefore strongest at intermediate supplier shares, where cost exposure is significant, but substitution remains feasible.
+Equilibrium Allocation.
+Figure 1: Equilibrium Allocations With Different Bargaining Power
 Notes: Each panel illustrates the negotiated price and quantity under a different value of $\phi$.
@@
 All panels assume $\theta<1$.
-2.3.2 Oligopsony Markdown
-Under full importer bargaining power ($\phi=1$), the bilateral markup simplifies to:
-$\mu_{ij}^{\text{oligopsony}}=\theta\cdot\frac{1-(1-x_{ij})^{\frac{1}{\theta}}}{x_{ij}}\leq 1,$ (2.10)
-which we refer to as an oligopsony markup, or markdown, since it lies weakly below one.
-[^10] Although $\mu_{ij}^{\text{oligopsony}}$ denotes a price–cost ratio, we refer to it as a markdown because it falls below one.
-This differs from standard oligopsony models, where markdowns typically reflect a wedge between input prices and marginal revenue product.
-See Section 2.5 for further discussion.
-Expression (2.10) characterizes the lowest price at which the exporter is willing to supply, with the importer extracting the exporter’s entire surplus through bargaining.
-[^11] Formally, the markdown is proportional to the percentage increase in quasi-rents attributable to the $i-j$ match.
-See equation (A.9) in Appendix A.1.
-The markdown declines with the buyer’s share $x_{ij}$: larger buyers induce greater quasi-rents and can negotiate lower prices.
-As $x_{ij}\to 0$, the importer behaves atomistically, generates no quasi-rents, and pays a price equal to marginal cost.
-This dependence of the price–cost ratio on $x_{ij}$ reflects a form of oligopsony power that arises only under decreasing returns to scale.
-When $\theta=1$, marginal and average costs coincide, quasi-rents vanish, and the importer can extract only oligopoly rents.
-In this case, the price cannot fall below marginal cost, and the markdown remains fixed at one.
-Panel (C) of Figure 1 illustrates the equilibrium under $\phi=1$.
-Compared to Panel (A), the price lies at the intersection of the downstream demand curve and the exporter’s residual average cost curve, leaving no surplus to the exporter.
-The result is a lower price and higher quantity relative to the oligopoly case.
+Figure 1 illustrates the equilibrium allocation for a single $i$-$j$ match under alternative bargaining weights $\phi$.
+The exporter faces a downward-sloping residual demand curve and its associated marginal revenue curve, shown in black and grey, respectively.
+The importer faces an upward-sloping residual marginal and average cost curves, shown in blue and red.
+Panel (A) shows the oligopoly benchmark ($\phi=0$).
+Quantity is pinned down by the intersection of marginal revenue and marginal cost, and price is then determined from residual demand.
+Exporter rents can be decomposed into two components: oligopoly rents, shown in dark blue, equal to (pMC)q(p-MC)q, and quasi-rents, shown in light blue, equal to the gap between marginal and average cost under decreasing returns, (MCAC)q(MC-AC)q.
+Panel (B) shows full importer bargaining power ($\phi=1$).
+The negotiated allocation lies at the intersection of residual demand and average cost curves, leaving the exporter with zero net gains from trade.
+In this case, the importer captures the full quasi-rents generated by decreasing returns.
+Relative to the oligopoly benchmark, the negotiated price is lower and the traded quantity is higher.
+Panel (C) shows the intermediate case ($\phi\in(0,1)$).
+The negotiated outcome lies between the two polar regimes, consistent with Proposition 1.
+The striped area highlights the corresponding range of price–marginal cost ratios, which may lie either above or below one depending on bargaining power and the magnitude of quasi-rents.
+The grey shaded region represents total exporter surplus, equal to (pAC)q(p-AC)q, as bargaining power varies between the two extremes.
+As $\phi$ rises, the importer captures a larger share of the surplus, so exporter rents decline and converge to zero at $\phi=1$.
+Markups and Market Shares.
+Both components of the bilateral markup vary systematically with bilateral market shares.
+A higher supplier share $s_{ij}$ raises the oligopoly markup, while a higher buyer share $x_{ij}$ lowers the oligopsony markdown.
+The co-movements between prices and bilateral shares are thus informative about the presence of the two forces, as summarized in the following proposition.
+Proposition 2.
+The bilateral markup $\mu_{ij}$ satisfies:
+  1. 1.
+    Supplier Share. If $\mu_{ij}$ increases in $s_{ij}$, then $\phi<1$.
+  2. 2.
+    Buyer Share. If $\mu_{ij}$ decreases in $x_{ij}$, then $\phi>0$ and $\theta<1$.
+Proof: See Appendix A.2.
+These cross-sectional predictions guide the empirical analysis in Section 3.4.
 Clarification of Terminology
-The terms “buyer power” and “oligopsony power” are often used interchangeably in the literature, but rarely precisely defined (Noll, 2005). We distinguish between these concepts in our framework.
-We use buyer power or importer market power to refer broadly to the importer’s ability to influence prices through bargaining (i.e., any $\phi>0$).
-In contrast, we reserve oligopsony power for the specific case when a higher buyer share $x_{ij}$ enables the importer to negotiate lower prices, holding $\phi$ fixed.
-Importantly, oligopsony power requires both $x_{ij}>0$ and $\theta<1$, as it depends on the scale-dependent gap between marginal and average cost.
-2.3.3 Two-Sided Market Power and Bargaining Weights
-Panel (B) of Figure 1 shows the intermediate case with $\phi\in(0,1)$.
-The equilibrium price lies on the thick blue curve and reflects a weighted average of the outcomes in Panels (A) and (C), consistent with Proposition 1.
-The resulting price–marginal cost ratio can exceed or fall below one, depending on the relative bargaining power of the exporter and importer.
-Determinants of $\omega_{ij}$.
-The bargaining weight $\omega_{ij}$ in equation (2.6), defined in equation (2.7), governs the balance between oligopoly and oligopsony forces in price setting.
-It depends both on the exogenous bargaining power $\phi$ and an endogenous term $\lambda_{ij}$, defined as:
-[…]
-To interpret $\lambda_{ij}$, we decompose it into two terms.
-The cost exposure component, $\lambda_{ij}^{C}=\frac{(\eta-1)s_{ij}}{\varepsilon_{ij}-1}$, captures how sensitive importer $j$’s profits are to price changes from supplier $i$ and increases with the supplier share $s_{ij}$.
-The network dependence term, $\lambda_{ij}^{N}=\left(1-(1-s_{ij})^{\frac{\eta-1}{\rho-1}}\right)^{-1}$, reflects how much of importer $j$’s profits rely on the match with $i$, and declines with $s_{ij}$ as reduced diversification weakens the importer’s outside option.
-Taken together, $\lambda_{ij}$ captures both the importer’s incentive to negotiate aggressively (through cost exposure) and the strength of its fallback option (through network dependence).
-It follows a hump-shaped pattern in $s_{ij}$, increasing at low values and declining at higher ones, and converges to 1 as $s_{ij}\to 0$ or $s_{ij}\to 1$.
-This interpretation clarifies the meaning of $\omega_{ij}$ as a measure of the importer’s effective bargaining power.
-While $\phi$ governs baseline influence in Nash bargaining, $\lambda_{ij}$ endogenously adjusts it based on the structure of the trading relationship.
-Testable Implications.
-Equation (2.6) writes the bilateral markup $\mu_{ij}$ as a convex combination of the oligopoly markup and the oligopsony markdown, with bargaining power as weight.
-Since these components scale with bilateral market shares, the model delivers clear predictions on how markups co-move with $s_{ij}$ and $x_{ij}$ under two-sided market power.
-Proposition 2.
-The bilateral markup $\mu_{ij}$ exhibits the following properties:
-  1. 1.
-    Markup and Exporter’s Share. If $\mu_{ij}$ increases with $s_{ij}$, then $\phi<1$.
-  2. 2.
-    Markup and Importer’s Share. If $\mu_{ij}$ decreases with $x_{ij}$, then $\phi>0$ and $\theta<1$.
-Proof: See Appendix A.2.
-Proposition 2 provides sufficient conditions for detecting two-sided market power.
-A positive relationship between $\mu_{ij}$ and supplier share $s_{ij}$ implies $\phi<1$, since exporters retain pricing power and $\mu_{ij}^{\text{oligopoly}}$ increases with $s_{ij}$.
-If $\phi=1$, prices are fully disciplined by buyers, and no such pattern should emerge.
-While the dependence of $\omega_{ij}$ on $s_{ij}$ may attenuate this effect, a positive slope remains inconsistent with full buyer power.
-Conversely, a negative relationship between $\mu_{ij}$ and buyer share $x_{ij}$ signals oligopsony power.
-This requires both positive bargaining leverage for importers ($\phi>0$) and decreasing returns to scale ($\theta<1$).
-When both conditions are met, markups reflect the joint influence of importer and exporter market power.
-2.4 Equilibrium: Tariff Pass-Through Elasticities
-Our pricing framework provides a basis for analyzing the short-run impact of an unanticipated tariff imposed on imports from country $c$.
+We use buyer power or importer market power (equivalently, countervailing power) to refer broadly to the importer’s ability to influence prices through bargaining (i.e., any $\phi>0$).
+In contrast, we reserve oligopsony power for the case in which a higher buyer share $x_{ij}$ allows the importer to negotiate lower prices, other things equal.
+Oligopsony power therefore requires both $\phi>0$ and $\theta<1$, as it relies on the presence of quasi-rents in the exporter’s technology.
+2.4 Pass-Through Characterization
+Our pricing framework allows for an analytical characterization of the short-run effects of unanticipated tariffs.
 Let $T_{c}$ denote the gross tariff rate.
-The (log) price that exporter $i$ from country $c$ charges importer $j$ is given by:
-$\ln p_{ij}=\ln\mu_{ij}+\ln c_{i}+\ln T_{c}.$
-While a tariff on country $c$ may, in principle, affect all trade relationships involving exporters from $c$, we focus on the short-run, direct effect at the relationship level.
-We interpret the tariff shock as a small, unanticipated perturbation around the initial equilibrium.
-Accordingly, we treat $T_{c}$ as a pair-specific shock and hold constant prices in all other relevant matches such that $dp_{rj}=0$ for all $r\in\mathcal{Z}_{j}\setminus{i}$ and $dp_{i\ell}=0$ for all $\ell\in\mathcal{Z}_{i}\setminus{j}$.
+We assume that the tariff enters multiplicatively into the price charged by exporter $i$ from country $c$ to importer $j$:
+$\ln p_{ij}=\ln\mu_{ij}+\ln MC_{i}+\ln T_{c(i)}.$
+While a tariff on country $c$ may affect all relationships involving its exporters, we treat it as a pair-specific shock by holding all other match prices fixed: $dp_{rj}=0$ for all $r\in\mathcal{Z}_{j}\setminus{i}$ and $dp_{i\ell}=0$ for all $\ell\in\mathcal{Z}_{i}\setminus{j}$.
 The following proposition characterizes this direct component of tariff pass-through.
@@
 The tariff pass-through elasticity into the bilateral import price $p_{ij}$, holding fixed all other prices in the network and general equilibrium variables, is:
-$\Phi_{ij}\equiv\frac{d\ln p_{ij}}{d\ln T_{c}}=\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}},$ (2.11)
+$\Phi_{ij}\equiv\frac{d\ln p_{ij}}{d\ln T_{c(i)}}=\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}},$ (2.15)
 where:
-$\Gamma_{ij}\equiv-\frac{d\ln\mu_{ij}}{d\ln p_{ij}}\quad\textit{and}\quad\Lambda_{ij}\equiv-\frac{d\ln c_{i}}{d\ln p_{ij}}$
-denote the partial elasticities of the equilibrium markup $\mu_{ij}$ and the exporter’s marginal cost $c_{i}$ to changes in the bilateral price $p_{ij}$, respectively.
+$\Gamma_{ij}\equiv-\frac{d\ln\mu_{ij}}{d\ln p_{ij}}\quad\textit{and}\quad\Lambda_{ij}\equiv-\frac{d\ln MC_{i}}{d\ln p_{ij}}$
+denote the partial elasticities of the equilibrium markup $\mu_{ij}$ and the exporter’s marginal cost $MC_{i}$ to changes in the bilateral price $p_{ij}$, respectively.
 Proof: See Appendix A.3.
-Proposition 3 highlights two distinct mechanisms shaping tariff pass-through in firm-to-firm trade: a markup channel, capturing strategic pricing responses, and a cost channel, capturing how the exporter’s marginal cost adjusts in response to price changes.
-In models with CES demand, monopolistic competition, and constant marginal costs, both $\Gamma_{ij}$ and $\Lambda_{ij}$ are zero, yielding full pass-through ($\Phi_{ij}=1$).
-More generally, $\Phi_{ij}$ may lie above, below, or at one depending on the magnitudes of $\Gamma_{ij}$ and $\Lambda_{ij}$.
-2.4.1 Markup Elasticity
-The markup elasticity $\Gamma_{ij}$ measures how the equilibrium markup $\mu_{ij}$ responds to changes in the bilateral price $p_{ij}$. It is given by:
-[…]
-where $\omega_{ij}^{\Gamma}\equiv\omega_{ij}\cdot\frac{\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\in[0,1]$, $\Gamma_{ij}^{\text{oligopoly}}$ denotes the oligopoly markup elasticity, $\Gamma_{ij}^{\text{oligopsony}}$ the oligopsony markup elasticity, and $\Gamma_{ij}^{\omega}$ the elasticity of the bargaining weight with respect to the bilateral price.
-The term in square brackets captures the direct elasticity of the markup, expressed as a convex combination of the oligopoly and oligopsony components.
-The second term reflects how the bargaining weight $\omega_{ij}$ itself responds to price changes.
-We discuss each component in turn.
-Oligopoly Markup Elasticity
-The oligopoly markup elasticity dominates when $\phi\to 0$, in which case $\Gamma_{ij}\to\Gamma_{ij}^{\text{oligopoly}}$. It is given by
-$\Gamma_{ij}^{\text{oligopoly}}\equiv-\frac{d\ln\mu_{ij}^{\text{oligopoly}}}{d\ln p_{ij}}=\frac{1}{\varepsilon_{ij}-1}\cdot\frac{\rho-\varepsilon_{ij}}{\varepsilon_{ij}}\cdot(\rho-1)(1-s_{ij})\geq 0.$
-This elasticity reflects the standard logic of strategic complementarities in price-setting: when a tariff raises the bilateral price $p_{ij}$, exporter $i$ reduces its markup to limit trade diversion, resulting in incomplete pass-through (see Amiti et al., 2014; Auer and Schoenle, 2016; Garetto, 2016; Amiti et al., 2019a).
-The response is strongest at intermediate supplier shares, giving rise to a U-shaped relationship between $\Gamma_{ij}^{\text{oligopoly}}$ and $s_{ij}$.
-[^12] Amiti et al.
-(2014) shows that, to a first-order approximation, pass-through decreases with the exporter’s share through strategic complementarities.
-However, we cannot rely on the same approximation, given our focus on bilateral markets, where both very low and very high market shares are observed in the data.
-Oligopsony Markdown Elasticity
-The oligopsony markdown elasticity dominates when $\phi\to 1$, in which case $\Gamma_{ij}\to\Gamma_{ij}^{\text{oligopsony}}$. It is given by
-$\Gamma_{ij}^{\text{oligopsony}}\equiv-\frac{d\ln\mu_{ij}^{\text{oligopsony}}}{d\ln p_{ij}}=\left(\frac{x_{ij}\cdot(1-x_{ij})^{\frac{1}{\theta}-1}}{\theta\cdot(1-(1-x_{ij})^{\frac{1}{\theta}})}-1\right)(1-x_{ij})\,\varepsilon_{ij}\leq 0.$
-This elasticity captures the logic of strategic substitutabilities among importers.
-[^13] Strategic substitutabilities arise because a decline in demand by other buyers lowers marginal costs and prices, encouraging importer $j$ to expand purchases.
-As the price $p_{ij}$ rises, importer $j$ reduces demand, shrinking the exporter’s quasi-rents.
-This weakens the importer’s bargaining position, reducing the markdown and amplifying the price response, potentially resulting in more-than-complete pass-through.
-The elasticity $\Gamma_{ij}^{\text{oligopsony}}$ depends on both the importer’s buyer share $x_{ij}$ and the exporter’s supplier share $s_{ij}$.
-It is U-shaped in $x_{ij}$: the elasticity vanishes when $x_{ij}\to 0$ (atomistic buyer) or $x_{ij}\to 1$ (monopsonist), and reaches its peak at intermediate values.
-It also declines with $s_{ij}$, since a higher supplier share reduces the demand elasticity $\varepsilon_{ij}$, limiting changes in $x_{ij}$ and weakening the markdown response.
-The Role of the Endogenous Bargaining Weight
-The elasticity $\Gamma_{ij}^{\omega}\equiv\frac{d\ln\omega_{ij}}{d\ln p_{ij}}$ captures how the bargaining weight $\omega_{ij}$ responds to price changes.
-Its sign and magnitude depend on the exporter’s supplier share $s_{ij}$ and are derived in Appendix A.3.
-For empirically relevant values of $s_{ij}$ and parameter ranges, this elasticity is typically small.
-In particular, as $s_{ij}$ approaches zero or one, where $\omega_{ij}$ converges to the importer’s exogenous bargaining strength $\phi$, the elasticity $\Gamma_{ij}^{\omega}$ goes to zero.
-2.4.2 Cost Elasticity
-The cost elasticity $\Lambda_{ij}$ measures how exporter $i$’s marginal cost $c_{i}$ responds to changes in the bilateral price $p_{ij}$, via adjustments in traded quantity.
-While prior work emphasizes the role of the cost channel in shaping pass-through (e.g., Burstein and Gopinath, 2015; Amiti et al., 2019a), our contribution is to show that this elasticity varies systematically across matches through both demand and supply channels.
-Formally:
-$\Lambda_{ij}\equiv-\frac{d\ln c_{i}}{d\ln p_{ij}}$ $=\frac{d\ln c_{i}}{d\ln q_{ij}}\cdot\left(-\frac{d\ln q_{ij}}{d\ln p_{ij}}\right)$
-$=\frac{1-\theta}{\theta}\cdot x_{ij}\cdot\varepsilon_{ij}\geq 0.$
-The elasticity declines with the exporter’s supplier share $s_{ij}$, which reduces the residual demand elasticity $\varepsilon_{ij}$, and increases with the importer’s buyer share $x_{ij}$, which amplifies the residual supply elasticity $\frac{d\ln c_{i}}{d\ln q_{ij}}$.
-Importantly, this mechanism operates independently of bargaining power ($\phi$), and instead reflects the concentrated nature of the trade network.
-2.4.3 Pass-Through Elasticity and Bilateral Market Shares
+Proposition 3 highlights two distinct mechanisms shaping tariff pass-through in firm-to-firm trade: a markup channel, capturing strategic markup responses, and a cost channel, capturing how the exporter’s marginal cost adjusts in response to price changes.
+[^14] Appendix B.4 extends equation (2.15) to incorporate indirect effects, for example shocks to exporter i that affect other prices and quantities and in turn $p_{ij}$.
+While we do not model these dynamics explicitly, the extension illustrates how spillovers across relationships can drive a wedge between reduced form and structural estimates.
+We characterize each channel in turn.
+The Markup Channel ($\Gamma_{ij}$).
+The elasticity $\Gamma_{ij}$ captures strategic markup responses to price changes. It can be written as
+$\Gamma_{ij}=\left[(1-\omega_{ij}^{\Gamma})\Gamma_{ij}^{\text{oligopoly}}+\omega_{ij}^{\Gamma}\Gamma_{ij}^{\text{oligopsony}}\right]+\left(1-\frac{\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\right)\Gamma_{ij}^{\omega},$
+where $\omega_{ij}^{\Gamma}\equiv\omega_{ij}\frac{\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\in[0,1]$, $\Gamma_{ij}^{\text{oligopoly}}$ and $\Gamma_{ij}^{\text{oligopsony}}$ denote the oligopoly and oligopsony markup elasticities, and $\Gamma_{ij}^{\omega}$ is the elasticity of the bargaining weight with respect to the bilateral price.
+The term in square brackets captures the direct markup adjustment, holding bargaining weights ($\omega_{ij})$ fixed. The second term reflects the endogenous response of these weights.
+When exporter bargaining power dominates ($\phi\to 0$), $\Gamma_{ij}$ converges to $\Gamma_{ij}^{\text{oligopoly}}$, given by
+$\Gamma_{ij}^{\text{oligopoly}}=-\frac{d\ln\mu_{ij}^{\text{oligopoly}}}{d\ln p_{ij}}=\frac{1}{\varepsilon_{ij}-1}\cdot\frac{\rho-\varepsilon_{ij}}{\varepsilon_{ij}}\cdot(\rho-1)(1-s_{ij})\geq 0.$
+In this regime, the markup channel operates through strategic complementarities, as is standard in oligopoly models.
+[^15] See, e.g., amiti2014importers; auer2016market; garetto2016firms; amiti2019international.
+An increase in the bilateral price induces the exporter to compress its markup in order to limit trade diversion, generating incomplete pass-through.
+The response is strongest at intermediate supplier shares, implying a U-shaped relationship between $\Gamma_{ij}^{\text{oligopoly}}$ and $s_{ij}$.161616amiti2014importers show that, in their model, a first-order approximation around small supplier shares implies that pass-through declines with exporter share.
+We cannot rely on this approximation in our bilateral pricing framework, where both low and high shares are observed.
+When importer bargaining power dominates ($\phi\to 1$), $\Gamma_{ij}$ converges to $\Gamma_{ij}^{\text{oligopsony}}$, given by
+$\Gamma_{ij}^{\text{oligopsony}}=-\frac{d\ln\mu_{ij}^{\text{oligopsony}}}{d\ln p_{ij}}=\left(\frac{x_{ij}(1-x_{ij})^{\frac{1}{\theta}-1}}{\theta\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]}-1\right)(1-x_{ij})\varepsilon_{ij}\leq 0.$
+In this case, the elasticity $\Gamma_{ij}$ is governed by markdown adjustments and reflect strategic substitutabilities among importers.
+As $p_{ij}$ rises, importer $j$ contracts demand, reducing quasi-rents and weakening its bargaining position.
+The resulting markdown compression amplifies the price response and can generate more-than-complete pass-through.
+The elasticity $\Gamma_{ij}^{\text{oligopsony}}$ depends on both the buyer share $x_{ij}$ and the supplier share $s_{ij}$.
+It is U-shaped in $x_{ij}$, vanishing as $x_{ij}\to 0$ (atomistic buyer) or $x_{ij}\to 1$ (monopsonist), and peaking at intermediate values.
+It declines in $s_{ij}$ because a higher supplier share reduces the demand elasticity $\varepsilon_{ij}$, limiting variation in $x_{ij}$ and weakening the markdown response.
+With two-sided market power $(\phi\in(0,1))$, the elasticity $\Gamma_{ij}$ reflects a convex combination of oligopoly and oligopsony markup elasticities, with weights $\omega_{ij}^{\Gamma}$ proportional to importer bargaining power.
+It also reflects the endogenous response of bargaining weights through $\Gamma_{ij}^{\omega}\equiv\frac{d\ln\omega_{ij}}{d\ln p_{ij}}$, whose sign and magnitude depend on $s_{ij}$.
+[^17] The exact expression for this term is provided in Appendix A.3.
+As $s_{ij}\to 0$ or $s_{ij}\to 1$, $\omega_{ij}$ converges to the exogenous parameter $\phi$, implying $\Gamma_{ij}^{\omega}\to 0$.
+The Cost Channel ($\Lambda_{ij}$).
+The cost elasticity $\Lambda_{ij}$ captures how marginal costs respond to price changes. An increase in prices reduces demand and, under decreasing returns, lowers marginal costs. Formally,
+$\Lambda_{ij}=\frac{1-\theta}{\theta}\,x_{ij}\,\varepsilon_{ij}\geq 0.$ (2.16)
+The first term, $(1-\theta)/\theta$, captures the slope of marginal cost or, equivalently, the inverse export supply elasticity at the exporter level.
+The buyer share $x_{ij}$ captures the exposure of the exporter’s total output to the relationship: importers that account for a greater fraction of the exporter’s output exert stronger scale effects on marginal cost.
+The residual demand elasticity $\varepsilon_{ij}$ captures how strongly bilateral quantity responds to a change in price.
+The product $(1-\theta)/\theta\cdot x_{ij}$ is the residual inverse export supply elasticity faced by buyer $j$: it scales with the buyer share, reflecting each buyer’s exposure to the exporter’s overall supply curvature.
+Together, $x_{ij}$ and $\varepsilon_{ij}$ determine the strength of the match-specific scale effect.
+This mechanism may appear similar to a standard cost channel driven by firm-level decreasing returns.
+While upward-sloping export supply can generate incomplete pass-through, a model with atomistic buyers attributes this entirely to supply curvature.
+As shown in Appendix B.5, this implies uniform pricing and identical pass-through across buyers sourcing from the same exporter.
+In contrast, our framework features non-atomistic buyers and produces relationship-specific residual cost elasticities that scale with buyer shares, generating cross-sectional heterogeneity in pass-through.
+Pass-Through Elasticity and Bilateral Market Shares
+Figure 2 plots the pass-through elasticity $\Phi_{ij}$ as a function of $s_{ij}$ and $x_{ij}$ for three values of $\theta$ (rows) and $\phi$ (columns).
+[^18] We set $\theta=0.9$ as an intermediate value of the returns-to-scale parameter, consistent with estimates in the production function literature, typically between 0.9 and 1; see, e.g., basu1997returns.
+The figure shows that the same primitives governing oligopsony power also shape pass-through.
+Variation in $\theta$ (vertical) captures how exporters’ returns to scale, via the inverse export supply elasticity, determine the balance between cost and markup channels, while variation in $\phi$ (horizontal) captures how bargaining power governs strategic markup adjustment.
 Figure 2: Pass-Through Elasticity and Bilateral Market Shares
-Notes: The figure presents heatmaps of the pass-through elasticity $\Phi_{ij}$ across combinations of $s_{ij}$ (x-axis) and $x_{ij}$ (y-axis), under alternative assumptions about $\phi$ and $\theta$.
-We set $\theta=0.5$ in the decreasing returns case and $\phi=0.5$ in the intermediate bargaining case.
+Notes: The figure presents heatmaps of the pass-through elasticity $\Phi_{ij}$ as a function of market shares $s_{ij}$ (x-axis) and $x_{ij}$ (y-axis), for $\phi\in\{0,0.5,1\}$ (columns) and $\theta\in\{0.5,0.9,1\}$ (rows).
 Other parameters are fixed at $\gamma=0.5$, $\varrho=1$, $\nu=4$, and $\rho=10$.
-The interaction between markup and cost channels generates rich heterogeneity in pass-through elasticities across matches, shaped by bargaining power ($\phi$), returns to scale ($\theta$), and bilateral market shares ($s_{ij}$ and $x_{ij}$).
-Figure 2 illustrates these patterns using heatmaps of the pass-through elasticity $\Phi_{ij}$ as a function of supplier share ($s_{ij}$) and buyer share ($x_{ij}$), under three bargaining regimes—$\phi\to 0$, $\phi\in(0,1)$, and $\phi\to 1$—with decreasing returns to scale ($\theta<1$) in the top row and constant returns to scale ($\theta=1$) in the bottom row.
-Several key insights emerge.
-First, pass-through $\Phi_{ij}$ increases with importer bargaining power $\phi$.
-As $\phi$ rises, the markup elasticity shifts toward the oligopsony markdown elasticity, increasing pass-through through strategic substitutabilities.
-This is most visible in the shift from the left to the right columns within each row.
-Second, when $\theta<1$, pass-through is incomplete across most of the $(s_{ij},x_{ij})$ space.
-This is because the cost elasticity $\Lambda_{ij}$, which is always non-negative, dominates the overall pass-through elasticity $(\Gamma_{ij}+\Lambda_{ij})$ when $\Gamma_{ij}$ is negative due to strong buyer power.
-Moreover, pass-through values under $\phi\in(0,1)$ and $\phi=1$ are nearly indistinguishable, indicating that pass-through is relatively insensitive to the exact level of $\phi$ in this regime.
-This reflects two forces: (i) the cost channel is strong when returns to scale are decreasing and importers’ bargaining power is high, and (ii) the markup elasticity $\Gamma_{ij}$ is either muted by offsetting effects of strategic complementarities and substitutabilities or generally low in values.
-[^14] For more details, see Appendix A.4.
-Third, pass-through $\Phi_{ij}$ depends on $x_{ij}$ only when $\theta<1$, in which case it declines with $x_{ij}$.
-This yields a sharp empirical prediction: a negative relationship between pass-through and buyer share $x_{ij}$ implies decreasing returns to scale, regardless of the level of bargaining power.
-This result is formalized below.
+Moving from top to bottom, $\theta$ rises from 0.5 to 1.
+At $\theta=0.5$, export supply is steep, so marginal cost is highly sensitive to output and the cost channel dominates across the $(s_{ij},x_{ij})$ space, rendering pass-through largely invariant to $\phi$ and the markup channel.
+As $\theta$ increases, the cost channel weakens and markup adjustment becomes more important.
+Markdown responses also attenuate, but the cost channel continues to dominate in $x_{ij}$ (Appendix A.4), so pass-through becomes less incomplete.
+At $\theta=1$, supply is flat, the cost channel vanishes, and pass-through is governed entirely by markup adjustment.
+Moving left to right, higher $\phi$ traces the role of bargaining power in shaping strategic markup adjustment.
+At $\phi=0$, the markup channel reflects strategic complementarities that dampen pass-through; at $\phi=1$, it reflects strategic substitutabilities associated with oligopsony power that amplify pass-through.
+This gradient is most pronounced in the bottom row ($\theta=1$), where only the markup channel operates, and attenuates as $\theta$ falls.
+At $\theta=0.5$, markup adjustment is second order.
+A key implication of Figure 2 is that pass-through is incomplete over most of the parameter space, with the sole exception of the constant-returns, full-importer-power case ($\theta=1$, $\phi=1$).
+Different $(\theta,\phi)$ pairs generate incomplete pass-through through distinct mechanisms and to varying degrees.
+In the lower-left region of the heatmaps–corresponding to the low-share range where most bilateral matches lie in the data (Table 1)–average pass-through is similar across specifications, making models difficult to distinguish on that margin alone.
+By contrast, the sign and magnitude of the pass-through gradient in $x_{ij}$ differ sharply across variants.
+The following proposition formalizes this contrast, which we exploit in Section 4.3 to discriminate among models using cross-sectional pass-through moments.
 Proposition 4.
-If pass-through $\Phi_{ij}$ decreases with the importer’s buyer share $x_{ij}$, then $\theta<1$.
-Proof: See Appendix A.4.
-By contrast, the relationship between $\Phi_{ij}$ and the supplier share $s_{ij}$ is less clear-cut.
-As shown in Figure 2, pass-through declines with $s_{ij}$ primarily when $\phi$ is low and $\theta$ is near one, a setting in which the markup channel dominates and cost-based adjustments are limited.
-In this region, strategic complementarities give rise to a U-shaped relationship between $\Phi_{ij}$ and $s_{ij}$.
-As either $\phi$ or $x_{ij}$ increases, the cost channel becomes more prominent, and pass-through tends to rise with $s_{ij}$.
-While we do not formally characterize the conditions under which this reversal occurs, we show below that $\phi$ is weakly increasing in $s_{ij}$.
-This monotonicity supports interpreting the empirical evidence as inconsistent with the low-$\phi$, high-$\theta$ case.
+When $\theta<1$, the pass-through elasticity $\Phi_{ij}$ decreases with the importer’s buyer share $x_{ij}$.
+Proof: See Appendix A.4. Intuitively, when $\theta<1$, the cost channel dominates the markdown channel. Equation (2.16) implies that pass-through falls with $x_{ij}$.
+The relationship between $\Phi_{ij}$ and $s_{ij}$ is, by contrast, ambiguous.
+Under strategic complementarities, pass-through is U-shaped in $s_{ij}$.
+Under strategic substitutabilities and through the cost channel, it is increasing in $s_{ij}$.
+The former pattern arises when bargaining power is low and returns to scale are close to constant, while the latter emerges as either force strengthens.
+Which effect dominates is ultimately an empirical matter.
 2.5 Discussion
-This section reviews key modeling assumptions and discusses potential extensions.
-2.5.1 Bargaining Protocol and Quantities
-The baseline model assumes demand-determined quantities, whereby the importer chooses input quantities to minimize total cost, taking the price as given.
-This assumption provides analytical tractability, yields closed-form solutions for markups and pass-through, and nests standard models of international trade.
-An important implication is that prices are allocative, a feature that aligns with our empirical findings and with firm-level evidence in related settings (Gopinath and Itskhoki, 2011).
-We consider two alternative bargaining protocols as benchmarks for understanding the role of quantity determination.
-The first is efficient bargaining, discussed more formally in Appendix B.1.
-This case corresponds to the vertically integrated benchmark, where importer and exporter jointly negotiate over both price and quantity to maximize total surplus.
-While theoretically appealing, this setup implies that prices are non-allocative transfers, inconsistent with the empirical evidence.
-Moreover, the vertically integrated case may be unrealistic in the context of arm’s-length firm-to-firm trade, where limited commitment and contracting frictions are prevalent (Antràs, 2020).
-The second alternative, detailed in Appendix B.2, is supply-driven bargaining: the exporter sets quantity for a given price, and price is then negotiated based on the resulting supply curve.
-As $\phi\to 1$, this nests the classic monopsony benchmark commonly used in labor markets (e.g., Berger et al., 2022).
-The key distinction between this setup and our baseline lies in their welfare implications: in the supply-driven case, buyer power lowers both prices and quantities, while supplier power helps restore efficiency.
-In contrast, under demand-driven bargaining, buyer power mitigates upstream distortions and improves efficiency.
-[^15] See Avignon et al.
-(2024) and Demirer and Rubens (2025) for recent discussions of how quantity-setting assumptions affect welfare outcomes.
-While both alternatives offer useful benchmarks, they are less suited to the goals of this paper.
-The supply-driven model, in particular, does not provide closed-form expressions for key variables, which makes it challenging to generate clear predictions or link the model to data compared to our baseline setup.
-Still, it remains a valuable direction for future work, especially for studying welfare effects when buyer power leads to inefficiencies.
-2.5.2 Outside Options
-In our model, each firm’s outside option reflects payoffs from trade with all other existing partners, excluding the focal match.
-This assumption allows us to express markups and pass-through in terms of market shares and a small set of parameters, facilitating structural estimation and counterfactual analysis.
-While analytically convenient, the assumption may appear restrictive if disagreement leads a firm to form or sever other relationships.
-Appendix B.3 explores a more flexible setup in which disagreement affects the importer’s cost and the exporter’s revenue non-parametrically.
-Although more realistic, this extension introduces an identification problem: the parameters $\phi$ and $\theta$ can no longer be separately identified from the outside option, which limits their interpretability and empirical tractability.
-2.5.3 General Equilibrium Forces
-We conclude the theoretical section by noting that all results are derived under a partial equilibrium approach, holding fixed general equilibrium variables such as wages, demand shifters, and competitor behavior.
-This approach simplifies the analysis but also raises questions about the role of general equilibrium adjustments in the results.
-General equilibrium considerations are critical for interpreting pass-through elasticities.
-The structural pass-through elasticity derived in Section 2.4 is a direct or partial elasticity, holding constant aggregate variables such as wages, demand conditions, and, importantly, competitors’ prices and sourcing decisions.
-In reality, tariff shocks may also affect variables such as foreign wages or domestic export prices.
-If these effects are not fully observed or controlled for, reduced-form estimates may conflate the direct impact with general equilibrium responses, making it harder to align empirical pass-through coefficients with model-based elasticities (Burstein and Gopinath, 2015).
-[^16] In Appendix B.4, we extend equation (2.11) to account for indirect effects, such as how a shock to exporter $i$ influences other prices and quantities, which may in turn affect $p_{ij}$.
-While we do not model full general equilibrium dynamics, this extension illustrates how spillovers across relationships may lead to reduced-form estimates diverging from structural ones.
-Therefore, our goal is not to recover structural elasticities from reduced-form regressions, which would be inappropriate in this context (Berger et al., 2022).
-Instead, we pursue a structural approach, which helps mitigate similar concerns.
-Our approach has three advantages.
-First, the theoretical relationship between bilateral shares, markups, and pass-through holds parametrically regardless of the specific general equilibrium environment.
-While the parameter values depend on how aggregate variables such as tariffs, demand shifters, and wages co-move in equilibrium, we do not need to model their relationships explicitly.
-Instead, identification is based on the cross-sectional variation of prices across importers within exporter–product–year cells, without making assumptions on the general equilibrium environment.
-Second, we can also isolate and test the cross-sectional predictions of the model independently of general equilibrium forces.
-In particular, even if aggregate variables shift over time, the model predicts specific patterns between firm-to-firm shares, markups, and pass-through at a given point in time.
-These predictions can be assessed empirically using within-period variation across matches, which we exploit in the next sections.
-Third, we can evaluate the performance of the estimated model by comparing its ability to replicate observed price changes in response to tariff shocks, thereby gauging the significance of general equilibrium forces.
-As shown in the empirical analysis, the estimated model fits the observed price changes well.
-This suggests that the short-run effects of tariffs on prices can be understood primarily through the lens of partial equilibrium mechanisms, and supports the usefulness of our framework for studying firm-to-firm pricing and pass-through.
+We briefly review our key modeling choices before turning to the empirical analysis.
+Bargaining Protocol and Quantities
+Our baseline model assumes demand-driven quantities, whereby the importer chooses quantities given negotiated prices.
+This ensures allocative pricing and nests the standard monopolistic-competition benchmark as a limiting case, consistent with firm-level evidence in international trade (e.g., gopinath2011search).
+We derive two alternative protocols in the Appendix.
+The first is efficient bargaining (Appendix B.1), which maximizes joint surplus but implies non-allocative prices, a prediction at odds with the data.
+The second is supply-driven bargaining (Appendix B.2), a monopsony-style benchmark, which implies that prices and quantities move together along the exporter’s supply curve.
+This implication is inconsistent with our empirical evidence that tariff-induced price increases are accompanied by quantity declines.
+One caveat of our framework is that when $\phi$ is sufficiently high, equilibrium prices can fall below the efficient level ($p<p^{\text{eff}}$), implying quantities above the competitive benchmark ($q>q^{\text{eff}}$).
+Recent theories of bilateral oligopoly rule out this case by imposing a supplier participation constraint that prevents quantities from exceeding the exporter’s supply (marginal cost) curve when $p<p^{\text{eff}}$ (e.g., demirer2025welfare; avignon2024markups).
+We abstract from this extension because, at high $\phi$, it places equilibrium on the upward-sloping supply curve, implying positive price-quantity comovement, contrary to our evidence.
+General Equilibrium and Network Spillovers
+Our theoretical results are derived in partial equilibrium, holding wages, aggregate demand shifters, and competitors’ prices and quantities fixed.
+The pass-through elasticity derived in Section 2.4 therefore isolates the direct, intensive-margin response of markups and marginal costs to a tariff shock.
+Embedding this framework in a general equilibrium environment introduces two main complexities.
+First, under decreasing returns, quantity adjustments along one link affect the exporter’s marginal cost across all links.
+Second, aggregate shocks or network spillovers may differentially affect outside options across matches of the same exporter.
+Appendix B.3 discusses how generalizing outside options modifies the bargaining problem.
+Fully accommodating these channels, however, requires committing to a specific macroeconomic closure for wages, final demand, and network formation, which is beyond the scope of this paper.
+As we show below, the partial equilibrium approach is supported by the close fit between model predictions and untargeted pass-through moments.
 3 Data and Stylized Facts
@@
 Section 3.3 details the sample selection and provides summary statistics.
-Finally, Section 3.4 presents evidence testing the model’s predictions.
+Finally, Section 3.4 presents reduced-form evidence on prices and pass-through through the lens of the model.
 3.1 Data Sources
@@
 Each observation corresponds to a shipment from a foreign exporter to a U.S.
-importer and includes the transaction date, product classification at the 10-digit Harmonized System (HS10) level, FOB import value in U.S.
+importer and includes the transaction date, product classification at the 10-digit Harmonized System (HS10) level, Free on Board (FOB) import value in U.S.
 dollars, physical quantity, transportation mode, and country of origin.
 Exporters are identified using a manufacturer ID (MID) constructed by the Census Bureau from the exporter’s name, street address, city, and country.
-[^17] The MID combines the country code, (elements from) the firm name, city, and address (Kamal and Monarch, 2018).
-Because the algorithm is not standardized, it may generate inconsistent identifiers due to misspellings or minor location changes, leading to one firm having multiple MIDs or several firms sharing one.
-Following Kamal and Monarch (2018), we construct a robustness version that truncates location fields to improve consistency.
-Our baseline uses the full MID, and results are robust to this alternative.
+[^19] The MID combines country code, elements of the firm name, city, and address (Kamal2018).
+Because the algorithm is not standardized, identifiers may be inconsistent: one firm may have multiple MIDs or multiple firms may share one.
+Following Kamal2018, we construct a robustness measure that truncates location fields to improve consistency.
+Results are robust.
 To focus on arm’s-length trade, we exclude related-party transactions from the baseline sample.
@@
 Customs forms, flagging relationships with ownership stakes of at least five percent.
-While widely used, this measure may misclassify firms due to its reliance on self-reporting and a low reporting threshold (Ruhl, 2015).
+While widely used, this measure may misclassify firms due to its reliance on self-reporting and a low reporting threshold (kimintrafirmAERPP).
 To improve accuracy, we construct an alternative indicator using ORBIS, which provides firm-level cross-border ownership links.
@@
 import tariffs introduced during the 2018 trade war.
-We use the dataset from Fajgelbaum et al.
-(2020), which records the timing, product coverage, and country-specific scope of these measures at the HS8–month level.
+We use the dataset from fajgelbaum2020return, which records the timing, product coverage, and country-specific scope of these measures at the HS8-month level.
 The tariffs averaged 25 percentage points and were imposed on top of existing rates, targeting selected goods.
@@
 This formulation implies that the elasticity of the importer’s marginal cost with respect to the price of foreign input $h$ is $\frac{d\ln c_{j}}{d\ln p_{jh}^{f}}=\alpha_{jh}\gamma\in(0,1]$.
+We also note that allowing this elasticity to vary across importer-product pairs in turn makes $\eta$ in equation (2.13) to vary across the same dimension, $\eta_{jh}$.
 We construct the exporter’s supplier share as $s_{ijh}\equiv\frac{p_{ijh}q_{ijh}}{\sum_{k\in\mathcal{Z}_{j}^{h}}p_{kjh}q_{kjh}}$, where $\mathcal{Z}_{j}^{h}$ denotes the set of firm $j$’s foreign suppliers of input $h$.
@@
 and thus cannot account for the full set of an exporter’s buyers.
-[^18] To address the possibility that the importer’s buyer share $x_{ij}$ may be overstated due to unobserved sales to other destinations, we replicate the analysis using only exporters from Canada and Mexico.
-These countries direct the majority of their exports to the U.S.
-(71% for Canada and 73% for Mexico in 2019), making the assumption of destination-specific production less restrictive.
-Reassuringly, our estimates remain stable in this subsample.
-Full results are available upon request.
+[^20] To address potential overstatement of buyer share $x_{ij}$ due to unobserved exports to other destinations, we replicate the analysis for Canada and Mexico, which send most exports to the U.S.
+(71% and 73% in 2019).
+Estimates are similar in this subsample.
+Full results available upon request.
 3.3 Sample Construction and Summary Statistics
-We apply a series of restrictions to the LFTTD to align the empirical sample with the model’s focus on decentralized bargaining over intermediate inputs.
-Full details are provided in Appendix C.2.
-The selection criteria are designed to ensure that we observe relationship-level price changes, exclude related-party transactions, and maintain sufficient variation in buyer–supplier matches for identification purposes.
-Appendix Table C.1 summarizes how the sample evolves with each restriction across key dimensions.
-We begin by restricting the sample to importer–exporter–product triples observed in two consecutive years, allowing us to compute relationship-level price changes, which are central to our analysis.
-While this requirement eliminates over half of the raw matches, it retains a substantial share of trade, covering roughly 88% of import value between 2001–2016 and 80% between 2017–2018, highlighting the importance of repeated relationships.
-Next, we restrict the sample to capital and intermediate inputs by excluding products classified as consumption goods under the Broad Economic Categories (BEC) system.
-This step reduces the number of suppliers and relationships, as shown in Table C.1.191919As a robustness check, we also consider a broader sample that includes consumption goods.
-Appendix Table C.2 reports the composition of this sample and Table C.3 reports the corresponding summary statistics.
-These statistics are broadly similar to the baseline, indicating that including consumption goods does not significantly alter the data composition.
-To further mitigate measurement error, we apply three filters: (i) exclude transactions involving energy goods, (ii) drop observations with unit values outside the 1st–99th percentile within product, and (iii) remove transactions with absolute log price changes above four.
-We then exclude related-party transactions, which are less likely to reflect decentralized bargaining and may involve internal pricing strategies such as transfer pricing.
-[^20] Bernard et al.
-(2006) document that prices in related-party trade differ systematically from arm’s-length transactions, with lower average prices and distinct pass-through behavior.
-In our baseline definition, a buyer–supplier pair is considered related if ORBIS identifies a shared corporate parent.
-[^21] We retain all observations not flagged as related in either ORBIS or the LFTTD.
-This approach preserves sample size while ensuring that limited ORBIS coverage does not unduly constrain the selection.
-See Appendix C.1 for details.
-For robustness, we consider two alternative definitions: one based solely on the LFTTD’s related-party flag, and another that combines this flag with ORBIS data identifying the U.S.
-importer as a multinational (domestic or foreign-owned).
-Lastly, we impose restrictions based on our identification strategy.
-Since the latter relies on observing the variation in prices across U.S.
-buyers for the same supplier-product-year combination, we restrict our sample to supplier–product pairs in which the supplier transacts with at least two U.S.
+We apply a sequence of restrictions to the LFTTD to construct the estimation sample.
+Some restrictions are required for identification and counterfactuals; others align the baseline sample with the model’s environment.
+Appendix C.2 provides details, and Appendix Table C.1 traces the sample under each restriction.
+Essential restrictions.
+Three restrictions are essential.
+First, we require importer-exporter-product triplets to be observed in two consecutive years, a necessary restriction for bilateral price to be computed.
+Second, because identification of the bargaining and returns-to-scale parameters relies on cross-sectional price dispersion across importers sourcing the same exporter-product-year, we restrict the sample to exporter-product pairs that transact with at least two U.S.
 buyers in consecutive years.
-Table C.1 reports that after these restrictions, the sample accounts for approximately \$160 billion in import value and 250 thousand buyer-supplier-product-year combinations between 2017-2018.
+Third, to limit measurement error and ensure that unit values more closely reflect prices rather than compositional changes within HS10 products, we follow standard practice and (i) exclude energy goods, (ii) drop unit values outside the 1st-99th percentile within product, and (iii) remove observations with absolute log price changes exceeding four.
+Other restrictions.
+We impose two additional restrictions to align the data with the model.
+First, we exclude related-party transactions, as prices in such relationships may reflect forces, such as transfer pricing, outside our framework (bernard2006transfer).
+In the baseline, pairs are classified as related if ORBIS identifies a common corporate parent; alternative definitions based on the LFTTD related-party flag or multinational status yield similar results.
+Second, we focus on capital and intermediate inputs by excluding consumption goods under the Broad Economic Categories (BEC) classification.
+This reflects the model’s focus on supply chains, where lock-in from relationship-specific investments limits short-run scalability and is more pronounced upstream (e.g., Nunn2007).
+In robustness exercises, however, we show that the main results extend to the full set of imported goods.
+After all restrictions, the baseline sample covers approximately \$160 billion in import value and 250,000 buyer-supplier-product-year observations in 2017-2018.
+This corresponds to about 20% of total U.S.
+imports by value and roughly 15% of buyer-supplier-product triplets.
+Although not representative of aggregate U.S.
+trade, the sample isolates repeated, arm’s-length production relationships where relationship-level price changes can be credibly measured and the model’s mechanisms credibly identified.
 Summary Statistics
 Table 1 reports summary statistics for our final sample.
-Panel A shows that the concentration of importers and exporters is substantial.
+[^21] Appendix Table C.2 provides additional statistics.
+Appendix Table C.3 and Table C.4 report sample composition and summary statistics for the pooled sample including all BEC categories.
+Panel A documents substantial concentration on both sides of the market.
 On average, an exporter supplies 32% of an importer’s total imports of a given HS10 product, with a median share of 15%.
-The average buyer share is lower, at 25%, with a median of 10%.
-The two shares are highly dispersed and largely uncorrelated, with a correlation of 0.04.
+The average buyer share is lower, at 25%, with a median of 10%, likely reflecting the exclusion of cases with buyer share equal to one.
+The two shares are highly dispersed and essentially uncorrelated (correlation 0.04).
 Table 1: Summary Statistics for Main Estimation Sample (2001–2018)
-Variable Mean Std. Dev. P25 Median P75
-Panel A: Characteristics of Trade Relationships
-$s_{ijh}$: Supplier share 0.32 0.35 0.03 0.15 0.57
-$x_{ijh}$: Buyer share 0.25 0.29 0.02 0.10 0.40
- Relationship length (product $h$) 4.00 2.80 2.50 3.50 5.50
- Relationship length (all products) 4.80 3.30 2.50 4.50 6.50
- # Transactions (product $h$) 120 1100 6.50 16 50
- # Transactions (all products) 360 3000 11 36 140
- # Products per pair 3.80 7.30 1.50 2.50 4.50
- Multi-HS10 dummy 0.59 0.49 0.00 1.00 1.00
- # Suppliers per buyer (HS10) 1.80 3.20 1.50 2.50 5.50
- Buyer tenure (all products) 9.90 5.00 6.50 10.00 14.00
- Buyer tenure (product $h$) 6.90 4.40 3.50 6.50 10.00
- # Buyers per supplier (HS10) 3.20 3.90 2.50 3.50 7.50
- Supplier tenure (all products) 8.00 4.60 4.50 8.50 12.00
- Supplier tenure (product $h$) 6.40 4.00 3.50 6.50 9.50
- Corr. between $s_{ijh}$ and $x_{ijh}$ 0.041
-Panel B: Prices
-$\log p$ (pre-duty) 3.50 2.80 1.40 3.10 5.40
-$\log p$ (pre-duty, excl. charges) 3.40 2.80 1.30 3.00 5.40
-$\log p^{\text{duty}}$ (post-duty) 3.50 2.80 1.40 3.10 5.40
-Notes: This table reports summary statistics for the estimation sample used in the empirical analysis.
-The data span 2001–2016 and include importer–supplier–product matches observed in two consecutive years.
-The sample excludes consumption goods (based on BEC), energy products, statistical outliers, and related-party trade, and is restricted to suppliers trading with at least two U.S.
-buyers in two consecutive years.
-This sample corresponds to the “+ Supplier Multi-Buyer” row in Panel B of Table C.1.
-Columns report the mean, standard deviation, and selected quantiles (25th, 50th, 75th percentile) for each variable.
-Prices in Panel B are log unit values (FOB value over quantity), with variants including charges or duties.
-$s_{ijh}$ denotes exporter $i$’s share in buyer $j$’s imports of product $h$; $x_{ijh}$ denotes buyer $j$’s share in exporter $i$’s U.S.
-exports of the same product.
-Relationship length and tenure are in years; concentration is measured at the HS10–year level.
-Counts of buyers, suppliers, and origin countries are per product per firm.
-Statistics are based on confidential LFTTD data and rounded to four significant digits per U.S.
+Variable Mean Median P25–P75
+Panel A: Trade Relationships
+$s_{ijh}$: Supplier share 0.32 0.15 0.03–0.57
+$x_{ijh}$: Buyer share 0.25 0.10 0.02–0.40
+Relationship length 4.0 3.5 2.5–5.5
+Relationship length (all products) 4.8 4.5 2.5–6.5
+# Transactions 120 16 6.5–50
+# Products per pair 3.8 2.5 1.5–4.5
+Buyer tenure 6.9 6.5 3.5–10.0
+Supplier tenure 6.4 6.5 3.5–9.5
+Panel B: Prices (log unit values)
+$\log p$ (pre-duty, excl. charges) 3.40 3.00 1.30–5.40
+$\log p$ (pre-duty) 3.50 3.10 1.40–5.40
+$\log p^{\text{duty}}$ (post-duty) 3.50 3.10 1.40–5.40
+Notes: This table reports summary statistics for the main estimation sample at the importer-supplier-product level.
+Full details on sample construction are provided in Appendix C.2.
+Variables are measured at the HS10 product level unless noted otherwise.
+Panel A reports trade relationship characteristics.
+$s_{ijh}$ is exporter $i$’s share in buyer $j$’s imports of product $h$; $x_{ijh}$ is buyer $j$’s share in exporter $i$’s U.S.
+exports of product $h$.
+Relationship length is the number of years between the supplier-buyer pair’s first transaction in product $h$ and year $t$; Relationship length (all products) is defined analogously, using the pair’s first transaction in any product.
+# Transactions is the cumulative number of transactions between the pair in product $h$, from their first transaction through year $t$.
+# Products per pair is the number of distinct HS10 products traded by the pair in year $t$.
+Buyer tenure (Supplier tenure) is the number of years between the buyer’s (supplier’s) first recorded transaction in the LFTTD and year $t$, pooling across partners and products.
+Panel B reports log unit values (value divided by quantity).
+$\log p$ (pre-duty, excl.
+charges) corresponds to the FOB (Free on Board) unit value, excluding shipping and insurance charges; $\log p$ (pre-duty) includes these charges to obtain the CIF (Cost, Insurance, and Freight) unit value; and $\log p^{\text{duty}}$ (post-duty) further incorporates import duties.
+Statistics are based on confidential LFTTD data and rounded according to U.S.
 Census Bureau disclosure guidelines.
 Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
-Long-term relationships are a hallmark of intermediate input trade (Antràs and Chor, 2013; Monarch, 2022).
-In our data, importer–exporter pairs trade the same product for an average of four years and remain connected across all products for nearly five years.
-Pairs transact frequently (median: 16 times per product), often spanning multiple products (mean: 3.8), and typically involve durable links, with average tenures ranging from 7 to 10 years.
-Panel B of Table 1 reports descriptive statistics for three bilateral (log) price measures based on FOB unit values: baseline pre-duty prices, pre-duty prices excluding insurance and other ancillary charges, and post-duty prices.
-All three exhibit wide dispersion across importer–exporter–product matches, with interquartile ranges exceeding four log points.
-To assess the sources of this variation, we perform a variance decomposition exercise, as described in Appendix D.1.
-Table D.1 shows that product-year fixed effects explain about 50% of the total variance, while match-specific residuals account for 4%.
-Crucially, when we focus only on variation within supplier–product–year cells, 77% of the remaining dispersion is explained by match-specific factors.
-This pattern holds across price definitions, underscoring the importance of relationship-specific forces in pricing.
-[^22] See Fontaine et al.
-(2020) for related evidence in French data.
-3.4 Test of Model Predictions
-We now examine how the comovements between markups, pass-through, and bilateral market shares align with the model’s predictions in Propositions 2 and 4.
-While not a formal test, this analysis provides supporting evidence for the mechanisms emphasized in the theory.
-Section 4 then develops a structural approach to quantify these forces more directly.
-3.4.1 Test of Proposition 2: Markups and Bilateral Market Shares
-We begin by testing Proposition 2, which predicts that with two-sided market power ($\phi\in(0,1)$ and $\theta<1$), bilateral markups increase with the exporter’s supplier share ($s_{ijht}$) and decrease with the importer’s buyer share ($x_{ijht}$).
-Since markups are not observed, we use log prices, equal to log markups plus log marginal costs, and include supplier–product–time fixed effects to absorb cost variation and isolate the markup component.
-We estimate the following specification:
+Trading relationships are also persistent: the typical importer-exporter pair has been transacting in a given HS10 product for about four years, and in any product for nearly five.
+Activity within the relationship is intensive: at the median, a pair has accumulated 16 transactions in a product to date and trades 2.5 distinct HS10 products in a given year.
+Buyer and supplier tenure across all partners and products average about seven and six years, respectively.
+Panel B of Table 1 reports descriptive statistics for three bilateral (log) price measures, each defined as transaction value divided by quantity: an FOB (Free on Board) pre-duty price that excludes shipping and insurance charges; a CIF (Cost, Insurance, and Freight) pre-duty price that adds those charges; and a post-duty price that further incorporates import duties.
+All three are highly dispersed across importer-exporter-product-year observations, with interquartile ranges of about four log points.
+To assess the sources of this variation, we conduct a variance decomposition following Fontaine2019 (Appendix D.1).
+Product-year fixed effects account for about 50% of the total variance, while match-specific residuals account for 4%.
+Focusing on within supplier–product–year variation, 77% of the remaining dispersion is explained by match-specific factors.
+This pattern holds across price definitions and underscores the importance of relationship-specific forces in pricing.
+3.4 Reduced-Form Evidence on Prices and Pass-Through
+As a preliminary exploration of mechanisms, we examine the empirical relationships between markups, pass-through, and bilateral market shares.
+We interpret these patterns through the lens of Propositions 2 and 4 as descriptive evidence consistent with two-sided market power and decreasing returns to scale, before turning to structural estimation in Section 4.
+Markups and Bilateral Market Shares
+Proposition 2 implies that under two-sided market power ($\phi\in(0,1)$) and decreasing returns to scale ($\theta<1$), bilateral markups rise in the exporter’s supplier share ($s_{ijht}$) and fall in the importer’s buyer share ($x_{ijht}$).
+To characterize these empirical patterns, we estimate
 $\ln p_{ijht}$ $=\alpha_{s}\,s_{ijht}+\alpha_{x}\,x_{ijht}+\mathbf{X}_{ijht}\boldsymbol{\gamma}+\mathbf{FE}+\upsilon_{ijht},$ (3.1)
-where the coefficients of interest are $\alpha_{s}$ and $\alpha_{x}$, which we expect to be positive and negative, respectively.
-To address endogeneity concerns, we construct leave-one-out instruments that isolate variation in market structure plausibly exogenous to the pricing decision of a given buyer–supplier pair.
-Specifically, we use the average supplier share (excluding $i$) among other buyers of exporter $i$ (excluding $j$) to instrument for $s_{ijh}$.
-Similarly, to instrument for $x_{ijh}$, we use the average buyer share (excluding $j$) among other suppliers to importer $j$ (excluding $i$).
-Table 2 shows the results.
-Columns (1)–(2) include exporter ($\text{FE}_{i}$), importer ($\text{FE}_{j}$), and product–year ($\text{FE}_{ht}$) fixed effects.
-Column (3)–(4) replaces exporter fixed effects with exporter–product–year ($\text{FE}_{iht}$) to account for unobserved marginal costs.
-Columns (5) and (6) further replace importer fixed effects with importer–product–year ($\text{FE}_{jht}$) to capture buyer-specific demand shocks.
-All regressions control for the relationship duration, measured as years since the first shipment of product $h$ between $i$ and $j$.
+where $\alpha_{s}$ and $\alpha_{x}$ are expected to be positive and negative, respectively.
+Because markups are not observed, we use log prices as the dependent variable, and include supplier–product–time fixed effects to absorb cost variation and isolate the markup component.
 Table 2: Prices and Bilateral Concentration
@@
 OLS IV OLS IV OLS IV
-$s_{ijht}$ 0.183 0.174 0.251 0.169 0.269 0.681
+$s_{ijht}$ 0.166 0.170 0.237 0.169 0.250 0.642
 (0.003) (0.029) (0.004) (0.029) (0.006) (0.015)
-$x_{ijht}$0.537 -0.077 -0.594 -0.186 -0.533 -0.777
+$x_{ijht}$ -0.559 -0.093 -0.617 -0.237 -0.549 -0.783
 (0.003) (0.0174) (0.003) (0.023) (0.006) (0.011)
@@
 Observations 1,200,000 1,200,000 1,200,000 1,200,000 1,200,000 1,200,000
-R-squared 0.957 0.010 0.976 0.032 0.991 0.032
+R-squared 0.958 0.012 0.976 0.041 0.991 0.038
 First-stage F stat. 5,270 3,485 19,760
@@
 SW F stat ($x_{ijht}$) 21,120 7,197 43,550
-Notes: This table reports OLS and IV estimates of equation (3.1), where the dependent variable is the log FOB unit value of product $h$ imported by buyer $j$ from supplier $i$ in year $t$.
+Notes: This table reports OLS and IV estimates of equation (3.1), where the dependent variable is the log CIF duty-excludive unit value of product $h$ imported by buyer $j$ from supplier $i$ in year $t$.
 Columns alternate between OLS and IV specifications.
@@
 Columns (3)–(4) use supplier–product–year ($\text{FE}_{iht}$) and buyer ($\text{FE}_{j}$) fixed effects.
-Columns (5)–(6) include fully interacted buyer–product–year and supplier–product–year fixed effects ($\text{FE}_{jht}$, $\text{FE}_{iht}$), flexibly controlling for sourcing and pricing patterns.
+Columns (5)–(6) include fully interacted buyer–product–year and supplier–product–year fixed effects ($\text{FE}_{jht}$, $\text{FE}_{iht}$).
 IV estimates (even-numbered columns) use leave-one-out instruments: $s_{ijht}$ is instrumented with the average share of other suppliers across buyers of $i$ (excluding $j$), and $x_{ijht}$ with the average share of other buyers across suppliers to $j$ (excluding $i$).
-Because the model includes multiple endogenous regressors, we report both first-stage and conditional F-statistics from Sanderson and Windmeijer (2016), which assess instrument strength for each endogenous regressor conditional on the others, addressing limitations of standard first-stage tests in multi-equation IV settings.
+Because the model includes multiple endogenous regressors, we report both first-stage and conditional F-statistics from SANDERSON2016212.
 Standard errors are robust.
@@
 Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
-We estimate $\alpha_{s}>0$ and $\alpha_{x}<0$, with both coefficients statistically and economically significant.
-A positive $\alpha_{s}$ indicates oligopoly power, requiring exporter bargaining power ($\phi<1$), while a negative $\alpha_{x}$ reflects oligopsony power, requiring importer bargaining power ($\phi>0$) and upward-sloping supply ($\theta<1$).
-These findings are consistent with the model’s core assumption of two-sided market power, i.e., $\phi\in(0,1)$ and $\theta<1$.
-3.4.2 Test of Proposition 4: Pass-Through and Bilateral Market Shares
-We next test Proposition 4, which links tariff pass-through to the importer’s buyer share.
-Under decreasing returns ($\theta<1$), the model predicts that pass-through declines with the buyer share via the cost channel, providing a direct test for $\theta<1$.
-For this analysis, we focus on 2017–2018, when U.S. imports experienced sharp and unanticipated tariff increases under the Trump administration. We estimate the following regression specification:
-$\Delta\ln p_{ijht}=$ $\alpha_{0}+\alpha_{1}\,\Delta\ln(1+\tau_{cht})+\alpha_{s}\,\Delta\ln(1+\tau_{cht})\cdot s_{ijh,t-1}+\alpha_{x}\,\Delta\ln(1+\tau_{cht})\cdot x_{ijh,t-1}$
-$+\alpha_{2}\,s_{ijh,t-1}+\alpha_{3}\,x_{ijh,t-1}+\textbf{X}_{ijht}\boldsymbol{\gamma}\,+\mathbf{FE}+\epsilon_{ijht}.$ (3.2)
-where we use the change in the duty-exclusive price as the dependent variable, defined as $\Delta\ln p_{ijht}=\Delta\ln p_{ijht}^{\text{duty}}-\Delta\ln(1+\tau^{\text{app}}_{cht})$, where $\tau^{\text{app}}_{cht}$ is the applied ad-valorem tariff.
-This transformation isolates price changes net of applied duties, mitigating measurement error.
+A concern is that market shares mechanically depend on prices, potentially inducing spurious correlation.
+To address this, we also report IV estimates using leave-one-out measures of market structure.
+We instrument $s_{ijht}$ with the average supplier share (excluding $i$) across other buyers of exporter $i$ (excluding $j$), and instrument $x_{ijht}$ with the average buyer share (excluding $j$) across other suppliers to importer $j$ (excluding $i$).
+Table 2 reports the results.
+Columns (1)-(2) include exporter, importer, and product-year fixed effects.
+Columns (3)-(4) replace exporter fixed effects with exporter-product-year fixed effects to absorb marginal costs.
+Columns (5)-(6) further include importer-product-year fixed effects to capture buyer-specific demand shocks.
+All specifications control for relationship duration (years since the first shipment of product $h$ between $i$ and $j$).
+Odd columns report OLS; even columns report IV estimates.
+Across specifications, we find $\alpha_{s}>0$ and $\alpha_{x}<0$, both statistically and economically significant.
+While the magnitudes are not directly interpretable in isolation, these patterns provide clear support for the model’s core mechanisms.
+Through the lens of Proposition 2, they are consistent with markup variation reflecting both oligopoly and oligopsony forces.
+Pass-Through and Decreasing Returns to Scale
+Proposition 4 establishes that, in our framework, a negative comovement between pass-through and buyer share arises if and only if production exhibits decreasing returns to scale ($\theta<1$).
+We test this implication using the reduced-form regression
+$\Delta\ln p^{*}_{ijht}=\$$\alpha_{0}+\alpha_{1}\,\Delta\ln(1+\tau_{cht})+\alpha_{s}\,\Delta\ln(1+\tau_{cht})\cdot s_{ijh,t-1}+\alpha_{x}\,\Delta\ln(1+\tau_{cht})\cdot x_{ijh,t-1}$
+$+\alpha_{2}\,s_{ijh,t-1}+\alpha_{3}\,x_{ijh,t-1}+\mathbf{X}_{ijht}\boldsymbol{\gamma}+\mathbf{FE}+\epsilon_{ijht},$ (3.2)
+The dependent variable is the change in the duty-exclusive price, denoted $p^{*}_{ijht}$.
+While the model is expressed in terms of statutory tariffs, observed duty-inclusive prices reflect applied duties and may therefore be subject to measurement error.
+Focusing on duty-exclusive prices isolates price adjustments net of duties and corresponds to the exporter’s received price in the model.
+[^22] We write the buyer’s effective price as $p_{ijh}=p^{*}_{ijh}(1+\tau)$, where $p^{*}_{ijh}$ denotes the exporter’s received price and $\tau$ the statutory tariff rate.
+We focus on the 2017-2018 period, when U.S.
+imports were subject to large and unexpected tariff increases.
 The interaction terms capture heterogeneity in pass-through with respect to bilateral market shares, $s_{ijh,t-1}$ and $x_{ijh,t-1}$, measured at the beginning of the period.
@@
 We estimate two specifications.
-The first includes product–time ($\text{FE}_{ht}$) and exporting country–sector ($\text{FE}_{cs}$) fixed effects, following standard practice in the pass-through literature.
+The first includes product-time ($\text{FE}_{ht}$) and exporting country-sector ($\text{FE}_{cs}$) fixed effects, following standard practice in the pass-through literature.
 This is our baseline.
-The second is more demanding, adding importer–time ($\text{FE}_{jt}$) and exporting country–time ($\text{FE}_{ct}$) fixed effects.
+The second is more demanding, combining product–time ($\text{FE}{ht}$) with importer–time ($\text{FE}{jt}$) and exporting country–time ($\text{FE}_{ct}$) fixed effects.
 Table 3: Pass-Through and Relationship Heterogeneity
@@
 (1) (2) (3) (4) (5) (6)
-$\Delta\ln(1+\tau_{cht})$0.151 -0.188 -0.171 -0.045 -0.066 -0.123
+$\Delta\ln(1+\tau_{cht})$ -0.126 -0.168 -0.138 -0.019 -0.032 -0.094
 (0.091) (0.105) (0.094) (0.097) (0.091) (0.103)
-$\Delta\ln(1+\tau_{cht})\cdot\ln\text{longevity}_{ijht}$ 0.026 0.042
+$\Delta\ln(1+\tau_{cht})\cdot\ln\text{longevity}_{ijht}$ 0.030 0.046
 (0.022) (0.019)
-$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.050 0.054 0.053
+$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.029 0.033 0.032
 (0.084) (0.080) (0.079)
-$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.403 -0.403 -0.411
+$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$ -0.407 -0.407 -0.416
 (0.112) (0.112) (0.112)
@@
 R-squared 0.04 0.04 0.04 0.04 0.04 0.04
-Notes: This table reports estimates of the pass-through of statutory tariffs, $\Delta\ln(1+\tau_{cht})$, to duty-exclusive prices at the exporter–importer–product–year level, $\Delta\ln p_{ijht}$.
-Columns (2) and (6) interact tariffs with the log of relationship longevity, measured as the number of years that buyer $j$ and supplier $i$ have transacted in product $h$.
+Notes: This table reports estimates of the pass-through of statutory tariffs, $\Delta\ln(1+\tau_{cht})$, to CIF duty-exclusive prices at the exporter-importer-product-year level, $\Delta\ln p_{ijht}$.
+Columns (2) and (6) interact tariffs with the log relationship length (in years) within a product $h$.
 Columns (3) and (5) interact tariffs with the lagged supplier share, $s_{ijht-1}$, defined as supplier $i$’s share in buyer $j$’s imports of product $h$.
 Columns (4) and (5) interact tariffs with the lagged buyer share, $x_{ijht-1}$, defined as buyer $j$’s share in supplier $i$’s exports of product $h$.
-All regressions include product–year and exporter country–sector fixed effects ($\text{FE}_{ht}+\text{FE}_{cs}$).
+All regressions include product-year and exporter country-sector fixed effects ($\text{FE}_{ht}+\text{FE}_{cs}$).
 Controls include: (i) $\ln\text{longevity}_{ijht}$; (ii) $\Delta\ln q_{i(-j)ht}$, the change in exporter $i$’s total sales of $h$ to U.S.
@@
 Census Bureau disclosure guidelines.
-See Table D.11 for results using an alternative definition of arm’s-length trade based on LFTTD.
 Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
 Table 3 presents the results using the baseline fixed effects.
-Column (1) shows that, on average, pass-through into duty-exclusive prices is incomplete: a 10% tariff increase reduces exporter prices by 1.5% , corresponding to 85% pass-through rate.
-Column (2) adds an interaction with relationship age to account for the role of match longevity, which has been shown to influence price adjustments.
-The results confirm that pass-through rises with relationship length, consistent with the results in Heise (2024).
-Columns (3)–(6) show that the coefficient on supplier share ($\alpha_{s}$) is positive but insignificant, suggesting limited pass-through heterogeneity on the exporter side.
-In contrast, the coefficient on buyer share ($\alpha_{x}$) is consistently negative and statistically significant, indicating that importers with greater buyer share face lower pass-through.
-Table D.2 in Appendix D.2 confirms that these patterns hold under more demanding fixed effects.
-Table D.3 shows they are robust to alternative price definitions and general equilibrium controls.
-To assess nonlinearities, we interact tariff changes with quartiles of lagged supplier and buyer shares.
-Figure D.1 shows no systematic pattern across supplier share quartiles but a clear, monotonic decline across buyer share quartiles, regardless of fixed effects.
-Together, these results suggest that pass-through is largely unresponsive to supplier concentration but declines strongly with buyer concentration.
-This pattern aligns with the model’s predictions under two-sided market power: the strong buyer share gradient provides direct evidence of decreasing returns, as formalized in Proposition 4, while the weak supplier share gradient reflects the dominance of cost-channel, which suggests a high-$\phi$, low-$\theta$ environment.
-4 Structural Estimation
-The patterns documented above are consistent with the model’s predictions under two-sided market power.
-However, as discussed, they do not warrant direct inference on the structural parameters.
-We now turn to a structural estimation approach to quantify the role of two-sided market power in shaping international prices.
-Our estimation targets two key parameters: the importer’s bargaining power, $\phi$, and the returns to scale parameter, $\theta$, which governs the elasticity of foreign export supply.
-The remaining parameters are taken from the literature or directly measured from the data.
-We set the elasticity of substitution across foreign varieties to $\rho=10$, consistent with Anderson and van Wincoop (2004) and Edmond et al.
-(2023), who adopt similar values to match observed U.S.
+Column (1) shows that, on average, pass-through into duty-exclusive prices is incomplete: a 10% tariff increase reduces exporter prices by 1.3% , corresponding to 87% pass-through rate.
+Column (2) shows that this finding is robust to controlling for relationship age.
+It also reveals that pass-through increases with the length of the relationship, consistent with heise2024firm.
+Columns (3), (5), and (6) show that the coefficient on supplier share ($\alpha_{s}$) is positive but statistically insignificant, suggesting limited heterogeneity in pass-through on the exporter side.
+In contrast, the coefficient on buyer share ($\alpha_{x}$) (Columns (4)-(6)) is negative and highly statistically significant, indicating that importers with greater buyer share experience lower pass-through.
+Table D.2 in the Appendix D.2 confirms that these patterns persist under more demanding fixed effects, while Table D.3 shows they are robust to alternative price definitions and the inclusion of general equilibrium controls.
+To assess for nonlinearities, we interact tariff changes with quartiles of supplier and buyer shares.
+Figure D.1 shows no systematic pattern across supplier share quartiles, but a clear, monotonic decline across buyer share quartiles, regardless of the fixed effects.
+Taken together, these results indicate that pass-through is largely insensitive to supplier concentration but declines sharply with buyer concentration.
+By Proposition 4, the negative gradient in $x$ implies $\theta<1$; these patterns are consistent with the first row of Figure 2.
+4 Structural Evidence
+The reduced-form patterns in Section 3.4 are consistent with two-sided market power and decreasing returns, but they cannot separately identify the structural parameters or convincingly isolate the model’s mechanisms from general-equilibrium effects.
+We therefore adopt a structural approach.
+Estimation focuses on $\phi$ and $\theta$, the two primitives that govern bargaining and decreasing returns; the remaining parameters are calibrated from the literature.
+We set the elasticity of substitution across foreign varieties to $\rho=10$, consistent with Anderson2004 and edmond2023costly, who use similar values to match observed U.S.
 markups.
-The downstream demand elasticity faced by importers is set to $\nu=4$, based on the estimates in Broda and Weinstein (2006).
-[^23] Appendix E.1 provides further discussion.
-The elasticity of importer $j$’s marginal cost with respect to the foreign input price index is set to $\gamma=0.5$, following Eldridge and Powers (2018), who document the share of imported inputs in total material costs for U.S.
-manufacturers.
-Input cost shares $\alpha_{jht}$ are directly measured from the data.
-Finally, we set the importer’s returns to scale parameter to $\varrho=1$ in the baseline specification.
-As discussed in Section 2.3.1, the parameters $\nu$, $\gamma$, and $\varrho$ jointly determine the elasticity $\eta$ of importer $j$’s foreign input bundle $q^{f}_{j}$ with respect to its price index $p^{f}_{j}$.
-Since $\varrho$ enters the model only through $\eta$, fixing it to one does not restrict generality.
-We verify that our results are robust to alternative values of $\eta$.
+The downstream demand elasticity is set to $\nu=4$, based on variety-level import demand estimates in Broda2006.232323Appendix E.1 provides additional details.
+The baseline model features a single foreign input, whereas firms in the data import many (Section 3.2).
+In the empirical implementation, we therefore replace the common elasticity $\gamma$ with the term $\alpha_{jh}\gamma$, where $\alpha_{jh}$ is the observed expenditure share of input $h$ in firm $j$’s imported intermediates.
+We set $\gamma=0.5$ equal to the share of imported inputs in total material costs for U.S.
+manufacturers from Eldridge2018.
+Finally, the importer’s returns-to-scale parameter $\varrho$ has no direct counterpart in the supply-chain production literature.
+We therefore adopt $\varrho=1$ as a parsimonious benchmark and assess robustness to alternative values, including $\varrho=0.5$.
+Because equilibrium outcomes depend on importer technology $(\nu,\gamma,\varrho)$ only through the composite elasticity $\eta_{jh}$, these exercises are equivalent to varying $\eta_{jh}$ directly.
 4.1 Identification and Estimation of the Parameters $\theta$ and $\phi$
-Let $\Omega_{ijt}$ denote the information set available to a given $i-j$ pair during negotiations.
-This includes observed market shares ($s_{ijht}$, $x_{ijht}$, $\alpha_{jht}$) and calibrated parameters ($\nu$, $\gamma$, $\rho$, $\varrho$).
-As shown in equation (2.6), the bilateral markup depends only on the model primitives $(\phi,\theta)$ conditional on $\Omega_{ijt}$, i.e., $\mu_{ij}=\mu\left(\phi,\theta\mid\Omega_{ijt}\right).$ The log price of product $h$ transacted between exporter $i$ and importer $j$ in year $t$ can thus be written as:
-$\ln p_{ijht}=\ln\mu\left(\phi,\theta\mid\Omega_{ijt}\right)+\ln c_{iht},$
-where $\ln c_{iht}$ denotes the exporter’s marginal cost. From equation (2.1), this is given by:
-$\ln c_{ijht}=\frac{1-\theta}{\theta}\ln q_{iht}+\ln k_{iht}.$
-In the theoretical model, the term $k_{iht}$ captured exporter-level cost shifters.
-In the empirical implementation, we generalize this term to allow for match-specific cost components, such as relationship-specific know-how or specialization, by letting it vary flexibly at the match-year level, $k_{ijht}$, thereby absorbing all $(i,j,h,t)$-specific variation.
-In contrast, the term $\frac{1-\theta}{\theta}\ln q_{iht}$ is constant across importer matches for a given supplier–product–year.
-Identification of $(\phi,\theta)$ thus relies on cross-sectional variation in bilateral market shares, as discussed next.
-Consider an exporter $i$ matched with two importers, $j$ and $\ell$.
-Conditional on the joint information set $\boldsymbol{\Omega}_{ij\ell t}\equiv(\Omega_{ijt},\Omega_{i\ell t})$, we assume that the unobserved component of marginal cost is mean-independent of the buyer identity: $\Delta k_{ij\ell ht}\equiv\mathbb{E}_{k}\left[k_{ijht}-k_{i\ell ht}\mid\boldsymbol{\Omega}_{ij\ell t}\right]=0.$ [^24] Omitted variables may induce $\Delta k_{ij\ell ht}\neq 0$, raising endogeneity concerns.
-One possible issue is endogenous network formation: unobserved factors may simultaneously affect both match formation and pricing, biasing estimates of $\hat{\phi}$ and $\hat{\theta}$.
-We address this by differencing across buyers of the same exporter, which removes exporter-level shocks common to all matches.
-To address remaining concerns, we further implement an instrumental variable strategy.
-Taking log price differences across buyers $j$ and $\ell$ served by the same exporter $i$ in year $t$ yields the following moment condition:
-$g(\phi,\theta\mid\boldsymbol{\Omega}_{ij\ell t})\equiv\mathbb{E}_{k}\Big{[}\ln p_{ijht}-\ln p_{i\ell ht}-\big{(}\ln\mu(\phi,\theta\mid\Omega_{ijt})-\ln\mu(\phi,\theta\mid\Omega_{i\ell t})\big{)}\Big{|}\boldsymbol{\Omega}_{ij\ell t}\Big{]}=0.$ (4.1)
-Identification requires that equation (4.1) does not hold for two pairs ($\phi,\theta)$ such that $(\phi^{A},\theta^{A})\neq(\phi^{B},\theta^{B})$.
-Since the oligopoly markup is independent of the returns to scale parameter $\theta$, identification of $\theta$ requires that the oligopsony channel plays a role in price determination, i.e., $\phi>0$, a condition supported by the reduced-form evidence in Tables 2 and 3.
-We therefore focus on the empirically relevant case of bilateral bargaining power $\phi\in(0,1)$.
-The markup function $\mu(\phi,\theta\mid\Omega)$ is strictly monotonic in both parameters and thus invertible in each.
-It follows that the moment condition in equation (4.1) is also invertible in $\phi$ and $\theta$.
-Identification then relies on observing multiple importer–exporter pairs in the same year, or multiple matches for a given exporter over time, under the assumption that bargaining weights remain constant across matches.
-[^25] Formally, identification relies on the nonlinearity of the markup equation (2.6) in $s_{ijht}$ and $x_{ijht}$.
-Consider moment conditions from two periods $t$ and $t-1$: the associated derivatives with respect to $(\phi,\theta)$ are not collinear, satisfying the full-rank condition.
-Similar variation across multiple matches for the same exporter in a given year (e.g., $i$$j$$k$ vs.
-$i$$j$$\ell$) also secures identification.
-Importantly, our identification strategy does not hinge on strong assumptions about the exogeneity of tariffs or other aggregate variables.
-While the estimated parameters may be shaped by general equilibrium forces, identification relies solely on cross-sectional variation across buyers within supplier–product–year cells.
-As a result, we do not need to specify or model the broader general equilibrium environment.
+Let $\Omega_{ijt}$ denote the information set available to pair $(i,j)$ at the time of negotiation.
+This includes observed market shares and calibrated parameters.
+As shown in equation (2.10), conditional on $\Omega_{ijt}$ the bilateral markup depends only on the structural parameters $(\phi,\theta)$, so that $\mu_{ij}=\mu\!\left(\phi,\theta\mid\Omega_{ijt}\right).$
+We write the log price of input $h$ sold by exporter $i$ to importer $j$ in year $t$ as
+$\ln p_{ijht}=\ln\mu\!\left(\phi,\theta\mid\Omega_{ijt}\right)+\ln MC_{ijht}\qquad\text{with}\qquad\ln MC_{ijht}=\frac{1-\theta}{\theta}\,\ln q_{iht}+\ln k_{ijht},$
+where $p_{ijht}$ is the CIF duty-exclusive unit value and $k_{ijht}$ denotes a match-year-specific cost component.
+In the theoretical model, the marginal cost term ($\ln MC_{ijht}$) is common across buyers of a given exporter-product-year.
+In the empirical implementation, we allow it to vary flexibly at the match-year level to absorb any $(i,j,h,t)$-specific cost differences in levels.
+The scale term $\frac{1-\theta}{\theta}\ln q_{iht}$ remains common across buyers within a supplier-product-year and captures the effect of decreasing returns.
+Consider exporter $i$ matched with two importers, $j$ and $\ell$.
+Conditional on the joint information set $\boldsymbol{\Omega}_{ij\ell t}\equiv(\Omega_{ijt},\Omega_{i\ell t})$, we assume mean independence of the unobserved cost component across buyers:
+$\mathbb{E}_{k}\!\left[\ln k_{ijht}-\ln k_{i\ell ht}\mid\boldsymbol{\Omega}_{ij\ell t}\right]=0.$ (4.1)
+Differencing log prices across buyers served by the same exporter yields the moment condition
+$\mathbb{E}_{k}\Big[\ln p_{ijht}-\ln p_{i\ell ht}-\big(\ln\mu(\phi,\theta\mid\Omega_{ijt})-\ln\mu(\phi,\theta\mid\Omega_{i\ell t})\big)\mid\boldsymbol{\Omega}_{ij\ell t}\Big]=0.$ (4.2)
+After differencing, exporter-level marginal cost components cancel out, so the remaining price differences across buyers reflect differences in bilateral markups.
+Since the markup function depends only on $(\phi,\theta)$ conditional on $\Omega_{ijt}$, variation in market shares identifies the structural parameters.
+Identification requires that equation (4.2) holds uniquely at the true $(\phi,\theta)$.
+Because the oligopoly component of the markup does not depend on $\theta$, identification of $\theta$ requires $\phi>0$, so that the oligopsony channel affects pricing.
+The empirical evidence in Tables 2 and 3 supports the presence of such bilateral effects.
+For $\phi\in(0,1)$, the markup function is strictly monotonic in both parameters over the empirically relevant range.
+This monotonicity implies local invertibility of the moment condition in $(\phi,\theta)$.
+Identification is therefore secured by observing multiple matches for the same exporter, either within a year or over time, under the assumption that bargaining weights are constant across matches.
+[^24] Identification follows from the nonlinearity of equation (2.10) in $s_{ijht}$ and $x_{ijht}$, which ensures that the derivatives with respect to $(\phi,\theta)$ satisfy a rank condition.
+Threats to Identification
+A first concern is endogenous match formation or omitted match characteristics.
+If such factors affect $\Delta k_{ij\ell ht}$, assumption (4.1) may be violated, potentially biasing estimates of $\hat{\phi}$ and $\hat{\theta}$.
+Our strategy mitigates this concern by differencing across buyers of the same exporter, which removes shocks common to all matches of that supplier.
+To further address residual concerns, we implement an instrumental variable strategy; see Appendix E.2 for details.
+A second concern is that price dispersion within supplier-product-year cells, even within narrowly defined HS10 categories, may reflect product or service differences rather than bargaining.
+In the model, such heterogeneity is captured by the idiosyncratic cost component $k_{ijht}$, which shifts price levels but does not drive identification.
+Identification instead comes from the systematic relationship between prices and bilateral market shares.
+To overturn our interpretation, unobserved match characteristics would need to replicate the nonlinear price-share relationship generated by bilateral bargaining under decreasing returns.
+Finally, identification does not rely on strong assumptions about the exogeneity of tariffs or other aggregate shocks.
+Although general equilibrium forces may influence the estimated parameters, identification is based solely on cross-sectional variation across buyers within supplier-product-year cells.
+We therefore require only that such aggregate forces do not differentially affect prices across matches of the same supplier.
 Estimation
@@
 $\min_{\{\mathbf{{\phi}},\theta\}}\,\,\mathbf{g}(\mathbf{{\phi}},\theta)\ \mathbf{Z}^{\prime}\ \mathbf{W}\ \mathbf{Z}\ \mathbf{g}(\mathbf{{\phi}},\theta)^{\prime},$ (4.3)
-where $\mathbf{g}(\mathbf{\phi},\theta)$ stacks all moment conditions in equation (4.1) across all $i$$j$$\ell$ pairs and years and $\mathbf{W}$ is the optimal weighting matrix.
-[^26] Appendix E.2 presents Monte Carlo simulations based on a data-generating process that mirrors the setup in Section 2.
-The results confirm that our estimators is consistent.
+[…]
 To address endogeneity concerns, we first include fixed effects by demeaning $\mathbf{g}({\phi},\theta)$ at the HS10 product, year, and buyer level.
@@
 In particular, the vector $\mathbf{Z}$ includes the total number of importers and exporters in each HS10 product-year, which we interpret as proxies for the pool of potential US buyers and foreign suppliers in a given variety.
-We also include in $\mathbf{Z}$ the mean and median of the distributions of the two bilateral shares within each year, excluding the focal pairs $i-j$ and $i-\ell$ to preserve over-identification.
+We also include in $\mathbf{Z}$ the mean and median of the distributions of the two bilateral shares within each year, excluding the focal pairs iji-j and $i$-$\ell$ to preserve over-identification.
 These instruments vary with the competitive structure within each HS10 product-year and are correlated with the endogenous variables through market structure, but, by construction, are not correlated with the idiosyncratic shocks affecting individual matches.
 Extension: pair-specific bargaining weights
-While our baseline assumes a constant bargaining weight $\phi$ across all importer–exporter pairs, we also consider an extension to allow $\phi$ to vary at the pair level.
+While our baseline assumes a constant bargaining weight $\phi$ across all importer-exporter pairs, we also consider an extension to allow $\phi$ to vary at the pair level.
 Given the large number of trade pairs in the data, estimating a separate $\phi_{ij}$ for each is computationally burdensome.
@@
 where $\boldsymbol{\kappa}$ is a parameter vector to be estimated and $\mathbf{X}_{ijht}$ includes covariates that plausibly influence bargaining outcomes but are not direct determinants of gains from trade in our model.
-Specifically, we include: (i) the longevity of the $i-j$ relationship, (ii) the number of transactions between $i-j$ in a year, (iii) the relative outside option of the two, measured by the ratio of the quantity of the exporter $i$’s sales to buyers other than $j$ in year $t-1$ over the quantity of the importer $j$’ purchases from suppliers other than $i$ in year $t-1$, and (iv) an indicator variable of whether the buyer and supplier transact multiple HS10 products.
+Specifically, we include: (i) the longevity of the iji-j relationship, (ii) the number of transactions between iji-j in a year, (iii) the supplier’s outside option relative to the importer’s, measured by the ratio of the quantity of the exporter $i$’s sales to buyers other than $j$ in year t1t-1 over the quantity of the importer $j$’ purchases from suppliers other than $i$ in year t1t-1, and (iv) an indicator variable of whether the buyer and supplier transact multiple HS10 products.
 4.2 Estimation Results
 We estimate equation (4.3) using data from 2001 to 2016.
-We exclude 2017 and 2018, as these years will be used to validate the model out-of-sample in Section 4.3, leveraging the tariff shocks that occurred during this period.
-To avoid convergence issues when $\phi$ is near one, we estimate the transformed parameter $\bar{\phi}\equiv\ln\frac{\phi}{1-\phi}$, which enters the markup equation linearly.
+We exclude 2017 and 2018, which we reserve for out-of-sample validation in Section 4.3 using the tariff shocks during that period.
+To avoid convergence issues when $\phi$ is close to one, we estimate the transformed parameter $\bar{\phi}\equiv\ln\frac{\phi}{1-\phi}$, which enters the markup equation linearly.
 Table 4: Estimated Model Primitives
 Panel A: Calibrated Parameters $\hat{\nu}$ $\hat{\gamma}$ $\hat{\rho}$ 4 0.5 10 Panel B: Estimated Parameters (GMM) (1) (2) (3) (4) Rel.
-bargaining power: $\ln\widehat{\frac{\phi}{1-\phi}}$ 1.565 0.863 (0.055) (0.043) Returns to scale ($\hat{\theta}$) 0.454 0.497 0.383 0.502 (0.004) (0.006) (0.006) (0.007) Constant 4.118 1.454 (0.428) (0.180) Longevity -0.360 0.332 (0.062) (0.064) Number of HS10 transactions -0.264 -0.003 (0.029) (0.014) Multiple HS10 dummy -0.180 0.131 (0.047) (0.034) Lagged outside option -0.235 -0.230 (0.031) (0.030) None Yes Yes No No $\text{FE}_{h}+\text{FE}_{t}+\text{FE}_{j}$ No No Yes Yes Observations 3,120,000 Panel C: Implied Bargaining Powers ($\hat{\phi}$) Mean 0.827 0.922 0.703 0.860 (0.008) (0.074) (0.009) (0.099) Median 0.945 0.886 (0.074) (0.099) Notes: This table presents model estimates based on our main estimation sample, which focuses on U.S.
-imports of intermediate inputs and capital goods for the period 2001-2016.
-Panel A reports calibrated parameters: the elasticity of demand ($\nu$), the elasticity of costs with respect to foreign input prices ($\gamma$), and the elasticity of substitution across foreign varieties ($\rho$).
-We set $\varrho=1$, so that $\eta=2.5$.
-Panel B presents GMM estimates.
-Columns (1) and (3) impose a constant $\phi$ across bilateral pairs, while Columns (2) and (4) estimate the full vector $\boldsymbol{\kappa}$ to allow for heterogeneity in bargaining power.
-Specifications differ in the inclusion of fixed effects.
-Controls include: (i) the log of relationship longevity between exporter $i$ and importer $j$; (ii) the log of the number of transactions between $i$ and $j$ in a given year; (iii) the log of the relative outside option, defined as the ratio of exporter $i$’s sales to other U.S.
-buyers (excluding $j$) over importer $j$’s purchases from other suppliers (excluding $i$), both in year $t{-}1$; and (iv) a dummy variable equal to one if the $i-j$ pair transacts in more than one HS10 product.
-Panel C reports the mean and median of the implied bargaining power.
-Standard errors are robust; those in Panel C are computed using the delta method.
-The set of instruments includes the number of exporters and importers at the HS10 level, as well as lagged bilateral shares (excluding the focal pair).
+bargaining power: $\ln\widehat{\frac{\phi}{1-\phi}}$ 1.563 1.168 (0.051) (0.043) Returns to scale ($\hat{\theta}$) 0.455 0.505 0.417 0.508 (0.004) (0.004) (0.005) (0.006) Constant 4.327 2.099 (0.361) (0.231) Longevity -0.216 0.476 (0.043) (0.090) Number of HS10 transactions -0.342 -0.115 (0.030) (0.024) Multiple HS10 dummy 0.004 0.113 (0.038) (0.039) Lagged outside option -0.285 -0.242 (0.028) (0.032) None Yes Yes No No $\text{FE}_{h}+\text{FE}_{t}+\text{FE}_{j}$ No No Yes Yes Observations 3,120,000 Panel C: Implied Bargaining Powers ($\hat{\phi}$) Mean 0.827 0.931 0.763 0.904 (0.007) (0.081) (0.008) (0.079) Median 0.957 0.927 (0.081) (0.079) Notes: This table presents model estimates using our main estimation sample of U.S.
+imports of intermediate inputs and capital goods, 2001–2016.
+Panel A reports calibrated parameters: demand elasticity ($\nu$), cost elasticity to foreign input prices ($\gamma$), and the elasticity of substitution across foreign varieties ($\rho$).
+Panel B reports GMM estimates.
+Columns (1) and (3) impose a constant $\phi$, while Columns (2) and (4) allow heterogeneous bargaining power via $\boldsymbol{\kappa}$.
+Specifications differ in fixed effects.
+Controls include relationship longevity, transaction intensity, the relative outside option (lagged), and a multi-product indicator.
+Panel C reports the mean and median implied bargaining power.
+Robust standard errors; Panel C uses the delta method.
+Instruments include HS10-level counts of exporters and importers and lagged bilateral shares (excluding the focal pair).
+All estimations use CIF duty-exclusive unit values.
 The number of observations is rounded to four significant digits in accordance with U.S.
@@
 Across specifications, U.S.
-importers appear to wield substantial bargaining power, with estimated values of $\phi$ ranging from 0.70 to 0.92.
+importers appear to wield substantial bargaining power, with estimated values of $\phi$ ranging from 0.7 to 0.9.
 Our preferred estimate, reported in Column (1), is $\hat{\phi}=0.83$, implying that U.S.
 importers have, on average, roughly four times the bargaining power of their foreign suppliers.
-[^27] These findings are consistent with evidence from related settings.
-Morlacco (2019) documents significant markdowns by French importers in input trade, while Atkin et al.
-(2024) show that Argentine importers often exercise considerable bargaining power.
-The returns to scale parameter $\hat{\theta}$ is consistently estimated below one, ranging from 0.40 to 0.50 across specifications, with a preferred estimate of 0.45.282828While standard estimates of returns to scale often cluster near one, $\theta$ here captures a short-run returns to scale elasticity, i.e., the slope of the exporter’s marginal cost curve during bargaining.
-This implies a residual export supply elasticity between 0.25 and 0.375 for the average importer, indicating relatively steep foreign supply curves.
-[^29] The implied residual supply elasticity is computed from the marginal cost slope $c^{\prime}_{q_{ijh}}=\frac{1-\theta}{\theta}\cdot x_{ijh}$, using the average buyer share of 0.25 from Table 1.
+The returns to scale parameter $\hat{\theta}$ is consistently estimated below one, ranging from 0.42 to 0.51 across specifications, with a preferred estimate of 0.46.252525While our estimate of $\theta$ lies below standard production-function estimates, which are typically closer to one, this reflects the short-run nature of the object we identify.
+Standard estimates rely on firm-level variation in inputs and outputs and need not recover the same object as our parameter, which is a strictly short-run parameter.
+This implies an inverse residual export supply elasticity of about 0.24-0.35 for the average importer.
+[^26] The implied elasticity is computed as $d\ln MC_{ih}/d\ln q_{ijh}=\frac{1-\theta}{\theta}\cdot x_{ijh}$, using the average buyer share of 0.25 from Table 1.
 These values are consistent with evidence from U.S.
-manufacturing under short-run constraints: Boehm and Pandalai-Nayar (2022) report median inverse elasticities around 0.3 at typical capacity levels, and Broda et al.
-(2008) document similarly low elasticities across many traded goods.
-Moving to the estimates of the vector $\hat{{\boldsymbol{\kappa}}}$, we find that the coefficients on relationship longevity, frequency of transactions, and the multiple-product indicator are highly significant, although their sign varies depending on the set of fixed effects included.
-By contrast, the coefficient on the relative outside option is stable across specifications: importers hold less bargaining power when their supplier has a stronger outside option.
+manufacturing under short-run constraints: boehm2022convex report median inverse elasticities around 0.3 at typical capacity levels, and broda2008optimal document similarly low elasticities across many traded goods.
+Moving to the estimates of the vector $\hat{{\boldsymbol{\kappa}}}$, the results are mixed across the relationship-level controls.
+Relationship longevity is significant in both specifications but changes sign when fixed effects are included, suggesting that its correlation with bargaining power is sensitive to how unobserved heterogeneity is handled.
+The coefficient on transaction frequency is consistently negative and significant, while the multiple-product indicator is positive but precisely estimated only in the fixed-effects specification.
+By contrast, the coefficient on the relative outside option is stable and precisely estimated across all specifications: importers hold less bargaining power when their supplier has a stronger outside option.
 Specifically, an increase in the supplier’s past sales to other buyers relative to the importer’s purchases from other suppliers is consistently associated with lower bargaining power for the importer.
 Robustness
-We assess the robustness of our structural estimates to alternative sample definitions and model calibrations.
-Appendix D.3 first considers a broader sample that includes all products in the BEC classification, notably extending the baseline by adding consumption goods.
-We also examine an alternative sample that uses related-party trade indicators from the LFTTD instead of ORBIS.
-On the calibration side, we consider a lower elasticity of substitution across foreign varieties ($\rho=5$ instead of 10) and introduce decreasing returns to scale in downstream production by setting $\varrho=0.5$ rather than 1.
-This choice aligns with the estimated returns to scale on the exporter side and allows us to test the robustness of the estimates to alternative values of the importer’s downstream demand elasticity, $\eta$.
-Across all variations, the structural estimates remain highly stable.
+We assess the robustness of the structural estimates to alternative sample definitions and model calibrations. Table 5 reports the results.
+Table 5: Robustness of Model Estimates
+Baseline Only Consumption All Products $\rho=5$ $\varrho=0.5$ (1) (2) (3) (4) (5) Bargaining power ($\phi$) 0.827 0.718 0.765 0.877 0.856 (0.007) (0.005) (0.004) (0.004) (0.007) Returns to scale ($\theta$) 0.455 0.558 0.508 0.465 0.470 (0.004) (0.003) (0.002) (0.002) (0.003) Observations 3,120,000 3,024,000 6,143,000 3,120,000 3,120,000
+Notes: This table reports robustness checks for the structural estimates.
+Columns vary either the sample definition or the calibrated parameters.
+“Only Consumption” restricts the sample to consumption goods.
+“All Products” includes all product categories.
+Columns $\rho=5$ and $\varrho=0.5$ modify the calibrated elasticities.
+All columns report GMM estimates without fixed effects.
+All estimations use CIF duty-exclusive unit values.
+Observation counts are rounded in accordance with U.S.
+Census Bureau disclosure guidelines.
+Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
+Restricting the sample to consumption goods only (column (2)) yields a slightly lower estimate of $\phi$ and a slightly higher estimate of $\theta$.
+Column (3) reports estimates for the pooled sample, including all products, which lie in between, as expected.
+The model’s core mechanisms, therefore, operate in consumption goods transactions as well.
+We nonetheless retain the baseline focus on capital and intermediate inputs, for which the framework is more naturally suited.
+[^27] Adapting the framework specifically to consumption goods would likely require modifications to the modeling of the importer’s technology and its calibration.
+On the calibration side, we consider two alternative parameter values.
+First, we lower the elasticity of substitution across foreign varieties to $\rho=5$ (instead of 10), at the lower end of the range reported in Anderson2004.
+Second, we introduce decreasing returns to scale in downstream production by setting $\varrho=0.5$, close to the supplier-side estimate.
+Both $\rho$ and $\varrho$ affect how cross-sectional price dispersion varies with the exporter’s supplier share, and thus with oligopoly power.
+Columns (4) and (5) show that the structural estimates remain highly stable.
+[^28] Results are robust to alternative related-party trade classifications.
+See Appendix D.3 and Table D.6 for details.
+This suggests that $\phi$ is primarily identified from variation in buyer share rather than supplier share.
 Implied Markups
-Using equation (2.6), we compute markups for all buyer–supplier–product matches given the estimated parameters and the observed distribution of market shares.
-The resulting markup distribution is clustered near the competitive benchmark.
-Our preferred estimates yield a mean markup of 0.94, with the median even closer to competitive levels.
-These low markups reflect the strong countervailing power of importers.
-The model implies that observed markups are a convex combination of oligopoly and oligopsony markups, which average 1.34 and 0.87, respectively.
-The large estimated bargaining power of buyers ($\hat{\phi}\approx 0.8$) shifts weight toward the oligopsony case, allowing importers to extract a substantial share of the surplus from exporters.
-[^30] The effective bargaining weight $\omega_{ijh}$ averages 0.77 (standard deviation 0.05), slightly below $\hat{\phi}$.
-This indicates that network effects, on average, dampen the importers’ effective bargaining power relative to $\phi$, although the gap is small.
+Using equation (2.10), we compute markups for all buyer-supplier–product matches given the estimated parameters and the observed market shares.
+The resulting markup distribution is clustered near the competitive benchmark, with a median of 1.01.
+This contrasts with the oligopoly benchmark, whose median is 1.21 and whose interquartile range spans 1.11 to 2.45.
+In the model, observed markups are a convex combination of oligopoly and oligopsony markups, where the median oligopsony markup is 0.95.
+The large estimated bargaining power of buyers ($\hat{\phi}\approx 0.8$) shifts weight toward oligopsony, allowing importers to capture a substantial share of the surplus from exporters.
 4.3 Model Validation
-We assess the model’s empirical validity by testing its ability to predict both the level and heterogeneity of price and quantity changes following the 2017–2018 tariff increases.
-These moments were not targeted in the estimation.
-Model-predicted price changes are computed as:
-$\widehat{\Delta\ln p_{ijht}}=\Phi_{ijht}(s_{ijht},x_{ijht}\mid\boldsymbol{\hat{\Theta}})\cdot\Delta\ln(1+\tau_{cht}),$ (4.4)
-where $\tau_{cht}$ denotes the ad-valorem tariff on product $h$ from country $c$, and $\Phi_{ijht}$ is the model-implied pass-through elasticity defined in equation (2.11), which depends on bilateral shares and the estimated parameter vector $\boldsymbol{\hat{\Theta}}$.
-Although the model is primarily designed to explain prices, it embeds a demand-driven allocation rule that links prices to quantities (and sales) via the importer’s demand curve.
-This mapping imposes a specific quantity-setting structure, enabling a direct comparison between predicted price responses and observed adjustments in quantities and trade values.
-Predicted quantity changes are given by $\widehat{\Delta\ln q_{ijht}}=-\hat{\varepsilon}_{ijht}\cdot\widehat{\Delta\ln p_{ijht}},$ where $\hat{\varepsilon}_{ijht}$ denotes the match-specific residual demand elasticity implied by the model.
-For sales, the corresponding mapping is $\widehat{\Delta\ln r_{ijht}}=-(1-\hat{\varepsilon}_{ijht})\cdot\widehat{\Delta\ln p_{ijht}}$.
+Proposition 2 yields a closed-form expression for the pass-through elasticity, which we use to assess the model’s out-of-sample performance.
+Specifically, we examine whether the model can predict the level and heterogeneity of price responses to the 2017–2018 tariff increases, which are excluded from estimation.
+Using observed shares in 2017 and the estimated parameters, the model-implied change in duty-exclusive prices is
+$\widehat{\Delta\ln p^{*}_{ijht}}=\Phi^{*}_{ijht}\!\left(s_{ijht},x_{ijht}\mid\boldsymbol{\hat{\Theta}}\right)\cdot\Delta\ln\left(1+\tau_{c(i)ht}\right),$ (4.5)
+where $\tau_{cht}$ denotes the statutory tariff rate and $\Phi^{*}_{ijht}$ is the model-implied pass-through elasticity into duty-exclusive prices.
+[^29] Proposition 2 characterizes pass-through into duty-inclusive prices; pass-through into duty-exclusive prices follows as $\Phi^{*}=\Phi-1$.
+Because the model delivers a partial-equilibrium mapping from tariffs to prices while the data reflect equilibrium outcomes, this exercise constitutes a joint test of the model’s partial-equilibrium structure and its underlying mechanisms.
+Table 6 reports pass-through elasticities in the data and in the model for the baseline specification (equation (3.2)) and three restricted variants, using baseline fixed effects.
+[^30] For the restricted variants, equation (4.5) is evaluated using the baseline parameter estimates from column (1) of Table 4 without re-estimation, except where parameters are restricted by construction (e.g., $\phi=0$ or $\theta=1$).
+Column (1) reports the empirical estimates.
+Column (2) corresponds to the baseline model.
+Column (3) shuts down both bargaining and decreasing returns by imposing $\phi=0$ and $\theta=1$, yielding a standard Nash–Bertrand model with constant marginal costs, as in kikkawa2019imperfect.
+Column (4) sets $\theta=1$, retaining bargaining but eliminating decreasing returns, as in gopinath2011search.
+Column (5) sets $\phi=0$, eliminating importer bargaining power by assigning all pricing power to exporters, while preserving decreasing returns.
 Table 6: Price Responses and Relationship Heterogeneity: Data vs. Model
-Panel A: Data
-(1) (2) (3) (4)
-$\Delta\ln(1+\tau_{cht})$0.151 -0.066 -0.223 -0.163
-(0.093) (0.093) (0.109) (0.107)
-$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.054 0.029
-(0.072) (0.157)
-$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.403 -0.271
-(0.113) (0.135)
-R-squared 0.04 0.04 0.31 0.31
-Panel B: Model
-(1) (2) (3) (4)
-$\Delta\ln(1+\tau_{cht})$0.248 -0.144 -0.249 -0.136
-(0.008) (0.009) (0.010) (0.011)
-$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.104 0.091
-(0.011) (0.010)
-$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.475 -0.486
-(0.026) (0.029)
-R-squared 0.32 0.46 0.50 0.59
-$\text{FE}_{ht}+\text{FE}_{cs}$ Yes Yes No No
-$\text{FE}_{ht}+\text{FE}_{ct}+\text{FE}_{jt}$ No No Yes Yes
+Panel A: Main Effect
+Data Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$
+(1) (2) (3) (4) (5)
+$\Delta\ln(1+\tau_{cht})$ -0.126 -0.222 -0.094 -0.013 -0.279
+(0.091) (0.008) (0.004) (0.001) (0.008)
+R-squared 0.04 0.31 0.22 0.21 0.34
+Panel B: Including Interactions
+Data Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$
+(1) (2) (3) (4) (5)
+$\Delta\ln(1+\tau_{cht})$ -0.032 -0.145 -0.049 -0.005 -0.219
+(0.091) (0.008) (0.009) (0.001) (0.009)
+$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.033 0.125 -0.079 -0.011 0.122
+(0.080) (0.013) (0.014) (0.003) (0.011)
+$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$ -0.407 -0.421 -0.046 -0.013 -0.366
+(0.112) (0.023) (0.023) (0.003) (0.022)
+R-squared 0.04 0.44 0.23 0.23 0.41
+$FE_{ht}+FE_{cs}$ Yes Yes Yes Yes Yes
 Observations 249,000
-Notes: This table reports the pass-through of tariffs to duty-exclusive prices at the exporter–importer–product level.
-Panel A presents reduced-form estimates from the data.
-Panel B shows corresponding pass-through estimates generated by the model.
-Columns (2)–(4) interact tariff changes with lagged supplier share ($s_{ijht-1}$) and lagged buyer share ($x_{ijht-1}$).
-Columns (1) and (2) use baseline fixed effects ($\text{FE}_{ht}+\text{FE}_{cs}$), while Columns (3) and (4) employ a more stringent specification with product–year, country–year, and buyer–year fixed effects ($\text{FE}_{ht}+\text{FE}_{ct}+\text{FE}_{jt}$).
+Notes: This table reports the response of duty-exclusive CIF prices to tariff changes at the exporter–importer–product level.
+Column (1) presents reduced-form estimates using observed price changes in the data.
+Columns (2)–(5) report the corresponding responses implied by model-predicted price changes under each specification.
+Panel A reports the main effect of tariff changes, while Panel B allows the response to vary with lagged supplier shares ($s_{ijht-1}$) and buyer shares ($x_{ijht-1}$).
+All specifications include product–time and country–sector fixed effects ($FE_{ht}+FE_{cs}$).
 Standard errors are clustered at the HS8 product and exporter-country level.
@@
 Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
-4.3.1 Price Predictions
-We begin with price outcomes.
-Table 5 compares tariff pass-through elasticities in the data (Panel A) and in the model (Panel B), in terms of average effects and their heterogeneity with respect to buyer and supplier shares.
-Columns (1) and (2) adopt a baseline specification with product–time and country–sector fixed effects, while Columns (3) and (4) adopt the more demanding specification with buyer–time and country–time fixed effects.
-The model predicts an average pass-through elasticity on duty-exclusive prices of $-0.25$, which translates to a 75% pass-through rate (Columns (1) and (3) of Panel B).
-These values fall within the 95% confidence intervals of the corresponding reduced-form estimates in the data (Columns (1) and (3) of Panel A).
-Columns (2) and (4) of Panel A, consistent with Column (6) of Table 3, show that pass-through declines significantly with buyer share ($x_{ijht}$) but exhibits no robust relationship with supplier share ($s_{ijht}$).
-The model closely replicates these patterns: as shown in Panel B, pass-through declines steeply in $x_{ijht}$ and increases mildly in $s_{ijht}$, with magnitudes comparable to those in the data.
-This alignment supports our interpretation of Table 3 as consistent with a low-$\theta$, high-$\phi$ environment, which is also consistent with our structural estimates.
-4.3.2 Goodness-of-fit Test
-Having shown that the model replicates average and heterogeneous price responses to tariff shocks, we formally test its predictive performance by evaluating how well model-implied price changes explain observed variation.
-This exercise complements Table 5 by providing a direct measure of goodness of fit relative to standard alternatives.
-Specifically, we estimate:
-$\Delta\ln p_{ijht}=\beta\,\,\widehat{\Delta\ln p_{ijht}}\,+\,\mathbf{FE}\,+\,u_{ijht},$ (4.5)
-where $\Delta\ln p_{ijht}$ is the observed change in the duty-inclusive price for product $h$ between exporter $i$ and importer $j$, and $\widehat{\Delta\ln p_{ijht}}$ is the corresponding model-predicted change.
-To benchmark the results, we compare the full model to three nested alternatives that sequentially shut down bilateral bargaining ($\phi=0$), decreasing returns to scale ($\theta=1$), or both.
-The fully restricted case ($\phi=0$, $\theta=1$) corresponds to a standard Nash-Bertrand model with constant marginal costs (e.g., Dhyne et al., 2022); the intermediate case with $\phi>0$ and $\theta=1$ mirrors the bargaining framework of Gopinath and Itskhoki (2011), which overlooks the heterogeneous buyer dimension.
-In all cases, we use the same parameter values from Column (1) of Table 4, without re-estimating.
-[^31] As discussed in Section 4.1, $\theta$ is not identified when $\phi=0$.
-Moreover, fixing $\theta=1$ and estimating equation (4.5) using the value of $\phi$ jointly obtained with $\theta$ from equation (4.2) yields an upper bound on the value of $\hat{\beta}$ attainable with a re-estimated $\phi$.
-Appendix E.3 provides formal proof and supporting simulation evidence.
-A key challenge is that observed price changes may reflect shocks unrelated to tariffs, increasing the risk of rejecting a valid model for reasons unrelated to its tariff predictions.
-In addition, while the model uses statutory tariff changes, observed duty-inclusive prices reflect actual duties paid, potentially introducing measurement error.
-To address both issues, we follow the logic of the IV-based goodness of fit test in Adão et al.
-(2023) and estimate equation (4.5) using two-stage least squares, instrumenting $\widehat{\Delta\ln p_{ijht}}$ with statutory tariff changes.
-This isolates variation in predicted prices that is directly attributable to tariff shocks, which is the one targeted by the model.
-Under the null that the model accurately captures pass-through, the IV coefficient $\hat{\beta}$ should equal one.
-Figure 3: IV-Based Goodness-of-Fit Test
-Notes: Each point reports the coefficient from an IV regression of observed log price changes on model-predicted changes $\widehat{\Delta\ln p_{ijht}}$, using statutory tariffs as instruments.
-Lines show 95% confidence intervals.
-Blue and red denote regressions with $\mathrm{FE}_{ht}+\mathrm{FE}_{cs}$ and $\mathrm{FE}_{ht}+\mathrm{FE}_{ct}+\mathrm{FE}_{jt}$, respectively.
-Standard errors are clustered by product and exporter-country.
-Observation counts (249,000) are rounded per Census disclosure guidelines.
-Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
-Figure 3 presents the results across the four model variants, each estimated under two alternative fixed effects specifications.
-Blue points and lines correspond to regressions with product–time and country–sector fixed effects, while red ones use the more stringent specification with product–time, country–time, and buyer–time fixed effects.
-In all cases, the baseline model provides the best fit: the estimated coefficient is very close to one and not statistically different from it under both specifications.
-Models with constant returns to scale ($\theta=1$), whether or not they include bargaining, perform noticeably worse, although their coefficients are not rejected at conventional significance levels.
-These variants lack a cost channel, and while the case with $\phi=0$ features strategic complementarities that generate some degree of incomplete pass-through, this mechanism alone does not produce sufficient variation to match the data.
-Similarly, the fourth model with decreasing returns but no bargaining ($\phi=0$, $\theta<1$), which was previously rejected based on price-level evidence (Table 2), also underperforms relative to the baseline model.
-In this case, markup responses driven by strategic complementarities amplify cost-based adjustments, but in a way that also fails to replicate the observed pass-through patterns.
-[^32] Appendix D.4 (Table D.6) presents additional robustness checks using alternative calibrations, including $\rho=5$ instead of 10 and $\varrho=0.5$ instead of 1.
-The model’s predictive performance remains stable across these variations, lending further support to our main findings.
-Overall, these results suggest that cost adjustments, along with weak strategic complementarities, are necessary to account for the empirical evidence.
-This is consistent with a setting in which oligopsony forces are dominant.
-4.3.3 Quantity Predictions
-Next, we assess the model’s ability to predict changes in bilateral quantities.
-Table D.7 in Appendix D.4 reports relationship-level quantity responses to tariff changes.
-Columns (3)–(4) of Panel A show that tariff increases reduce traded volumes, with larger declines under more stringent fixed effects.
-However, the interaction terms with supplier and buyer shares are statistically imprecise, suggesting inconclusive evidence of heterogeneity in the data.
-[^33] In robustness exercises (not shown), the signs on the interaction coefficients with supplier and buyer shares occasionally flip, though they largely remain statistically insignificant.
-We therefore conclude that the data do not offer robust evidence of heterogeneous quantity responses by relationship structure.
-Panel B shows that the model generates sizable average quantity declines and predicts heterogeneity across relationships.
-In particular, quantity responses become less negative with higher buyer shares, consistent with the model’s allocative logic: stronger buyers face smaller price increases and thus smaller quantity reductions.
-In contrast, the positive interaction with supplier share is not explained by price adjustments because the model predicts little variation in pass-through along this dimension.
-Instead, it reflects the curvature of the importer’s demand curve embedded in the model.
-Table D.8 in Appendix D.4 evaluates model fit for quantities (Panel A) and sales (Panel B) across the four alternative parameterizations discussed above.
-Since all models share the same demand system, performance differences reflect variation in price predictions rather than differences in quantity mechanisms.
-While all specifications are formally rejected, the baseline model and the specification with decreasing returns to scale perform best, mirroring their superior performance in prices.
-Despite differences in magnitude and precision, the model captures the broad directional patterns in quantity responses, lending support to its core allocative mechanism.
-However, its weaker quantitative fit suggests that additional forces beyond price-based allocation shape the observed quantity adjustments.
-The model’s strong performance on prices highlights this asymmetry and points to potential gains from extending the framework to incorporate supply-side responses.
-5 Aggregate Implications of Bargaining in Firm-to-Firm Trade
-This section examines the impact of the Trump tariffs on aggregate import prices through the lens of our bargaining model.
-We begin by situating our analysis within the broader literature on tariff pass-through.
-We then use the model to simulate the aggregate effects of the tariffs and assess the mechanisms driving price adjustments.
-5.1 Comparison with Existing Pass-Through Estimates
-Several recent studies find near-complete pass-through of the 2018 Trump tariffs to U.S.
-import prices (e.g., Fajgelbaum et al., 2020; Amiti et al., 2019b, 2020).
-These analyses primarily rely on product-level data.
-In contrast, our match-level estimates indicate incomplete pass-through.
-For instance, in a standard specification similar to those used in the literature, Column (1) of Table 5 (Panel A) reports a pass-through elasticity on duty-exclusive prices of $-0.15$, implying an 85% pass-through rate.
-This divergence partly reflects differences in sample composition.
-As discussed in Section 3.3, our estimates are based on repeated firm-to-firm relationships where bilateral prices can be observed over time.
-They therefore capture within-relationship price changes net of compositional shifts across firms or products, rather than on changes in product-level unit values.
-In addition, we restrict attention to arm’s-length transactions involving intermediate goods and suppliers with two or more buyers.
-Within this sample, we estimate pass-through rates ranging from 78 to 85%, depending on the fixed effects used, as shown in Table 5.
-The model’s predicted pass-through closely matches these empirical patterns.
-To further assess the role of sample composition, Table D.9 in Appendix D.5 reports pass-through estimates across alternative sample definitions.
-Expanding the baseline to include single-buyer relationships raises the estimate to approximately 83%.
-Adding matches involving related parties, energy goods, or extreme price changes increases it to roughly 90%.
-The most inclusive specification, which further incorporates final consumption goods, yields pass-through elasticities between 93% and 95%, depending on the fixed effects.
-Although these broader samples imply higher pass-through, they still fall short of full pass-through, in contrast to product-level studies.
-[^34] Using the monthly LFTTD data, we show in Table D.10 in Appendix D.5 that pass-through remains incomplete in product-level regressions when restricting to the firms and products used in our baseline sample, confirming that this is a feature of repeated firm-to-firm relationships.
-These product-level estimates are based on aggregated unit values that combine within-match price changes with shifts in the composition of transactions, particularly the inclusion of one-off (spot) exchanges.
-Such transactions may differ systematically from repeated firm-to-firm relationships.
-In particular, incomplete pass-through may be a feature of relational trade, where prices reflect bilateral bargaining and upward-sloping residual supply.
-Spot exchanges, by contrast, are more likely to reflect marginal cost pricing and thus exhibit full pass-through.
-While speculative, this interpretation is consistent with Cajal-Grossi et al.
-(2023), who find lower markups and more competitive pricing in spot relationships.
-5.2 Tariffs and Aggregate Import Prices
-We use the estimated model to assess the impact of the 2018 tariffs on aggregate import prices and decompose the underlying contributions of markup and marginal cost changes.
-A full evaluation of the tariffs’ macroeconomic effects would require a general equilibrium framework that incorporates changes in expenditure, wages, export prices, and retaliation, as in Fajgelbaum et al.
-(2020).
-While such elements are essential for welfare analysis, they lie beyond the scope of our model.
-Instead, we focus on one central component of tariff incidence: the change in aggregate import prices within ongoing firm-to-firm relationships.
-We compute model-implied bilateral price changes using equation (4.4) and estimate aggregate pass-through by regressing these predicted changes on the tariff shock, weighting each observation by its initial-period import value.
-The approach connects the model’s micro-level predictions to aggregate outcomes and provides a framework for decomposing the contribution of distinct adjustment channels.
-To that end, we separately isolate the roles of markup and marginal cost adjustments.
-For each mechanism, we compute the model-implied price changes while holding the other channel constant, and re-estimate the weighted regression.
-The predicted price change due solely to markup adjustment is defined as:
-$\widehat{\Delta^{\Gamma}\ln p_{ijht}}=\Phi^{\Gamma}_{ijht}(s_{ijht},\,x_{ijht}\,\mid\,\hat{\boldsymbol{\Theta}})\cdot\Delta\ln T_{cht},\quad\text{where}\quad\Phi^{\Gamma}_{ijht}\equiv\frac{1}{1+\Gamma_{ij}}.$
-Similarly, the predicted price change driven solely by cost adjustment is given by:
-$\widehat{\Delta^{\Lambda}\ln p_{ijht}}=\Phi^{\Lambda}_{ijht}(s_{ijht},\,x_{ijht}\,\mid\,\hat{\boldsymbol{\Theta}})\cdot\Delta\ln T_{cht},\quad\text{where}\quad\Phi^{\Lambda}_{ijht}\equiv\frac{1}{1+\Lambda_{ij}}.$
+Panel A of Table 6 reports average pass-through of about 87% in the data.
+The baseline model predicts a lower pass-through, around 78%, yet remains close to the data with overlapping confidence intervals.
+The restricted variants also generate incomplete pass-through, to varying degrees: shutting down both bargaining and decreasing returns ($\phi=0$, $\theta=1$) yields about 91%; eliminating importer bargaining alone ($\phi=0$) produces the lowest pass-through, around 72%; removing decreasing returns alone ($\theta=1$) raises pass-through substantially.
+While all variants generate incomplete pass-through, confidence intervals do not reject any of them, limiting the informativeness of these differences for model fit.
+As illustrated in Figure 2, pass-through reflects the balance of the cost and markup channels, and different $(\phi,\theta)$ combinations can generate similar averages through different mechanisms, particularly in the relevant range of buyer and supplier shares.
+In the baseline, strategic substitutabilities partly offset the cost channel (Column (2)); setting $\phi=0$ shifts the markup channel toward complementarities, further lowering pass-through and moving the model farther from the data (Column (5)); eliminating the cost channel alone still yields incomplete pass-through under strategic complementarities (Column (3)); and removing both the cost channel and the strategic complementarities produces near-complete pass-through (Column (4)).
+Average pass-through is therefore a weak discriminator across models, with more informative evidence coming from cross-sectional moments.
+Panel B of Table 6 interacts tariff changes with lagged supplier and buyer shares.
+In the data, pass-through declines sharply with buyer share and increases modestly with supplier share.
+The baseline model closely replicates both gradients, with magnitudes comparable to those in the data.
+This alignment also sheds new light on Figure D.1: the approximately monotonic decline in pass-through across buyer-share quartiles, which we documented but did not interpret structurally in Section 3.4, finds a natural counterpart in the model, where the cost channel scales linearly with $x_{ij}$ (Equation 2.16).
+Eliminating decreasing returns ($\theta=1$, Columns (3)–(4)) substantially attenuates the buyer-share gradient and counterfactually predicts lower pass-through for exporters with larger supplier shares.
+By contrast, setting $\phi=0$ (Column (5)) yields cross-sectional patterns similar to the baseline.
+This is expected, given Figure 2: when decreasing returns are strong, pass-through heterogeneity is largely governed by the cost channel, leaving limited scope for bargaining power to shape the cross-sectional gradient.
+Taken together, the baseline is the only specification consistent with both panels.
+Column (3) comes closest to the data on average (91%), but is strongly rejected on cross-sectional grounds.
+Column (5) matches the cross-sectional patterns but noticeably overstates price absorption relative to both the data and the baseline.
+The model’s ability to reproduce observed patterns of tariff pass-through lends support to its underlying mechanisms and suggests that the partial equilibrium framework captures key forces governing price adjustment in firm-to-firm trade.
+Additional Validation.
+Table 6 uses baseline fixed effects throughout; Table D.7 in Appendix D.4 repeats the exercise under more stringent fixed effects and yields the same conclusions, both in terms of average pass-through and cross-sectional patterns.
+One concern is that observed price changes may reflect shocks unrelated to tariffs, potentially leading the data to reject (or accept) a correctly (or incorrectly) specified model for spurious reasons.
+To address this, we formalize the comparison using the IV goodness-of-fit test of adao2023putting, which isolates tariff-driven variation in observed prices.
+Figure D.2 in Appendix D.4 reports the results.
+Consistent with Panel A of Table 6, none of the variants can be formally rejected based on average pass-through alone.
+Among the non-rejected specifications, the baseline provides a closer fit than Column (5), consistent with Table 6.
+Table D.8 shows that these results are robust to alternative calibrations of the model parameters and remain stable across various fixed-effects specifications.
+Finally, we assess the model’s ability to predict bilateral quantity and sales responses.
+[^31] Predicted quantity and sales changes are given by $\widehat{\Delta\ln q_{ijht}}=-\hat{\varepsilon}_{ijht}\widehat{\Delta\ln p_{ijht}}$ and $\widehat{\Delta\ln r_{ijht}}=-(1-\hat{\varepsilon}_{ijht})\widehat{\Delta\ln p_{ijht}}$, where $\hat{\varepsilon}_{ijht}$ is the match-specific residual demand elasticity computed from observed shares and estimated parameters.
+Table D.9 in Appendix D.4 reports the results for quantities.
+Across specifications, tariff increases are associated with declines in traded volumes in both the data and the model, with larger effects under more demanding fixed effects.
+Interaction coefficients with buyer and supplier shares are imprecisely estimated in the data and sensitive to specification, suggesting limited power to detect heterogeneous effects.
+By contrast, the model predicts systematic patterns.
+Specifically, quantity responses become less negative at higher buyer shares, reflecting the smaller price increases faced by high-share buyers.
+Appendix Table D.10 extends the IV tests to quantities and sales jointly.
+Consistent with our price results, the baseline and decreasing-returns specifications provide the best fit to the data, although these tests formally reject all model variants.
+This comparatively weaker fit for quantities does not necessarily undermine the model’s mechanism.
+Quantities are measured with significantly greater error in customs data, and the mapping from prices to quantities relies on estimated demand elasticities that may amplify this noise.
+Despite these measurement challenges, the results remain qualitatively consistent with the model’s core implication: prices govern the allocation of trade, with quantities adjusting along the demand curve.
+5 Aggregate Pass-Through in Firm-to-Firm Trade
+We use the estimated model to quantify the price impact of the 2018 tariffs within continuing buyer-supplier pairs and its underlying mechanisms.
+To aggregate the model’s micro-level predictions, we compute model-implied bilateral price changes and estimate a weighted pass-through for this subset of trade by regressing predicted price changes on the tariff shock, weighting observations by their initial import shares.
+We then use the model’s expression for cost and markup elasticities to decompose predicted price changes into its two underlying determinants.
+For each channel, we recompute counterfactual price changes while holding the other component fixed, and re-estimate the weighted regression.
 Table 7: Aggregate Tariff Pass-Through and Decomposition
@@
 (1) (2)
-Panel A: Aggregate Passthrough ()
-Agg. pass-through elasticity 70.7 64.6
-Cost channel only: $1/(1+\Lambda_{ij})$ 69.3 62.7
-Markup channel only: $1/(1+\Gamma_{ij})$ 97.1 88.0
+Panel A: Aggregate Pass-through (% )
+Agg. pass-through: $1/(1+\Lambda_{ij}+\Gamma_{ij})$ 72.7 66.5
+  Cost channel only: $1/(1+\Lambda_{ij})$ 71.5 64.8
 Panel B: Variance Decomposition of $\Lambda_{ij}+\Gamma_{ij}$
-Cost Elasticity: $\Lambda_{ij}$ 1.01 0.99
-Markup Elasticity: $\Gamma_{ij}$0.01 0.01
+Cost Elasticity: $\Lambda_{ij}$ 1.023 0.997
+Markup Elasticity: $\Gamma_{ij}$ -0.023 0.003
 Notes: This table reports model-implied aggregate pass-through estimates following the 2018 U.S.
 tariff increases.
-Column (1) includes product–time and country–sector fixed effects.
-Column (2) includes product–time, country–time, and buyer–time fixed effects.
+Column (1) includes product-time and country-sector fixed effects.
+Column (2) includes product-time, country-time, and buyer-time fixed effects.
 The overall pass-through elasticity is computed as $1+$ the estimated coefficient on $\Delta\ln(1+\tau)$, and decomposed into contributions from the cost channel ($\Lambda_{ij}$) and the markup channel ($\Gamma_{ij}$).
-The counterfactual “Cost channel only” row shows the predicted pass-through when markup elasticities are set to zero, while the “Markup channel only” row sets cost elasticities to zero.
+The counterfactual “Cost channel only” row shows the predicted pass-through when markup elasticities are set to zero.
 Panel B reports the relative contribution of each channel to the cross-sectional variance of $\Lambda_{ij}+\Gamma_{ij}$, the total elasticity governing pass-through.
@@
 Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
-Panel A of Table 6 reports the aggregate tariff pass-through predicted by the model, which incorporates both markup and marginal cost adjustments: 71% under the baseline specification (Column (1)) and 65% under the alternative using product–time, country–time, and buyer–time fixed effects (Column (2)).
-These estimates are lower than the average pass-through of 75% obtained from unweighted regressions (Table 5), reflecting the more incomplete pass-through associated with larger, high-share buyers who account for a greater share of aggregate trade.
-Nearly all of the predicted price response is driven by the cost channel.
-When markup elasticities are shut down (i.e., setting $\Gamma_{ij}=0$), the model yields pass-through rates between 69 and 63%, closely matching the full-model estimates.
-By contrast, shutting down the cost elasticities results in much higher pass-through, ranging from 88 to 97%.
-Panel B offers an alternative approach to quantify the contribution of each channel by decomposing the variance of the term $\Lambda_{ij}+\Gamma_{ij}$, which governs the pass-through elasticity $1/(1+\Lambda_{ij}+\Gamma_{ij})$.
-This decomposition attributes the cross-sectional variation in pass-through to cost and markup elasticities, with their shares summing to one by construction.
-The results reinforce the earlier findings: nearly all of the variation is explained by the cost elasticity $\Lambda_{ij}$, underscoring its dominant role in shaping price responses.
-Together, the results suggest a clear conclusion: tariff pass-through is incomplete, reflecting the combined effects of cost-side adjustments and bargaining.
-The underlying mechanism is intuitive.
-When tariffs reduce demand from large U.S.
-buyers, exporters move down their marginal cost curves, dampening the price impact of the shock.
-Although markups adjust endogenously, the opposing forces of strategic complementarities and substitutabilities limit their quantitative contribution to aggregate price change.
-This is a robust implication of our framework.
-As illustrated in Figure 2, when returns to scale are decreasing ($\theta<1$) and buyer power is high (large $\phi$), most of the variation in pass-through across the $(s_{ij},\,x_{ij})$ space is driven by cost adjustments, with a comparatively smaller role for markup responses.
-5.3 The Importance of Bargaining in Firm-to-Firm Trade
-Although bargaining plays a central role in our model, it appears to have limited influence on tariff pass-through, which primarily reflects cost-side adjustments.
-This is not because bargaining is unimportant, but because it endogenously mutes the markup channel.
-When buyer power is strong ($\phi$ high), the markup elasticity $(\Gamma_{ij})$ tends toward zero.
-As a result, pass-through is driven almost entirely by cost adjustments and appears relatively unresponsive to markup variation.
-In this sense, the weak role of the markup channel in shaping pass-through is itself a consequence of strong importer market power.
-In addition, bargaining remains essential to understand price levels, markups, and the welfare consequences of market power.
-As shown in Table 2, two-sided market power is critical to explaining the cross-sectional variation in prices and markups.
-Moreover, as discussed in Section 2.5, $\phi$ governs how markup dispersion maps into misallocation.
-When exporters hold bargaining power, markup heterogeneity leads to underproduction.
-But when bargaining power shifts toward importers, the same dispersion can improve efficiency by reallocating output toward lower-cost suppliers.
-Ignoring bargaining would therefore yield misleading conclusions about the allocative effects of firm-to-firm trade.
+Panel A of Table 7 reports aggregate pass-through predicted by the model, alongside a counterfactual that isolates the cost channel by shutting down markup adjustments.
+The model predicts aggregate pass-through around 73% under the baseline specification (Column (1)) and 67% under the specification with product–time, country–time, and buyer–time fixed effects (Column (2)), somewhat below the unweighted average of 78% reported in Column (2) of the top panel of Table 6, reflecting the greater weight placed on large, high-share buyers, which exhibit more incomplete pass-through.
+Marginal cost adjustments account for most of the aggregate response.
+Shutting down markup elasticities ($\Gamma_{ij}=0$) yields pass-through rates between 71.5% and 64.8%, very close to the aggregate estimates.
+This indicates that the cost channel accounts for most of the aggregate response.
+Markup adjustments operate in the opposite direction, exerting only a small offsetting effect and therefore raising pass-through only slightly relative to the cost-only benchmark.
+Panel B provides a complementary decomposition by analyzing the variance of $\Lambda_{ij}+\Gamma_{ij}$, which governs pass-through through $1/(1+\Lambda_{ij}+\Gamma_{ij})$.
+By construction, the contributions of cost and markup elasticities sum to one.
+Consistent with Panel A, nearly all cross-sectional variation in aggregate pass-through is accounted for by the cost elasticity $\Lambda_{ij}$, with markup elasticities playing only a minor role.
+5.1 Comparison with Existing Pass-Through Estimates
+Several influential studies document near-complete pass-through of the 2018 U.S.
+tariffs at the product level (e.g., fajgelbaum2020return; amiti2019impact; amiti2020s).
+Within our baseline sample of repeated firm-to-firm relationships, the implied average (unweighted) pass-through is bounded between 81 and 87 percent, substantially incomplete and closely matched by the model (see Appendix Tables 6 and D.7).
+[^32] In a specification comparable to those in the literature, Column (1) of Table 3 yields a duty-exclusive pass-through elasticity of 0.13-0.13 (s.e.
+0.09); a more stringent set of fixed effects yields 0.19-0.19 (Appendix Table D.2).
+Two features distinguish our setting from product-level studies.
+First, our baseline imposes sample restrictions: we focus on repeated, arm’s-length links trading intermediate and capital goods, with suppliers serving at least two buyers.
+Second, we measure prices at the buyer-supplier–product level and at annual frequency, whereas product-level studies aggregate across relationships and use monthly data.
+To assess how each contributes to the gap with the literature, we evaluate them in turn.
+Sample restrictions.
+To isolate the role of sample composition, we progressively relax our baseline restrictions while keeping the unit of observation at the relationship level.
+Appendix Table D.11 shows the results.
+Including consumption goods raises pass-through from 82% to 88%; allowing single-buyer suppliers raises it further to about 92%; and the broadest specification, which also restores related-party trade, energy goods, and extreme price changes, reaches 95%.
+Each restriction contributes, but none on its own closes the gap with the literature: within the sample of repeated supplier–buyer links, pass-through remains incomplete regardless of which other restrictions are relaxed.
+Aggregation.
+To isolate the role of aggregation, we aggregate our baseline firm-to-firm prices to the product–source-country level and use monthly rather than annual data, following the literature.
+Appendix D.5 provides more details on the data construction, and Table D.12 reports the results.
+Within our baseline sample of repeated links, pass-through remains incomplete even at the product level, confirming that aggregation alone does not close the gap.
+Similarly, pass-through remains incomplete even once all sample restrictions are lifted at the product level (Columns (1)–(4)).
+The remaining difference is accounted for by the final step: once we add non-consecutive relationships (Column (5)), that is, transactions not observed in both years, pass-through reaches completeness, matching the literature.
+The near-complete pass-through in product-level studies is thus driven primarily by the inclusion of non-consecutive, one-off relationships.
+Within the population of ongoing buyer-supplier links, pass-through is systematically incomplete.
+Understanding aggregate pass-through therefore requires accounting for the extensive margin of trade, that is, how the set of active trading relationships responds to tariffs, beyond pricing within existing relationships.
 6 Conclusions
-Firm-to-firm relationships are a central feature of international trade.
-These relationships often involve market power on both sides, with prices determined through bilateral negotiations rather than market-clearing conditions.
-We develop a framework that departs from standard models of price-taking buyers by allowing both importers and exporters to influence price formation through bargaining.
-This approach yields analytical and empirical tools to study how market power and network frictions shape prices, markups, and the transmission of shocks in firm-to-firm trade.
-Using transaction-level data, we show that U.S.
-importers wield substantial bargaining power and face upward-sloping residual supply, consistent with oligopsonistic behavior.
-Within ongoing firm-to-firm relationships, tariff pass-through during the 2018 trade war was incomplete, with exporters absorbing much of the tariff through cost-side adjustments.
-This muted price response is primarily driven by cost-side adjustments: while markups do adjust, their contribution to pass-through is limited because strong buyer power endogenously flattens the markup elasticity.
-In this sense, the limited role of markups in shaping tariff responses is itself a consequence of strong importer power.
-Although not the primary focus of our analysis, the results highlight the importance of bargaining for understanding the allocative implications of market power in firm-to-firm trade.
-In our model, when supplier power dominates, markup dispersion reflects inefficiencies and underproduction.
-By contrast, when buyer power is strong, the same dispersion can lead to efficient reallocation and higher aggregate output.
-These findings suggest that markup heterogeneity does not uniformly signal misallocation, and that its welfare consequences depend critically on the distribution of bargaining power.
-Extending the framework to study these welfare implications more formally remains a promising avenue for future work.
-More broadly, our study offers a foundation for analyzing how market power shapes the incidence and transmission of shocks through the price channel.
-While we focus on detailed price-setting within firm-to-firm relationships, Acemoglu and Tahbaz-Salehi (2025) emphasize the general equilibrium implications of market power in production networks in shaping aggregate fluctuations.
-A promising direction for future research is to integrate these perspectives by combining micro-level bargaining dynamics with macro-level spillovers to study the broader implications of buyer and supplier power, including in domestic supply chains.
+We develop and estimate a structural model of two sided market power in firm-to-firm trade to shed light on tariff pass-through in global supply chains.
+The framework features oligopoly power on the exporter side and oligopsony power on the importer side, and yields closed-form expressions for bilateral markups and pass-through as functions of two sufficient statistics: the exporter’s supplier share and the importer’s buyer share, facilitating the mapping from theory to data, structural estimation, and the decomposition of pass-through mechanisms.
+The two primitives governing oligopsony power, namely the inverse export supply elasticity and bargaining power, also govern the determinants of pass-through: the inverse export supply elasticity determines whether pass-through reflects cost or markup adjustments, while bargaining power shapes the nature of strategic markup adjustment.
+The framework therefore nests as special cases both the competitive benchmark under decreasing returns and the standard oligopoly pass-through model under constant returns.
+Applied to U.S.
+imports in the 2018 tariff episode, our estimates reveal substantial oligopsony power and steep export supply curves.
+At these parameters, the cost channel dominates, explaining why pass-through within continuing buyer-supplier relationships is substantially incomplete and declines systematically with buyer share.
+This specific result likely reflects the relationship specificity and supply-side frictions that characterize U.S.
+firm-to-firm trade.
+In other settings, with weaker relationship specificity or flatter export supply, oligopsony power is lower and the framework accommodates a larger role for strategic markup adjustment.
+Incomplete pass-through through the cost channel is thus a feature of short-run pricing in ongoing firm-to-firm relationships.
+Over longer horizons, as firms reallocate across partners and capacity constraints relax, effective returns to scale rise toward constant and the markup channel prevails, in line with the firm-level pass-through literature that attributes incomplete pass-through to strategic markup adjustment (e.g., amiti2014importers).
+Even there, our framework provides a richer characterization of how bargaining power shapes pass-through across bilateral relationships.
+Sample composition matters too: product-level studies pool repeated arm’s-length links with one-off and newly formed relationships, and we show that this extensive-margin variation accounts for the bulk of the near-complete pass-through documented at the product level (amiti2019impact; fajgelbaum2020return).
+Our analysis is explicitly partial equilibrium.
+While we show that this approach yields accurate predictions of pass-through within firm-to-firm relationships, a full account of tariff incidence requires accounting for extensive-margin adjustments in the set of active relationships, as well as general-equilibrium responses in wages and export prices.
+Whether the cost-channel dominance we document at the intensive margin survives once these forces are taken into account is a key open question for future research.
 Appendix A Mathematical Derivations
-This appendix derives the key equations from the main text: the bilateral markup $\mu_{ij}$ (equation (2.6)) and the bilateral pass-through elasticity $\Phi_{ij}$ (equation (2.11)).
 A.1 Proof of Proposition 1
@@
 Solving the outer CES aggregator for total input demand yields:
-$q_{j}^{f}$ $=\gamma c_{j}q_{j}\left(p_{j}^{f}\right)^{-1},$ (A.1)
-$q_{j}^{d}$ $=(\varrho-\gamma)c_{j}q_{j}\left(p_{j}^{d}\right)^{-1},$ (A.2)
-where $c_{j}$ is the unit cost of output $q_{j}$, given by:
-$c_{j}=\left[\varphi_{j}^{-1}\left(\frac{p_{j}^{f}}{\gamma}\right)^{\gamma}\left(\frac{p_{j}^{d}}{\varrho-\gamma}\right)^{\varrho-\gamma}\right]^{\frac{1}{\varrho}}q_{j}^{\frac{1-\varrho}{\varrho}}\equiv k_{j}q_{j}^{\frac{1-\varrho}{\varrho}}.$ (A.3)
+$q_{j}^{f}$ $=\gamma MC_{j}q_{j}\left(p_{j}^{f}\right)^{-1},$ (A.1)
+$q_{j}^{d}$ $=(\varrho-\gamma)MC_{j}q_{j}\left(p_{j}^{d}\right)^{-1},$ (A.2)
+where $MC_{j}$ is the unit cost of output $q_{j}$, given by:
+$MC_{j}=\left[\varphi_{j}^{-1}\left(\frac{p_{j}^{f}}{\gamma}\right)^{\gamma}\left(\frac{p_{j}^{d}}{\varrho-\gamma}\right)^{\varrho-\gamma}\right]^{\frac{1}{\varrho}}q_{j}^{\frac{1-\varrho}{\varrho}}\equiv k_{j}q_{j}^{\frac{1-\varrho}{\varrho}}.$ (A.3)
 This form illustrates symmetry with exporter-side technology. Finally, note that the cost share of foreign inputs is constant and equal to
 $\frac{p_{j}^{f}q_{j}^{f}}{p_{j}^{f}q_{j}^{f}+p_{j}^{d}q_{j}^{d}}=\frac{\gamma}{\varrho}.$
-A relevant object for our derivations will be $\frac{d\ln c_{j}}{d\ln p_{j}^{f}},$ namely, the elasticity of the marginal cost $c_{j}$ with respect to $p_{j}^{f}.$ To find this elasticity, we first use the demand function downstream to write $q_{j}$ as
-$q_{j}=\left(\frac{\nu}{\nu-1}\right)^{-\nu}c_{j}^{-\nu}D_{j},$ (A.4)
+A relevant object for our derivations will be $\frac{d\ln MC_{j}}{d\ln p_{j}^{f}},$ namely, the elasticity of the marginal cost $MC_{j}$ with respect to $p_{j}^{f}.$ To find this elasticity, we first use the demand function downstream to write $q_{j}$ as
+$q_{j}=\left(\frac{\nu}{\nu-1}\right)^{-\nu}MC_{j}^{-\nu}D_{j},$ (A.4)
 where $D_{j}$ is the firm-level demand shifter. Substituting equation (A.4) into equation (A.3) and rearranging, we can write:
-$c_{j}$ $=\left[z_{j}^{-1}\left(\frac{p_{j}^{f}}{\gamma}\right)^{\gamma}\left(\frac{p_{j}^{d}}{\varrho-\gamma}\right)^{(\varrho-\gamma)}\right]^{\frac{1}{\varrho+\nu-\nu\varrho}}\left(\frac{\nu}{\nu-1}\right)^{-\nu\frac{1-\varrho}{\varrho+\nu-\nu\varrho}}\left(D_{j}\right)^{\frac{1-\varrho}{\varrho+\nu-\nu\varrho}},$
+$MC_{j}$ $=\left[z_{j}^{-1}\left(\frac{p_{j}^{f}}{\gamma}\right)^{\gamma}\left(\frac{p_{j}^{d}}{\varrho-\gamma}\right)^{(\varrho-\gamma)}\right]^{\frac{1}{\varrho+\nu-\nu\varrho}}\left(\frac{\nu}{\nu-1}\right)^{-\nu\frac{1-\varrho}{\varrho+\nu-\nu\varrho}}\left(D_{j}\right)^{\frac{1-\varrho}{\varrho+\nu-\nu\varrho}},$
 which implies
-$\frac{d\ln c_{j}}{d\ln p_{j}^{f}}=\frac{\gamma}{\varrho+\nu-\nu\varrho}.$
+$\frac{d\ln MC_{j}}{d\ln p_{j}^{f}}=\frac{\gamma}{\varrho+\nu-\nu\varrho}.$
 Armed with these equations, we proceed to find the elasticity of interest. Given the log demand:
 $\ln q_{ij}=\ln q_{j}^{f}+\rho\ln\varsigma_{ij}-\rho\left(\ln p_{ij}-\ln p_{j}^{f}\right),$
-and equation (A.1), we find equation (2.9):
+and equation (A.1), we find:
 $\varepsilon_{ij}=-\frac{d\ln q_{ij}}{d\ln p_{ij}}$ $=-\left(\frac{d\ln q_{j}^{f}}{d\ln p_{j}^{f}}+\rho\right)\frac{d\ln p_{j}^{f}}{d\ln p_{ij}}+\rho$
 $=\left((\nu-1)\frac{d\ln c_{j}}{d\ln p_{j}^{f}}+1-\rho\right)s_{ij}+\rho$
-$=\eta s_{ij}^{f}+(1-s_{ij}^{f})\rho,$ (A.5)
+$=\eta s_{ij}+(1-s_{ij})\rho,$ (A.5)
 where we defined
@@
 Equilibrium Price
-The problem of the $i-j$ pair is to choose a bilateral price $p_{ij}$ that solves the following problem:
+The problem of the iji-j pair is to choose a bilateral price $p_{ij}$ that solves the following problem:
 $\max_{p}\Big(\underbrace{\pi_{i}\left(p\right)-\tilde{\pi}_{i\left(-j\right)}}_{GFT_{ij}^{i}\left(p\right)}\Big)^{1-\phi}\Big(\underbrace{\pi_{j}\left(p\right)-\tilde{\pi}_{j\left(-i\right)}}_{GFT_{ij}^{j}\left(p\right)}\Big)^{\phi},$ (A.6)
@@
 The FOC associated with the problem (A.6) can be written as:
-00 0 $=\frac{d\ln\pi_{i}}{d\ln p_{ij}}+\frac{\phi}{1-\phi}\cdot\frac{GFT_{ij}^{i}}{\pi_{i}}\frac{\pi_{j}}{GFT_{ij}^{j}}\cdot\frac{d\ln\pi_{j}}{d\ln p_{ij}},$ (A.7)
+$0$ $=\frac{d\ln\pi_{i}}{d\ln p_{ij}}+\frac{\phi}{1-\phi}\cdot\frac{GFT_{ij}^{i}}{\pi_{i}}\frac{\pi_{j}}{GFT_{ij}^{j}}\cdot\frac{d\ln\pi_{j}}{d\ln p_{ij}},$ (A.7)
 where we used the fact that $\frac{dGFT_{ij}^{k}}{dp}=\frac{d\pi_{k}}{dp}$ for $k=\{i,j\}.$ In what follows, we derive expressions for $\frac{d\ln\pi_{i}}{d\ln p_{ij}},\frac{d\ln\pi_{j}}{d\ln p_{ij}},GFT_{ij}^{i}$ and $GFT_{ij}^{j}.$
 Exporter $i$’s Profits and Gains from Trade–Firm $i$’s profit under a successful negotiation can be expressed as
-$\pi_{i}=p_{ij}q_{ij}+\sum_{k\neq j}p_{ik}q_{ik}-\theta c_{i}q_{i},$
-where $c_{i}=k_{i}q_{i}^{\frac{1-\theta}{\theta}}.$ The elasticity of the profit $\pi_{i}$ with respect to $p_{ij}$ can be found as:
-$\frac{d\ln\pi_{i}}{d\ln p_{ij}}$ $=\frac{q_{ij}}{\pi_{i}}(\varepsilon_{ij}-1)\left(\mu_{ij}^{\text{oligopoly}}c_{i}-p_{ij}\right),$ (A.8)
+$\pi_{i}=p_{ij}q_{ij}+\sum_{k\neq j}p_{ik}q_{ik}-\theta MC_{i}q_{i},$
+where $MC_{i}=k_{i}q_{i}^{\frac{1-\theta}{\theta}}.$ The elasticity of the profit $\pi_{i}$ with respect to $p_{ij}$ can be found as:
+$\frac{d\ln\pi_{i}}{d\ln p_{ij}}$ $=\frac{q_{ij}}{\pi_{i}}(\varepsilon_{ij}-1)\left(\mu_{ij}^{\text{oligopoly}}MC_{i}-p_{ij}\right),$ (A.8)
 where $\varepsilon_{ij}\equiv-\frac{d\ln q_{ij}}{d\ln p_{ij}}$ is defined in equation (A.5) and
@@
 The “outside” profit of firm $i$ are
-$\tilde{\pi}_{i\left(-j\right)}=\sum_{k\neq j}p_{ik}q_{ik}-\theta\widetilde{c_{i}q_{i}}.$
+$\tilde{\pi}_{i\left(-j\right)}=\sum_{k\neq j}p_{ik}q_{ik}-\theta\widetilde{MC_{i}q_{i}}$.
 For a constant $k_{i}$ and given $q_{i}=\sum_{j\in\mathcal{Z}_{i}}q_{ij}$, the total cost in case of failed agreement can be found as:
-$\theta\widetilde{c_{i}q_{i}}$ $=\theta c_{i}q_{i}\left(\frac{\widetilde{q_{i}}}{q_{i}}\right)^{\frac{1}{\theta}}.$
-$=\theta c_{i}q_{i}\left(1-x_{ij}\right)^{\frac{1}{\theta}}.$
+$\theta\widetilde{MC_{i}q_{i}}$ $=\theta MC_{i}q_{i}\left(\frac{\widetilde{q_{i}}}{q_{i}}\right)^{\frac{1}{\theta}}.$
+$=\theta MC_{i}q_{i}\left(1-x_{ij}\right)^{\frac{1}{\theta}}.$
 The exporter’s gains from trade $GFT_{ij}^{i}$ are thus given by:
 $GFT_{ij}^{i}\left(p_{ij}\right)$ $\equiv\pi_{i}\left(p\right)-\tilde{\pi}_{i\left(-j\right)}$ (A.9)
-$=q_{ij}\left(p_{ij}-c_{i}\mu_{ij}^{\text{oligopsony}}\right),$
+$=q_{ij}\left(p_{ij}-MC_{i}\mu_{ij}^{\text{oligopsony}}\right),$
 where $\mu_{ij}^{\text{oligopsony}}:=\theta\left[\frac{\Delta_{ij}^{x}}{x_{ij}}\right]$ and $\Delta_{ij}^{x}:=\left[1-\left(1-x_{ij}\right)^{\frac{1}{\theta}}\right].$
 Importer $j$’s Profits and Gains from Trade–Firm $j$’s profit under a successful negotiation can be expressed as
-$\pi_{j}=p_{j}q_{j}-\varrho c_{j}q_{j}=\frac{\gamma}{\eta-1}c_{j}q_{j},$
-where $c_{j}$ is defined in equation (A.3), and $q_{j}$ in equation (A.4). The elasticity of the profit $\pi_{j}$ with respect to $p_{ij}$ can be found as:
-$-\frac{d\ln\pi_{j}}{d\ln p_{ij}}=(\eta-1)s_{ij}.$ (A.10)
+$\pi_{j}=p_{j}q_{j}-\varrho MC_{j}q_{j}=\frac{\gamma}{\eta-1}MC_{j}q_{j},$
+where $MC_{j}$ is defined in equation (A.3), and $q_{j}$ in equation (A.4). The elasticity of the profit $\pi_{j}$ with respect to $p_{ij}$ can be found as:
+$\frac{d\ln\pi_{j}}{d\ln p_{ij}}=(\eta-1)s_{ij}.$ (A.10)
 The outside profit of firm $j$ under a failed negotiation is
-$\tilde{\pi}_{j\left(-i\right)}=\frac{\gamma}{\eta-1}\widetilde{c_{j}q_{j}},$
-where $\varrho\widetilde{c_{j}q_{j}}$ denotes the total “outside” cost of importer $j$. The assumptions on technology and demand downstream imply that we can write:
-$\widetilde{c_{j}q_{j}}$ $=\left(1-s_{ij}\right)^{\frac{\eta-1}{\rho-1}}c_{j}q_{j}.$
+$\tilde{\pi}_{j\left(-i\right)}=\frac{\gamma}{\eta-1}\widetilde{MC_{j}q_{j}}$,
+where $\varrho\widetilde{MC_{j}q_{j}}$ denotes the total “outside” cost of importer $j$. The assumptions on technology and demand downstream imply that we can write:
+$\widetilde{MC_{j}q_{j}}$ $=\left(1-s_{ij}\right)^{\frac{\eta-1}{\rho-1}}MC_{j}q_{j}.$
 Putting things together, the importer’s gains from trade can be written as:
@@
 We’re now ready to solve for the bilateral price $p_{ij}.$ Substituting equations (A.8), (A.9), (A.10), and (A.11) into the FOC in equation (A.7), we obtain:
-$\mu_{ij}:=\frac{p_{ij}}{c_{i}}=(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}+\omega_{ij}\mu_{ij}^{\text{oligopsony}}$ (A.12)
+$\mu_{ij}:=\frac{p_{ij}}{MC_{i}}=(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}+\omega_{ij}\mu_{ij}^{\text{oligopsony}}$ (A.12)
 where the weighting factor is
@@
 A.2 Proof of Proposition 2
-Proposition 2 characterizes how the bilateral markup $\mu_{ij}$ co-moves with the exporter’s and importer’s bilateral market shares, $s_{ij}$ and $x_{ij}$.
-Part (i): $\mu_{ij}$ and Exporter’s supplier share $s_{ij}$
-To study how $\mu_{ij}$ responds to the exporter’s supplier share $s_{ij}$, note that the oligopoly markup $\mu^{\text{oligopoly}}_{ij}$ is strictly increasing in $s_{ij}$ under standard assumptions ($\eta<\rho<\infty$).
-Differentiating the overall markup $\mu_{ij}$ in equation (A.12) with respect to $s_{ij}$ yields:
-$\frac{\partial\mu_{ij}}{\partial s_{ij}}=(1-\omega_{ij})\cdot\underbrace{\frac{\partial\mu^{\text{oligopoly}}_{ij}}{\partial s_{ij}}}_{>0}+\frac{\partial\omega_{ij}}{\partial s_{ij}}\cdot\underbrace{\left(\mu^{\text{oligopsony}}_{ij}-\mu^{\text{oligopoly}}_{ij}\right)}_{<0}.$
-The first term is strictly positive when $\omega_{ij}<1$ (i.e., $\phi<1$), reflecting the direct effect of $s_{ij}$ on the oligopoly markup.
-The second term captures how changes in $s_{ij}$ affect the bargaining weight $\omega_{ij}$ and thus the relative influence of the oligopsony markdown.
-Although $\omega_{ij}$ is hump-shaped in $s_{ij}$, increasing at low values and decreasing at high values, its sensitivity to $s_{ij}$ is limited in most of the parameter space.
-As a result, $\frac{\partial\omega_{ij}}{\partial s_{ij}}$ is typically small, and the first (positive) term generally dominates.
-In contrast, when $\phi=1$ (full importer bargaining power), the bilateral markup reduces to the oligopsony markdown: $\mu_{ij}=\mu^{\text{oligopsony}}_{ij}$, and $\omega_{ij}=1$. In this case,
-$\frac{\partial\mu_{ij}}{\partial s_{ij}}=0.$
-Thus, we have:
-$\frac{\partial\mu_{ij}}{\partial s_{ij}}=\begin{cases}\gtrless 0&\text{if }\phi<1\text{ (typically positive)},\\ =0&\text{if }\phi=1.\end{cases}$
-Thus, $\mu_{ij}$ can increase with $s_{ij}$ in our theory only when $\phi<1$, which proves the first part of Proposition 1.
-Part (ii): $\mu_{ij}$ and Importer’s buyer share $x_{ij}$
-The effect of $x_{ij}$ on $\mu_{ij}$ is limited to the oligopsony markdown term and is therefore easier to characterize. Differentiating the bilateral markup with respect to $x_{ij}$ yields:
-$\frac{\partial\mu_{ij}}{\partial x_{ij}}=\omega_{ij}\cdot\underbrace{\frac{\partial\mu^{\text{oligopsony}}_{ij}}{\partial x_{ij}}}_{<0\ \text{if}\ \theta<1}.$
-The markdown $\mu^{\text{oligopsony}}_{ij}$ decreases with $x_{ij}$ when marginal cost is increasing (i.e., $\theta<1$), as larger buyers elicit stronger cost reductions and thus negotiate lower prices.
-This implies:
-$\frac{\partial\mu_{ij}}{\partial x_{ij}}<0\iff\omega_{ij}>0\text{ and }\theta<1.$
-Therefore, observing a negative slope of $\mu_{ij}$ with respect to the importer’s share $x_{ij}$ provides evidence of both importer bargaining power ($\phi>0$ so that $\omega_{ij}>0$) and decreasing returns to scale ($\theta<1$).
-This completes the proof of Proposition 2.
+Proposition 2 characterizes how the bilateral markup $\mu_{ij}$ varies with bilateral market shares.
+Part I states that a positive relationship between $\mu_{ij}$ and the exporter’s supplier share $s_{ij}$ requires $\phi<1$.
+Part II states that a negative relationship between $\mu_{ij}$ and the importer’s buyer share $x_{ij}$ can arise only if $\phi>0$ and $\theta<1$.
+Proof of Part I
+The Nash bargaining solution implies that the bilateral markup satisfies
+$\mu_{ij}=(1-\omega_{ij})\mu^{\text{oligopoly}}(s_{ij})+\omega_{ij}\mu^{\text{oligopsony}}(x_{ij}),$ (A.15)
+with terms defined above. Differentiating with respect to $s_{ij}$ yields
+$\frac{\partial\mu_{ij}}{\partial s_{ij}}=(1-\omega_{ij})\frac{\partial\mu^{\text{oligopoly}}_{ij}}{\partial s_{ij}}+\frac{\partial\omega_{ij}}{\partial s_{ij}}\bigl(\mu^{\text{oligopsony}}_{ij}-\mu^{\text{oligopoly}}_{ij}\bigr).$ (A.16)
+Under the maintained assumption $\eta<\rho$, the oligopoly markup is strictly increasing in supplier share, so $\partial_{s}\mu^{\text{oligopoly}}_{ij}>0$, while the oligopsony markdown is strictly lower than the oligopoly markup, $\mu^{\text{oligopsony}}_{ij}-\mu^{\text{oligopoly}}_{ij}<0$.
+Consider first the case $\phi=1$, which implies $\omega_{ij}=1$ for all $s_{ij}$. Hence, $(1-\omega_{ij})=0$ and $\partial\omega_{ij}/\partial s_{ij}=0$. Both terms in (A.16) therefore vanish, so that
+$\frac{\partial\mu_{ij}}{\partial s_{ij}}=0\qquad\text{when }\phi=1.$
+Thus, a strictly positive relationship between $\mu_{ij}$ and $s_{ij}$ cannot arise under full importer bargaining power.
+Now suppose $\phi<1$.
+In this case $\omega_{ij}(s_{ij})$ is hump-shaped in $s_{ij}$ with $\omega_{ij}(0)=\omega_{ij}(1)=\phi$, implying that there exists $\bar{s}\in(0,1)$ such that $\partial\omega_{ij}/\partial s_{ij}<0$ for all $s_{ij}>\bar{s}$.
+For such values of $s_{ij}$, the second term in (A.16) is nonnegative, since it is the product of two negative terms.
+Together with $(1-\omega_{ij})>0$ and $\partial_{s}\mu^{\text{oligopoly}}_{ij}>0$, this implies
+$\frac{\partial\mu_{ij}}{\partial s_{ij}}>0\qquad\text{for all }s_{ij}>\bar{s}.$
+For $s_{ij}<\bar{s}$, the sign of $\partial\mu_{ij}/\partial s_{ij}$ depends on the relative strength of the two terms in (A.16).
+In particular, $\partial\mu_{ij}/\partial s_{ij}>0$ if and only if the direct oligopoly effect dominates the bargaining-weight effect, i.e.,
+$(1-\omega_{ij})\frac{\partial\mu^{\text{oligopoly}}_{ij}}{\partial s_{ij}}>\left|\frac{\partial\omega_{ij}}{\partial s_{ij}}\bigl(\mu^{\text{oligopsony}}_{ij}-\mu^{\text{oligopoly}}_{ij}\bigr)\right|>0.$
+We conclude that $\partial\mu_{ij}/\partial s_{ij}>0$ can arise only when $\phi<1$. This establishes Part I of Proposition 2. ∎
+Proof of Part II.
+From (A.15), the bilateral markup $\mu_{ij}$ depends on the importer’s buyer share $x_{ij}$ only through the oligopsony markdown $\mu^{\text{oligopsony}}_{ij}(x_{ij})$.
+Differentiating with respect to $x_{ij}$:
+$\frac{\partial\mu_{ij}}{\partial x_{ij}}=\omega_{ij}\frac{\partial\mu^{\text{oligopsony}}_{ij}}{\partial x_{ij}}.$
+If $\phi=0$, then $\omega_{ij}=0$ for all $x_{ij}$, implying $\partial\mu_{ij}/\partial x_{ij}=0$. Hence, a negative relationship between $\mu_{ij}$ and $x_{ij}$ cannot occur if $\phi=0$.
+For $\phi>0$, the sign of $\partial_{x}\mu_{ij}$ is equal to the sign of $\partial_{x}\mu^{\text{oligopsony}}_{ij}$. This is given by:
+$\frac{\partial\mu^{\text{oligopsony}}_{ij}}{\partial x_{ij}}=\frac{x_{ij}(1-x_{ij})^{\frac{1}{\theta}-1}-\theta\bigl[1-(1-x_{ij})^{\frac{1}{\theta}}\bigr]}{x_{ij}^{2}}.$
+We establish the following lemma.
+Lemma 1.
+For any $x\in(0,1)$ and $\theta>0$,
+$\operatorname{sign}\!\left(x(1-x)^{\frac{1}{\theta}-1}-\theta\bigl[1-(1-x)^{\frac{1}{\theta}}\bigr]\right)=\operatorname{sign}(\theta-1).$
+Proof.
+Let $t=(1-x)^{1/\theta}\in(0,1)$. Then the expression in the lemma can be written as
+$t^{1-\theta}-t-\theta+t\theta.$
+Express the above as a function of $\theta$, $f\left(\theta\right)$.
+One can show that $f\left(\theta\right)$ is strictly convex in $\theta$ ($\partial^{2}f\left(\theta\right)/\partial\theta^{2}>0$), and $f\left(0\right)=f\left(1\right)=0$.
+This implies
+[…]
+with equality at $\theta=1$. ∎
+By the lemma, $\partial\mu^{\text{oligopsony}}_{ij}/\partial x_{ij}<0$ if and only if $\theta<1$ since $x_{ij}^{2}>0$ for all $x\in(0,1]$. Since $\omega_{ij}>0$ when $\phi>0$, this implies
+$\frac{\partial\mu_{ij}}{\partial x_{ij}}<0\quad\Longleftrightarrow\quad\phi>0\text{ and }\theta<1.$
+Therefore, observing a negative relationship between the bilateral markup and the importer’s buyer share implies positive importer bargaining power and decreasing returns to scale, establishing Part II of Proposition 2.
 ∎
@@
 The log (tariff-inclusive) price is given by:
-$\ln p_{ij}=\ln\mu_{ij}+\ln c_{i}+\ln T_{c},$
-where $c_{i}$ and $\mu_{ij}$ are as in equations (2.1) and (2.6), respectively.
+$\ln p_{ij}=\ln\mu_{ij}+\ln MC_{i}+\ln T_{c},$
+where $MC_{i}$ and $\mu_{ij}$ are as in equations (2.1) and (2.10), respectively.
 Taking a full log-differential and rearranging terms yields:
@@
 $\frac{d\ln p_{ij}}{d\ln T_{c}}$ $=$ $\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}},$
-where $\Gamma_{ij}\equiv-\frac{d\ln\mu_{ij}}{d\ln p_{ij}}$ and $\Lambda_{ij}\equiv-\frac{d\ln c_{i}}{d\ln p_{ij}}$ are the partial markup and cost elasticities, respectively.
+where $\Gamma_{ij}\equiv-\frac{d\ln\mu_{ij}}{d\ln p_{ij}}$ and $\Lambda_{ij}\equiv-\frac{d\ln MC_{i}}{d\ln p_{ij}}$ are the partial markup and cost elasticities, respectively.
 The Cost Elasticity
 Taking the logarithm of equation (2.1), we obtain:
-$\ln c_{i}$ $=\ln k_{i}+\frac{1-\theta}{\theta}\ln q_{i}.$
+$\ln MC_{i}$ $=\ln k_{i}+\frac{1-\theta}{\theta}\ln q_{i}.$
 It immediately follows that:
-$\Lambda_{ij}\equiv-\frac{d\ln c_{i}}{d\ln p_{ij}}$ $=\frac{1-\theta}{\theta}\frac{d\ln q_{i}}{d\ln q_{ij}}\left(-\frac{d\ln q_{ij}}{d\ln p_{ij}}\right)$
+$\Lambda_{ij}\equiv-\frac{d\ln MC_{i}}{d\ln p_{ij}}$ $=\frac{1-\theta}{\theta}\frac{d\ln q_{i}}{d\ln q_{ij}}\left(-\frac{d\ln q_{ij}}{d\ln p_{ij}}\right)$
 $=\frac{1-\theta}{\theta}\cdot x_{ij}\cdot\varepsilon_{ij}\geq 0.$
@@
 $\quad-\left(\frac{x_{ij}(1-x_{ij})^{\frac{1}{\theta}-1}}{\theta\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]}-1\right)\Bigg]\varepsilon_{ij}$
-[…]
+$={}$ $\left[\frac{\left((1-x_{ij})^{\frac{1-\theta}{\theta}}-\frac{1-\theta}{\theta}x_{ij}(1-x_{ij})^{\frac{1}{\theta}-2}\right)\theta\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]-x_{ij}(1-x_{ij})^{\frac{2}{\theta}-2}}{\theta^{2}\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]^{2}}\right](1-x_{ij})$
 $\quad-\left(\frac{x_{ij}(1-x_{ij})^{\frac{1}{\theta}-1}}{\theta\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]}-1\right)$
-[…]
+$={}$ $\left[\left(\frac{1-\frac{1-\theta}{\theta}\cdot\frac{x_{ij}}{1-x_{ij}}}{\theta\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]}-\frac{(1-x_{ij})^{\frac{1}{\theta}-1}}{\mu_{ij}^{\text{oligopsony}}\theta\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]}\right)(1-x_{ij})^{\frac{1}{\theta}}\right$.
 $\quad\left.-\left(\frac{x_{ij}(1-x_{ij})^{\frac{1}{\theta}-1}}{\theta\left[1-(1-x_{ij})^{\frac{1}{\theta}}\right]}-1\right)\right]$
@@
 A.4 Proof of Proposition 4
-We now prove Proposition 4, which characterizes the equilibrium comovement between pass-through $\Phi_{ij}$ and the bilateral market shares $x_{ij}$ and $s_{ij}$.
-Comparative Statics with Respect to $x_{ij}$
-We begin by analyzing how pass-through $\Phi_{ij}$ responds to the importer’s buyer share $x_{ij}$.
-As discussed in the main text, $\Phi_{ij}$ depends on $x_{ij}$ through two channels: (i) the oligopsony markdown elasticity $\Gamma_{ij}^{\text{oligopsony}}$, which is non-monotonic (U-shaped) in $x_{ij}$, and (ii) the cost elasticity $\Lambda_{ij}$, which increases linearly in $x_{ij}$ when $\theta<1$.
-When $\theta=1$, both elasticities are zero: $\Gamma_{ij}^{\text{oligopsony}}=\Lambda_{ij}=0$. Therefore, pass-through is constant in $x_{ij}$:
-$\left.\frac{\partial\Phi_{ij}}{\partial x_{ij}}\right|_{\theta=1}=0.$
-When $\theta<1$, however, both elasticities are active. The pass-through elasticity can be written as:
-$\Phi_{ij}=\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}}=\frac{1}{1+\cdots+\omega_{ij}\cdot\frac{\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\cdot\Gamma_{ij}^{\text{oligopsony}}+\Lambda_{ij}},$
-where $\cdots$” denotes terms that do not depend on $x_{ij}$. Differentiating with respect to $x_{ij}$:
-$\frac{\partial\Phi_{ij}}{\partial x_{ij}}=-\frac{1}{(1+\Gamma_{ij}+\Lambda_{ij})^{2}}\cdot\frac{\partial}{\partial x_{ij}}\left(\omega_{ij}\cdot\frac{\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\cdot\Gamma_{ij}^{\text{oligopsony}}+\Lambda_{ij}\right).$
-Let:
-$T_{1}\equiv\omega_{ij}\cdot\frac{\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\cdot\Gamma_{ij}^{\text{oligopsony}},\quad T_{2}\equiv\Lambda_{ij}.$
-We can expand these as:
-$T_{1}$ $=\omega_{ij}\cdot\left(\frac{(1-x_{ij})^{\frac{1}{\theta}-1}-\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\right)(1-x_{ij})\varepsilon_{ij},$
-$T_{2}$ $=\frac{1-\theta}{\theta}\cdot x_{ij}\cdot\varepsilon_{ij}.$
-While $T_{1}$ is non-monotonic in $x_{ij}$, $T_{2}$ increases linearly in $x_{ij}$ and dominates $T_{1}$ for all $x_{ij}$. This can be verified numerically; for instance, when $\theta=\frac{1}{2}$:
-$T_{1}+T_{2}=x_{ij}\cdot\varepsilon_{ij}\left(1-\frac{\omega_{ij}}{\mu_{ij}}\cdot\left(\frac{5}{2}-\frac{3}{2}x_{ij}\right)\right),$
-which increases in $x_{ij}$ as the term in parentheses remains positive.
-Hence, the term in the numerator of $\frac{\partial\Phi_{ij}}{\partial x_{ij}}$ increases in $x_{ij}$, implying:
-$\operatorname{sign}\left(\frac{\partial\Phi_{ij}}{\partial x_{ij}}\right)=-\operatorname{sign}\left(\frac{\partial\Lambda_{ij}}{\partial x_{ij}}\right),$
-so that:
-$\frac{\partial\Phi_{ij}}{\partial x_{ij}}\leq 0.$
-This proves that, under $\theta<1$, pass-through decreases in $x_{ij}$, a robust and testable implication of the model.
-Comparative Statics with Respect to $s_{ij}$
-We now turn to how pass-through $\Phi_{ij}$ varies with the exporter’s supplier share $s_{ij}$.
-Unlike the case of $x_{ij}$, this relationship is more complex, as all components, namely, $\Gamma_{ij}^{\text{oligopoly}}$, $\Gamma_{ij}^{\text{oligopsony}}$, $\Gamma_{ij}^{\omega}$, and $\Lambda_{ij}$, may vary with $s_{ij}$.
-Specifically,
-$\Phi_{ij}=\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}}=\frac{1}{1+\frac{(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopoly}}+\frac{\omega_{ij}\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopsony}}+\left(1-\frac{\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\right)\Gamma_{ij}^{\omega}+\Lambda_{ij}}.$
-Differentiating with respect to $s_{ij}$ gives:
-$\frac{\partial\Phi_{ij}}{\partial s_{ij}}$ $=-\frac{\partial}{\partial s_{ij}}\bigg{(}\frac{(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopoly}}+\frac{\omega_{ij}\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopsony}}$
-$\quad+\left(1-\frac{\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\right)\Gamma_{ij}^{\omega}+\Lambda_{ij}\bigg{)}\cdot\left(1+\Gamma_{ij}+\Lambda_{ij}\right)^{-2}.$
-When $\theta=1$, both $\Gamma_{ij}^{\text{oligopsony}}=0$ and $\Lambda_{ij}=0$, and the expression simplifies to:
-$\frac{\partial\Phi_{ij}}{\partial s_{ij}}=-\frac{\partial}{\partial s_{ij}}\left(\frac{(1-\omega_{ij})\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopoly}}+\left(1-\frac{\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\right)\Gamma_{ij}^{\omega}\right)\cdot(1+\Gamma_{ij})^{-2}.$
-To build intuition, consider the limit $s_{ij}\to 1$, where $\omega_{ij}\to\phi$ and $\Gamma_{ij}^{\omega}\to 0$. Then:
-$\frac{\partial\Phi_{ij}}{\partial s_{ij}}\propto\begin{cases}-\frac{\partial}{\partial s_{ij}}\left(\frac{(1-\phi)\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopoly}}+\frac{\phi\mu_{ij}^{\text{oligopsony}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopsony}}+\Lambda_{ij}\right)&\text{if }\theta<1,\\ -\frac{\partial}{\partial s_{ij}}\left(\frac{(1-\phi)\mu_{ij}^{\text{oligopoly}}}{\mu_{ij}}\Gamma_{ij}^{\text{oligopoly}}\right)&\text{if }\theta=1.\end{cases}$
-The term involving $\Gamma_{ij}^{\text{oligopoly}}$ is hump-shaped in $s_{ij}$, while the remaining components, particularly $\Lambda_{ij}$, decline monotonically with $s_{ij}$ due to $\frac{\partial\varepsilon_{ij}}{\partial s_{ij}}=-(\rho-\eta)<0$.
-Quantitatively, these latter terms dominate in most parameter ranges.
-Thus, we have:
-$\operatorname{sign}\left(\frac{\partial\Phi_{ij}}{\partial s_{ij}}\right)\propto\begin{cases}(+)&\text{if }\theta<1\text{ or }s_{ij}\text{ sufficientlylarge},\\ (-)&\text{if }\theta=1\text{ and }s_{ij}\text{ sufficiently small}.\end{cases}$
-Due to the non-monotonicity and interaction of multiple channels, the relationship between $\Phi_{ij}$ and $s_{ij}$ does not yield a clean, general prediction.
-While pass-through may increase with $s_{ij}$ under decreasing returns to scale and sufficient buyer power, the sign of this relationship depends on the relative strength of strategic complementarities, oligopsony markdowns, and cost elasticity.
-As a result, we omit this relationship from Proposition 3 and focus instead on the more robust prediction involving $x_{ij}$.
-Appendix B Theory: Discussion and Extensions
+Proposition 4 states that if the tariff pass-through elasticity $\Phi_{ij}$ decreases with the importer’s buyer share $x_{ij}$, then the exporter’s technology exhibits decreasing returns to scale, i.e.
+$\theta<1$.
+Proof.
+The bilateral pass-through elasticity is
+$\Phi_{ij}=\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}},$
+where $\Gamma_{ij}$ and $\Lambda_{ij}$ denote the markup and marginal cost elasticities, respectively.
+Since $\partial\Phi_{ij}/\partial x_{ij}$ has the opposite sign of $\partial(\Gamma_{ij}+\Lambda_{ij})/\partial x_{ij}$, it suffices to show that
+$\partial(\Gamma_{ij}+\Lambda_{ij})/\partial x_{ij}>0\quad\text{for}\quad\theta<1$
+to prove the claim.
+We first consider the case $\theta=1$.
+Under constant returns to scale, the exporter’s marginal cost is constant and $\Lambda_{ij}=0$.
+Moreover, the oligopsony markdown equals one, implying $\Gamma^{\text{oligopsony}}_{ij}=0$.
+As $\Gamma_{ij}$ depends on $x_{ij}$ only through $\Gamma^{\text{oligopsony}}_{ij}$, it follows that $\partial\Gamma_{ij}/\partial x_{ij}=0$ and hence $\partial\Phi_{ij}/\partial x_{ij}=0$.
+We now consider $\theta<1$.
+The cost elasticity is
+$\Lambda_{ij}=\frac{1-\theta}{\theta}x_{ij}\varepsilon_{ij},$
+which is strictly increasing in $x_{ij}$ and independent of the bargaining parameter $\phi$.
+The markup elasticity depends on $x_{ij}$ only through the oligopsony component $\Gamma^{\text{oligopsony}}_{ij}$.
+Varying $\phi$ therefore only scales the contribution of this term.
+Consequently, if
+$\left.\frac{\partial}{\partial x_{ij}}(\Gamma_{ij}+\Lambda_{ij})\right|_{\phi=1}>0,$
+where the contribution of $x$ is maximized, then the same must hold for all $\phi\in[0,1]$. It is therefore sufficient to study the case $\phi=1$.
+When $\phi=1$, the oligopsony markdown elasticity can be written as
+$\Gamma^{\text{oligopsony}}_{ij}=f(x_{ij})\,\varepsilon_{ij},\qquad f(x)=\left(\frac{x(1-x)^{\frac{1}{\theta}-1}}{\theta\left[1-(1-x)^{\frac{1}{\theta}}\right]}-1\right)(1-x).$
+Hence,
+$\frac{\partial}{\partial x_{ij}}\bigl(\Gamma^{\text{oligopsony}}_{ij}+\Lambda_{ij}\bigr)=\varepsilon_{ij}\left[f^{\prime}(x_{ij})+\frac{1-\theta}{\theta}\right].$
+The function $f$ is smooth on $(0,1)$ and U-shaped: $f^{\prime}(x)<0$ for small $x$, $f^{\prime}(x)>0$ for large $x$, and $f^{\prime}(x)$ is strictly increasing in $x$.
+Thus the minimum of $f^{\prime}(x)$ over $(0,1)$ is attained at the boundary as $x\to 0$.
+A first-order expansion around $x=0$ yields
+$f(x)=\frac{\theta-1}{2\theta}x+O(x^{2}),\qquad\lim_{x\to 0}f^{\prime}(x)=\frac{\theta-1}{2\theta}.$
+When $\theta<1$, this implies
+$f^{\prime}(x)\geq-\frac{1-\theta}{2\theta}>-\frac{1-\theta}{\theta}\qquad\forall x\in(0,1).$
+Therefore,
+$f^{\prime}(x_{ij})+\frac{1-\theta}{\theta}>0\quad\forall x_{ij}\in(0,1),$
+and hence
+$\frac{\partial}{\partial x_{ij}}\bigl(\Gamma_{ij}+\Lambda_{ij}\bigr)>0\qquad\forall\phi\in[0,1]\quad\&\quad\theta<1.$
+Appendix B Theory: Extensions
 B.1 Efficient Bargaining
-In the efficient bargaining setup, the importer and exporter negotiate over a two-part tariff $(p,q)$ by maximizing a Nash product of generalized firm-specific gains from trade (GFT):
-$\max_{p,q}\left[GFT_{ij}^{i}(p,q)\right]^{1-\phi}\left[GFT_{ij}^{j}(p,q)\right]^{\phi}\quad\text{s.t.}\quad GFT_{ij}^{i}\geq 0,\quad GFT_{ij}^{j}\geq 0$
-where:
-$GFT_{ij}^{i}(p_{ij},q_{ij})$ $\equiv\pi_{i}(p)-\tilde{\pi}_{i(-j)}=p_{ij}q_{ij}-\left(\theta c_{i}q_{i}-\theta\widetilde{c_{i}q_{i}}\right),$
-$GFT_{ij}^{j}(p_{ij},q_{ij})$ $\equiv\pi_{j}(p)-\tilde{\pi}_{j(-i)}=\left(p_{j}q_{j}-\tilde{p}_{j}\tilde{q}_{j}\right)-p_{ij}q_{ij}.$
-This formulation corresponds to a setting in which the firm pair first selects the input quantity $q_{ij}$ to maximize joint surplus, and then negotiates over the price $p_{ij}$ to determine how surplus is split.
+In the efficient bargaining setup, the importer and exporter jointly choose the price $p_{ij}$ and quantity $q_{ij}$ by maximizing the Nash product:
+$\max_{p_{ij},\,q_{ij}}\left(GFT_{ij}^{i}\right)^{1-\phi}\left(GFT_{ij}^{j}\right)^{\phi}\quad\text{s.t.}\quad GFT_{ij}^{i}\geq 0,\quad GFT_{ij}^{j}\geq 0,$
+where $GFT_{ij}^{k}\equiv\pi^{k}_{ij}-\tilde{\pi}^{k}_{(-\cdot)}$, $k\in\{i,j\}$, are as defined in Section 2.
+This corresponds to a setting in which the firm pair first selects $q_{ij}$ to maximize joint surplus, and then negotiates over $p_{ij}$ to split it.
+The first-best quantity is determined where the exporter’s marginal cost equals the importer’s marginal revenue product
+$q_{ij}^{*}:MR_{j}(q^{*}_{ij})=MC_{i}(q^{*}_{ij}).$
+Given $q_{ij}^{*}$, the price $p_{ij}$ solves:
+$\max_{p_{ij}}\left(p_{ij}q^{*}_{ij}-\Delta C_{i}(q^{*}_{ij})\right)^{1-\phi}\left(\Delta R_{j}(q^{*}_{ij})-p_{ij}q^{*}_{ij}\right)^{\phi},$
+where $\Delta R_{j}(q_{ij})\equiv R_{j}(q_{j})-\tilde{R}_{j}(\tilde{q}_{j})$ and $\Delta C_{i}(q_{ij})\equiv TC_{i}(q_{i})-TC_{i}(\tilde{q}_{i})$ are the incremental revenue for the importer and the incremental cost for the exporter from the match.
 Figure B.1: Efficient Bargaining
-The efficient quantity solves:
-$q_{ij}^{*}$ $:\quad\arg\max_{q}\left[GFT_{ij}^{i}(q)+GFT_{ij}^{j}(q)\right]$
-$:\quad\frac{d(p_{j}q_{j})}{dq_{ij}}=\frac{d(\theta c_{i}q_{i})}{dq_{ij}}$
-That is, the efficient quantity $q^{*}_{ij}$ equates the exporter’s marginal cost to the importer’s marginal revenue, similar to the outcome under vertical integration. Solving yields:
-$q_{ij}^{*}:\quad p_{ij}=c_{i}(q_{ij}^{*}),$
-so that the negotiated price equals the exporter’s marginal cost.
-Given $q_{ij}^{*}$, the price $p_{ij}$ solves the Nash bargaining problem:
-$\max_{p_{ij}}\left(p_{ij}q^{*}_{ij}-\Delta C_{i}(q^{*}_{ij})\right)^{1-\phi}\left(\Delta R_{j}(q^{*}_{ij})-p_{ij}q^{*}_{ij}\right)^{\phi},$
-where $\phi\in(0,1)$ denotes the importer’s bargaining power.
-The terms $\Delta R_{j}(q_{ij})$ and $\Delta C_{i}(q_{ij})$ represent the additional revenue for the importer and the additional cost for the exporter attributable to the match:
-$\Delta R_{j}(q_{ij})$ $=p_{j}q_{j}-\widetilde{p_{j}q_{j}},$
-$\Delta C_{i}(q_{ij})$ $=\theta c_{i}q_{i}-\theta\widetilde{c_{i}q_{i}}.$
 Standard derivations lead to the equilibrium price:
 $p_{ij}=(1-\phi)\cdot\frac{\Delta R_{j}(q_{ij})}{q_{ij}}+\phi\cdot\frac{\Delta C_{i}(q_{ij})}{q_{ij}}$
-$=(1-\phi)\cdot MC_{i}(q^{*}_{ij})+\phi\cdot AC(q^{*}_{ij})$
 a weighted average of the per-unit downstream revenue gain and per-unit upstream cost increase from the match, with bargaining weights given by $\phi$.
-Figure B.1 illustrates the set of feasible equilibria. The efficient quantity is $q^{*}$, while the negotiated price ranges between $MC(q^{*})$ and $AC(q^{*})$ depending on $\phi$.
+Figure B.1 illustrates this result.
+The efficient quantity $q^{*}$ is determined where the exporter’s marginal cost $\mathrm{MC}(q_{i})$ equals the importer’s marginal revenue product $p(q_{i})$.
+At $q^{*}$, the negotiated price lies between $\Delta C_{i}/q^{*}=\mathrm{AC}(q^{*})$ and $\Delta R_{j}/q^{*}=p(q^{*})$.
+The weight $\phi$ determines the exact price within this interval: as $\phi\to 0$ the price approaches $\Delta R_{j}/q^{*}$, and as $\phi\to 1$ it approaches $\Delta C_{i}/q^{*}$.
 B.2 Supply-Driven Quantity Bargaining
@@
 $p_{ij}(q_{ij})=k_{i}q_{i}^{\frac{1-\theta}{\theta}},\quad\text{where}\quad q_{i}=q_{i(-j)}+q_{ij},$
-as in equation (2.1). Solving the bargaining problem using derivations similar to those in Section A.1 of this Appendix yields the following expression for the bilateral price:
-$p_{ij}=(1-\omega_{ij}^{S})\cdot c_{i}+\omega_{ij}^{S}\cdot\psi_{ij}^{-1}MRP_{j},$
+as in equation (2.1).
+Since $p_{ij}=MC_{i}(q_{ij})$ holds on the supply curve, solving the Nash bargaining problem yields the following first-order condition characterizing the equilibrium quantity (and hence price):
+$\psi_{ij}\,p_{ij}-MRP_{j}\;=\;\frac{1-\omega_{ij}^{S}}{\omega_{ij}^{S}}\,(\psi_{ij}-1)\,p_{ij},$
 where
-$\psi_{ij}=1+c^{\prime}_{i,q_{ij}}=1+\frac{1-\theta}{\theta}x_{ij}.$
-Here, $\omega_{ij}^{S}$ is given by:
-$\omega_{ij}^{S}\equiv\frac{\frac{\phi}{1-\phi}\cdot\lambda^{S}_{ij}}{1+\frac{\phi}{1-\phi}\cdot\lambda^{S}_{ij}}\quad\text{and}\quad\lambda^{S}_{ij}\equiv\frac{GFT_{ij}^{i}}{GFT_{ij}^{j}}\psi_{ij}=\frac{p_{ij}q_{ij}-\theta c_{i}q_{i}\Delta_{ij}^{x}}{p_{j}q_{j}\Delta_{ij}^{s}-p_{ij}q_{ij}}\psi_{ij}.$
+$\psi_{ij}=1+c^{\prime}_{i,q_{ij}}=1+\frac{1-\theta}{\theta}x_{ij},$
+and $\omega_{ij}^{S}$ has the same functional form as $\omega_{ij}$ in Proposition 1:
+$\omega_{ij}^{S}\equiv\frac{\phi\,\psi_{ij}\,GFT_{ij}^{i}}{(1-\phi)\,GFT_{ij}^{j}+\phi\,\psi_{ij}\,GFT_{ij}^{i}}\in(0,1),$
+where $GFT_{ij}^{i}=p_{ij}q_{ij}-\theta MC_{i}q_{i}\Delta_{ij}^{x}$ and $GFT_{ij}^{j}=p_{j}q_{j}\Delta_{ij}^{s}-p_{ij}q_{ij}$ are as in Section 2.
+The formula has the same structure as $\omega_{ij}$ in Proposition 1: $\omega^{S}_{ij}$ is a ratio of the importer’s weighted gains $\phi\,\psi_{ij}\,GFT^{i}_{ij}$ to total weighted gains, where the supply-slope factor $\psi_{ij}$ plays the role of $1+e_{i}$ (the labor-supply elasticity factor) in their model.
+The key difference from the baseline is that $\omega^{S}_{ij}$ depends on equilibrium gains-from-trade levels rather than on observable market shares alone, which is why the equilibrium condition does not reduce to a closed-form convex combination.
+For $\phi=0$, $\omega_{ij}^{S}=0$: the exporter captures the full importer surplus ($GFT_{ij}^{j}=0$), and the equilibrium price is pinned by $\Delta R_{j}(q_{ij})=p_{ij}\,q_{ij}$.
+For $\phi=1$, $\omega_{ij}^{S}=1$: the first-order condition collapses to $\psi_{ij}\,p_{ij}=MRP_{j}$, recovering the standard oligopsony benchmark where the input price is a markdown factor $\psi_{ij}^{-1}$ below the importer’s marginal revenue product.
+Here $\psi_{ij}=1+$ the inverse of the importer’s residual supply elasticity, reflecting the cost of raising the unit price on all infra-marginal units when the supply curve is upward-sloping.
 Figure B.2: Supply-Driven Quantity Bargaining
-Hence, the bilateral price is a convex combination of the exporter’s marginal cost ($c_{i}$) and a term reflecting the markdown ($\psi_{ij}$) below the importer’s marginal revenue product ($MRP_{j}\equiv\frac{dp_{j}q_{j}}{dq_{ij}}$).
-As in the demand-driven quantity bargaining case, the weight $\omega_{ij}^{S}$ has an intuitive interpretation as the importer’s effective bargaining power.
-For $\phi=0$, $\omega_{ij}^{S}=0$, and the allocation is efficient with $p_{ij}=c_{i}$.
-For $\phi=1$, $\omega_{ij}^{S}=1$, and we obtain the standard oligopsony benchmark where the input price is a markdown below the importer’s marginal revenue product, with the markdown equal to one plus the importer’s residual supply elasticity.
-This reflects the importer’s view, where the cost of increasing quantity is driven by raising the unit price for all inframarginal units when supply curves are upward-sloping.
-Figure B.2 plots the set of feasible allocations. The red line indicates the pure oligopsony markdown, and the pink area represents the set of markdowns for $\phi\in(0,1)$.
-While this case is intuitive, it is more complex to characterize due to the term $\lambda^{S}_{ij}$, which is itself a function of prices and quantities and thus no longer a simple function of market shares as in the baseline case.
-As a result, bringing this model to the data is less straightforward than our baseline model.
+Figure B.2 illustrates the equilibrium.
+The blue curve is the exporter’s marginal cost $MC_{i}$ (equivalently, the supply curve), and the black curve is the importer’s marginal revenue product $MRP_{j}$.
+The red curve is the importer’s marginal factor cost $\psi_{ij}\,MC_{i}$, which exceeds $MC_{i}$ whenever $\theta<1$.
+At $\phi=1$ the equilibrium lies at point A, where $MRP_{j}=\psi_{ij}\,MC_{i}$: the importer acts as a pure monopsonist, restricting quantity below the efficient level $q^{*}$ (where $MC_{i}=MRP_{j}$).
+As $\phi$ decreases, bargaining power shifts toward the exporter, moving the equilibrium up the supply curve toward $q^{*}$.
+The shaded region on the supply curve indicates the feasible price range for $\phi\in(0,1)$.
+A key implication of this protocol is that prices and quantities positively co-move in equilibrium.
+Because the equilibrium always lies on the exporter’s supply curve $p_{ij}=MC_{i}(q_{ij})$, which is upward-sloping under decreasing returns ($\theta<1$), any shock that increases the equilibrium quantity also raises the equilibrium price.
+This stands in sharp contrast to the demand-driven baseline, where the equilibrium lies on the importer’s downward-sloping demand schedule and price increases are accompanied by quantity reductions.
+The data support the demand-driven prediction: tariff-induced price increases are accompanied by quantity declines ($p\uparrow,\,q\downarrow$).
 B.3 Generalized Outside Option
-In the baseline model, we assume that the importer’s (exporter’s) gains from trade are given by the firm’s total payoff from trading with all partners, minus the payoff from trading with all partners except exporter $i$ (importer $j$).
-In simpler terms, each importer (exporter) treats itself as the marginal buyer (supplier).
-This implies that, in case of disagreement, neither party considers the possibility of forming new relationships, which substantially influences the definition of outside profits for both importers and exporters.
-We now consider a more general specification that imposes less structure on outside options. Let
-$\Delta_{ij}^{c_{i}}=\frac{\tilde{c}_{i}}{c_{i}}\qquad\text{and}\qquad\Delta_{ij}^{c_{j}}=\frac{\tilde{c}_{j}}{c_{j}}$
+In the baseline model, the importer’s (exporter’s) gains from trade are the difference between its payoff when trading with all current partners and its payoff when trading with all partners except exporter $i$ (importer $j$).
+This rules out two forces that may matter in richer environments: after a failed negotiation, firms may form or reallocate relationships, and broader equilibrium adjustments may shift disagreement payoffs by changing costs or demand elsewhere in the network.
+We now consider a more general specification that places less structure on outside options and allows these effects to enter through disagreement payoffs.
+For tractability, we assume $\varrho=1$ throughout this section (constant returns to scale for the importer), so that $\eta-1=(\nu-1)\gamma$.
+Let
+$\Delta_{ij}^{MC_{i}}=\frac{\widetilde{MC}_{i}}{MC_{i}}\qquad\text{and}\qquad\Delta_{ij}^{c_{j}}=\frac{\widetilde{MC}_{j}}{MC_{j}}$
 denote the percentage change in exporter $i$’s and importer $j$’s marginal cost in the event of a failed negotiation.
 Under this generalization, the gains from trade for firms $i$ and $j$ are given by:
-$GFT_{ij}^{i}(p_{ij})$ $=q_{ij}\left(p_{ij}-c_{i}\mu_{ij}^{\text{oligopsony},G}\right),$
+$GFT_{ij}^{i}(p_{ij})$ $=q_{ij}\left(p_{ij}-MC_{i}\mu_{ij}^{\text{oligopsony},G}\right),$
 $GFT_{ij}^{j}(p_{ij})$ $=\pi_{j}\Delta_{ij}^{s,G},$
 where we define:
-$\mu_{ij}^{\text{oligopsony},G}$ $:=\theta\left[\frac{\Delta_{ij}^{x}}{x_{ij}}\right],\qquad\Delta_{ij}^{x}:=\left[\frac{1-\Delta_{ij}^{c_{i}}(1-x_{ij})}{x_{ij}}\right],$
+$\mu_{ij}^{\text{oligopsony},G}$ $:=\theta\left[\frac{\Delta_{ij}^{x}}{x_{ij}}\right]\qquad\text{with}\qquad\Delta_{ij}^{x}:=\left(1-\Delta_{ij}^{MC_{i}}(1-x_{ij})\right)$
+and
 $\Delta_{ij}^{s,G}:=1-\left(\Delta_{ij}^{c_{j}}\right)^{1-\nu}.$
 The first-order condition under this generalized outside option implies:
-$\mu_{ij}:=\frac{p_{ij}}{c_{i}}=(1-\omega^{G}_{ij})\mu_{ij}^{\text{oligopoly}}+\omega^{G}_{ij}\mu_{ij}^{\text{oligopsony},G},$ (B.1)
+$\mu_{ij}:=\frac{p_{ij}}{MC_{i}}=(1-\omega^{G}_{ij})\mu_{ij}^{\text{oligopoly}}+\omega^{G}_{ij}\mu_{ij}^{\text{oligopsony},G},$ (B.1)
 where the weighting factor is given by:
-$\omega^{G}_{ij}$ $:=\frac{\frac{\phi}{1-\phi}\lambda_{ij}^{G}}{1+\frac{\phi}{1-\phi}\lambda_{ij}^{G}}\in(0,1),$
-$\lambda_{ij}^{G}$ $:=\frac{\gamma s_{ij}}{\Delta_{ij}^{s,G}}\cdot\frac{\nu-1}{\varepsilon_{ij}-1}>0.$
-Equation (B.1) shares the structure of the markup equation in equation (2.6), but with two key differences.
-First, the term $\lambda_{ij}$ now depends on $\Delta_{ij}^{s,G}$, which cannot be expressed solely as a function of the supplier share.
-Second, the oligopsony markdown depends on $\Delta_{ij}^{x}$, which likewise cannot be written as a function of the buyer share alone.
-To summarize, generalizing the structure of outside options preserves the overall structure of the markup equation and yields similar comparative statics.
-However, it introduces an identification problem: the effective bargaining power and oligopsony markdown now depend on unobserved terms, such as $\Delta_{ij}^{s,G}$ and $\Delta_{ij}^{x}$, that cannot be expressed solely as functions of market shares.
-In the absence of external data to estimate these terms, empirical implementation becomes infeasible without further assumptions.
-Imposing the more restrictive baseline assumption enables us to express both components as functions of observable bilateral market shares and a small set of structural parameters, allowing us to implement the model using available data and avoid excessive computational complexity.
+$\omega^{G}_{ij}=\frac{\phi\,\xi^{G}(GFT_{ij}^{j},p_{ij})}{\phi\,\xi^{G}(GFT_{ij}^{j},p_{ij})+(1-\phi)(\varepsilon_{ij}-1)}\in(0,1).$
+Here
+$\xi^{G}(GFT_{ij}^{j},p_{ij})\equiv-\frac{d\ln GFT_{ij}^{j}}{d\ln p_{ij}}=\frac{(\eta-1)s_{ij}}{\Delta_{ij}^{s,G}}$
+is the generalized elasticity of the importer’s gains from trade with respect to the bilateral price.
+This expression is equivalent to the baseline elasticity up to the denominator.
+In the baseline model, that denominator is pinned down by the supplier share alone.
+Here, instead, it also depends on how the importer’s disagreement payoff changes after a failed negotiation, as summarized by $\Delta_{ij}^{s,G}$.
+Since $\Delta_{ij}^{s,G}$ reflects unobserved changes in marginal cost and outside options, effective bargaining power is no longer identified from $s_{ij}$ alone.
+Equation (B.1) shares the weighted-average structure of equation (2.10), but with two key differences.
+First, the generalized elasticity $\xi^{G}(GFT_{ij}^{j},p_{ij})$ now depends on $\Delta_{ij}^{s,G}$, so effective bargaining power is no longer pinned down by the supplier share alone.
+Second, the oligopsony markdown depends on $\Delta_{ij}^{x}$, so it is no longer determined by the buyer share alone.
+The main implication is that outside options can absorb both extensive-margin reallocation and broader equilibrium forces, but once they do, the pricing equation depends on unobserved objects.
+In particular, $\Delta_{ij}^{s,G}$ and $\Delta_{ij}^{x}$ are not recoverable from bilateral market shares alone.
+Without additional data or stronger assumptions, the model is therefore not identified in a way that is useful for our empirical exercise.
+The baseline specification imposes more structure precisely to recover markup components from observable bilateral shares and a small set of parameters.
 B.4 Full Pass-Through Elasticity
@@
 This section extends the analysis to incorporate certain indirect effects by capturing how a shock to exporter $i$ influences prices and quantities in other relationships, and how those changes feed back into the bilateral price $p_{ij}$.
-While we continue to abstract from full general equilibrium forces, this exercise clarifies how network spillovers may cause reduced-form pass-through estimates to deviate from structural ones, as emphasized by Berger et al.
-(2022).
-We generalize the cost shock to an exporter-level shock, denoted $\vartheta_{i}$, and re-derive the relevant elasticities to allow for cross-relationship spillovers.
+While we continue to abstract from full general equilibrium forces, this exercise clarifies how network spillovers may cause reduced-form pass-through estimates to deviate from structural ones, as emphasized by berger2022labor.
+We consider an exporter-level price shifter, denoted $\vartheta_{i}$, and re-derive the relevant elasticities to allow for cross-relationship spillovers.
 First, the impact on the importer’s buyer share becomes:
@@
 The effect on the exporter’s marginal cost is:
-$\frac{d\ln c_{i}}{d\ln\vartheta_{i}}$ $=\frac{1-\theta}{\theta}\left(-\varepsilon_{ij}x_{ij}\frac{d\ln p_{ij}}{d\ln\vartheta_{i}}-\sum_{z\in\mathcal{Z}_{i},\,z\neq j}x_{iz}\varepsilon_{iz}\frac{d\ln p_{iz}}{d\ln\vartheta_{i}}\right).$
+$\frac{d\ln MC_{i}}{d\ln\vartheta_{i}}$ $=\frac{1-\theta}{\theta}\left(-\varepsilon_{ij}x_{ij}\frac{d\ln p_{ij}}{d\ln\vartheta_{i}}-\sum_{z\in\mathcal{Z}_{i},\,z\neq j}x_{iz}\varepsilon_{iz}\frac{d\ln p_{iz}}{d\ln\vartheta_{i}}\right).$
 We also account for the fact that $p_{ij}$ affects rival suppliers’ shares and prices. The elasticity of supplier share with respect to the shock is:
@@
 In the main text, we focus on the direct pass-through elasticity $\Phi_{ij}$, which is tightly grounded in our model’s Nash-in-Nash structure and match-level pricing assumptions.
+B.5 Buyer Power, Decreasing Returns, and Pass-Through
+We compare a perfectly competitive benchmark with decreasing returns to scale to our bilateral bargaining framework in order to clarify the source of pass-through heterogeneity.
+Consider one supplier $i$ and two buyers $j\in\{1,2\}$.
+The supplier’s marginal cost is
+$MC_{i}=k_{i}q_{i}^{\frac{1-\theta}{\theta}},$
+which implies a marginal cost elasticity
+$mc_{q}\equiv\frac{d\ln MC_{i}}{d\ln q_{i}}=\frac{1-\theta}{\theta}>0.$
+Total output is $q_{i}=q_{i1}+q_{i2}$ and buyer shares are defined as $x_{ij}\equiv q_{ij}/q_{i}$. For simplicity, both buyers have constant demand elasticity $\sigma\equiv-d\ln q_{ij}/d\ln p_{ij}$.
+Under perfect competition and in the absence of price discrimination, prices are uniform and equal to marginal cost, $p_{ij}=p_{i}=MC_{i}$.
+A tariff shock $T$ enters multiplicatively into marginal cost, so that $d\ln p_{i}=d\ln MC_{i}+d\ln T$.
+Using $d\ln MC_{i}=mc_{q}\,d\ln q_{i}$ together with $d\ln q_{i}=-\sigma\,d\ln p_{i}$, we obtain
+$d\ln p_{i}=-mc_{q}\sigma\,d\ln p_{i}+d\ln T,$
+which implies
+$\Phi^{\text{PC}}\equiv\frac{d\ln p_{i}}{d\ln T}=\frac{1}{1+mc_{q}\sigma}=\frac{1}{1+\frac{1-\theta}{\theta}\sigma}.$
+Pass-through is therefore incomplete due to decreasing returns and identical across buyers.
+In the bargaining model, prices are set bilaterally according to $p_{ij}=\mu_{ij}MC_{i}$. Pass-through becomes relationship-specific:
+$\Phi^{\text{B}}\equiv\frac{d\ln p_{ij}}{d\ln T}=\frac{1}{1+\Gamma_{ij}+\Lambda_{ij}},$
+where $\Gamma_{ij}$ captures markup adjustments and
+$\Lambda_{ij}\equiv mc_{q}\,x_{ij}\,\sigma$
+captures the elasticity of marginal cost induced by buyer $j$’s share of output.
+Quantitatively, under our estimate of $\theta$, the markup channel is dominated by the cost channel (see Figure 2), so that $\Gamma_{ij}\approx 0$. Pass-through is then well approximated by
+$\Phi^{\text{B}}\approx\frac{1}{1+\Lambda_{ij}}=\frac{1}{1+\frac{1-\theta}{\theta}\sigma x_{ij}}.$
+The two environments share the same underlying mechanism: incomplete pass-through is governed by the slope of marginal cost, $mc_{q}$.
+The difference lies in the unit of adjustment.
+Under perfect competition, the relevant elasticity is aggregate and pass-through is uniform.
+Under bargaining, the effective elasticity is share-weighted, generating heterogeneity across relationships.
+Large buyers internalize a steeper effective supply curve and exhibit lower pass-through, whereas small buyers approximate the constant-returns benchmark.
+Bargaining thus provides a micro-foundation for decreasing-returns-type pass-through while delivering the relationship-level heterogeneity observed in the data.
 Appendix C Data Appendix
@@
 A location score is computed based on the match between city names in the MID and in ORBIS.
-A product match score is computed by comparing the NAICS6 industry code listed in ORBIS to the HS6 product code recorded in the customs data, using the concordance developed by Pierce and Schott (2009).
-We retain only those matches where both location and product scores exceed 90%.
-In addition, we drop from the matched dataset any ORBIS firm with fewer than five transactions to filter out spurious exporters and potential noise.
+A product match score is constructed at the MID–firm level by comparing the foreign firm’s NAICS6 industry code in ORBIS to the distribution of HS6 products shipped under that MID in the customs data, using the concordance developed by Pierce2009.
+We retain only those matches where both the location and product match scores exceed 90%.
+In addition, we drop from the matched LFTTD–ORBIS dataset any ORBIS-linked firm that is associated with fewer than five customs transactions.
+This filter is based on the number of shipment records matched to each ORBIS identifier in the merged data and is intended to eliminate weak or potentially noisy MID–firm matches.
 Another concern is that the MID may sometimes refer to intermediaries rather than manufacturers.
@@
 Taken together, these steps yield a linked dataset that offers a more transparent and conservative definition of related-party trade, based on majority ownership (at least 50%).
-In contrast to the standard related-party trade (RPT) flag in customs data, which applies a lower threshold of 6% ownership for imports, this approach reduces false positives and more precisely captures transactions where ownership ties are likely to influence pricing.
+In contrast to the standard related-party trade (RPT) flag in customs data, which applies a lower threshold of 5% ownership for imports, this approach reduces false positives and more precisely captures transactions where ownership ties are likely to influence pricing.
 By combining MID-based matching with firm-level ownership structures, the final dataset is well-suited for analyzing pricing behavior in cross-border transactions.
 C.2 Data Cleaning, Sample Construction and Summary Statistics
-We construct the analysis sample in several steps to align the data with the model’s structure and requirements.
-We begin by removing observations that are incomplete or inconsistent with the modeling framework.
-Specifically, we drop transactions with missing or zero values for import value or quantity, invalid exporter identifiers (e.g., strings with fewer than three characters or beginning with a number), or U.S.
+We construct the analysis sample in several steps to align the data with the model and the identification strategy.
+As in the main text, it is useful to distinguish between restrictions that are essential for identification and those that primarily ensure consistency with the model’s environment.
+Preliminary restrictions.
+We begin by aggregating transaction-level customs records to the annual importer–exporter–product (HS10) level.
+Because relatively few relationships appear in adjacent months or in the same month across years, annual aggregation provides a natural unit of observation for relationship-level price changes.
+To ensure reliable price measurement, we apply several data-quality filters.
+We drop transactions with missing or zero import value or quantity, invalid exporter identifiers, or U.S.
 importers that cannot be linked to the Longitudinal Business Database (LBD).
-We also exclude transactions associated with special provisions or temporary classifications (HS codes 98–99).
-Next, we restrict our attention to trade in intermediate and capital goods by removing HS10 products, which are classified as consumption goods under the BEC system.
-Although the customs data are reported at the transaction level, we aggregate them to the annual level for each buyer–supplier–product triplet.
-We choose annual aggregation because few relationships appear in adjacent months or in the same month across years.
-We then retain only those triplets that are active in at least two consecutive years and where the supplier transacts the same HS10 product with more than one U.S.
-buyer.
-This ensures a panel structure with repeated observations, supporting the identification of the model’s parameters.
-We also exclude related-party trade, which is less likely to reflect decentralized bargaining and more likely to involve internal pricing practices such as transfer pricing.
-[^35] Bernard et al.
-(2006) shows that related-party prices differ systematically from arm’s-length transactions, including lower average prices and different pass-through behavior.
-In our baseline, a buyer–supplier pair is flagged as related if ORBIS identifies a shared corporate parent.
+We exclude HS codes 98–99 (special provisions), remove observations with log unit values below the 1st or above the 99th percentile of the HS10–country distribution, and trim extreme year-on-year log price changes outside the interval [4,4][-4,4].
+[^33] These filters are applied to the estimation sample but retained when computing tenure and relationship length.
+We also exclude energy goods (HS chapter 27), which fall outside the scope of the model.
+Identification-driven restrictions.
+First, we require importer–exporter–product triplets to be observed in at least two consecutive calendar years.
+This panel structure is necessary to compute year-on-year price changes and to estimate pass-through.
+Second, our identification strategy relies on cross-sectional price variation across buyers sourcing the same product from the same supplier.
+We therefore restrict attention to supplier–product pairs that transact with at least two distinct U.S.
+buyers in consecutive years.
+This condition is essential for identifying the bargaining and returns-to-scale parameters, as outlined in Section 4.
+Model-alignment restrictions.
+The remaining restrictions ensure that the empirical setting matches the model’s focus on decentralized bargaining over intermediate inputs.
+We exclude products classified as consumption goods under the Broad Economic Categories (BEC) system, restricting attention to capital and intermediate inputs.
+We also exclude related-party trade, as such transactions may reflect intra-firm transfer pricing rather than arm’s-length bargaining.343434bernard2006transfer document systematic differences between related-party and arm’s-length prices.
+In the baseline, a buyer–supplier pair is classified as related if ORBIS identifies a common corporate parent.
 To preserve coverage, we retain all observations not flagged as related by either ORBIS or the LFTTD.
-For robustness, we also consider two alternative definitions: one based solely on the LFTTD flag, and another combining it with ORBIS data to identify the U.S.
-importer as a multinational.
-See Appendix C.1 for further details.
-To address outliers, we follow Heise (2024) and apply two filters.
-First, we drop observations with extreme price levels, defined as log unit values below the 1st or above the 99th percentile of the HS10 product–country distribution.
-Second, we trim extreme price changes by excluding year-on-year log price differences smaller than $-4$ or greater than $+4$.
-These outliers are removed from estimation but retained when computing tenure and relationship length.
-We also exclude all HS10 products under HS chapter 27 (energy-related goods), which lie outside the scope of the model.
-After applying these restrictions, which exclude outliers, energy products, and related-party trade, and focusing on repeated, arm’s-length relationships, the sample retains over 20% of U.S.
+For robustness, we consider alternative definitions based on the LFTTD related-party flag and on multinational status (see Appendix C.1).
+After applying these restrictions, the final sample isolates repeated, arm’s-length buyer–supplier relationships in intermediate goods with sufficient cross-sectional variation for identification; it retains over 20% of U.S.
 imports by value and nearly 15% of buyer–supplier–product triplets.
-Finally, we restrict attention to suppliers that sell the same HS10 product to more than one U.S.
-buyer.
-This condition is essential for the identification of the model’s firm-level parameters and ensures that the empirical setting aligns with the model’s structure oulined in Setion 4.
 Table C.1 summarizes the cumulative impact of these steps.
@@
 Subsequent rows show the effect of each restriction in turn, including the requirement that buyer–supplier pairs trade the same product in two consecutive calendar years, the exclusion of consumption and energy goods, and the restriction to arm’s-length relationships with sufficient variation for identification.
-The final rows show the estimation samples used in the analysis.
-For 2001–2016 (Panel A), the data include approximately \$880 billion in import value, 480,000 buyer–supplier pairs, and 630,000 buyer–supplier–HS10 triplets.
-For 2017–2018 (Panel B), the final sample includes \$160 billion in imports, 190,000 pairs, and 250,000 triplets.
-[^36] Panel A omits intermediate sample steps and shows only the full and final samples for 2001–2016, as these are not reported in the analysis nor disclosed under Census data policies.
+The final rows report the estimation samples used in the analysis.
+For 2001–2016 (Panel A), the sample includes approximately \$880 billion in import value, 480,000 buyer–supplier pairs, and 630,000 buyer–supplier–HS10 triplets.
+For 2017–2018 (Panel B), the corresponding sample comprises \$160 billion in imports, 190,000 pairs, and 250,000 triplets.
+[^35] Panel A omits intermediate sample steps and reports only the full and final samples for 2001–2016, as intermediate counts are neither used in the analysis nor releasable under Census disclosure policies.
 The same applies to Table C.3.
-These samples correspond exactly to the data used in estimation and post-estimation analysis.
-Table C.2 reports a similar analysis using a broader sample that includes capital, intermediate, and consumption goods (“BEC – All (Consec.)”), rather than limiting to consumption goods only.
-Analogous to Table 1, Table C.3 reports the summary statistics of this broader sample.
+These samples correspond exactly to those used in the structural estimation and post-estimation analysis.
+Table C.2 reports summary statistics for the main estimation sample, analogous to Table 1 in the main text.
+Tables C.3 and C.4 present analogous results for a broader sample that includes capital, intermediate, and consumption goods (“BEC – All (Consec.)”), rather than restricting to non-consumption goods.
 Table C.1: Sample Composition by Period – Excluding BEC-Classified Consumption Goods
@@
 Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
+Table C.2: Summary Statistics for Main Estimation Sample (2001–2018)
+Variable Mean Std. Dev. P25 Median P75
+Panel A: Trade Relationships
+   $s_{ijh}$: Supplier share 0.32 0.35 0.03 0.15 0.57
+   $x_{ijh}$: Buyer share 0.25 0.29 0.02 0.10 0.40
+   Relationship length (product $h$) 4.00 2.80 2.50 3.50 5.50
+   Relationship length (all products) 4.80 3.30 2.50 4.50 6.50
+   # Transactions (product $h$) 120 1100 6.50 16 50
+   # Transactions (all products) 360 3000 11 36 140
+   # Products per pair 3.80 7.30 1.50 2.50 4.50
+   Multi-HS10 dummy 0.59 0.49 0.00 1.00 1.00
+   # Suppliers per buyer (HS10) 1.80 3.20 1.50 2.50 5.50
+   Buyer tenure (all products) 9.90 5.00 6.50 10.00 14.00
+   Buyer tenure (product $h$) 6.90 4.40 3.50 6.50 10.00
+   # Buyers per supplier (HS10) 3.20 3.90 2.50 3.50 7.50
+   Supplier tenure (all products) 8.00 4.60 4.50 8.50 12.00
+   Supplier tenure (product $h$) 6.40 4.00 3.50 6.50 9.50
+   Corr. between $s_{ijh}$ and $x_{ijh}$ 0.041
+Panel B: Prices (log unit values)
+   $\log p$ (pre-duty, excl. charges) 3.40 2.80 1.30 3.00 5.40
+   $\log p$ (pre-duty) 3.50 2.80 1.40 3.10 5.40
+   $\log p^{\text{duty}}$ (post-duty) 3.50 2.80 1.40 3.10 5.40
+Notes: This table extends Table 1 with additional variables for the main estimation sample at the importer-supplier-product level.
+Full details on sample construction are provided in Appendix C.2.
+Panel A reports trade relationship characteristics.
+$s_{ijh}$ is exporter $i$’s share in buyer $j$’s imports of product $h$; $x_{ijh}$ is buyer $j$’s share in exporter $i$’s U.S.
+exports of product $h$.
+Relationship length (product $h$) is the number of years between the supplier-buyer pair’s first transaction in product $h$ and year $t$; Relationship length (all products) is defined analogously, using the pair’s first transaction in any product.
+# Transactions (product $h$) is the cumulative number of transactions between the pair in product $h$, from their first transaction through year $t$; # Transactions (all products) is defined analogously, summing across all products.
+# Products per pair is the number of distinct HS10 products traded by the pair in year $t$, and Multi-HS10 dummy equals one when this number exceeds one.
+# Suppliers per buyer (HS10) is the number of distinct exporters from which buyer $j$ sources product $h$ in year $t$; # Buyers per supplier (HS10) is defined analogously, counting distinct buyers per exporter.
+Buyer tenure (product $h$) is the number of years between the buyer’s first recorded transaction in product $h$ (with any supplier) and year $t$; Buyer tenure (all products) is defined analogously, using the buyer’s first recorded transaction in any product.
+Supplier tenure is defined analogously for the exporter.
+Corr.
+between $s_{ijh}$ and $x_{ijh}$ is the sample correlation between the two shares.
+Panel B reports log unit values (value divided by quantity).
+$\log p$ (pre-duty, excl.
+charges) corresponds to the FOB (Free on Board) unit value, excluding shipping and insurance charges; $\log p$ (pre-duty) includes these charges to obtain the CIF (Cost, Insurance, and Freight) unit value; and $\log p^{\text{duty}}$ (post-duty) further incorporates import duties.
+Statistics are based on confidential LFTTD data and rounded according to U.S.
+Census Bureau disclosure guidelines.
+Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
 Table C.3: Sample Composition by Period – All BEC Categories
@@
 Variable Mean Std. Dev. P25 Median P75
-Panel A: Characteristics of Trade Relationships
+Panel A: Trade Relationships
    $s_{ijh}$: Supplier share 0.27 0.32 0.02 0.10 0.43
@@
    Corr. between $s_{ijh}$ and $x_{ijh}$ 0.053
-Panel B: Prices
+Panel B: Prices (log unit values)
+   $\log p$ (pre-duty, excl. charges) 3.30 2.50 1.40 3.00 5.00
    $\log p$ (pre-duty) 3.40 2.50 1.50 3.10 5.10
-$\log p$ (pre-duty, excl. charges) 3.30 2.50 1.40 3.00 5.00
    $\log p^{\text{duty}}$ (post-duty) 3.40 2.50 1.50 3.10 5.20
-Notes: This table reports summary statistics for a sample that covers all BEC product categories except energy goods, and excludes statistical outliers and related-party trade.
-It further restricts to suppliers that trade with at least two different U.S.
-buyers in consecutive years.
-This corresponds to the cumulative sample underlying the “+ Supplier Multi-Buyer” row in Panel B of Table C.2.
-Columns report the mean, standard deviation, and selected quantiles (25th percentile, median, and 75th percentile) for each variable.
-Prices in Panel B are log unit values (FOB value over quantity), with variants including charges or duties.
-$s_{ijh}$ denotes exporter $i$’s share in buyer $j$’s imports of product $h$; $x_{ijh}$ denotes buyer $j$’s share in exporter $i$’s U.S.
-exports of the same product.
-Relationship length and tenure are in years; concentration is measured at the HS10–year level.
-Counts of buyers, suppliers, and origin countries are per product per firm.
-Statistics are based on confidential LFTTD data and rounded to four significant digits per U.S.
+Notes: This table reports summary statistics for the pooled sample, defined analogously to the main estimation sample (Table 1) but without restricting on BEC end-use category, at the importer-supplier-product level.
+Full details on sample construction are provided in Appendix C.2.
+Panel A reports trade relationship characteristics.
+$s_{ijh}$ is exporter $i$’s share in buyer $j$’s imports of product $h$; $x_{ijh}$ is buyer $j$’s share in exporter $i$’s U.S.
+exports of product $h$.
+Relationship length (product $h$) is the number of years between the supplier-buyer pair’s first transaction in product $h$ and year $t$; Relationship length (all products) is defined analogously, using the pair’s first transaction in any product.
+# Transactions (product $h$) is the cumulative number of transactions between the pair in product $h$, from their first transaction through year $t$; # Transactions (all products) is defined analogously, summing across all products.
+# Products per pair is the number of distinct HS10 products traded by the pair in year $t$, and Multi-HS10 dummy equals one when this number exceeds one.
+# Suppliers per buyer (HS10) is the number of distinct exporters from which buyer $j$ sources product $h$ in year $t$; # Buyers per supplier (HS10) is defined analogously, counting distinct buyers per exporter.
+Buyer tenure (product $h$) is the number of years between the buyer’s first recorded transaction in product $h$ (with any supplier) and year $t$; Buyer tenure (all products) is defined analogously, using the buyer’s first recorded transaction in any product.
+Supplier tenure is defined analogously for the exporter.
+Corr.
+between $s_{ijh}$ and $x_{ijh}$ is the sample correlation between the two shares.
+Panel B reports log unit values (value divided by quantity).
+$\log p$ (pre-duty, excl.
+charges) corresponds to the FOB (Free on Board) unit value, excluding shipping and insurance charges; $\log p$ (pre-duty) includes these charges to obtain the CIF (Cost, Insurance, and Freight) unit value; and $\log p^{\text{duty}}$ (post-duty) further incorporates import duties.
+Statistics are based on confidential LFTTD data and rounded according to U.S.
 Census Bureau disclosure guidelines.
@@
 estimated over the period 2001-2016.
-We consider three alternative prices: prices that exclude both duties and charges ($\ln p_{ijht}$), prices that include charges but exclude duties ($\ln p^{c}_{ijht}$), and prices that include duties but exclude charges ($\ln p^{\text{duty}}_{ijht}$).
+We consider three alternative prices: prices that exclude duties ($\ln p_{ijht}$), prices that exclude duties and charges ($\ln p^{c}_{ijht}$), and prices that include duties and charges ($\ln p^{\text{duty}}_{ijht}$).
 Table D.1 finds that controlling for product and year fixed effects explains approximately 50% of the overall price dispersion, while 4% is attributed to match-specific residuals.
@@
 Panel A: Overall price dispersion
-$FE_{ht}$ 0.483 0.485 0.486
-$FE_{i}$ 0.427 0.424 0.423
-$FE_{j}$ 0.0452 0.0464 0.0463
-Match residual 0.0444 0.0441 0.0442
+$FE_{ht}$ 0.485 0.483 0.486
+$FE_{i}$ 0.424 0.427 0.423
+$FE_{j}$ 0.0464 0.0452 0.0463
+Match residual 0.0441 0.0444 0.0442
 Panel B: Within exporter–product dispersion
-$FE_{j}$ 0.231 0.233 0.233
-Match residual 0.768 0.765 0.765
-Notes: The columns correspond to alternative price definitions: $\ln p_{ijht}$ excludes both duties and charges; $\ln p^{c}_{ijht}$ includes charges but excludes duties; $\ln p^{\text{duty}}_{ijht}$ includes duties but excludes charges.
+$FE_{j}$ 0.233 0.231 0.233
+Match residual 0.765 0.768 0.765
+Notes: The columns correspond to alternative price definitions: $\ln p_{ijht}$ excludes duties; $\ln p^{c}_{ijht}$ exclude both duties and charges; $\ln p^{\text{duty}}_{ijht}$ includes duties.
 The estimation sample includes importer–exporter–product matches observed in two consecutive calendar years, and applies the following restrictions: (i) excludes transactions involving consumption goods (based on the BEC classification), energy-sector products, statistical outliers, and related-party trade; and (ii) retains only suppliers that trade with at least two distinct U.S.
@@
 The control vector $\boldsymbol{X}_{ijht}$ includes the log of transaction value, the log of relationship longevity (years since the exporter first supplied the buyer with the given HS10 product), and the log of the relative number of partners (the supplier’s number of HS10-level buyers divided by the buyer’s number of HS10-level suppliers).
-The sample includes 1,2000,000 importer–exporter–product–year observations, which have been rounded to four significant digits per U.S.
+The sample includes 1,200,000 importer–exporter–product–year observations, which have been rounded to four significant digits per U.S.
 Census Bureau disclosure guidelines.
@@
 D.2 Pass-Through Heterogeneity
-Table D.2 examines the heterogeneity in tariff pass-through through specification (3.2) by including buyer-by-year fixed effects ($FE_{jt}$).
-This more demanding specification accounts for time-varying shocks at the buyer level, while also controlling for product–year and exporter country–year fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$).
-Findings replicate the results of incomplete tariff pass-through reported in Table 3, suggesting that exporters adjust marginal costs in response to demand shifts from dominant importers, thereby absorbing a substantial fraction of tariff shocks.
-These results demonstrate the critical role of the cost channel as, by Proposition 3, the pass-through decreases with the buyer share $x_{ij}$.
+Table D.2 examines heterogeneity in tariff pass-through using specification (3.2) with a more demanding set of fixed effects.
+In addition to product–time fixed effects ($FE_{ht}$), the specification includes importer–time and exporting country–time fixed effects ($FE_{jt}+FE_{ct}$), accounting for time-varying shocks at the buyer level while flexibly controlling for exporter-side conditions.
+Findings replicate the incomplete tariff pass-through estimates reported in Table 3, consistent with exporters adjusting marginal costs in response to demand shifts from large buyers and thereby absorbing a substantial fraction of tariff shocks.
+In line with Proposition 3, pass-through declines with the buyer share $x_{ij}$, highlighting the central role of the cost channel.
 Table D.2: Pass-Through and Relationship Heterogeneity, Stringent Fixed Effects
@@
 (1) (2) (3) (4) (5) (6)
-$\Delta\ln(1+\tau_{cht})$0.223 -0.342 -0.230 -0.154 -0.163 -0.292
+$\Delta\ln(1+\tau_{cht})$ -0.184 -0.306 -0.183 -0.113 -0.114 -0.247
 (0.107) (0.155) (0.093) (0.126) (0.104) (0.145)
-$\Delta\ln(1+\tau_{cht})\cdot\ln\text{longevity}_{ijht}$ 0.086 0.097
+$\Delta\ln(1+\tau_{cht})\cdot\ln\text{longevity}_{ijht}$ 0.088 0.101
 (0.047) (0.047)
-$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.023 0.029 0.018
+$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ -0.003 0.004 -0.007
 (0.149) (0.165) (0.165)
-$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.271 -0.271 -0.280
+$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$ -0.282 -0.280 -0.289
 (0.125) (0.134) (0.128)
@@
 R-squared 0.31 0.31 0.31 0.31 0.31 0.31
-Notes: This table reports estimates of the pass-through of statutory tariffs, $\Delta\ln(1+\tau_{cht})$, to duty-exclusive prices at the exporter–importer–product–year level, $\Delta\ln p_{ijht}$.
+Notes: This table reports estimates of the pass-through of statutory tariffs, $\Delta\ln(1+\tau_{cht})$, to duty-exclusive CIF prices at the exporter–importer–product–year level, $\Delta\ln p_{ijht}$.
 Columns (2) and (6) interact tariffs with the log of relationship longevity, measured as the number of years that buyer $j$ and supplier $i$ have transacted in product $h$.
@@
 Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
-Nonlinear Effects
+Nonlinear Effects.
 To explore nonlinearities in tariff pass-through along the distribution of bilateral concentration, we interact the tariff change with quartile dummies of the lagged supplier share ($s_{ijh,t-1}$) and buyer share ($x_{ijh,t-1}$).
@@
 Panel (A) of Figure D.1 includes product–time and exporting country–sector fixed effects, while Panel (B) features product–time, importer–time, and exporting country–time fixed effects.
-$\Delta\ln p_{ijht}=$ $\alpha_{0}+\alpha_{1}\Delta\ln(1+\tau_{cht})+\sum_{q=2}^{Q}\alpha_{s,q}\cdot\Delta\ln(1+\tau_{cht})\cdot\mathbf{1}\{s_{ijh,t-1}\in Q_{q}\}$
+$\Delta\ln p_{ijht}=\$$\alpha_{0}+\alpha_{1}\Delta\ln(1+\tau_{cht})+\sum_{q=2}^{Q}\alpha_{s,q}\cdot\Delta\ln(1+\tau_{cht})\cdot\mathbf{1}\{s_{ijh,t-1}\in Q_{q}\}$
 $+\sum_{q=2}^{Q}\alpha_{x,q}\cdot\Delta\ln(1+\tau_{cht})\cdot\mathbf{1}\{x_{ijh,t-1}\in Q_{q}\}+\gamma^{\prime}\textbf{X}_{ijht}+\textbf{FE}+\epsilon_{ijht}.$ (D.1)
 The results reveal no evidence of nonlinearities in pass-through with respect to supplier shares (in purple).
-Interaction coefficients across the upper quartiles of $s_{ijht}$ are uniformly positive, but small in magnitude and statistically insignificant, indicating that supplier concentration does not materially affect the degree of pass-through.
-In contrast, buyer shares exhibit a strong, monotonic relationship (in green): pass-through declines significantly at higher quartiles of $x_{ijht}$, consistent with the model’s prediction that dominant buyers constrain suppliers’ ability to shift cost shocks.
+Interaction coefficients across the upper quartiles of $s_{ijht}$ are uniformly positive, but small and statistically insignificant, indicating that supplier concentration does not materially affect the degree of pass-through.
+In contrast, buyer shares display a strong monotonic pattern (in green): pass-through declines significantly across higher quartiles of $x_{ijht}$, consistent with the model’s prediction that dominant buyers limit suppliers’ ability to transmit cost shocks.
 Figure D.1: Pass-Through by Bilateral Market Share Quartiles
@@
 Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
-Additional Robustness: GE Controls and Price Definitions
+Additional Robustness: GE Controls and Price Definitions.
 Table D.3 presents robustness checks using alternative price definitions and specifications.
-Columns (1)-(2) exclude the general equilibrium controls from our baseline model.
-Columns (3)-(4) use duty-exclusive prices that include charges, while Columns (5)-(6) use tariff-inclusive prices.
-In all cases, we interact tariff changes with lagged supplier and buyer shares, and hold fixed effects constant across specifications for comparability.
-Across all variations, pass-through estimates and their interaction effects with bilateral market shares remain stable in sign and magnitude, supporting the robustness of the main findings.
-Table D.3: Additional Robustness: GE Controls and Price Definitions
-Dependent variable: $\ln p_{ijht}$ (excl. GE controls) $\ln p^{c}_{ijht}$ (before duty, incl. charges) $\ln p^{\text{duty}}_{ijht}$ (tariff inclusive)
+Columns (1)-(2) exclude the general equilibrium controls included in the baseline model.
+Columns (3)-(4) use duty-exclusive prices that exclude charges, while Columns (5)-(6) use tariff-inclusive prices.
+In all cases, tariff changes are interacted with lagged supplier and buyer shares, and the set of fixed effects is held constant to ensure comparability across specifications.
+Estimates remain stable in sign and magnitude across all variants, confirming the robustness of the main results.
+Table D.3: Additional Robustness: General Equilibrium Controls and Price Definitions
+Dependent variable: $\ln p_{ijht}$ (excl. GE controls) $\ln p^{c}_{ijht}$ (before duty, excl. charges) $\ln p^{\text{duty}}_{ijht}$ (tariff inclusive)
 (1) (2) (3) (4) (5) (6)
-$\Delta\ln(1+\tau_{cht})$0.060 -0.153 -0.032 -0.114 0.506 0.401
+$\Delta\ln(1+\tau_{cht})$ -0.025 -0.103 -0.066 -0.163 0.506 0.401
 (0.090) (0.103) (0.093) (0.107) (0.099) (0.126)
-$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.048 0.036 0.033 0.004 0.017 0.007
+$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.027 0.011 0.054 0.029 0.017 0.007
 (0.076) (0.170) (0.072) (0.157) (0.068) (0.143)
-$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.399 -0.278 -0.407 -0.280 -0.421 -0.287
+$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$ -0.403 -0.287 -0.403 -0.271 -0.421 -0.287
 (0.112) (0.135) (0.113) (0.135) (0.121) (0.145)
@@
 Notes: This table reports robustness checks on tariff pass-through specifications using alternative price definitions and control sets.
-Columns (1)–(2) exclude general equilibrium controls; Columns (3)–(4) use pre-duty prices including charges; and Columns (5)–(6) use tariff-inclusive prices.
+Columns (1)–(2) exclude general equilibrium controls; Columns (3)–(4) use pre-duty prices excluding charges; and Columns (5)–(6) use tariff-inclusive prices.
 In each case, we report specifications using either baseline fixed effects ($FE_{ht}+FE_{cs}$) or a more stringent set of fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$).
@@
 We assess the robustness of our structural estimates in Section 4.2.
-We first replicate the main GMM estimation using an expanded sample that includes all Broad Economic Categories (BECs), rather than excluding consumption goods as in the baseline.
-This extended product scope allows us to test whether our key parameter estimates, namely, returns to scale ($\theta$) and relative bargaining power ($\phi$), are sensitive to the exclusion of consumer-oriented products.
-As shown in Table D.4, the results remain fairly stable, suggesting that the baseline findings are not driven by product composition.
-We then examine how sensitive the estimates are to alternative values of calibrated model parameters.
-First, we vary the elasticity of substitution across foreign varieties ($\rho$) by setting $\rho=5$ instead of 10, which is consistent with the lower end of the estimates in the literature.
-Second, we relax the assumption of constant returns to scale for the importers and set $\varrho=0.5$ instead of 1.
-We report the results in Table D.5.
-Column (1) shows the estimated values when setting $\rho=5$, and Column (2) shows the estimated values when setting $\varrho=0.5$.
-Furthermore, in Table D.5, we also estimate the parameters using an alternative sample constructed by utilizing the RPT indicator from LFTTD.
-[^37] See Section 3.3 for the discussion on selection through RPT indicators.
-Columns (3) and (4) report the estimated values for this set of sample.
-Throughout these alternative setups,the resulting estimates remain robust, suggesting that the estimated values are not sensitive to particular values of other parameters or set of sample.
+We first re-estimate the main GMM specification using an expanded sample that includes all Broad Economic Categories (BECs).
+We also estimate the structural parameters on a sample restricted to consumption goods only.
+Results are reported in Tables D.4 and D.5, respectively
+Estimated returns to scale are systematically higher for consumption goods.
+Restricting the estimation to consumption goods raises $\hat{\theta}$ from 0.42–0.51 in the baseline to 0.54–0.66.
+Expanding the sample to include all BEC categories also increases estimates, with $\hat{\theta}$ ranging from 0.48 to 0.60, consistent with the higher values observed for consumption goods.
+This pattern suggests that consumption goods are less relationship-specific and therefore face more elastic supply conditions, implying greater substitutability across suppliers and higher estimated returns to scale in the model.
+Finally, Table D.6 examines the robustness of our structural estimates to an alternative definition of related-party trade.
+Instead of using our baseline ORBIS-based classification, we re-estimate the model excluding transactions identified as related-party trade using the standard RPT indicator in the U.S.
+Census data.
+Appendix C.1 explains the difference between the two indicators.
+Reassuringly, the estimated bargaining-power and returns-to-scale parameters remain very close to the baseline results, indicating that our main conclusions are not driven by the particular way related-party relationships are classified.
 Table D.4: Estimated Model Primitives - All BEC Categories
 Panel A: Calibrated Parameters $\hat{\nu}$ $\hat{\gamma}$ $\hat{\rho}$ 4 0.5 10 Panel B: Estimated Parameters (GMM) (1) (2) (3) (4) Rel.
-bargaining power: $\ln\widehat{\frac{\phi}{1-\phi}}$ 1.162 0.558 (0.026) (0.022) Returns to scale ($\hat{\theta}$) 0.505 0.573 0.432 0.586 (0.003) (0.005) (0.004) (0.005) Constant 2.780 0.742 (0.190) (0.083) Longevity 0.061 0.742 (0.023) (0.069) Number of HS10 transactions -0.240 -0.019 (0.019) (0.011) Multiple HS10 dummy 0.080 0.190 (0.026) (0.025) Lagged outside option -0.188 -0.237 (0.017) (0.021) None Yes Yes No No $FE_{h}+FE_{t}+FE_{j}$ No No Yes Yes Observations 6,143,000 Panel C: Implied Bargaining Powers ($\hat{\phi}$) Mean 0.762 0.896 0.636 0.848 (0.005) (0.068) (0.005) (0.112) Median 0.913 0.878 (0.068) (0.112) Notes: This table presents model estimates based on a sample that includes all Broad Economic Categories (BEC), including consumption goods, for the period 2001-2016.
-Panel A reports calibrated parameters: the elasticity of demand ($\nu$), the elasticity of costs with respect to foreign input prices ($\gamma$), and the elasticity of substitution across foreign varieties ($\rho$).
-We set $\varrho=1$, so that $\eta=2.5$.
-Panel B presents GMM estimates.
-Columns (1) and (3) impose a constant $\phi$ across bilateral pairs, while Columns (2) and (4) estimate the full vector $\boldsymbol{\kappa}$ to allow for heterogeneity in bargaining power.
-Specifications differ in the inclusion of fixed effects.
-Controls include: (i) the log of relationship longevity between exporter $i$ and importer $j$; (ii) the log of the number of transactions between $i$ and $j$ in a given year; (iii) the log of the relative outside option, defined as the ratio of exporter $i$’s sales to other U.S.
-buyers (excluding $j$) over importer $j$’s purchases from other suppliers (excluding $i$), both in year $t{-}1$; and (iv) a dummy variable equal to one if the $i–j$ pair transacts in more than one HS10 product.
-Panel C reports the mean and median of the implied bargaining power.
-Standard errors are robust; those in Panel C are computed using the delta method.
-The set of instruments includes the number of exporters and importers at the HS10 level, as well as lagged bilateral shares (excluding the focal pair).
+bargaining power: $\ln\widehat{\frac{\phi}{1-\phi}}$ 1.180 0.940 (0.023) (0.021) Returns to scale ($\hat{\theta}$) 0.508 0.585 0.483 0.595 (0.002) (0.003) (0.003) (0.003) Constant 3.355 1.299 (0.196) (0.103) Longevity 0.115 0.902 (0.026) (0.090) Number of HS10 transactions -0.315 -0.090 (0.021) (0.016) Multiple HS10 dummy 0.137 0.199 (0.028) (0.029) Lagged outside option -0.267 -0.272 (0.019) (0.023) None Yes Yes No No $FE_{h}+FE_{t}+FE_{j}$ No No Yes Yes Observations 6,143,000 Panel C: Implied Bargaining Powers ($\hat{\phi}$) Mean 0.765 0.923 0.719 0.899 (0.004) (0.075) (0.004) (0.095) Median 0.946 0.928 (0.075) (0.095)
+Notes: This table presents model estimates based on a sample that includes all Broad Economic Categories (BEC), including consumption goods, for the period 2001-2016.
+Panel A reports calibrated parameters: demand elasticity ($\nu$), cost elasticity to foreign input prices ($\gamma$), and the elasticity of substitution across foreign varieties ($\rho$).
+Panel B reports GMM estimates.
+Columns (1) and (3) impose a constant $\phi$, while Columns (2) and (4) allow heterogeneous bargaining power via $\boldsymbol{\kappa}$.
+Specifications differ in fixed effects.
+Controls include relationship longevity, transaction intensity, the relative outside option (lagged), and a multi-product indicator.
+Panel C reports the mean and median implied bargaining power.
+Robust standard errors; Panel C uses the delta method.
+Instruments include HS10-level counts of exporters and importers and lagged bilateral shares (excluding the focal pair).
+All estimations use CIF duty-exclusive unit values.
 The number of observations is rounded to four significant digits in accordance with U.S.
@@
 Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
-Table D.5: Robustness of Model Estimates
-$\rho=5$ $\varrho=0.5$ Related Party Trade LFTTD (1) (2) (3) (4) Rel.
-bargaining power: $\ln\widehat{\frac{\phi}{1-\phi}}$ 1.455 1.838 1.618 0.892 (0.038) (0.052) (0.062) (0.047) Returns to scale ($\theta$) 0.427 0.473 0.453 0.381 (0.004) (0.003) (0.005) (0.007) Mean $\hat{\phi}$ 0.811 0.863 0.835 0.709 (0.006) (0.006) (0.009) (0.010) Median $\hat{\phi}$ 0.811 0.863 0.835 0.709 (0.000) (0.000) (0.000) (0.000) None Yes Yes Yes No $FE_{h}+FE_{t}+FE_{j}$ No No No Yes Observations 3,120,000
-Notes: This table reports robustness checks for the main model estimates.
-The columns explore sensitivity to changes in key calibrated parameters and sample definitions.
-Column (1) varies the elasticity of substitution across foreign varieties ($\rho$); Column (2) changes the downstream returns to scale parameter ($\varrho$).
-Columns (3)–(4) use a sample based on related-party indicators from LFTTD (RPT), without and with fixed effects for buyer, product, and time.
+Table D.5: Estimated Model Primitives - BEC Consumption Categories
+Panel A: Calibrated Parameters $\hat{\nu}$ $\hat{\gamma}$ $\hat{\rho}$ 4 0.5 10 Panel B: Estimated Parameters (GMM) (1) (2) (3) (4) Rel.
+bargaining power: $\ln\widehat{\frac{\phi}{1-\phi}}$ 0.936 0.729 (0.023) (0.021) Returns to scale ($\hat{\theta}$) 0.558 0.646 0.539 0.664 (0.003) (0.004) (0.003) (0.004) Constant 2.501 0.568 (0.213) (0.093) Longevity 0.628 1.414 (0.070) (0.198) Number of HS10 transactions -0.386 -0.137 (0.037) (0.030) Multiple HS10 dummy 0.452 0.344 (0.056) (0.052) Lagged outside option -0.266 -0.299 (0.027) (0.038) None Yes Yes No No $FE_{h}+FE_{t}+FE_{j}$ No No Yes Yes Observations 3,024,000 Panel C: Implied Bargaining Powers ($\hat{\phi}$) Mean 0.718 0.919 0.674 0.891 (0.005) (0.077) (0.005) (0.111) Median 0.942 0.928 (0.077) (0.111)
+Notes: This table presents model estimates based on a sample that only includes consumption goods based on Broad Economic Categories (BEC), for the period 2001-2016.
+Panel A reports calibrated parameters: demand elasticity ($\nu$), cost elasticity to foreign input prices ($\gamma$), and the elasticity of substitution across foreign varieties ($\rho$).
+Panel B reports GMM estimates.
+Columns (1) and (3) impose a constant $\phi$, while Columns (2) and (4) allow heterogeneous bargaining power via $\boldsymbol{\kappa}$.
+Specifications differ in fixed effects.
+Controls include relationship longevity, transaction intensity, the relative outside option (lagged), and a multi-product indicator.
+Panel C reports the mean and median implied bargaining power.
+Robust standard errors; Panel C uses the delta method.
+Instruments include HS10-level counts of exporters and importers and lagged bilateral shares (excluding the focal pair).
+All estimations use CIF duty-exclusive unit values.
+The number of observations is rounded to four significant digits in accordance with U.S.
+Census Bureau disclosure guidelines.
+Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
+Table D.6: Estimated Model Primitives under Alternative Related-Party Classification
+(1) (2) (3) (4) Implied Mean Bargaining Power ($\hat{\phi}$) 0.837 0.928 0.761 0.899 (0.008) (0.078) (0.008) (0.082) Returns to scale ($\hat{\theta}$) 0.455 0.500 0.410 0.504 (0.004) (0.005) (0.005) (0.006) None Yes Yes No No $FE_{h}+FE_{t}+FE_{j}$ No No Yes Yes
+Notes: Model estimates are based on U.S.
+arm’s-length imports of intermediate inputs and capital goods over 2001–2016, excluding transactions identified as related-party trade using the RPT indicator in the U.S.
+Census data; see Appendix C.1 for details.
+Columns (1) and (3) impose a constant $\phi$, while Columns (2) and (4) allow for heterogeneous bargaining power through $\boldsymbol{\kappa}$, parameterized as a function of relationship longevity, transaction intensity, the lagged relative outside option, and a multi-product indicator.
+The demand elasticity is set to $\nu=4$, the cost elasticity to foreign input prices to $\gamma=0.5$, and the elasticity of substitution across foreign varieties to $\rho=10$.
+Specifications differ in the fixed effects included.
+Robust standard errors are reported in parentheses.
+All estimations use CIF duty-exclusive unit values.
+All regressions include 3,120,000 observations.
 The number of observations is rounded to four significant digits in accordance with U.S.
@@
 D.4 Additional Results on Model Fit
-Sensitivity to alternative parameter values.
-Table D.6 presents IV-based goodness-of-fit tests as in equation (4.5) for an alternative set of calibrated parameters of the model and two alternative sets of fixed effects.
-The results show that models incorporating bargaining and decreasing returns to scale (Columns (1) and (5)) have coefficients closer to one, indicating empirical alignment, consistent with the results in Figure 3.
-Quantity responses and relationship heterogeneity: data vs. model.
+Pass-through and Relationship Heterogeneity: Data vs. Model.
+Table D.7 reports tariff pass-through in the data and in the model (baseline and restricted variants) under a more demanding set of fixed effects.
+Panel A shows that the model captures the main pass-through effect in the data very closely.
+Panel B shows how pass-through varies with buyer and supplier concentration, measured by lagged buyer and supplier shares.
+The baseline model closely replicates these patterns: pass-through declines sharply with buyer concentration ($x_{ijht-1}$) and increases modestly with supplier concentration ($s_{ijht-1}$), with magnitudes comparable to those observed in the data.
+IV-Based Goodness-of-Fit Test.
+We formally assess model fit using an IV regression following adao2023putting:
+$\Delta\ln p_{ijht}=\beta\,\widehat{\Delta\ln p_{ijht}}+\mathbf{FE}+u_{ijht},$ (D.2)
+where $\Delta\ln p_{ijht}$ is the observed change in the duty-inclusive price and $\widehat{\Delta\ln p_{ijht}}$ is the corresponding model-implied change.
+Because observed price changes may reflect shocks unrelated to tariffs, we instrument $\widehat{\Delta\ln p_{ijht}}$ with statutory tariff changes to isolate tariff-driven variation.
+[^36] Unlike the pass-through analysis in the main text, which uses duty-exclusive prices to mitigate measurement error in applied duties, the IV strategy isolates variation induced by statutory tariffs and evaluates the model against duty-inclusive price changes.
+Under the null that the model correctly captures tariff-induced price movements, the IV coefficient $\hat{\beta}$ equals one.
+Figure D.2 reports the IV estimates of $\beta$ across the four model variants under the two fixed-effects specifications.
+Blue markers correspond to regressions with product–time and country–sector fixed effects, while red markers use product–time, country–time, and buyer–time fixed effects.
+The model generates price changes closely aligned with those observed in the data.
+Formally, however, we cannot reject any of the variants at conventional levels of significance.
+Sensitivity of Model Fit to Alternative Parameter Values.
+In our baseline estimation we set $\varrho=1$, $\nu=4$, and $\gamma=0.5$.
+As all these parameters enter the model only though $\eta$, fixing the buyer’s returns to scale parameter $\varrho$ to one does not restrict generality.
+Nevertheless, we conduct a robustness exercise where we set $\varrho=0.5$, close to the returns to scale parameter we estimate for the supplier side.
+Further, we also consider a value of $\rho=5$, the lower end of the range provided in Anderson2004.
+Table D.8 presents IV-based goodness-of-fit tests as in equation (D.2) for these alternative set of calibrated parameters of the model and two alternative sets of fixed effects.
+The results show that models incorporating bargaining and decreasing returns to scale (Columns (1) and (5)) have coefficients closer to one, indicating empirical alignment, consistent with the results in Figure D.2.
+Model Fit: Quantity and Sales Responses.
 Table D.9 compares how quantities respond to tariffs in the data (Panel A) and in the model (Panel B).
@@
 In contrast, the interaction terms with supplier and buyer shares for the quantity responses in the data are statistically imprecise, suggesting inconclusive evidence of heterogeneity in the data.
-Testing the model-predicted quantity changes.
-Panel A in Table D.8 presents IV-based tests comparing observed quantity changes to those predicted by the model under alternative parameterizations.
-The model-predicted change in quantity is given by:
+Table D.10 reports IV-based tests comparing observed quantity (Panel A) and sales (Panel B) changes to those implied by the model under alternative parameterizations. The model predicts
 $\widehat{\Delta\ln q_{ijht}}=-\varepsilon_{ijht}\,\widehat{\Delta\ln p_{ijht}}$,
-where $\varepsilon_{ijht}$ is the match-specific demand elasticity and $\widehat{\Delta\ln q_{ijht}}$ is as in equation (4.5).
-Although formal tests reject all models, the baseline model with bargaining and decreasing returns (Columns (1) and (2)) demonstrates the strongest fit, indicating that this specification best captures the underlying mechanisms of tariff-induced quantity adjustments.
-Testing the model-predicted sales changes.
-Panel B in Table D.8 provides analogous IV-based tests for observed sales changes computed as:
+where $\varepsilon_{ijht}$ denotes the match-specific demand elasticity and $\widehat{\Delta\ln p_{ijht}}$ is defined in equation (D.2). For sales changes, we use:
 $\widehat{\Delta\ln r_{ijht}}=(1-\varepsilon_{ijht})\,\widehat{\Delta\ln p_{ijht}}$.
-The baseline specification (Columns (1) and (2)) shows superior alignment between predicted and observed sales, reinforcing the conclusion that the bargaining model with decreasing returns most effectively matches empirical patterns in the data.
+Although the formal tests reject all specifications, the specification that combines bargaining with decreasing returns and the specification with decreasing returns only provide the best quantitative fit, consistent with the results for prices.
+Table D.7: Price Responses and Relationship Heterogeneity: Data vs. Model
+Panel A: Main Effect (Stringent Fixed Effects)
+Data Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$
+$\Delta\ln(1+\tau_{cht})$ -0.184 -0.226 -0.095 -0.014 -0.286
+(0.107) (0.009) (0.008) (0.001) (0.011)
+R-squared 0.31 0.48 0.42 0.41 0.51
+Panel B: Including Interactions (Stringent Fixed Effects)
+Data Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$
+(1) (2) (3) (4) (5)
+$\Delta\ln(1+\tau_{cht})$ -0.114 -0.137 -0.030 -0.002 -0.200
+(0.104) (0.011) (0.007) (0.001) (0.014)
+$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.004 0.116 -0.120 -0.022 0.073
+(0.165) (0.013) (0.022) (0.003) (0.013)
+$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$ -0.280 -0.433 -0.078 -0.013 -0.372
+(0.134) (0.027) (0.020) (0.004) (0.025)
+R-squared 0.31 0.57 0.45 0.45 0.56
+$FE_{ht}+FE_{ct}+FE_{jt}$ Yes Yes Yes Yes Yes
+Observations 249,000
+Notes: This table reports the response of duty-exclusive CIF prices to tariff changes at the exporter–importer–product level.
+Column (1) presents reduced-form estimates using observed price changes in the data.
+Columns (2)–(5) report the corresponding responses implied by model-predicted price changes under each specification.
+Panel A reports the main effect of tariff changes, while Panel B allows the response to vary with lagged supplier shares ($s_{ijht-1}$) and buyer shares ($x_{ijht-1}$).
+All specifications include product–time, country–time, and buyer-time fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$).
+Standard errors are clustered at the HS8 product and exporter-country level.
+Observation counts are rounded to four significant digits per U.S.
+Census Bureau disclosure guidelines.
+Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
+Figure D.2: IV-Based Goodness-of-Fit Test
+Notes: Each point reports the coefficient from an IV regression of observed log duty-inclusive price changes on model-predicted changes $\widehat{\Delta\ln p_{ijht}}$, using statutory tariffs as instruments.
+Lines show 95% confidence intervals.
+Blue and red denote regressions with $\mathrm{FE}_{ht}+\mathrm{FE}_{cs}$ and $\mathrm{FE}_{ht}+\mathrm{FE}_{ct}+\mathrm{FE}_{jt}$, respectively.
+Standard errors are clustered by product and exporter-country.
+Observation counts (249,000) are rounded per Census disclosure guidelines.
+Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
 Table D.8: IV-Based Goodness-of-Fit Test, Alternative Parameters
@@
 (1) (2) (3) (4) (5) (6) (7) (8)
-$\widehat{\Delta\ln p_{ijht}}$ 1.211 0.781 0.735 1.368 1.247 0.871 0.740 1.544
+$\widehat{\Delta\ln p_{ijht}}$ 1.158 0.889 0.771 1.391 1.291 0.936 0.776 1.619
 (0.331) (0.253) (0.220) (0.397) (0.368) (0.267) (0.221) (0.4763)
@@
 Dependent variable: $\Delta\ln p_{ijht}$
+$\rho=5$ $\varrho=0.5$
 Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$ Baseline $\phi=0$, $\theta=1$ $\theta=1$ $\phi=0$
 (1) (2) (3) (4) (5) (6) (7) (8)
-$\widehat{\Delta\ln p_{ijht}}$ 1.000 0.648 0.612 1.132 1.047 0.720 0.615 1.302
+$\widehat{\Delta\ln p_{ijht}}$ 0.998 0.776 0.667 1.215 1.128 0.805 0.670 1.432
 (0.341) (0.265) (0.227) (0.415) (0.386) (0.275) (0.229) (0.490)
@@
 Observations 249,000
-Notes: Each column reports the coefficient from an IV regression of the observed change in log price on the corresponding model-predicted change $\widehat{\Delta\ln p_{ijht}}$, using statutory tariff changes as instruments.
+Notes: Each column reports the coefficient from an IV regression of the observed change in log duty-inclusive price on the corresponding model-predicted change $\widehat{\Delta\ln p_{ijht}}$, using statutory tariff changes as instruments.
 Columns (1)-(4) use predicted prices using a $\rho=5$ instead of $\rho=10$ as in the baseline.
@@
 Table D.9: Quantity Responses and Relationship Heterogeneity: Data vs. Model
-Panel A: Data
+Baseline FE Stringent FE
+Data Model Data Model
 (1) (2) (3) (4)
-$\Delta\ln(1+\tau_{cht})$0.568 0.006 -1.021 -0.761
+Panel A: Main effects
+$\Delta\ln(1+\tau_{cht})$ -0.568 -1.710 -1.021 -1.921
 (0.249) (0.075) (0.185) (0.106)
-$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$0.511 -0.278
-(0.323) (0.276)
-$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.201 -0.130
-(0.229) (0.242)
-R-squared 0.06 0.10 0.32 0.36
-Panel B: Model
-(1) (2) (3) (4)
-$\Delta\ln(1+\tau_{cht})$1.962 -3.967 -2.169 -3.984
+R-squared 0.06 0.25 0.32 0.40
+Panel B: Including Interactions
+$\Delta\ln(1+\tau_{cht})$ 0.006 -3.683 -0.761 -3.705
 (0.331) (0.203) (0.237) (0.229)
-$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 2.441 2.420
-(0.409) (0.551)
-$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$ 2.993 3.270
-(0.229) (0.316)
-R-squared 0.28 0.38 0.44 0.51
+$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ -0.511 2.755 -0.278 2.816
+(0.323) (0.461) (0.276) (0.617)
+$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$ -0.201 2.453 -0.130 2.703
+(0.229) (0.242) (0.242) (0.349)
+R-squared 0.10 0.36 0.36 0.49
 $FE_{ht}+FE_{cs}$ Yes Yes No No
@@
 Observations 249,000
-Notes: This table reports the pass-through of tariffs to quantities at the exporter–importer–product level.
-Panel A presents reduced-form estimates from the data.
-Panel B shows corresponding pass-through estimates generated by the model.
-Columns (2)–(4) interact tariff changes with lagged supplier share ($s_{ijht-1}$) and lagged buyer share ($x_{ijht-1}$).
-Columns (1) and (2) use baseline fixed effects ($FE_{ht}+FE_{cs}$), while Columns (3) and (4) employ a more stringent specification with product–year, country–year, and buyer–year fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$).
+Notes: Panel A reports main effects of tariff changes on quantities.
+Panel B introduces interactions between tariff changes and lagged supplier share ($s_{ijht-1}$) and lagged buyer share ($x_{ijht-1}$).
+Columns (1) and (2) use baseline fixed effects ($FE_{ht}+FE_{cs}$), while Columns (3) and (4) use a more demanding specification with product–year, country–year, and buyer–year fixed effects ($FE_{ht}+FE_{ct}+FE_{jt}$).
 Standard errors are clustered at the HS8 product and exporter-country level.
@@
 (1) (2) (3) (4) (5) (6) (7) (8)
-$\widehat{\Delta\ln q_{ijht}}$ 0.349 0.525 0.205 0.311 0.185 0.282 0.420 0.627
+$\widehat{\Delta\ln q_{ijht}}$ 0.401 0.592 0.242 0.361 0.213 0.320 0.496 0.726
 (0.165) (0.161) (0.100) (0.097) (0.087) (0.086) (0.205) (0.198)
@@
 (1) (2) (3) (4) (5) (6) (7) (8)
-$\widehat{\Delta\ln(p\cdot q)_{ijht}}$ 0.176 0.426 0.102 0.250 0.092 0.228 0.212 0.509
+$\widehat{\Delta\ln(p\cdot q)_{ijht}}$ 0.206 0.489 0.121 0.292 0.107 0.262 0.255 0.598
 (0.184) (0.177) (0.108) (0.105) (0.095) (0.094) (0.228) (0.216)
@@
 D.5.1 Pass-Through Across Samples
-Panel A and Panel B in Table D.9 evaluate how tariff pass-through estimates vary under different sample restrictions using baseline fixed effects (product–time, country–sector) and more stringent fixed effects (including buyer-time), respectively.
-For ease of exposition, the first two columns of Table D.9 replicate the numbers of Columns (1) and (3) of Table 5.
-These two columns show that the baseline sample yields a pass-through estimate of around 78-85%, closely matched by the model-implied prediction with no statistically significant difference.
-As the sample is progressively broadened–from including single-buyer matches (Column (3)), to adding related-party transactions, energy goods, and outliers (Column (4)), and finally to the most inclusive specification (Column (5))–estimated pass-through increases steadily, reaching up to 93-95%.
-This pattern underscores the sensitivity of reduced-form estimates to sample composition and the role of relationship filtering in uncovering pricing patterns consistent with bilateral bargaining.
+Table D.11 examines how tariff pass-through estimates vary across alternative sample restrictions under baseline fixed effects (product–time and country–sector; Panel A) and more demanding specifications that additionally include country-time (instead of country-sector) and buyer–time fixed effects (Panel B).
+For reference, the first two columns replicate Columns (1) and (2) of Table 6.
+These estimates imply pass-through of roughly 82–88 percent in the data, closely aligned with the model’s predictions, with no statistically significant differences.
+As the sample is progressively broadened–from including final goods (column 3), to incorporating single-buyer matches (column 4), and finally to the most inclusive specification, which adds related-party transactions, energy goods, and outliers (column 5)–estimated pass-through rises monotonically, reaching 94–96 percent.
+These results indicate that sample composition plays an important role in measured pass-through.
+At the same time, when attention is restricted to stable buyer–supplier relationships, pass-through remains incomplete and below the near-100 percent estimates often reported in the literature.
+To further reconcile our findings with this evidence, we next aggregate the data to the product level and exploit monthly variation.
 Table D.11: Tariff Pass-Through Across Different Samples
-Dependent variable: Model Baseline +Suppliers + Energy/ + Final
-$\Delta\ln p_{ijht}$ w/ RPT/ Goods
-$<$ 2 Buyers Outliers
+Dependent variable: Model Baseline + Final +Suppliers + Energy/
+$\Delta\ln p_{ijht}$ Goods w/ RPT/
+<< 2 Buyers Outliers
 (1) (2) (3) (4) (5)
 Panel A: Baseline Fixed Effects
-$\Delta\ln(1+\tau_{cht})$0.248 -0.151 -0.168 -0.099 -0.066
+$\Delta\ln(1+\tau_{cht})$ -0.222 -0.126 -0.151 -0.138 -0.060
 (0.008) (0.091) (0.089) (0.034) (0.044)
 $FE_{ht}+FE_{cs}$ Yes Yes Yes Yes Yes
-R-squared 0.32 0.04 0.02 0.02 0.02
+R-squared 0.31 0.04 0.04 0.02 0.02
 Panel B: Stringent Fixed Effects
-$\Delta\ln(1+\tau_{cht})$0.249 -0.223 -0.171 -0.123 -0.047
+$\Delta\ln(1+\tau_{cht})$ -0.226 -0.184 -0.118 -0.083 -0.036
 (0.009) (0.107) (0.075) (0.032) (0.042)
 $FE_{ht}+FE_{ct}+FE_{jt}$ Yes Yes Yes Yes Yes
-R-squared 0.50 0.31 0.18 0.14 0.13
-Observations 249,000 249,000 732,000 945,000 1,768,000
-Notes: This table reports tariff pass-through estimates to duty-exclusive prices at the exporter–importer–product level.
+R-squared 0.48 0.31 0.26 0.14 0.12
+Observations 249,000 249,000 473,000 1,490,000 1,768,000
+Notes: This table reports tariff pass-through estimates to duty-exclusive CIF prices at the exporter–importer–product level.
 Panel A uses baseline fixed effects: product–time and country–sector ($FE_{ht}+FE_{cs}$).
@@
 Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
-D.5.2 Pass-Through Using Data Aggregated at the Product-Level
+D.5.2 Product-Level Pass-Through Estimates
 To complement our main analysis, we replicate a standard pass-through specification using data aggregated at the product-country-month level.
@@
 Data
-We measure price changes at the HS10–country–month level, using the same subset of buyer–supplier–product links as in the baseline analysis.
-These are links with at least one transaction in both 2017 and 2018.
-For each product and country, we construct monthly prices by aggregating trade values and quantities.
-We then relate monthly price changes to changes in statutory tariffs, controlling for product-month, country-month, and country-sector fixed effects.
-Standard errors are clustered at the HS8–country level.
-Models
-Table D.10, reports results separately for all products and for the subset that excludes consumption goods.
-Within each group, columns reflect increasingly selective samples.
-Columns (1) and (4) includes buyer–supplier–product pairs observed in two consecutive calendar years.
-The next specification, Columns (2) and (5), excludes consumption goods and applies additional filters: energy-sector goods are dropped; transactions with extreme price levels and price changes are excluded; and related-party trade is removed.
-The final sample, Columns (3) and (6), is restricted to suppliers trading with at least two U.S.
-buyers for the same product in consecutive years.
+We measure price changes at the HS10–source-country–month level using the same set of buyer–supplier–product links as in the baseline analysis, restricted to relationships with at least one transaction in both 2017 and 2018.
+For each relationship, we retain the month in which transactions occur in both years.
+We then aggregate values and quantities across these transactions to the product–source-country–month level, construct unit values, and compute month-to-month price changes.
+Relative to Table D.11, this procedure introduces two changes: prices are aggregated across trading relationships to the product–source-country level, and price changes are measured at the monthly frequency rather than annually.
+The objective of Table D.12 is to evaluate how pass-through estimates change when prices are aggregated across firms within a product–source-country–month cell, while maintaining the same underlying set of buyer–supplier links observed in both 2017 and 2018.
+Column (1) reports results for the baseline sample.
+The sample is then progressively broadened along the same margins as in Table D.11: Column (2) adds consumption goods, Column (3) allows suppliers serving only one buyer, Column (4) further includes related-party transactions, energy goods, and extreme price changes, and Column (5) additionally incorporates non-consecutive buyer–supplier relationships.
 Results
-Across specifications, we find consistent evidence of incomplete tariff pass-through to U.S.
+Across specifications, we continue to find incomplete tariff pass-through to U.S.
 import prices.
-As in our firm-level regressions, the degree of pass-through incompleteness increases as we move to more selective samples, particularly those that condition on firms with multiple trading partners over consecutive years.
-The estimated effects are generally larger in magnitude than at the match level, suggesting that relationship-level frictions may be amplified when observed in aggregated trade flows.
-While this exercise remains suggestive, it helps connect the mechanisms explored in the main analysis to pricing patterns in product-level data.
-Table D.10: Tariff Pass-Through Using Data Aggregated at the Product-Level
-Dependent Variable: $\Delta\ln p_{cht}$
-All Products Excl. Consumption Goods
-(1) (2) (3) (4) (5) (6)
-$\Delta\ln(1+\tau_{cht})$0.115 -0.136 -0.321 -0.072 -0.119 -0.700
-(0.078) (0.058) (0.186) (0.136) (0.101) (0.164)
-Consecutive obs only Yes Yes Yes Yes Yes Yes
-(-) Energy/RPT/Outliers No Yes Yes No Yes Yes
-(-) Suppliers with $<$2 buyers No No Yes No No Yes
-Observations 800,000 540,000 180,000 510,000 320,000 100,000
-R-squared 0.13 0.14 0.22 0.13 0.14 0.22
-Notes: This table reports regressions of month-over-month changes in log unit values (FOB, excluding charges) on corresponding changes in statutory tariffs ($\Delta\tau_{hct}$), measured at the HS10–country–month level.
-The data are aggregated from the firm-level sample used in our main analysis, retaining only buyer–supplier–product triplets with consecutive transactions.
-Columns (1)-(3) refer to all products; Columns (4)-(6) exclude consumption goods, based on the Broad Economic Categories (BEC) classification.
-Each column reflects a progressively more restricted sample: Columns (1) and (4) include all consecutive transactions; Columns (2) and (5) drop energy goods, extreme price levels and changes, and related-party trade; and Columns (3) and (6) restrict to suppliers with multiple buyers in consecutive years.
-All regressions include product–month, country–month, and country–sector fixed effects.
-Standard errors are clustered at the HS8–country level and reported in brackets.
-The number of observations is rounded to four significant digits in accordance with U.S.
-Census Bureau disclosure guidelines.
-Source: FSRDC Project Number 2109 (CBDRB-FY25-P2109-R12520).
-D.5.3 Pass-Through Results Using LFTTD RPT Indicators
-As a robustness check, we re-estimate the baseline specifications using the related-party transaction (RPT) indicator provided in the LFTTD.
-Table D.11 shows that the pass-through estimates remain stable, and the interaction effects with supplier and buyer shares are qualitatively similar to those reported in Table 3.
-Table D.11: Pass-Through and Relationship Heterogeneity–Alternative Related Party Trade
-Dependent variable: $\Delta\ln p_{ijht}$
-(1) (2) (3) (4)
-$\Delta\ln(1+\tau_{cht})$0.163 -0.087 -0.304 -0.253
-(0.096) (0.100) (0.111) (0.114)
-$\Delta\ln(1+\tau_{cht})\cdot s_{ijht-1}$ 0.043 0.011
-(0.070) (0.174)
-$\Delta\ln(1+\tau_{cht})\cdot x_{ijht-1}$0.374 -0.222
-(0.110) (0.148)
-$FE_{ht}+FE_{cs}$ Yes Yes No No
-$FE_{ht}+FE_{ct}+FE_{jt}$ No No Yes Yes
-Observations 249,000 249,000 249,000 249,000
-R-squared 0.04 0.05 0.32 0.32
-Notes: This table reports estimates of the pass-through of statutory tariffs, $\Delta\ln(1+\tau_{cht})$, to duty-exclusive prices at the exporter–importer–product–year level, $\Delta\ln p_{ijht}$.
-Columns (2) and (4) include interactions between tariffs and lagged bilateral characteristics: supplier share ($s_{ijht-1}$), defined as supplier $i$’s share in buyer $j$’s imports of product $h$, and buyer share ($x_{ijht-1}$), defined as buyer $j$’s share in supplier $i$’s exports of product $h$.
-All regressions include the following controls: (i) $\ln\text{longevity}_{ijht}$, the number of years $i$ and $j$ have transacted in $h$; (ii) $\Delta\ln q_{i(-j)ht}$, exporter $i$’s sales of $h$ to U.S.
-buyers other than $j$; and (iii) $\Delta\ln p_{(-i)jht}$, the average price change charged by other suppliers of $h$ to buyer $j$, using lagged shares as weights.
-Columns (1)–(2) include product–year and exporter country–sector fixed effects ($FE_{ht}+FE_{cs}$), while Columns (3)–(4) include a more demanding set of fixed effects: product–year, importer–year, and exporter country–year ($FE_{ht}+FE_{ct}+FE_{jt}$).
-Standard errors are clustered at the HS8 product and exporter-country level.
-Arm’s length transactions are defined using LFTTD related party trade indicator.
-The number of observations is rounded to four significant digits in accordance with U.S.
+When the aggregation is performed using the same set of repeated buyer–supplier relationships as in the baseline analysis, pass-through remains well below one, consistent with the match-level evidence.
+As the sample is progressively broadened to include consumption goods, single-buyer suppliers, related-party transactions, energy goods, and eventually non-consecutive relationships, estimated pass-through increases steadily, approaching the near-complete transmission typically found in the literature.
+Differences between match-level and product-level pass-through therefore primarily reflect the set of relationships included in the data rather than aggregation itself.
+Incomplete pass-through remains a robust feature when attention is restricted to persistent buyer–supplier links.
+Table D.12: Product-Level Tariff Pass-Through Across Alternative Samples
+Dependent variable: Baseline + Final + Suppliers + Energy/ + Non-Consecutive
+$\Delta\ln p_{ijht}$ Goods w/ RPT/ Relationships
+<< 2 Buyers Outliers
+(1) (2) (3) (4) (5)
+$\Delta\ln(1+\tau_{cht})$ -0.654 -0.353 -0.063 -0.054 -0.017
+(0.245) (0.156) (0.073) (0.068) (0.068)
+$FE_{ht}+FE_{cs}$ Yes Yes Yes Yes Yes
+R-squared 0.23 0.22 0.14 0.13 0.11
+Observations 124,000 213,000 665,000 805,000 1,071,000
+Notes: This table reports product-level pass-through estimates.
+Product-level observations are constructed from firm-buyer-HS relationships that transact in both years t and t-1.
+Values and quantities are aggregated to the product–source-country–month level, unit values are constructed at that frequency, and month-to-month changes in log CIF prices are computed.
+Column (1) reports the baseline sample, restricted to repeated arm’s-length relationships in intermediate goods among multi-buyer suppliers.
+Column (1) aggregates the baseline sample used in Table 3 to the product–source-country level.
+Column (2) broadens the sample to include consumption goods (all BEC categories).
+Column (3) further adds suppliers that trade with only one U.S.
+importer.
+Column (4) further includes related-party transactions, energy commodities, and observations with extreme price levels or changes.
+Column (5) removes the consecutive-year restriction and includes all exporter–importer–product combinations.
+Standard errors are clustered at the HS8 product and country level.
+Observation counts are rounded in accordance with U.S.
 Census Bureau disclosure guidelines.
@@
 Appendix E Estimation Appendix
-E.1 Downstream Demand Elasticity $\left(\nu\right)$
-Consider a model where importer $j$ sells its output $q_{j}$ to downstream customers in different countries.
-A representative consumer in each country maximises utility by choosing a composite of domestic and imported goods.
-The sub-utility derived from the composite imported good will be given by a CES aggregation across imported varieties with a good-importer specific elasticity of substitution given by $\sigma_{g}$.
-Broda and Weinstein (2006) provide estimates of the elasticity $\sigma_{g}$ at the HS10 good $g$-level in U.S.
-import data.
-The plot below shows the distribution of these elasticities.
-We base the calibration of the elasticity $\nu$ in our model on these estimates.
-We consider a value of 4 for $\nu$, close to the mean value of 3.85, which we see as a conservative choice.
+E.1 Calibration of Downstream Demand Elasticity $\left(\nu\right)$
+Consider an extension to our model in which importer $j$ sells its output $q_{j}$ to downstream consumers in different countries.
+In each country, a representative consumer allocates expenditure between domestic and imported goods, and the sub-utility over imported varieties is CES with elasticity of substitution $\sigma_{g}$ at the HS10 good level.
+Broda2006 (BW) estimate these import demand elasticities for the United States.
+In our framework, the parameter $\nu$ governs the own-price elasticity of demand faced by importer $j$’s output in the downstream market.
+We interpret BW’s $\sigma_{g}$ estimates as measuring this variety-level demand elasticity.
+Accordingly, we calibrate $\nu$ using their estimates.
+The distribution of $\sigma_{g}$ is shown in Figure E.1.
+Given a mean value of 3.85 and a median of 2.8, we set $\nu=4$, which we view as a conservative choice.
 Figure E.1: Downstream Demand Elasticity
-Notes: The figure displays the estimates of the import demand elasticity $\sigma_{g}$ from Broda and Weinstein (2006).
-The mean and median value of $\sigma_{g}^{US}$ is 3.85 and 2.8, respectively.
+Notes: The figure displays the estimates of the import demand elasticity $\sigma_{g}$ from Broda2006.
+The mean and median value of $\sigma_{g}^{US}$ are 3.85 and 2.8, respectively.
 Estimates are truncated above at 20, and below at 1.
-E.2 Monte Carlo Simulation
-Data for one replicate. Each exporter $i\in\{1,\dots,200\}$ belongs to a block with exactly two importers, labeled $j(i)$ and $\ell(i)$. Store the log‐price difference
-$\Delta p_{i}\;=\;\ln p_{ij}\;-\;\ln p_{i\ell}\quad\text{and the pair }(j,\ell).$
-We set the marginal cost to 1 for all pairs for simplicity and use the parameters of $\varrho=1,\nu=4,\gamma=0.5,\rho=10$.
-$\phi^{*}=0.827$ and $\theta^{*}=0.454$.
-$s_{ij}$ and $x_{ij}$ are drawn from a $U[0,1]$ so that all shares within a block sum to 1.
-All Monte-Carlo exercises use 501 random replicas.
-Joint estimation of $\phi$ and $\theta$.
-  1. 1.
-    Candidate markups. For any $(\phi,\theta)\in(0,1)\times(0,1)$ compute the bilateral markup $\mu_{ij}(\phi,\theta)$ from the structural formula (2.6).
-  2. 2.
-    Model-implied gap for exporter $i$: ${\Delta\mu}_{i}(\phi,\theta)\;=\;\ln\mu_{ij}(\phi,\theta)\;-\;\ln\mu_{i\ell}(\phi,\theta).$
-  3. 3.
-    Non-linear least squares criterion.
-    $Q(\phi,\theta)\;=\;\sum_{i=1}^{200}\Bigl{[}\,\Delta p_{i}-{\Delta\mu}_{i}(\phi,\theta)\Bigr{]}^{2}.$
-  4. 4.
-    Estimation. Minimize $Q(\phi,\theta)$ subject to the simple box constraints
-    $0.01\leq\phi\leq 0.99,\qquad 0.01\leq\theta\leq 1$
-    We record the resulting estimates $\hat{\phi},\hat{\theta}$ for each of the 500 Monte-Carlo replicates.
-Results.
-Figure E.2 reports the frequencies for the jointly estimated $(\phi,\theta)$ under the parameters noted above.
-The left panel refers to $\hat{\phi}$ and the right panel to $\hat{\theta}$.
-Similarly, Figure E.3 shows the analogous results for an alternative set of parameters of $\varrho=1,\nu=2.5,\gamma=0.5,\rho=5$.
-Across the figures, the true and average estimates are numerically close, and the distributions of the estimates are centered around the true parameters, showing that our estimator is consistent.
-Figure E.2: Estimated $(\phi,\theta)$ when $\varrho=1,\nu=4,\gamma=0.5,\rho=10$
-Figure E.3: Estimated $(\phi,\theta)$ when $\varrho=1,\nu=2.5,\gamma=0.5,\rho=5$
-E.3 Estimation Under $\theta=1$
-In Section E.3.1, we first demonstrate that assuming $\theta=1$ in the estimation leads to overestimating $\phi$ when the true parameters are $(\phi^{*},\theta^{*})\in(0,1)^{2}$.
-Further, in Section E.3.2, we show that using $\phi^{*}$ (instead of the overestimated $\phi$) in the model validation exercise in Section 4.3 where $\theta$ is imposed to be $1$ would yield an upper bound on the attainable correlation between model-predicted and observed price changes.
-E.3.1 Estimation Bias When Fixing $\theta=1$ in the Estimation
-In what follows, we maintain the following simplification assumptions.
-We assume $Cor(s_{ij},x_{ij})=0$, mirroring the low correlation between the two bilateral shares (Table 1).
-We also impose marginal costs to be constant across firms, implying $p_{ij}=\mu_{ij}$ (can be relaxed by assuming a distribution for $\Delta k_{ij\ell}$).
-Finally, we assume that $\omega(\phi)=\phi$ to keep the notation clean.
-This assumption is without loss since $\omega$ is increasing in $\phi$.
-The objective function when jointly estimating $\phi$ and $\theta$ is:
-$\arg\min_{\phi,\theta}R(\phi,\theta)$ $=\mathbb{E}\big{[}p_{ij}-p_{i\ell}-\mu_{ij}(\phi,\theta)+\mu_{i\ell}(\phi,\theta)\big{]}^{2}$ (E.1)
-$=\mathbb{E}\big{[}\Delta p_{i}-\Delta\mu_{i}(\phi,\theta)\big{]}^{2}$
-where $j$ and $\ell$ are two importers to exporter $i$ and we dropped the $\ln$ sign to simplify the notation.
-Under the full-rank condition, minimization of (E.1) leads to the estimation of the true parameters $(\theta^{*},\phi^{*})$.
-Now define $\tilde{R}(\phi)=R(\phi,\theta)|_{\theta=1}$. When fixing $\theta=1$, the objective function (E.1) becomes:
-$\arg\min_{\phi}\tilde{R}(\phi)=\mathbb{E}\big{[}\Delta p_{i}-\Delta\mu_{i}(\phi;\theta=1)\big{]}^{2}.$ (E.2)
-To study the bias, we can replace $p_{ij}=(1-\phi^{*})\mu^{oligopoly}_{ij}+\phi^{*}\mu^{oligopsony}_{ij}(\theta^{*})$ and rearrange equation (E.2) as follows:
-$\tilde{R}(\phi)=\mathbb{E}\left[(\phi-\phi^{*})\,(\mu^{oligopoly}_{ij}-\mu^{oligopoly}_{i\ell})+\phi^{*}\,(\mu^{oligopsony}_{ij}(\theta^{*})-\mu^{oligopsony}_{i\ell}(\theta^{*}))\right]^{2},$
-where, in a slight abuse of notation, we denote the oligopsony markdown computed at the true $\theta^{*}$ as $\mu^{oligopsony}_{ij}(\theta^{*})$ to distinguish it from the oligopsony markdown when $\theta=1$, which is $\mu^{oligopsony}_{ij}(\theta=1)=1$.
-We can use the price equation to transform this into a function of only prices (data) and $\mu^{oligopsony}(\theta^{*})$:
-$\tilde{R}(\phi)=$ $\mathbb{E}\Bigg{[}(\phi-\phi^{*})\,\left(\frac{p_{ij}-\phi^{*}\mu^{oligopsony}_{ij}(\theta^{*})}{1-\phi^{*}}-\frac{p_{i\ell}-\phi^{*}\mu^{oligopsony}_{i\ell}(\theta^{*})}{1-\phi^{*}}\right)$
-$\hskip 28.45274pt+\phi^{*}\,\left(\mu^{oligopsony}_{ij}(\theta^{*})-\mu^{oligopsony}_{i\ell}(\theta^{*})\right)\Bigg{]}^{2}$
-$=$ $\mathbb{E}\Bigg{[}\frac{\phi-\phi^{*}}{1-\phi^{*}}\,(p_{ij}-p_{i\ell})-\frac{\phi-\phi^{*}}{1-\phi^{*}}\phi^{*}\left(\mu^{oligopsony}_{ij}(\theta^{*})-\mu^{oligopsony}_{i\ell}(\theta^{*})\right)$
-$\hskip 28.45274pt+\phi^{*}\,\left(\mu^{oligopsony}_{ij}(\theta^{*})-\mu^{oligopsony}_{i\ell}(\theta^{*})\right)\Bigg{]}^{2}$
-$=$ 1($\frac{1}{(1-\phi^{*})^{2}}\mathbb{E}\left[(\phi-\phi^{*})\,(p_{ij}-p_{i\ell})+\left(\mu^{oligopsony}_{ij}(\theta^{*})-\mu^{oligopsony}_{i\ell}(\theta^{*})\right)\,\Big{(}(1-\phi)\,\phi^{*}\Big{)}\right]^{2}$
-$=$ 1($\frac{1}{(1-\phi^{*})^{2}}\mathbb{E}\left[(\phi-\phi^{*})\,\Delta p_{i}+\Delta\mu^{oligopsony}_{i}\Big{(}(1-\phi)\,\phi^{*}\Big{)}\right]^{2}.$
-Notice that when $\phi=\phi^{*}$, the first term is zero, but the second term is not. The last equation above can be further rewritten as
-$\tilde{R}(\phi)=$ $\left(\phi-\phi^{*}\right)^{2}\cdot\left(\sigma_{p}^{2}+\overline{\Delta p}^{2}\right)+\big{(}\left(1-\phi\right)\cdot\phi^{*}\big{)}^{2}\cdot\left(\sigma_{\Delta\mu^{oligopsony}}^{2}+\overline{\Delta\mu}_{oligopsony}^{2}\right)$ (E.3)
-$+2\,\left(\phi-\phi^{*}\right)\cdot\left(1-\phi\right)\cdot\phi^{*}\cdot\bigg{(}\operatorname{Corr}\left(\Delta p_{i},\Delta\mu^{oligopsony}_{i}\right)\cdot\sigma_{p}\cdot\sigma_{\Delta\mu^{oligopsony}}\,+\overline{\Delta p}\cdot\overline{\Delta\mu}_{oligopsony}\bigg{)},$
-where we denote the standard deviation and average of variable $z$ by $\sigma_{z}$ and $\overline{z}$ respectively.
-Taking derivative with respect to $\phi$ and setting it equal to zero yields:
-$\phi=\frac{A\cdot\phi^{*}+B\cdot(\phi^{*})^{2}-\phi^{*}\cdot C\cdot(1+\phi^{*})}{A+B\cdot(\phi^{*})^{2}-2\cdot\phi^{*}C}=\phi^{*}\cdot\,\underbrace{\frac{A+B\cdot\phi^{*}-C\cdot(1+\phi^{*})}{A+B\cdot(\phi^{*})^{2}-2\cdot\phi^{*}C}}_{>1}>\phi^{*}$
-where $A=\sigma_{p}^{2}+\overline{\Delta p}^{2}>0,\;B=\sigma_{\mu^{oligopsony}}^{2}+\overline{\Delta\mu}_{oligopsony}^{2}>0,\;C=\operatorname{Corr}\left(\Delta p_{i},\Delta\mu^{oligopsony}_{i}\right)\cdot\sigma_{p}\cdot\sigma_{\mu^{oligopsony}}+\overline{\Delta p}\cdot\overline{\Delta\mu}_{oligopsony}>0$.
-Therefore, the objective (E.2) would estimate $\phi>\phi^{*}$ when setting $\theta=1$.
-This argument can be extended to heterogeneous marginal costs across pairs.
-In that case, we would replace $\Delta\mu_{i}$ for $\Delta p_{i}$.
-Simulation
-We repeat the simulations from Appendix E.2, now estimating only $\phi$ (with $\theta$ fixed at 1).
-We draw $s_{ij}\sim U[0,1]$, injecting enough within-exporter variation to identify $\phi$ from differences in markups.
-If $s_{ij}$ varies very little for exporter $i$, then $\ln\mu(s_{ij},\phi)-\ln\mu(s_{i\ell},\phi)\approx 0,$ so $\phi$ cannot be identified as the markup difference cancels out in the moment (4.1).
-Panel (A) of Figure E.4 maintains 2 importers and exporters in each market, showing a large bias.
-In Panel (B) of Figure E.4, we increase the number of buyers and suppliers to create more cross-sectional variation.
-The figure confirms our argument above: The estimated value of $\phi$ is larger than the true $\phi^{*}$ when the estimation imposes $\theta=1$.
-Figure E.4: Estimation of the parameter $\phi$ while fixing $\theta=1$
-(a) 2 importers and exporters
-(b) 10 importers and exporters
-Notes: In each similuation there are either 2 (Panel a) or 10 (Panel b) buyers and suppliers in each market.
-There are 100 markets.
-501 simulations.
-We set $\varrho=1,\nu=4,\gamma=0.5,\rho=10$ and compute $\phi$ by minimum distance.
-E.3.2 Correlation between Pass-Through and Price Changes
-In the previous section we proved that the estimated $\phi$ is necessarily larger than the true value of $\phi^{*}$, when one imposes $\theta=1$ in the estimation.
-In the exercise of Section 4.3, we do not re-estimate $\phi$ and use $\phi^{*}$ when testing the model under $\theta=1$.
-In this section, we show that the coefficients reported for $\theta-1$ of Figure 3 are upper bounds on the attainable correlations between predicted and observed price changes.
-In what follows, we drop $\ln$ and subscripts to simplify the notation.
-Call the re-estimated $\phi$ when setting $\theta=1$ as $\phi^{R}$.
-Then, combining equations (4.5) and (4.4) in Section 4.3, the pass-through regression for this model is
-$\Delta p\;=\;\beta^{R}\,\cdot\,\Phi^{R}\,\cdot\,\Delta T\,+\,u^{R},$ (E.4)
-where we denoted $\Phi^{R}=\Phi(s,x;\phi^{R},\theta=1)$.
-[^38] We disregard the fixed effects and consider each variable as demeaned for simplicity.
-We also assume that the residual is independent from $\Phi^{R}\,\cdot\,\Delta T$.
-From Appendix E.3.1, setting $\theta=1$ and re-estimating $\phi$, we would find $\phi^{R}>\phi^{*}$, so that
-$\tilde{\Phi}\,=\,\Phi(s,x;\phi^{*},1)\,<\,\Phi(s,x;\phi^{R},1)\,=\,\Phi^{R},$
-since the pass-through increases in $\phi$.
-Using $\tilde{\Phi}$ as an independent variable instead of $\Phi^{R}$ in the OLS regression (E.4),
-$\Delta p=\tilde{\beta}\,\cdot\,\tilde{\Phi}\,\cdot\,\Delta T+\tilde{u}$
-yields the following estimated coefficient:
-$\tilde{\beta}=\frac{\operatorname{Cov}(\tilde{\Phi}\,\Delta T,\;\Delta p)}{\operatorname{Var}(\tilde{\Phi}\,\Delta T)}=\beta^{R}\,\frac{\operatorname{Cov}(\tilde{\Phi}\,\Delta T,\;\Phi^{R}\,\Delta T+u^{R})}{\operatorname{Var}(\tilde{\Phi}\,\Delta T)}.$
-From independence of $\Delta T$ from $\Phi^{R}$ and $\tilde{\Phi}$ we know that
-$\operatorname{Cov}(\tilde{\Phi}\,\Delta T,\;\Phi^{R}\,\Delta T)=\operatorname{Var}(\tilde{\Phi}\,\Delta T)+\operatorname{Cov}(\Phi^{R},\tilde{\Phi})\,\cdot\,\operatorname{Var}(\Delta T),$
-which implies
-$\frac{\operatorname{Cov}(\tilde{\Phi}\,\Delta T,\;\Phi^{R}\,\Delta T)}{\operatorname{Var}(\tilde{\Phi}\,\Delta T)}=1+\frac{\operatorname{Cov}(\Phi^{R},\tilde{\Phi})\,\cdot\,\operatorname{Var}(\Delta T)}{\operatorname{Var}(\tilde{\Phi}\,\Delta T)}>1$
-since $\operatorname{Cov}(\Phi^{R},\tilde{\Phi})>0$ and $\operatorname{Var}(\cdot)>0$. Hence,
-$\tilde{\beta}=\beta^{R}\,\cdot\,\Bigl{(}1+\frac{\operatorname{Cov}(\Phi^{R},\tilde{\Phi})\,\operatorname{Var}(\Delta T)}{\operatorname{Var}(\tilde{\Phi}\,\Delta T)}\Bigr{)}\,>\,\beta^{R}.$
-Thus, since $\tilde{\beta}$, estimated with $\phi^{*}$ and $\theta=1$, exceeds the pass-through coefficient $\beta^{R}$ that would be obtained by re-estimating with $\phi=\phi^{R}$ and $\theta=1$ (i.e.
-$\tilde{\beta}>\beta^{R}$), $\tilde{\beta}$ constitutes an upper bound on the pass-through coefficient in equation (E.4).
-Consequently, the estimates for $\theta-1$ of Figure 3—obtained under $\phi=\phi^{*}$ and $\theta=1$ via the pass-through formula—can be interpreted as upper bounds on the coefficients one would recover by re-estimating with $\phi=\phi^{R}$ at $\theta=1$.
+E.2 Potential Endogeneity from Linkage Formation
+The marginal cost, from equation (2.1), can be written as
+$\ln MC_{iht}=\frac{1-\theta}{\theta}\ln q_{iht}+\ln k_{iht}.$
+As noted in the main text, we generalize $k_{iht}$ to allow for match-specific cost components and denote as $k_{ijht}$.
+This buyer–supplier–time specific term captures heterogeneity in effective marginal cost arising from match-specific productivity, quality differences, or other unobserved relationship-level factors.
+Accordingly, we interpret $\ln(k_{ijht})$ as a structural component of marginal cost, rather than an econometric residual.
+Although we do not explicitly model the determinants of $k_{ijht}$, in estimation we treat it nonparametrically as a buyer–supplier–product–time varying term.
+This approach captures idiosyncratic price differences not explained by bilateral markups and mitigates concerns related to selection into matches.
+Here we discuss how selection operating through $k_{ijht}$ could affect identification of the structural parameters. We model the value of a potential match between exporter $i$ and importer $j$ as
+$V(X_{ijht})=f(X_{ijht})+\Xi_{ijht},$
+where $X_{ijht}$ denotes observed match characteristics, $f(\cdot)$ is a deterministic function, and $\Xi_{ijht}$ captures unobserved, idiosyncratic variation.
+A match is observed if and only if $V(X_{ijht})>0$, which implies $\Xi_{ijht}>-f(X_{ijht})$.
+Hence, the observed transaction $(q_{ijht},p_{ijht})$ satisfies
+$q^{*}_{ijht}=\begin{cases}0,&\text{if }\Xi_{ijt}\leq-f(X_{ijt}),\\ q_{ijh}(p_{ijht}),&\text{otherwise},\end{cases}$ (E.1)
+where $q_{ijh}(\cdot)$ corresponds to the demand function defined in equation (2.6). When $\Xi_{ijht}\leq-f(X_{ijht})$, no match is formed, and bilateral prices and quantities are not observed.
+This framework clarifies how selection may bias estimation.
+If the unobserved component $\Xi_{ijht}$ is correlated with key determinants of bilateral markups ($s_{ijht}$ and $x_{ijht}$) then selection induces endogeneity by generating correlation between the regression error and the regressors.
+In that case, estimates of $\phi$ and $\theta$ may be biased, with the sign and magnitude of the bias depending on the underlying correlation structure.
+Even if $\Xi_{ijht}$ is uncorrelated with the regressors in the population, selection based on $\Xi_{ijht}$ can induce correlation between the regression error and the regressors whenever $\Xi_{ijht}$ is correlated with the selection rule $k_{ijht}$, a standard truncation problem.
+The absence of bias therefore requires the stronger condition, that is, selection must be random conditional on observables.
+In practice, if the unobserved match surplus $\Xi_{ijht}$ primarily reflects exporter-specific fundamentals or shocks that are stable across a supplier’s buyers, differencing across buyers within the same exporter mitigates this source of endogeneity.
+Our estimation strategy exploits cross-sectional variation in prices across buyers of the same supplier and is designed to reduce, though not eliminate, concerns about selection bias.
+When $\text{Cor}(k_{ijht},\Xi_{ijht})\neq 0$, and $k_{ijht}$ is also correlated with the bilateral market shares that determine the markup, the resulting bias can be positive or negative depending on the sign and strength of these correlations.
+However, if the unobserved match surplus $\Xi_{ijht}$ reflects exporter-specific fundamentals or shocks that are relatively stable across a supplier’s buyers, then differencing across matches within the same exporter helps mitigate this source of endogeneity.
+Our estimation strategy, which leverages cross-sectional variation in prices across buyers of the same supplier, is therefore designed to partially address this endogeneity concern.
+We further address the potential endogeneity of $k_{ijht}$ using instrumental variables.
+Specifically, we instrument for bilateral shares using the number of importers and exporters active in product $h$, which proxy for the broader population of potential foreign suppliers and U.S.
+buyers.
+These variables are plausibly exogenous to the match-specific surplus $\Xi_{ijht}$, and are not determined by the selection equation (E.1) above.
+The exclusion restriction is that $\Xi_{ijht}$ is uncorrelated with the overall size of the trading population in product $h$.
+Intuitively, a larger number of potential partners shifts the distribution of the order statistics of $\Xi_{ijht}$, influencing the likelihood of match formation without directly affecting whether a given match exceeds the threshold $f(X_{ijht})$.
+We acknowledge that alternative network formation processes could give rise to different endogeneity concerns.
+For instance, in dynamic formation games, $k_{ijht}$ may correlate with the expected continuation value of future matches, which would introduce additional selection channels.
+In such cases, the direction of the bias on $\phi$ and $\theta$ is harder to characterize.
+Nonetheless, the static selection framework outlined above provides a useful benchmark that aligns with our theoretical setup and clarifies the potential sources of endogeneity relevant for our identification strategy.